A Complete Bibliography of the International Journal of Number Theory

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

24 June 2022 Version 1.03

Title word cross-reference

(2, 2, 2) [AZTM15]. (2am − 1)^{x} + (2m)^{y} = (2am + 1)^{z} [MT12]. (3, 3) [ATT^{+}16]. (9, 3) [AITA19]. (A.2) [WY20]. (modr) [HL11]. (ℓ ≠ p) [Pal14]. (k, l) [Baj09]. (n − 1, 1) [RZ22]. (ϕ, Γ) [Ber11]. (q^{b}, q^{p} − b; q^{p})_{∞}^{3}(q^{jb}, q^{2p−jb}; q^{2p})_{∞} [MZ22]. (w + x + y + z)(1/w + 1/x + 1/y + 1/z) = n [BX18]. (x − d)^{4} + x^{4} + (x + d)^{4} = y^{n} [Zha17d]. (x_{1} + ⋯ + x_{n})^{2} = ax_{1} ⋯ x_{n} [Bao08]. −23 [GY13]. 0 [Gri11, Kaa11]. 1 [AK14, Ala14, AK15b, AK17, CWR16, Dub14, Fu11, HM06, Li14, Mem20a, Mor08, RSS18, SS10, Tou09]. 1/2 [Kan22]. 1/a + 1/b + 1/c + 1/d = 1 [BU11a]. 1/π [Gui21, Liu12b]. 1/π^{k} [Gui20]. 10 [PSY18]. 11 [AB15b, Ber07, Jam12, JZ17, Pen19]. 12 [KE17, FY17, PSY18]. 13 [JZ17]. 14 [NV10]. 16 [GJS14, SS10, Xio16]. 2 [AK14, AAA15, AK17, AZTM15, ARTZ19, Bro10a, CV17, CK20, Cop20, Cre18, CG19b, GK18a, Gla09, HZZ12, INST14, Ito18, KM12b, Kum13, LW19, LR19b, LL11, Liu13b, Mat08, Mem17, MY21, Mol12, OLG19, OT05, PSY17, QH19, RSS18, RT17, RSW14, Roz18, Sch12, SW06, Tay17, Ula19, ZH16a]. 23 [BD15a]. 24 [UW14]. 27/64 [CC21a]. 2Z [Tsu15]. 2x^{2} + 3y^{2} + 6z^{2} + 6t^{2} [AW12]. 3 [AK14, Ala14, AK15b, AK17, AITA19, Awt12, CM16, DL20, FN14, Gol16, JZ17, Li14, Lin13b, MM13a, Mor16, Pac19, RSS18, Roe14, Wan16, XY14]. 32 [GJS14, SS10]. 4 [Dai14, Gir16a, HR11, HJ14, JO20, Lin14, RSS18]. 4Z [Tsu15]. 4n^{2}p = x^{2} + qy^{2} [Sun15b]. 5 [AB15b, Ber07, CGM15, Han17, JZ17, LMW16, Mat19, NH17, RS11, Sak14]. 5(mod24) [KK12]. $520 [RS13d]. 5^{A} [Rob11]. 5x^{2} [KK15]. 6 [AK14, AK17, Hua17, RSS18]. 7 [AB15b, BS12, BD15a, Ber07, Dai16b, Jam12, Lin14, Mah19, RS11, Sak14]. 8 [BCH08, JO20, Mil18, PSY18, XY11]. 9 [Ala14, AK15b, BS12, DW16, Par18b, She17, XY14, Yao15]. _{1}ψ_{1} [Vil18b]. _{2}ψ_{2} [BSM16]. 3F_{2} [McC10, WW18]. _{3}F_{2}(1) [Tau18]. _{m} [GR14]. _{p} [Hyo15]. a [Kim16b]. A^{4} + 2^{δ}B^{2} = C^{n} [BEN10]. A_{0} [Val14]. A_{4} [RS14b]. ax^{2} + by^{2} + cz^{2} + dxy = 0 [LR19a]. ax^{y} + by^{z} + cz^{x} = 0 [ZY12]. B [BD21, CDHS15, DE21]. ā + b̄ ≡ c̄ mod p [Cha21]. Q̄^{∗} [Pot21]. β [LLZ18, TYZ16, ZLL16]. β_{n} = ∑_{k=0}^{n} \binom{n}{k}^{2}\binom{n+k}{k} [CMS20]. \binom{3k}{k} [Sun16a]. \binom{n−i−1}{i−1} [She07]. mod [Kön20, Tsu15]. mod3 [Lin13b]. mod4 [IIO20]. modZ [Gir14b]. modn [SC16]. modp [AM09, Amo21, BP11, Nţ21, Álv14]. c [BB21b]. C_{2}^{r} ⊕ C_{n} [FZ16]. C_{p} ⊕ C_{p^{n}} [QH17]. χ [Edd16]. χ(2) = 1 [Edd16]. Co_{2} [KRLT20]. D [HK18, AsMS20, Baš12]. d(n!)/m! [BFLT10]. D_{1}x^{2} − D_{2}y^{2} = λk^{z} [YF15]. D_{m}^{+} [KM06]. Δ [Bai16]. Δ(x) [Li17, ZZ11]. Δx [TZ09]. e [Has13b]. E(Q)_{tor} ≅ Z/3Z [BKY19]. E(x) [TZ09]. ℓ [JZ17, Pen08]. EQ_{tor} ≅ Z/2Z ⊕ Z/2Z, Z/2Z ⊕ Z/4Z [BKY20]. ϵ(k) = c(k + 1) [ZMS15]. ϵ(k_{n}) = ck_{n}^{τ} + ck_{n}^{τ−1} [TZ17]. η [CC16]. f(n) [GV13]. f(x) [GV13]. f(x)x^{n} + g(x) [DFV13]. F_{n} = P (x) [TU20]. F_{n} ± p^{a} [Šiu16]. F_{p} = u^{2} + pv^{2} [GBL15]. \frac{520}{π} [RS13d]. \frac{L′}{L} (1, χ_{D}) [MM13b]. \frac{F_{p}[x]}{⟨f(x)⟩} [WZWQ16]. G

[Kön20, Nat11, ST19]. Gal(k_{3}^{(2)}|k) [ATT^{+}16]. Gal(Q_{p}/Q_{p}) [Roz18]. Γ [BS10a, Bar13a]. Γ^{3} \ H [SS07a]. Γ^{∗}(10, p) [VR20]. Γ_{0}(2) [Got20, HY18]. Γ_{0}(4) [Moo19]. Γ_{0}(N) [Gil17]. Γ_{0}(p) [Coo09]. Γ_{0}(pN) [BP11]. Γ_{1}(q)

[Dja11]. Γ_{1}(t) [BV18]. Γ \ H [SS07b]. gcd [Abo08, KE18]. GL(2) [Ass21, JK21]. GL(3) [Agg21]. GL(3) × GL(2) [Str22]. GL(n) [Wal21]. h [BC18, ST19]. i [She07]. j [Sha14, XLD22]. j(z) [DM13]. K [Kit13, BK11, BL13, CG13, DL20, FGT15, HLT20, Hu13a, Kei13, Kei21, lLW21, Liu09, xMgC22, MW16, Mol12, MS21, QGX22, RS13a, RB18, Rib11, RL18, SD20, Tre15b]. K3 [Ito18, Tre15a]. k = Q(\sqrt{−3}, \sqrt{d}) [ATT^{+}16]. L [Aga10, Agg21, And12, Ass21, AMPS17, Aym22, BT18b, Bou11, Bro10b, Bru18, BM11, Bui12, Bui13, CY18, Cho13a, CK17, Dah18, DM19, Dja11, Dja13, – DLV20, Dum09, FW10, FRcT20, FG22, FJ20, GK13, Gil13, GY16, HRL11, Hah21, HM16, Hay14, Hou21, Hun18, KP14, KMT11, Kug22, Lan19, Lei12, LLZ16, LZ14, MNZ19, MN13, MU18, MSV18, Mis16, MY13, Mui12, NMZJ22, OS17, OO17, PT14, łPS13, Pat19a, Pat10, Pey20, Pi20, RGHK20, San09, Sas15, SSS22, Str22, SY19b, Suz05, Tam14, Tan12, TX16, Tem10, Van12b, Van16a, Yam15, You12, BARCVS13, HS21a, She16, Wit07, ZH16c]. L(1, χ) [Jun14]. Λ [She14, BS10a, Bar13a, Kum21]. lcm(a, b, c) = 7, 8 [Par18b]. lcm(a_{1}, a_{2}, a_{3}, a_{4}) ≤ 4 [LP17]. ln(1 + x^{n}) [AKMR12]. log [Kas19]. m [MS16b, Toh08]. M(2, K) [Tu11]. m^{2} + n^{2}x + 2mn^{y} = m + n^{2z} [hY20]. A^{n} [AS09]. F_{1} [LLM18]. F_{p}[x] [Tho13]. F_{q} [BK13, Tür11]. F_{q}(t) [She22]. F_{q}[T] [AC15, Yam16, RGHK20]. F_{q}[x] [WHZ19]. F_{ζ} [LM15b]. P^{1} [GL19, GG22]. P^{1} − {0, 1, ∞} [Tan18]. P_{Q}^{1} [Haj15]. Q [BCF16, BCF21, BF19, FK11, PZ18, Par11, SG17, Kit13]. Q(\sqrt{−4pq}) [Mil18]. Q(\sqrt{−8pq}) [Mil18]. Q(\sqrt{−p}) [XLD22]. Q(\sqrt{8pq}) [Mil18]. Q( ^{3} \sqrt{d}, ζ_{3}) [AITA19]. Q(\sqrt{pq}) [Kum21]. Q(\sqrt{pq}, \sqrt{2 + \sqrt{2}}) [Kum21]. Q_{p} [BP15b]. R_{an,exp} [JQ21]. Z [BGW12, Ot20, Shl12, Tsu15]. Z/2Z ⊕ Z/6Z [BKY20]. Z[i] [BB18a]. Z[x] [GV13]. Z_{2} [FK11, Mat17a, Kum21]. Z_{m} × Z_{n} [NT14]. Z_{n} [TW19]. Z_{p} [IMO13, Mat18, Kit13, Mem20a]. Z_{p}^{2} [LZ14]. Z_{p}^{n} [BS10a, Bar13a]. Z_{p}^{r}(1) [Cob21]. M_{H}(G) [Lim15b]. S_{n} [Lou16]. GCF_{ϵ} [TZ17, ZMS15]. GL(2) × GL(3) [Pi20]. GL(3) [AP17, Gue15, SY19b]. GL(3) × GL(2) [CY18]. Gl(3, E) [FZ12]. GL(3, Z) [AD16b]. Gl(4) [FZ06]. GL(n, R) [Gol07]. GL_{2} [Wu17, Kid16]. GSp(4) [Fli11]. GSp_{4} [You12]. GSp_{4} × GL_{2} [Bro10b]. GU(2, 2) [Cau20]. PGL_{2}(F_{ℓ}) [Rob18]. pod [RS11]. SL(2, Z) [Wei15a]. SL(3, r) [Ehr09]. SL_{2}(F) [GS16]. SL_{2}(K) [AMPS17]. SL_{2}(F_{3}) [DR16]. SL_{2}(Q(\sqrt{−D})) [Pet16]. SL_{2}(R) [Zem17]. SL_{3}(Z) [AP08]. SO(2, 3) [GK18a]. SO(4) [Fli11]. Sp(2n) [Zha17c]. Sp_{2}(F) [Zor11]. spt [GJS14]. U(1, 1) [Hof13, Zha17b]. U(n, n) [Zha17c]. μ [Ari13]. N [Liu14, GC17, HLS11, Kim16b, Len19, MY21, PQSW14, San17, Sin09, Sun19a, WS16, WS17]. n(n + d) ⋯ (n + (k − 1)d) = by^{2} [FLS12]. n^{2} [HLS11]. n ≥ 1 [Bro10a]. Ω [SZZ18]. ω(q) [Gar08]. {ωn} [SK13]. P [Liu19a, Ade18, Álv14, BS10a, Bar13a, BS15, Ber11, Bud20, BV11, CF22, CR18, Dor20, DA19, Edi05, FT15, FJ20, Gar18b, Gir18a, GS16, Gra18, HP11, HL12, HK14, Hir22, Hun18, INST14, Jar22, Jen05, Kam08, Kas19, Ked05, Kid16, KH11, KT20, Kön20, LMS10, Lei12, Len14, Len17b, LZ14, LR06, May12, May14, McC12, Mem20b, Miz08, Mor22, MOS14b, N18, OLG19, Pot18, Pri09, Pum20, Sal09, SA09, Sch14, Som22, Tak08, Tho06, Van12b, Van16a, Vie10, You16b, You17]. P (b) ≤ Ck [FLS12]. p(n) [Nic06]. p(n, k) [Kar19]. p^{2} [Che22a, PT14]. p^{α}q^{β} [CSJ17]. p_{k}(n) [DLZ19]. p ≡ 3(mod8) [Ska17a]. p ≡ 3 mod 4 [WY20]. p ≡ 3 (mod 4) [Guo21]. ϕ [VB19]. ϕ(n) [Sin09]. ϕ(n)/n [Tou09, Tou09]. ϕ(X^{m} − 1) = X^{n} − 1 [FL15]. π [ABCM14, PS17]. π(x) > li(x) [CP10a]. π^{±2} [CC21a]. ±1 [BSK17]. pq [EG12, WZ12]. ψ [SW08]. q [Alo19, ACS09, ES20, BW17, BMW06, Bun08, BV09, BV11, CP08, CC21a, DA19, Gor19, Guo15, GL17b, Guo19b, GN19, Guo19a, Guo21, HM16, JM10, Kam08, KP16, hKS21, KKL21, LLMA16, LP20, Liu13a, Liu17a, Mad12, MAM06, Mer11a, Miz08, NP18, Sau15, SSU21, Sch18a, Tan19, WC19, WY20, ZL18b]. Q(\sqrt{−5}) [Vul06]. q^{[p/8]}(modp) [Sun15b]. r [Gib14, Liu18, SD20, Tol06]. r_{s}(n) [CL11b, CL13]. Res_{x}(P(x), x^{2} + sx + t) = a [AGL18]. RΓ(N, T) [Cob21]. ρ, q [DA18]. S [GR11, GM19, Sch21b, SZ15, Zie11]. s(a, b) [CDHS15]. s s + 1 [NY21]. s = −1 [San09]. S = {2, q} [SZ15]. S_{3} [TT14]. S_{k}(Γ_{1}(4)) [FY13]. S_{n}(t) [CQh22]. σ(n!)/m! [BFLT10]. σ(n) ≡ a(modn) [APP13]. ∑_{2l+5m=n} σ(l)σ(m) [CY14]. ∑_{3i+j=n} σ(i)σ_{3}(j) [YX14]. ∑_{3l+5m=n} σ(l)σ(m) [RS13b]. ∑_{4l+5m=n} σ(l)σ(m) [CY14]. ∑_{4l+9m=n} σ(l)σ(m) [Ye15a]. ∑_{ak+bl+cm=n} σ(k)σ(l)σ(m) [Par18b]. ∑_{i+25j=n} σ(i)σ(j) [XTY14]. ∑_{i+3j=n} σ(i)σ_{3}(j) [YX14]. ∑_{l+15m=n} σ(l)σ(m) [RS13b]. ∑_{l+20m=n} σ(l)σ(m) [CY14]. ∑_{l+27m=n} σ(l)σ(m) [AK16]. ∑_{l+32m=n} σ(l)σ(m) [AK16]. ∑_{l+36m=n} σ(l)σ(m) [Ye15a]. ∑_{m<n/9} σ(m)σ(n − 9m) [Wil05]. ∑_{n≤x} f(n){x/n}^{k} [WS20b]. ∑ a_{1}m_{1} + a_{2}m_{2} + a_{3}m_{3} + a_{4}m_{4} = nσ(m_{1})σ(m_{2})σ(m_{3})σ(m_{4}) [LP17]. t [Agg21, BN14]. θ [Mok20]. U [Alk15]. u(u + a)(u + 2a) = v(v + 1) [LL17]. u, v [HMST16]. U_{t} [BV18]. v [Fra21, She22]. φ(n!)/m! [BFLT10]. φ(n) = φ(n + k) [Kim21b]. φ(p − 1) = φ(q − 1) [Kim21b]. ∥(4/3)^{k}∥ [Pup15]. S̃p(2n) [Zha17c]. S̃p_{2}(F) [Zor11]. X [EGL21, CP10a, Gou18]. X(4) [JKM09]. X(4p) [JKM09]. x^{2} + 2^{a} · 5^{b} = y^{n} [LT08]. x^{2} + 2y^{2} + 2z^{2} + 6t^{2} [AW12]. x^{2} + 3y^{2} + 3z^{2} + 6t^{2} [AW12]. x^{2} + C = 2y^{n} [MLST09]. x^{2} + p^{a}q^{b} = y^{q} [GN21]. x^{2} + xy + 7y^{2} + z^{2} + zt + 7t^{2} [Ye16]. x^{2} + xy + y^{2} + z^{2} + zt + t^{2} [Cha08]. x^{2} + y^{2} + 2z^{2} + 3t^{2} [AW12]. x^{2} + y^{2n} = z^{3} [Dah11]. x^{2} − (a^{2} − 1)y^{2} = 1 [YF19]. x^{2} − ay^{2} = 1 [FY21]. x^{2} − dy^{2} [Din09]. x^{3} + y^{3} + z^{3} = q [Huo18]. x^{3} + y^{3} = 2 [Jed14]. x^{6} + Ax + B [BS10b]. x^{a} ± y^{b} ± z^{c} ± w^{d} = 0 [BU11a]. x^{d} + ax + b [BK13]. x^{d} + c [Pan22]. x^{n} + cx^{n−1} + d [Har12]. x^{n} − 1 [Tho13]. X^{n} − 1 = BZ^{n} [BM17]. x^{p^{r}} − m [BE21]. x^{pq^{b}} − 1 [Wan15b]. x^{y} + y^{x} = z^{2} [YF18]. x_{1}x_{2} + x_{2}x_{3} + x_{3}x_{4} + x_{4}x_{1} = n [DT22]. y^{2} − bz^{2} = v_{1}^{2} [FY21]. y^{2} − pz^{2} = 1 [YF19]. y^{2} = x^{3} + Ax [Dra11, VY13]. y^{2} = x^{5} + ax [Jed22]. Z [Yas16]. |L(1, χ)| [Edd16].

-additive [Mad12]. -Adic [Awt12, HZZ12, Kum13, Mat08, OT05, QH19, Ula19, Ade18, BS10a, Bar13a, BS15, Ber11, Bud20, BV11, CF22, Cop20, Dor20, DA19, Edi05, Fra21, FJ20, Gir18a, GS16, Gra18, HP11, HL12, HK14, Hir22, Hun18, INST14, Jar22, Jen05, Kam08, Kas19, Ked05, Kid16, KH11, KM12b, KT20, LMS10, Lei12, Len14, Len17b, LZ14, LR06, McC12, Mem17, Miz08, Mor22, Mor16, MOS14b, Nat11, PS17, Pot18, Pum20, Roe14, SA09, She22, Som22, ST19, Tak08, Van12b, Van16a, Vie10, You16b, You17, She14]. -analogue [Guo19b, Guo21, JM10, Liu17a, Mer11a, WY20]. -Analogues [CC21a, CP08, HM16, LP20]. -ary [DE21, Sau15]. -aspect [Agg21]. -automorphic [Dja11]. -Bernoulli [Kam08, N18]. -binomial [GL17b, KP16, SSU21]. -bracket [Sch18a]. -Calkin [HMST16]. -class [AITA19, AZTM15, ARTZ19, May12, ZH16c]. -coefficients [SW08]. -color [Kei21]. -Colored [Lin13b, Lin14, xMgC22]. -congruence [GN19, Guo19a, NP18]. -congruences [Gor19]. -congruent [Mok20]. -conjecture [Lim15b]. -continued [BMW06]. -coordinates [EGL21]. -Core [BS12, Gol16, NY21]. -cores [BN14, Wan16]. -curves [BF19]. -cyclotomic [INST14]. -décomposé [Tho06]. -decomposed [Tho06]. -degree [Gar18b]. -derivative [Liu13a]. -descent [Cre18]. -Diamond [Fu11, Mor08, AB15b, DW16, Dai16b, LMW16]. -digit [Baš12]. -digital [Alo19]. -Dimensional [Sch12, Len19, Roz18, ZH16a]. -Diophantine [SZ15]. -dissection [XY11]. -Euler [DA18]. -expansion [LLZ18]. -expansions [TYZ16, ZLL16]. -exponential [LLMA16]. -extension [FK11, Kum21, Mat17a]. -extensions [BARCVS13, Sch14, DR16, Kit13, Kön20, LZ14, Mat18]. -Fermat [Guo15]. -Fibonacci [BL13, RL18]. -finite [HK18]. -fold [BC18]. -free [FGT15, Gib14, Kei13, Liu18]. -function [CC16, GK13, GJS14, Hay14, FG22, LLZ16, OS17, You12]. -functions [Van16a, And12, Bou11, BM11, Bui12, Bui13, Cho13a, Dja13, Dum09, FW10, Gil13, HRL11, KP14, KMT11, MN13, MY13, Mui12, łPS13, Pat10, Suz05, Tam14, Tan12, Tem10, Agg21, Ass21, BT18b, Bro10b, CY18, CK17, Dah18, DM19, Dja11, – DLV20, FJ20, GY16, Hah21, HM16, Hou21, Hun18, Lan19, Lei12, LZ14, MNZ19, MU18, MSV18, NMZJ22, OO17, Pey20, Pi20, RGHK20, San09, SSS22, Str22, SY19b, TX16, Van12b]. -generalized [DL20]. -geometry [LM15b]. -gonal [JO20]. -groups [OLG19, Sal09, Kit13]. -hypergeometric [ZL18b]. -identities [hKS21]. -integral [GR11]. -integrals [DA19]. -invariant [Kum21, Sha14, BV18]. -invariants [Ari13, BS10a, Bar13a, XLD22]. -Isogenies [MM13a]. -Lattices [KM06]. -Lehmer [MW16]. -Lipschitz [Mem20a]. -marked [BK11]. -modules [Ber11]. -numbers [Alk15]. -operator [BV18]. -orders [Lou16].

-ordinarity [FT15]. -packets [AMPS17, FRcT20, Mis16]. -pairs [BD21]. -parts [Wit07]. -periods [Fli11]. -power [Mol12]. -powers [Mol12]. -prime [SD20]. -primitive [CK20]. -Rank [HR11, Álv14, Mil18, Pri09]. -Ranks [Gla09, IMO13]. -Regular [Pen19, XY14, BD15a, HS21a, JZ17, NH17, Pen08, RB18, She16]. -restricted [GM19]. -result [SZZ18]. -ring [May14]. -Selmer [FN14, Li14]. -series [ACS09, BW17, Bun08, BV09, BV11, KKL21, Liu13a, MAM06, Tan19, PT14, Pat19a, WW18]. -sextic [TT14]. -Sidon [Pac19]. -sum [KE18]. -sum-free [Baj09]. -supercongruences [pFG22]. -th [AsMS20, FY17, lLW21, San17, Toh08, Tre15b]. -torsion [Pri09]. -transformation [WC19]. -transforms [BS10a, Bar13a]. -trigonometric [ES20]. -tuple [Dai14]. -tuples [Liu09, MS16b, Rib11, Tol06]. -Unit [Zie11, Sch21b]. -Value [Aga10]. -Values [Aga10, Bru18, Kug22, Sas15, Yam15]. -Volkenborn [DA18]. -Wieferich [BB21b]. -wise [Hu13a, SD20]. -zero-sum [PQSW14].

1 [Par20].

2-adic [Ade21]. 2-Class [CeM21]. 2-part [Kum21]. 2-Salem [BK22].

=th [QGX22].

A-motives [Fra21]. ABC [Pas15]. Abel [Ada12, BP18]. Abelian [SS10, Baj09, BGm22, Bru18, Cop20, Gam14, Gil17, Her16, Hol19, Kim16a, Nom14, Orr17, Pal14, PHLS19, Pol14a, Sca17, Sug15, Tho06, Bal08, Bos09, Bou11, Cul12, FP10, Lev06, LMP10, Mam10, Mei18, Paz13, Per12, Pic10, RT17, Via10, Vol10]. abelien [DP08, DP08, Tho06]. abéliennes [Gil17, SS10]. absolute [Fla19]. absolutely [Sah16]. Abundant [Kob14, Kob16].

Acknowledgment [GHK^{+}15a, HKN11]. Action [Col12, Wal06, Wal08, Cop20]. Actions [Kab10, KN09a]. Adaptation [Mok20]. addition [SC16]. Additive [Ell15, GHHP06, Koc08, Kum09, LMP10, Par20, QH17, Roy10, Vau15, WC11, Zie11, BG15b, hF22, KS16, Mad12, Nat11, QH16, Zhu18]. Additively [LL18]. Adelic [Ass21, Mag13]. Adic [Awt12, HZZ12, Kum13, Mat08, OT05, QH19, Ula19, Ade18, Ade21, BS10a, Bar13a, BS15, Ber11, Bud20, BV11, CF22, Cop20, Dor20, DA19, Edi05, Fra21, FJ20, Gir18a, GS16, Gra18, HP11, HL12, HK14, Hir22, Hun18, INST14, Jar22, Jen05, Kam08, Kas19, Ked05, Kid16, KH11, KM12b, KT20, LMS10, Lei12, Len14, Len17b, LZ14, LR06, McC12, Mem17, Miz08, Mor22, Mor16, MOS14b, Nat11, PS17, Pot18, Pum20, Roe14, SA09, She14, She22, Som22, ST19, Tak08, Van12b, Van16a, Vie10, You16b, You17]. Affine [Kab10, Alv21, Gho11, Sar22]. after [cC21b]. Ahlgren [TB06]. Airy [HRL11]. Akashi [Lim15b, Lim16]. Alex [CCS13]. Algebra [Li13, GR14, Hyo15]. Algebraic [ABP09, BZ15, BK12, Bun12, BO19, GH14, GMR11, Gur11, Höh11, JP14, Mah12, PS11, Roy05, SSD11, TT18a, Tan18, Vää16, Veg11, Abo08, BK22, Cha20, Eld19, Fla19, JK16, Kal18, Kha19, Kid16, KV19, Mor16, mPP22, PS19, Pot18, Tre15a, WJ20]. Algebras [BFCC09, KL12, Ter13, Amo21, AB18, GK18a, Hou17, LS17b, PP18]. Algorithm [AS06, CS10, Sch12, Roz18]. algorithms [Wei15b]. Aliquot [BBS08, Par15a]. Almost [Che18c, MT17, Wu18, Cha15, Cha18b, Kim22a, LWC19, RY16, ZL18a]. almost-primes [LWC19]. also [VB19]. Alternating [GZ11, Sof18, Gir16a, HP11]. alternative [CC16]. Amicable [LP16, Par15a]. Among [Moy13, Cha18a, JL20, Klu16, Zhu22]. Ample [FP10, San18b]. analog [AMM17, AMSV21, Cha10a]. Analogs [DWW09]. Analogue [DS07, Dic21, Guo19b, Guo21, JL20, JM10, Kua15, Liu17a, Mer11a, Ska17a, Wan20, WY20]. Analogues [ASD16, BBCZ05, CK14, CC21a, Dix11, UW14, WY21, CP08, DE21, EMS21, HM16, LP20, You17]. Analysis [BH10, SS21, Ade18, Ade21, CWW15, LMW16, Wu17]. Analytic [CL11a, JVW20, Mui12, Tay09, Zor13, BO19, Kid16]. Analytical [Mah19]. Anderson [MP22a]. Andrews [AB15a, CHS15, Che22c, Gar10, GJS14, MMO08, NS18]. Andrzej [US13]. angles [BPZ14]. Ankeny [IN22, YF18]. Annihilating [Rüh10]. anti [Cla08]. anti-Hasse [Cla08]. anticyclotomic [Hun18, KPW17, Mat18]. Any [Bra14, PV12, WY20]. Apagodu [Liu17a]. Apéry [CMS20, CCS10, GZ12, PPT12]. Apéry-Like [CCS10, PPT12]. appearance [LP14]. Appell [CM12, FWX21, HN08, LLM18, TSB20]. Appendix [CCS13, US13]. Application [Bal08, CL11a, Kab10, Sak11, BV11, BŻ19, Jar22, RS13b]. Applications [Bou05, Cao11, Cul12, KTT06, Lov05, OT05, Riv09, Won18, AjW21, BSM16, CeM21, DA19, Gor19, He20, LM15a, LZ15, MS16a, OS21, lWfS22, ZH16b]. Approach [Mol05, Toh13, BP17, Cho17, Dah11, JMV16, Kol15, Mah19, SR19]. approximable [BLW21, BK15b]. Approximate [EH08]. approximating [MP16a, MP16b]. Approximation [Gir14a, KMW10, Roy05, Vul10a, Adi15, Bud20, BP15b, GL18, HYZ17, KM22, LLMA16, Leo15, Liu17b, Mar14, MQ18, NRS20, Poë20, Ram17, Vul06, Yam16, You16a]. Approximations [Mer11b, Phi11, Boc08, Vää15, Vää18]. Arakawa [You16b]. Arakelov [Hou17, Tra17]. Arakelov-modular [Hou17]. Arbitrary [Buc11, Mau17, Moh19, Wal17a]. Archimedean [Tu11]. arcs [Li16]. arctangents [ABCM14]. Area [BG06]. areas [Cho20b]. argument [Pey20].

Arguments

[Pil09, EL18, Hof17, NJ22, Pas15]. Arising [Ngu11, ACS09, HK20, PSZ16, SVY20, Tsu18]. Arithmetic [Ada12, AG13, Boy10, CT13, Che14, CHS15, El 11, El 12, Fre12, Has13a, HHM21, Jah10, KK10, Kna08, KN09b, MN13, MC13, MP10, Min12, Mis17, MS09a, NH17, Oh11, OT05, Pap11, Pen08, PT15, SSD11, Şen12, She16, SZZ13, Ula12, Wan15a, Wan16, Yao15, Adi15, AAAW17, BN14, BF10, Bor18, BS16b, BS19, CHL19, CGZ21, CL22, DGM19, EK20, ETT17, Gib14, Got20, GR14, Haj15, HKL^{+}21, Has13b, HLN19, KM12b, Kou15, KP18, MT21, Pas15, RSS20, Rei21, RGHK20, San14, San15, SS15b, Suz17, Szc15, Van21b, Vil18a, ZJ19, Zhe15, Zho21, Zoe19, Riv09]. Arithmetical [AZZ05, Bun08, BV09, Kab10, Sch10, CTZ16, Hau18, Sah16]. arithmétiques [BF10, Riv09]. Artin [IN22, YF18, CK17, FM12, Ren19, Var22, Won18, WS10]. ary [DE21, Sau15]. Asai [Gro20, JK21]. Aspect [GGW11, Agg21, Ass21, Sun15a]. aspects [Dic15, Hit18]. Associated [LT14, Our09, Wil10, Agn22, Bor22, CJKM18, Che18a, DA18, DA19, Eve22, HL11, Hsu20, Hyo15, IIO20, Kas19, hKS21, LW20, MP15, RGHK20, SA09, Tay17, Van12b, Van16a]. associées [Tay17]. Asymptotic [BW17, BP15a, Bel09, FF15, Fol21, GK18b, Has13a, HL11, Kan13, KO20, KN09a, Tem10, Tou19, Xu19a, BP18, BB20c, CJKM18, DM13, Lin18, MNZ19, Mol12, Sed15, ST19, Wei20, WS20b]. Asymptotically [Has10, IFD08]. Asymptotics [Cio20, HM12, KK19, Mao14, Nob12, O’S16, Wal13]. Atkin [BV18, DWW09, Pat15, Sak11, Sak14]. atoms [SY14]. Attached [KM06, Tan12, AD16b, BT18b, Las17, LLZ16, Mil13, Pat19a]. Author [Ano05, Ano06, Ano07a, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13, Ano14, Ano15a, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21]. Automatic [ZZ21]. automaticity [HWh20]. Automorphic [BKK06, Bui13, Gil13, AB18, AP08, Dau14, Dja11, DS21, GY16, GR14, JL20, LLZ16, TX16, Wei15a, Wu17]. Automorphisms [Gla09, Neb13, MP22b, Zem17]. aux [JM10]. Average [Alk07, DG13, Kim16a, Dra17, Kim16b, NH20, NT14, Par15a, Peh16, Zho18]. Averages [BS19, FW10, CGZ21, CWW15, KE18]. avoiding [EK20]. away [RT17].

bêta [HM06]. bêta-développement [HM06]. Bad [Yas13, Rob15]. badly [BK15b]. Báez [BD10]. Bailey [BSM16, CZ10, ZH16b]. Balanced [Bro12, NQ08]. Balancing [KS14]. Barban [Smi11]. Barker [BE13]. Barnes [BB14a, JMV16]. Base [BD08, CF09, PV12]. based [Meš16]. Bases [FY13, GHHP06, LMP10, MS12b, Pic10, ANH14, Baš12, QH16, QH17, SPY18, ST19]. basic [Fer22]. basis [JK16]. Bateman [Ano15c]. Bauer [BKB13]. be [Viñ19]. Beautiful [Cha10a]. Beck [And21c, Che18b, Che22c, LW20]. Beck-type [LW20]. Behavior [BKK06, Ono08, Zha14, Alk20, And16, BB20c, CL22, Cow22, HL11, Kle17, LP14, Mol12, WS20b, Tou09]. behaviors [Pon16]. Beilinson [Sca17]. being [Ska17b, Ska19]. Beiter [MR12]. Bell [KwK21]. below [BG17]. Belyi [AT20]. Benford [CHL19, JTY16]. Berndt [AAW08b, CK14, TB06]. Bernoulli [Ade18, Ade21, BB14a, Che22b, CP18b, DJ22a, HN08, JMV16, Kam08, KL17, KY18, LY12, N18, OS21, Ros18, Sun15c, Sun16b, XC10, You16b]. Bessel [BK15a, MV14b]. Beta [Hic14, HM06, Sto21]. beta-development [HM06]. Between [ACH05, Bou11, GL11, Paz13, RS13c, Suz05, Van07, Bao19, BCSX20, CFTZ14, Dud15, EMS21, FGT15, Fri16, KK18, MU18, Orr17, XY19]. Beurling [GK18b, NH20, Zha15a]. Bhargava [AC15]. Bianchi [LŞ20, Won20]. bias [Lam16]. Biased [Alk20]. Bicyclic [GML12, Sch13]. Bielliptic [BB21a]. Bijective [Kim10b, Yee09, Gol16]. Bilateral [CZ10, Gui20, Che18a]. Bilu [Pet05]. Binary [CC07, CEO06, DS13, EV07, Has13a, Kan14, Kum09, Liu11, Our09, Res09, Vul10c, APW14, And16, BHPR17, BCF21, EF16, Liu19a, Pat20, Xio16, Ye15b, Zhu18, Zub20]. Binomial [CZ09, GZ11, ST11, ASD16, AT22, Apa18, CP18b, DL22, GL17b, KP16, KS14, LM16, Len17b, Len17a, Len18, Ma17, sMT21, San18a, SSU21, Sof18, Sun17a, ZY20]. Biography [Ano15b, Ano15c]. biparametric [SSU21]. bipartitions [NH17]. Biquadratic [GML12, AZTM15, Bae19, CeM21, SS18]. bivariate [BSK17]. blocks [SS15b, Ska19]. Böcherer [RT11]. Böcherer-type [RT11]. Bodies [DK06, Now16]. body [HM06]. Bombieri [Yao18]. Bombieri-type [Yao18]. Bootstrapped [MR15]. Borcherds [Ehl10, Hof13, MPY13]. Borel [Zha21]. bornée [Hua17, Mig15]. Borwein [Li20b]. Bound [Hal12, Kna08, Mue12, Agg21, Amo07, BG17, BLW21, BŻ19, CP10a, Che22a, DP22, Edd16, JsKL21, wJShY21, Pac19, Pup15, YF15]. boundaries [dS18a]. Boundary [BKK06, Roe14]. Bounded [BB09b, LRL10, Thu08, DS17, Hua17, KV19, Mig15]. Bounding [Sha14]. Bounds [Akh12, BD08, EK07, Leh08, LRL10, Pon06, Via10, Wat08, AV19, Cre18, Ivi16, KM18a, Kim17a, Moh19, Ste16, SY19b, Zel19, Zhe15]. Boyd [Lal10, LW19]. bracelet [CG13, RS13a, Yao15]. bracket [Sch18a]. Brackets [Meh12, JS19, Lan19]. Brahmagupta [Cho20b, CBJ22]. branches [ACS16, ACB19]. Brauer [Bal19, Bis15, Car11b, Ga20, Ito18, PS22]. Breaking [Sch13]. Briggs [Bha20]. Brill [KM18a]. Brocot [Len19, Sch12]. Broken [Fu11, Mor08, AB15b, DW16, Dai16b, LMW16]. Bruhat [ACS16, ACB19]. Bruiner [IW16]. Brumer [DR16, KRY09, KRS10, Nom14]. Brun [Deb19]. build [RC17]. Bumby [BTW06]. bundles [Tür11]. Burgess [Alk12, BG17, JsKL21, Tre15b, WHZ19]. Burnside [Mil13]. Busche [Tót13].

Cadoret [Tür18]. calculate [HK21]. calculation [IIO20]. Calkin [HMST16]. cancellations [cC21b]. Canonical [Car11a, Fin12, LW13, Tem10, Bat17b, Fin20, Vs21]. Cantor [LM15a, WLWY22]. capitulation [Pit11, Pit11]. cardinality [JL18b, ZLL16]. Carlitz [Bao05, Koc08]. Carmichael [Har08, McN13, MW16]. Cartan [Mr22]. cas [Tho06]. Case [Bou11, BKY19, BKY20, Hir06, ZY08, Chi19, Ge18, Lu20, Sch21a, Tho06, Vie10]. Cases [VB19]. Casimir [Mau12]. Cassels [FN14, GHHP06]. Catalan [Koy22, LLS21, Pan22, Sun17b, Zha14]. Cauchy [Kom13, KL17, KY18]. ceiling [MCW19]. Centers [FS11]. Central [BFCC09, BM11, Bui12, Bui13, CK17, CZ09, FW10, GZ12, Pi20, Rod14, San18a, Dja11, LS17b, sMT21, MR15, PT14]. Certain [AALW08, AALW09, AK14, BB09a, Bao10, BK12, BS15, Boy10, BGW11, Bru05, CK14, Cha10b, CD11, DD09, GML12, Gun06, JSS14, Kim10a, Kök13a, MS05, Mer11b, Min12, Mui11, MM07, MW11, OT05, SS08, XY13, ANH14, Ala14, AAW16, AAAW17, AK15c, BE21, Bor18, BB14b, Bun08, BV09, Cha20, Cho18b, EW15, GM18a, GL16, HLT20, Hin15, Jar22, JS19, KK19, KE17, KW14, Kug22, LS18, Miy15, Oga14, PR21, RS14a, RSS18, Ray21, SS18, Sha14, Tay17, Ula19, Vää15, Vää16, XZ21, Ye15b]. certaines [LR22, Tay17]. Cesàro [CGZ21]. Châtelet [Bal19, Rom19]. Chabauty [CDc20]. chains [Cho20c]. challenge [RS13d]. Champaign [Ell15]. Champions [GL11]. Chan [FWX21]. Change [CF09, HY18]. Changes [RSW14, GS17, KM14, KM18b, MM14]. chaos [Mah20]. Character [BH10, Bla11, CK14, FZ06, FZ12, OSW11, Sad16, Wat08, EW15, Wal17a]. Characteristic [Ada12, AT20, BZ13, Koc08, BJ19, BK22, Ito18, Koy22, Lim16, Pal14, Sar22, She14]. Characterization [Fol09, JKS17, Mem17, Wei15b, Leo18]. Characterizing [ABP09]. Characters [BZ05, Kat10, KM09, KM12c, AH19, DJ15, LHK18, Pol14a, Tót18, Val14, Won17]. Chasing [cC21b]. Chebotarev [ApKK22, Deb19, Kan13, VW17, Wan20]. Chebyshev [AT22, BST10, Gir22, IT10, PT19]. Chebyshev-Like [BST10]. Chen [Yan13]. chiffres [AST22]. Chinese [CC16]. Chowla [Bor22, Cha12, Ham18, IN22, YF18]. circle [Bzd17, DFV13, Kel17]. circles [Cha15, Ska21]. Circulant [LS12]. Circulants [Wil10]. Circular [Bel09]. Class [Bao14, CLM08, CeM21, GML12, Har07, Has13a, Hoe10, JM11, Mat12, MM13a, MS07, OSW11, łPS13, Sin10, Vig12, Wil12, YXJ13, ZY08, AAA15, AAW16, AITA19, AMO17, AZTM15, ARTZ19, BJ19, Bao15, CHL19, CTZ16, CGPY15, Dai16a, DM19, Fuj20, FK11, GL19, GG22, GM19, Ga20, GT17, JK16, Kum21, LM15a, LS19, LM15c, May12, Mil18, Mon14, Mor16, OS11, Oga14, Poë20, Rau16, RSY18, Rei21, Ric13, SSS22, Tsa17, Viñ19, Wit07, XLD22, ZH16c]. classe [Ric13]. Classes [AAW08a, DS13, GH14, Kab10, Kan13, KN09b, Lez12, MS10, SS10, Amo07, AST22, AMO17, Cau20, Gro20, Hsu16, HK20, Kom17, LS18, PR21, Şia15, Tsa17]. classic [cC21b]. classical [BP17, FT18, Isa21, ZH16a]. Classification [HKN12]. Classifying [Hin18, KRS10, SS07a]. closed [Zhu22]. Closest [BO12]. closures [AB19, Szc15]. CM [Ari13, Bou11, Hsu20, KK18, KY20, Kim16a, Nom14, Sch21a, Sua09]. CM-extensions [Nom14]. CM-types [KY20]. co [Pit11, Sto21]. co-factors [Sto21]. co-kernels [Pit11]. Coates [KT20]. Cocompact [Vul09]. Codes [GNS06, Mah12, Our09, SZ07, Mou17, Nar22]. Coefficient [Gek11, ASD16, Bat17b, Dud14]. Coefficients [AK14, BGW11, BR11, CK13, CZ09, Dai16a, EG09, Gri11, GZ11, HKKL12, Kaz08, Koh10, MZ22, MS13, Moy13, MM07, Raj09, RSW14, ST11, WZ09, YXJ13, Ala14, AAA15, AK15b, AAAW17, AK17, And16, Apa18, BSK17, Che19, DM13, DL22, Fin20, Gar08, GS17, GL17b, Hou21, Hsu20, HY18, JTY16, JV21, JL17a, yK21, KK18, KP16, KM14, KS14, Kri16, KMV20, KM18b, LM16, LS14, Len17b, Len17a, Len18, LR06, LX19, Ma17, sMT21, MP22a, MM14, MSV18, Mem20a, Mem20b, MR12, OS11, Oli14, OO17, RSS18, San18a, SSU21, SPY18, SW08, Tan19, Ula19, Vep17, Wei15a, Xio17, ZY20]. Cohen [JS19, Kri16, Lan19, Meh12, RS14b, YZ22]. cohomological [HL12]. Cohomology [Bel09, Gim13, Klo13, KLR09, KR10, Raj11, Wie09, AP08, Cau20, Sca17]. coincidences [BK16]. Collatz [Koh08]. Collatz-Type [Koh08]. color [AAR16, Kei21]. Colored [Lin13b, Lin14, Lov05, xMgC22]. colors [CG19b]. combination [AT22, Sun19a, WS16, WS17]. combinations [BHPR17, CK21, KTZ16, KTZ18, Pol14a]. Combinatorial [BM21a, Fu11, HR11, JKK16, Kim10a, All09, Bat17a, Kim21a, Kur21, LR06, ZS18]. Combinatorics [Bur21, SS21]. coming [Oli22]. Common [AK11, BG06, Cha18a, Bre19, FF15, FZ18, NJ22, RC17, SA09]. communtative [Wit20]. commutable [Par18a]. commutative [Lei12, SSTW14, SH08]. Compact [MMW11]. companion [Rob18]. compatibility [Orr17]. compatible [Cau20]. Complementary [BLS07]. Complete [Ngu11, YF19, Mer11a]. completely [BZ15, HL11]. completions [Kid16]. Complex [Baj14, CCS13, HKN12, Olo09, WZ09, Zue18, Kan14, Kat17, LY21, Lei21, Nat17, Tre15a, Zub20]. Components [FG14, Bis15]. Comportement [Tou09]. composite [JsKL21]. Composition [ZLZ13, SZZ18]. Compositions [HHP09, GM19]. Compositum [WS10, Kha19]. Computation [CL19, FZ12, BT18b, HS20]. Computational [Hit18, CDc20]. Computations [Fin14]. Computing [ACB19, Dor20, FN14, Sam10, Sch21b, Tra17, Roz18]. concatenation [Lin22]. Concavity [LY12, SSU21]. Concentration [Zum11]. Concerned [Li13]. Concerning [BHV11, PPT12, Sun13, Sun14, BB20a, Boc08, Har15, HLT20, sMT21, RS13d, Sun15c, Tót18]. Condition [MP12, Sun15b, Yam16]. Conditional [Baz11, Lan09, Sar22]. Conditions [McN13, LZ18b, NRS20]. Conductor [DW09, VS10]. Configurations [Kom09]. Congruence [DLZ19, Fu11, KK10, KN09b, LS14, Lin14, Mac16, Meš12, MPY13, RS11, SW13, And21c, BKS16, CSJ17, Cha21, Che22b, Eld19, FWX21, GS18, GS17, GL17b, GN19, Guo19a, HW18, LZ18b, Lim16, Liu17a, Meš16, NRS20, NP18, RSY18, YC17, Zem17]. Congruences [And13, BBS08, BO11, Ber07, Bon08, Bou11, BL08, Bri09, CMS20, CCS10, Cha10b, CM12, CE07, CWW08, CG13, CGM15, CGH18, CG19b, Dai14, DR10, DS21, EG07, Gar10, GJS14, Gir16a, GZ12, Jam12, Lag10a, Lag10b, Moy13, NS18, PPT12, RS13a, She17, Sin10, Str15, ST11, Sun12, Sun13, Sun14, Sun15b, Sun18, Ver10, XC10, ZW14b, ACX18, AB15a, APP13, Apa18, BP11, BB20b, BD22, Che22c, Gor19, Gui20, Guo15, JZ17, LLW18, MS16a, Mao17, MW19, sMT21, MTWZ17, MV16, MS18a, RB18, Ros18, SW08, Sun15c, Sun16a, Sun17b, Wan16]. Congruent [Ben13, Rol11, CL11b, CL13, CG19a, KK12, Mok20, RSY18]. Congruential [EMW06, EMG08, NW05]. Conic [AG13, Tür11]. conjectural [FWX21, HS19b, Wal21, Wei19]. Conjecture [BHV11, Cha12, Chi19, FM12, LS06, Lai10, Lal10, LS05b, Miy11, Mos07, SX14, Xia13, XY14, Yan13, Zyw11, ATIA20, Bor22, BKY19, BKY20, Che18b, CGPY15, DR16, DR22, Do17, pFG22, FW20, GF18, Ge18, HKL^{+}21, Har18, Hui18, Hum14, IW16, IN22, Kat17, Li20b, Lim15b, Nom14, Pas15, PT19, RT11, San18b, Tak21, Tho06, Var22, VY13, Vse14, Wit20, YF18, KT20, Kuo09]. Conjectured [Meš12, Guo19a, Mao17, NP18]. Conjectures [HKN10, Toh13, And21c, Gug21, GL16, HKN11, hKS21, LV17, PS15, Won18, XY19]. Conjugate [BS12, Roy05]. Conjugates [HS21b, Höh11, Koh08]. connected [MV14b]. connection [Vää15]. connections [Kal18]. Conoyaux [Pit11]. Consecutive [BK16, CP10b, GL11, Hu13b, JL22, LS06, Tan09, TZ12a, TZ12b, AM17, Cha18a, LY19, Lin22, PT15, RY16, TYZ16]. Considerations [BHV11]. consisting [AAW16]. Constant [HJW14, HW18, Kop08, Mue12, Oza17, RST08, FZ16, Gra18, MOS14a, MV16, OLG19, WZWQ16, WHZ19, Wei20, Xu19a]. Constant-Perturbed [HJW14]. Constants [Bac08, Bar14, GGW11, Bha20, CP17, CP18a, MORS16, OS17]. Constrained [LMO^{+}19, Nit09]. constructed [Mou17].

Constructing [CF22, STW10].

Construction

[Coo09, Har11a, Hou17, JLSW15, Pic10, Kur21, Lin18, Mon14, Vie10]. Constructions [BCU14, FM20]. containing [KW18]. Continuation [Mui12, Tay09, JVW20]. Continued [BO11, BKB13, Cha10b, Gir11, Gir12b, Lin13a, SXJ13, Zha12, Zha05, ZT13, BI21, BMW06, Cha10a, CH05, DFG^{+}14, GM18b, HYZ17, HWh20, Jam16, MAM06, Mer22, Oye16, Par22, Rob20, SW14, Van16b, XY11, Zha15b]. Continuing [CLMR09]. Continuous [PSZ16, Haj20]. Control [BL09, Kum13, Liu22, LV12]. convergence [CC21a, Len17b, TLZ22]. convergent [Sah16]. converse [Pet16, Zha17b]. convex [LTZ20]. Convolution [Cho20a, Hua14, Roy07, Sch13, Wil05, AK16, AH19, CP18b, CY14, Hau18, JV21, KE17, Lan19, LP17, Par18b, RS13b, XTY14, YX14, Ye15a]. Convolutional [SZ07]. Convolutions [KS13, LS12, Lei21, Vie10, You16b]. Conway [KRLT20]. coordinates [EGL21]. coprime [JLl21, MS16b, San18a]. Core [BS12, Gol16, NY21]. cores [BN14, Wan16]. corps [Amo07, HM06, Kra07, Mai11]. Corput [HKLP09]. Corrections [BD08]. Correlation [Smi13, BPZ14, GY16]. correlations [LX19]. Correspondence [Fre12, RS13c, Bir19, Dau14, PT14]. Corresponding [DD09]. Counter [EMG08]. Counter-Dependent [EMG08]. Counterexamples [Ram17]. Counting [DM21, Hub11b, KT15, KM05, KV19, MP14, Tür11, Bor20, Che18c, Deb19, HH21, May14, TT14, Tou19, Zoe19]. Courbes [Kra07, Bil11, Phi11, Ric13]. courtes [LR22]. Covering [San14, Har15, JW17, Ram17]. Coverings [Kop08, LM11, GS16]. Covers [BZ13, GG22, Mei18, Pet16]. Crank [Kaa11, AsMS20, CM16, Che22c, EMS21, JS15, LRS14]. cranks [MS21]. critère [BD10]. Critères [Bil11]. Criteria [BBCM13, FT15, Bil11, Hat16, WY21]. Criterion [TK10, BD10, hC21, Gir14b, Meš16, Viñ19]. Critical [Far08, KS13, Lev06, Bru18, CV17, FG22, Lan19, Pat19b, WZ12]. Cryptanalysis [Nit09]. Crystalline [Cho13b, Roz18]. Cubefull [Cha13]. Cubes [RR11, LZ18c, Rez21, ZG20]. Cubic [BO11, BEH10, Cha10b, EJ11, GY13, Har11b, Hub11a, JR13, KMS21, Koz09, MR06, SW06, Sun16a, Xio11, ZLZ10, BCF16, BS16b, Cha10a, Hu22, JLSW15, Jed14, JS20, Kat17, KW14, LL17, Mor16, MOS14b, Mur15, Sof18, Tak15, TT18b, BF10]. cubiques [BF10]. Curious [AAW08b, Pat10]. Curve [CL11a, Hub11b, LM14, AM09, CPS18, Dra11, DP22, FN14, FY17, LZ22, Shl12, Tür11]. Curves [AW13, ABP09, BL09, BK12, Ben13, Bou11, BU11b, BM12, CCS13, DW09, Dum05, FG12, FS08, Fre12, Gla09, GY13, Kow06, Koz09, Kum13, LRL10, Mat08, McC10, Naj12, Pál07, Sai10, SSD11, Ula10, Veg11, VZ14, WZ09, Yas13, Abo08, Ari13, BB21a, BJ19, Bai16, BCF16, BCF21, Bil11, BF19, CF17, Cen16, Cop20, DG17, EGP21, FT15, GR11, Gar18a, Gar18b, II16, JS20, Kal18, Kra07, LW19, LY21, LL22, Li14, LR19b, Mat17a, Mr22, MP22b, Nţ21, Phi11, Pri09, Pum20, Ric13, Sad16, Sar22, Sch21a, Sha14, SS17, Shp19, Tak08, Tak15, Tro17, Van18, VY13, Wei19, Yas15, dS18a]. Cusp [Das13, EG07, EH08, Fli11, Gan10, HKKL12, RS13c, RSW14, Sah11, AB18, BV18, BP11, CJKM18, CL22, GS17, JL17a, Kim22a, KM14, KMV20, LX19, MM14, SPY18, Wei15a]. cuspforms [Bro10a, Lan19, XZ21]. Cuspidality [Zan16]. Cycles [GH14, MR06, Par15a]. Cyclic [BS10b, CF09, Koz09, Lag10a, Lag10b, LO12, LP13, McG12, SW06, BARCVS13, Gor19, Liu18, PQSW14, Sch14, SY14, TT18b, Wan21, Wit07, YL16]. Cyclotomic [AB09, HJW14, Hoe10, Leo15, LM11, Rob10, Bat17b, Bzd17, CeM21, Cho18a, DFV13, Dud14, Fuj20, FK11, FJ20, INST14, Jar22, Kit13, Kum21, LL22, Mat17a, MR12, Mou17, Ngu19, SW08, Zha17a].

dans [AST22, HM06, Tho06]. Davenport [Smi11, Che15, MOS14a, PZ18, RST08, WZWQ16]. Decay [Gek11]. Decomposable [Liu08]. décomposé [Tho06]. decomposed [Tho06]. Decomposition [Pur13, ELO16]. Decompositions [ZLZ10]. Dedekind [BS16a, BH14, CC16, CDHS15, Gir12a, Gir14a, Gir14b, Gir15, Gir16b, Gir17, Gir18a, Gir18b, Gir19b, Gir19a, Ham13, Ham16, Hea14, Isa21, JRW11, KW16, Kad12, Kan22, KK08, Li09, LS19, Lou15, Mac17, MP21, SVY20, Tsu15, ZCX22]. deduced [LQ21]. Deducible [BO11]. definable [JQ21]. defined [BE21, BCF21, DWW17, HWh20, Tür11]. Defining [Sin09, PR17]. Definite [CCL13, CEIK07, Oh11, CL19, Hou17, JKO18, Ot20]. Deformation [KL12, Hat16]. Deformations [Pan11, Ray21, Ray22, Miy15]. Deformed [EOY05]. degeneracy [KY20]. Degenerate [LY12]. Degeneration [Anc17]. degré [Hua17]. Degree [Bru05, BD08, Cao11, CF09, Das13, Klo13, Kop08, LO12, McG12, Rey13, Roy05, SXJ13, CL22, Cho18b, CWR16, Dic15, GK18a, Gar18b, GS17, Hua17, II16, Les16, Mah19, PST21, Pum20, Qua19, RGHK20, Sca17, Shi16, Shp19, Sto21, Wal17a, Wit07]. Degrees [Ula12, mPP22]. Del [Car11b, Hua17]. Delannoy [GZ12, LLW18, Nob12]. Deligne [Sca17]. Denominators [Gir12a, Ham13, Kam08]. Dense [Via10, Def15]. Densities [BC11, Das13, Hir22, Jon21, Wan20]. Density [Cho13b, GT10, HHP09, Kob14, LR19b, LL11, Mam10, MP12, MS07, Mun10, Sah11, Deb19, Kob16, MR15, MS16b, Mou17, Rei21]. density-like [Rei21]. dent [Now16]. depend [AV19]. Dependent [EMG08, MS13]. depth [Ass21, AMPS17, Mal20]. Derivation [Hau18, Mur17, BT18a]. Derivative [Chu13, CY18, Hun18, KV19, Liu13a, SA09]. Derivatives [BM11, LZ11, Tan12, KTZ16, KTZ18, TT18a]. Derived [CW12]. descent [Cre18]. described [CWW15]. Description [Ngu11, HP11]. designated [ACX18]. detect [CS21]. determinant [HS19a]. determinantal [II16]. determinants [CK21, DJ22a, HW15, ZS18]. Determination [ZLZ10, Sun15a]. Determinations [BBCZ05]. Determining [Lan19]. deux [AST22]. development [HM06, PT19]. développement [HM06]. d’extensions [SS10]. Diagonal [Bao10, Pum20, BCU14, GL20, HK21, Jon21, LR19a, Var22, Vep17]. Diamond [Fu11, Mor08, AB15b, DW16, Dai16b, LMW16]. Diatomic [Coo10]. Difference [Bya06, Bya09, MC13, CGG15, Dud15, MU18]. Differences [Hay10, Shk18, Bre19, CFTZ14, Fol21, Len14]. Different [Suz05, WS10, ZT13, Baš12, JS20, Oli22, Tsa17]. différente [Sod21]. differentiable [Mem20b]. Differential [Hub11a, KN09a, Mas07, Mat19, May19, Gra20, Ked05, Tan18]. Digit [SSTW14, Baš12, DWW17, MS15]. digital [Alo19]. digits [Eve22, HLS11, LLL22, Sau15, SYZ14, SFM17, TYZ16]. Digraph [MM11]. dihedral [Ber11, May14, Mil13]. Dimension [BK15b, Gan10, IFD08, NV10, TZ21, BLW21, BV11, NR16, NSS15, TLZ22]. Dimensional [AS06, MR06, NV10, Sch12, All17, Cob21, HR21, Hol19, Len19, Per17, Roz18, SW14, ZH16a]. dimensions [FW20, Tra17]. diophantine [Cho18b, CBJ22, GN21, LL17, MT17, YF18, BTW06, BEN10, Boc08, Cai18, CFM18, Dah11, GL20, GL18, Har11b, Huo18, Jah10, KMW10, Kre17, Li19, Liu17b, LT08, Mar14, Miy11, MT12, Mu17, MQ18, MLST09, NRS20, Rey13, RLT22, SS08, Shl12, SZ15, TU20, Van16b, Vul06, Vul10a, Yam16, hY20, YF15, ZY12, Zha17d, ZL19]. Diophantus [DFL08, DJ10]. Dirichlet [AH19, BB20c, BM11, Bui12, Dah18, DJ15,

– DLV20, DS18b, Eve22, Far08, FP12a, GK13, HM16, Hau18, Ivi16, JS19, LHK18, MNZ19, MP12, MN13, NMZJ22, Ono17, Pol14a, Sch13, Tam14, Tót18]. d’irréductibilité [Bil11]. discovery [ZZ21]. discrepancy [Che15, Lin22]. Discrete [BH10, GK13]. Discriminant [BD08, GY13, Rob11, VS10, FY17, JLSW15, Kha19]. Discriminants [BF10, DU10, Lou16, Mam10, Kim17a, hKS21, Kön20, KK12, Rau16]. disjoint [Ska19]. d’isogénie [Ric13]. dissection [XY11]. Dissections [AK15a]. Distance [GM12, AMM17, Fri16]. distinct [BN14, BB18b, CGM15, Dai14, GZZ15, LP16, Ska17b, Wan15a]. distinction [Lu20]. distinguished [Smi21]. Distinguishing [CG15, Ghi11, VX18]. Distributed [Gra07a, Gra07b]. Distribution [CW12, EMW06, EMG08, Gek11, HKLP09, Hu13b, Hua17, Kel10, Klü12, Kna08, LO12, Liu12a, MCW19, Mei18, Pat19a, Shp14, Smi13, SW06, Dra17, GK18b, Gib14, Kön20, MU18, NW05, Rau16, Ric13, SK13, She22, Shp19, SSS22, Tol06]. distributions [PSZ16]. Dividing [MS09a, JKS17, San17]. divisibilities [ZY20]. Divisibility [DJ22b, Gri11, yK21, US13, WS19, WG11, Zha16, Bel22, BO19, CG18, DP22, GL16, She07, Sun17a, Yam16, Zhu22]. divisible [San18a]. Division [BFCC09, DP22, GR14, Szc16]. Divisor [AG12, Fel12, Roy07, Vil18a, AB19, Ary17, Bor22, CP18b, Cho20a, Def15, FZ18, Ivi16, Kim17b, LZ18b, RGHK20, Wit07]. Divisors [AK11, BBS08, BLS07, DS13, GT10, Gra07a, Gra07b, Hu13b, LS05b, Min12, MS13, Pol14b, Rho09, Dra17, FF15, Kim21a, LH16, Tho13, Tra17, Wan15b]. d’Iwasawa [AMO17, Pit11]. Dodecic [Awt12]. domains [Ala16, Čes16, LS19, ZCX22]. Domb [MW19]. Dominant [Dub11]. d’ordre [SS10]. d’Ostrowski [AST22]. dots [CG13, RS13a, Yao15]. Double [EOY05, KMT11, LM11, SY17, Sou10, Wak17, BP15a, BK15a, Dah18, DM20, GL16, Har18, KK19, LQ21, Ma17, MS18a, Pet16, Shp18, ZL18b]. doubled [BN14]. doubling [HWh20, Zor11]. Doubly [WLWY22]. Dream [And07]. Drinfel’d [DP08, DWW09, BV18, hC21, EG14, Gek11, Kat21, Kaz08, KL09, MP22a, Pál07, PP18, ZJ19]. Dual [NV10, Pic10, Che18a, Lim15a]. Duality [BC10, KO10, SY19a]. Duarte [BD10]. Ducci [LTZ20]. d’un [Amo07, JM10]. d’une [Gil17, Ric13]. Durfee [BK11, pCG21]. Dwork [Goo17, Koc08, MV16, Shp14]. Dyadic [Bri11]. Dynamical [FHS11, San18b, NSS15]. Dynamics [FG14, KMW10, Kru16, LMS^{+}22, ZJ19]. Dyson [And21c, Bri09, Gar10, Zha22b].

Each [IFD08, Gir18b, RSY18]. Effective [Kna08, RD13, ZLZ10, Bud20, DP22, Sun15a]. efficiency [HYZ17]. Efficient [CS10, Hub11b]. EGZ [OLG19]. EGZ-constant [OLG19]. Eichler [AC13, Gim13, KLR09, KR10, Raj11, SG17]. Eigenclasses [DD09]. eigencurve [AP17, Roe14]. eigenform [Hsu20, JV21, MP15, Pen19]. Eigenforms [Ghi11, Meh12, STW10, Bao19, DM15, KK18, Van12b, VX18, Van16a]. Eigenvalue [IKS07, Bro10a, HL11]. Eigenvalues [Das13, Sah11, CL22, Kim22a, MP15, Wal17a]. Eight [ZG20, RSY18]. Eisenstein [BP17, Coo09, Ehr09, GNS06, GP12, HN21, JM16, KN09a, KM12a, Kri16, MOR21, MO20a, Miz08, Mon14, Moo19, Our09, Oza17, Pál07, Pet16, SVY20, Su16, Tay09, Wal17a, Wal17b, Wil19]. Elementary [Apa18, Mol05, VS10, SST19]. elements [BK22, CK20, Col12, Din21, Dub18, Lim16, Pal14, mPP22]. Elite [CLMR09, Wit10]. ellipsoids [JL20]. Elliptic [AW13, Bai16, BL09, Bou11, BM12, CLN05, CPS18, CCS13, Cop20, DW09, GL19, GML12, GY13, Kow06, Koz09, LM14, LRL10, Mat08, McC10, MM13a, Mun10, Naj12, RS13c, Ros08, Shl12, Tak15, Ula10, Vig12, WZ09, Yas13, AM09, Ari13, Bil11, CF17, Cen16, DG17, Dra11, DP22, FN14, FT15, Fri16, FY17, Gar18b, JS20, LY21, LL22, Li14, LZ22, LR19b, Mat17a, Nţ21, Ric13, Sar22, SSU21, Sch21a, Sch15, SS17, Shp19, Tak08, Tro17, VY13, Wei19, Yas15, dS18a]. elliptical [Kra07]. elliptiques [Bil11, Kra07, Ric13]. embedding [ACB19, PS22, SS10]. embeddings [Bat17b, Mai11]. endomorphisms [Sug15]. Engel [HYZ17]. engendré [Amo07]. Enlarged [Li16]. ensemble [Jun14]. entiers [JM10]. Entire [Pil09]. Entry [BI21]. Enumerating [PS17]. Epipelagic [FRcT20]. Epstein [Cou08]. Equal [JRW11, Rez21, CCH^{+}19, Cho20b, Gir16b, ZL18a]. Equality [Tsu15, Chi19, Gir14b]. Equation [Bao06, EH08, Har07, JR13, Kum09, Mol05, Suz05, AGL18, Bao08, BM17, BEN10, BX18, Cha21, CBJ22, Dah11, DT22, FL15, GN21, Huo18, JP14, Koy22, Kra15, LR19a, Li19, LL17, LT08, LM15c, MT12, Mor16, MLST09, Pal14, TU20, hY20, YF15, YF18, ZY12, Zha17d, Kra15]. Equations [Akh12, Bao10, Bao14, BTW06, Gün12, Hub11a, JZ06, KP14, Kan13, KMT11, Miy11, Ngu11, Rey13, Rif19, SS08, Wak12a, BZ15, Bao15, Bel22, BCDY15, Cho18b, EGL21, FY21, GJR19, GL20, Gra20, HS22, Hei18, Hin15, Ked05, Kim21b, Kre17, Mah19, Mat19, MOS14b, Pan22, SS18, Sch15, Sze19, Tan18, YF19]. Equidistribution [DS13, PS11, Pet05, Dic15, IW16]. Equivalence [Bar14, HK20]. Equivalent [MS12a, DS18b]. Equivariant [ES12, SS14, Bru18]. Erdos [AMM17, Boc08, Bre19, FZ16, LLZ18, Pat19a, Tan14, US13, WHZ19, Yan13, ZW14a, Zha21]. ergodic [Mem17]. Erratum [CL13, Gra07a, Lag10b, MP16a, Van16a]. Error [EH08, HN07, Bor20, CFTZ14, CTZ16, LZ18b, MP14, TT14]. Escott [Cho17]. espace [Mig15]. espaces [Phi11]. Estimate [MS07, Bud20, BV11, Li22, Lin22, NR16, TT14]. Estimates [AB18, JV21, Roy10, SS08, Zor13, BK15b, CG22, KW16, Kei13, KMPS20, MR15, SS19, SPY18, VY13]. Estimating [FZ18, SA09]. Eta [CEP12, Sin10, Wil12, YXJ13, AAA15, Oga14, Ye15b, Zue18]. Eta-Function [CEP12, Zue18]. eta-quotient [Oga14]. eta-quotients [Ye15b]. Euclidean [BFCC09, CCL13, CL19, Gra11, Hsu16, Lez12, McG12]. Euler [BK16, BKW13, Bha20, CW12, Che17, CP18b, DJ22a, Dja13, DA18, EOY05, ELY05, GC17, KW12, Li09, OS17, OEL08, She14, Sof18, Suz05, TK10, Vää18, XYZ17, dS18a]. Euler-type [Vää18]. Eulerian [KK19]. Evaluating [CP18b, EOY05, Roy07]. Evaluation [AK16, CY14, ELY05, KP16, KE17, LP17, OEL08, Par18b, RS13b, XTY14, YX14, Ye15a, HW18, XYZ17]. Evaluations [BEH10, HW15, Zha05, BP15a, BP18, Xu19b]. Even [BB09a, ELY05, HKN12, Kom09, Neb13, AD18, BJ19, Dai14, Edd16, FY13, Hof17, HL11, JL18a, Kit13, Len17a, RSY18, Sun17c]. Eventually [JL17b]. Every [Far08]. Everywhere [AP08, BCF16, BCF21, HK21, Tak15]. Exact [BB18b, BR11, FW10, Len14, SK13, VR20]. Examples [Lez12, SS07b]. exceeding [Liu14]. Exceptional [Lan09, LLL22, Liu19b, Naj12, Pol11, BG15a, CS21, LLZ18, Zha15b, Zha21]. excludant [BM21a]. Existence [BZ13, Yu11, Ben15, DS18b]. Expanders [Vin14]. Expansion [Alk07, Chu13, Lin13a, CJKM18, CH05, DA19, Hic14, KK20, LZ15, LLZ18, SYZ14, SFM17, TZ17, WLWY22, Xu19a, Zha15b, ZMS15]. Expansions [EG09, KN09a, Luc10, Mor11, SXJ13, BW17, Fol21, KM22, LLL22, Liu13a, Mor22, Sah16, Sau15, Sto21, TZ21, TLZ22, Tan19, TYZ16, Won20, ZLL16]. Explicit [Che22a, Deb19, DGM19, IIO20, JsKL21, Kad12, Kat21, KR16, Li12, OEL08, Pea22, PST20, PRVS08, Sze19, Van07, XYZ17, YXJ13, Zha05, Zor11, AP17, DGM16, Edd16, Fer22, HKL^{+}21, Li15]. Exploration [Row10]. Exponent [FS08, Amo07]. Exponential [AS09, Alo19, BBS08, EH08, Gün12, HRL11, Koc08, KMPS20, MS05, MS10, Min12, Miy11, NW05, Wei15a, GL17a, Hou21, Kei13, LLMA16, SY17, YF15, YF18]. Exponents [CLM08, MPY13, Fer22]. exposant [Amo07]. expression [Mac17]. extended [Wea22]. Extending [McC12]. Extension [SW13, All17, CH15, FK11, KT15, Kum21, Mat17a, SSU21, Yas15]. Extensions [Bos09, Bri11, Bru05, BD08, DD10, Koz09, LRL10, Mam10, Pic10, WS10, Zha15a, Amo18, ACB19, BS20, BARCVS13, Cas12, CFM18, CK20, DR16, Kit13, Kön20, LL22, LZ14, Mai11, Mat18, MP22a, Mon14, Nom14, PS17, PST21, San09, SS10, Sch14, She14, Shp19]. Extensive [LMW16]. Exterior [MY13]. Extra [Gla09]. extraspecial [SS10, SS10]. Extremal [Kom09, Neb13]. Extreme [łPS13].

Factor [Bar14, BL13, Bre19, CM20, CW19, Kim16a, WHZ19, Wei20, Zel19]. Factorial [BB09b, Ala16, You16a]. Factorials [CD11, AC15, LS17a]. Factoring [BGW12, Ozd21]. Factorization [Gil13, KM05, GV13, Har12, SS15a, SH08]. Factorizations [GGW11, DL22, JLl21, SS15a]. Factors [AK12a, AK11, Cha15, GZ11, Bad17, Čes16, Cha18a, JK21, LS05a, LP16, May19, RSY18, RL18, Sto21, Tro17, Zha17c]. faible [Do17, Tho06]. failures [Rom19]. faithful [BZ15]. False [CZ10]. Families [Bar13b, GLZ15, JV08, LZ11, LV12, Mas08, RSY18, Zie08, ACX18, Bal19, BN14, BD22, FT15, Hin15, Hol19, Liu19a, Pha22, RB18, Ray22, Shp19, ZW14a]. Family [BK12, HRL11, IFD08, JZ06, JSS14, KL12, KPT08, Ber11, DFG^{+}14, Dja11, GM18b, Gue15, Haj20, JP14, Jed22, Kom17, Liu14, LL17, Mal20, MTWZ17, Mon14, NSS15, Rom19, Ros18, Sad16, Van21a]. Farey [Hay10, LM16, Sch12]. Fast [MS15]. fastest [Gui21]. Fermat [BCDY15, CW12, FG12, Guo15, II16, IIO20, Jed14, JR13, KO20, Kra15, LS05a, LM15c, Mor16, Pan22, Ska17a, VZ14]. Few [BC18, LP16, Tro17]. Fibers [Mun10]. Fibonacci [HKN11, BL13, CCZ15, DL20, HKN10, KK15, Lee10, LX21, LP14, RL18, San21]. Fibonacci-norm [HKN11, HKN10]. Fibrations [Mun10]. Field [AK11, Bao06, CCL13, GY13, Hea14, HM10, JM11, Mam10, Mat12, Rob11, SXJ13, Tam14, Tu11, Amo07, ABK19, ARTZ19, Bao08, Cop20, Dor20, Gir18a, KT15, KM18a, KKK16, Kua15, Mai11, May12, MP15, MO20b, Oga14, Oza17, Par11, Viñ19, WJ20]. Fields [AW13, And12, AB09, AITA19, AC13, Awt12, BL09, Bao10, Bos09, Bou05, Bri11, Bru05, BD08, Cao11, CCL13, CLM08, Coh06, FG12, FP10, FG14, FS08, GML12, GP12, Ham13, Has10, Hoe10, JR13, Klo13, Klü12, Kow06, LZ11, LO12, LM11, LS09a, Mag13, MS12a, McG12, Naj12, Pap11, PS11, Pic10, Pot21, Rho09, Rob10, Roy05, Rüh10, SW06, Sua09, Tan09, Veg11, Vin14, VZ14, Vul10a, Wan12, Wid11, ZLZ10, Zie11, Zie08, AMM17, ANH14, AK15c, AZTM15, ARTZ19, Bae19, BS20, Bal19, BARCVS13, BH14, BE21, BGm22, BS16b, Cen16, CL19, CeM21, Dai16a,

– DLV20, DP22, FT15, Fuj20, GJR19, GM18a, GZ22, GL19, GG22, Ham16, Ham18, Har18, Hsu16, HSW14, Hum14, INST14, JLSW15, JS20, JK16, KO20, KMS21, Kat17, Kha19]. fields [KW18, KM12b, KN19, KMPS20, KW14, Kom17, Kra07, Kug22, Leb10, LS17a, LLM18, Li22, LM15c, MP16a, MP16b, May14, May19, Mei18, Mér20, Moh19, Mon14, Mor16, MOS14b, Nţ21, Ngu19, Ozd21, Pal14, PR17, PS17, PS19, PST21, PZ16, Pol14a, Pum20, QQH15, Rob18, Ros17, RS14b, Sal09, San09, Sch14, Shi16, Shl12, Shp19, Tak21, Tak15, TT14, Tra17, TT18b, TSB20, Wan17, Wit07, ZH16c]. Fifth [Liu12d, LW20, RL19, ZZ11]. fifth-order [LW20, RL19]. fifth-power [ZZ11]. Figurate [Toh13]. Filtrations [MO20b, Las17]. Finding [Bad17, Baj14, Cha18b, DD09, SS17]. Fine [Mat18, Lim15a, All09, And07, BW21, Yee09]. Finite [AW13, Ada12, Bao06, Bao10, BH10, Cao11, FG12, FG14, FS08, GML12, Hin19, LP13, Nat09, Pic10, RS06, Veg11, Vin14, VZ14, Yas16, Zie08, AMM17, Bae19, Bao08, BSM16, BGm22, Din21, EMS21, Fri16, HK18, KM18a, KKK16, LM18, LS17a, LLM18, Liu18, MP16a, MP16b, MP15, Mei18, Mér20, Moh19, Mur17, MS18b, Ozd21, Pot21, QQH15, Shp19, TSB20, WZ12, Wan21, YL16]. finitely [Peh16]. Finiteness [CEIK07, IT10, Ito18, Sam10, Wei15b]. First [IKS07, PsR21, Bro10a, CY18, GC17, Kim16a, MS15, QH19]. Fishburn [Str15]. fitting [San09]. five [LZ18c, Mal21]. Fixed [BZ13, Fri16, RPR17]. Flach [Gro20]. flatness [Zha17a]. Fleck [CH15]. Fleck-type [CH15]. flows [Mer22]. fold [BC18]. Fomenko [FM12]. fonction [JM10, Nic06, Tou09]. fonctionnelles [Phi11]. fonctions [LR22]. forest [HMST16, Nat15]. Forests [Nat17, Zub20]. Form [Das13, EG09, Har11b, Liu08, RSW14, Sah11, Agn22, APP13, AMPS17, Bel22, Cha08, Din09, Do17, FZ16, Hsu20, KK15, KK19, LR17, LN19, LH16, LR06, MZ22, SA09, Šiu16, Ye16]. Formal [MY13, Sai10, SXJ13, BGW12, HM06, Hyo15, Ma17, Ula19]. forme [Do17]. formelles [HM06]. formes [Tay17]. Forms [AALW08, AALW09, AAW10a, AAW10b, AK14, Alk07, Aza09, BG11, Boy10, Bra14, BKK06, BR11, CW07, CEIK07, Chi09, CC07, CK13, CWW08, Cou08, DR10, DH13, DS13, EG07, ES12, EH08, EV07, Fuk08, Gan10, Gek11, Gim13, Gri11, Gun06, GMR11, GY13, HJ14, Has13a, HKKL12, HK13, JSS14, Kan10, Kaz08, KL12, KMW10, KLR09, KM09, KR10, KM12c, Koh10, Kök13a, Kök13b, Kök13c, LV12, Mah12, MS12a, Mas07, Mor11, MS13, Moy13, Mui11, MM07, Oh11, OT05, Raj09, Raj11, RS13c, Res09, Roy07, Rüh10, Sha09, Tan12, Ver10, Vul10c, XY13, Zum11, AAR16, AA20, AW12, APW14, Ala14, AK15b, AAW16, AK17, AB18, BG15a, BV18, BP11, BSM16, Bir19, BDTT16, BT18b, BM21b, BB14b, BK15b, CF17, CJKM18, CL22, Chi19, CWR16]. forms [Dai16a, Dic15, DM20, EF16, EK20, Fli11, Fol21, FM20, Got20, Gra20, GR14, Gue15, GS17, Han17, Hei18, HL12, Hir22, Hu22, Hun18, JS19, JL17a, JL20, JLR17, Jon21, JKO18, Kan14, KPW17, KwK21, Kim22a, Kim22b, KM14, KE17, KW14, KMV20, KM18b, LS14, LLZ16, LX19, Mal20, Mal21, Mat19, MM14, Mor22, Ot20, Oli14, PR21, Pat20, PsR21, PP18, PSY17, PSY18, Pum20, RS14a, RSS18, Rez21, Rob18, RZ22, RT11, SS14, SG17, SPY18, SV15, She14, Shp18, Ste16, Su16, Tay17, Var22, Vep17, Wei15a, Won20, Xio16, Ye15b, Zha16]. Formula [AAW08b, Chu13, Dix11, Li12, LS09b, Mag13, MMW11, SZ07, VS10, BP18, CbL21, CC16, CG19a, DM13, Gir19a, GBL15, Ham16, Lam16, Li15, LY09, Pat20, RR20, Sch14, SK13, UW14, Vil18b, Wan20, Yam15]. Formulae [PPT12, BSM16, Sed15, WW18]. Formulas [BR11, El 11, GML12, Has13a, LY05, Mac12, Mas08, OEL13, SS07b, YXJ13, CFTZ14, ELO17, EL18, FY13, GLZ15, KR16, Kur21, LZ15, OS21, San15, WC19]. Four [Cai10, LP13, MR06, OPY08, SX14, Wit10, Xia13, BP11, FM20, Shi16, Sun19c, WS16, WS17, ZL19]. Fourier [AAA15, Alk07, BH10, BGW11, CK13, EG09, GS17, Hsu20, HY18, HKKL12, JV21, JL17a, yK21, KK18, Koh10, KM14, Kri16, KMV20, KM18b, LX19, MM14, MM07, Oli14, Raj09, SPY18, Wei15a, WZ09, Wil12, Won20, YXJ13]. Fourteen [AAW10a]. Fourth [Tam14, WG11, LZ18b, TZ09, Van21b]. Fourth-Order [WG11]. fractal [LM15a]. fractals [NSS15]. Fraction [BO11, Cha10b, Lin13a, SXJ13, BI21, BMW06, Cha10a, CH05, XY11, Zha15b]. Fractional [BS16a, Buc11, Dub09, Kum09, Fre19, Gir15, Min22, Nat15, Sch18b, Zhu18]. Fractions [BKB13, Gir11, Gir12b, Hay10, Zha12, Zha05, ZT13, DFG^{+}14, GM18b, HYZ17, HWh20, Jam16, LM16, MAM06, Mer22, O’S16, Oye16, Par22, Rob20, SW14, Van16b]. Franel [Sun18, WS19, ZY20]. Free [DU10, HN11, Kad12, Aym22, Bai16, Baj09, Dud14, FGT15, Gib14, JLSW15, JL22, Kei13, Liu18, Mer11a, PQSW14, PHLS19, Ros17, Wan21]. freeness [KPW17]. French [Amo07, AMO17, BD10, BF10, Bil11, DP08, Do17, Gil17, HM06, Hua17, JM10, Kra07, Kra15, Mai11, Mig15, Nic06, Phi11, Pit11, Ric13, Riv09, SS10, Sun17b, Tay17, Tho06, Tou09, Zha14]. fresh [BP17]. friable [Dra17]. Fricke [AAW10b, Sak11, Sak14]. Fringe [AMMS17]. Frobenius [GHK^{+}15a, BJ19, Bae19, CG19b, DU10, GHK^{+}15b, GT17, Lin14, RPR17]. Fu [RS13a, Yao15]. Fuchs [Tan14]. Fuchsian [Fri16, Vul09]. Fujii [PT19]. full [Col12, Har18, MT18, Sha17]. fully [LDSM19]. Function [AALW08, AZ05, Alk12, And12, BB09a, BBS08, BL09, Ber07, Bet10, CEP12, Cha10b, Cou08, EK07, Far08, GK13, Gar10, GJS14, Ham13, HN07, Hay14, Hea14, Ivi05, Kaa11, KW12, Kow06, Lau13, LZ12, Li09, Lin13b, LS09a, Liu12c, Pap11, Rho09, Rod14, Roy07, RSW14, Tam14, Tan09, Wan12, Xio11, ZLZ10, AK15a, AD18, Bae19, BS20, BARCVS13, BH14, BK16, BCSX20, BJMV14, CC16, CHJL21, Che14, Cho20a, CH15, DK21, – DLV20, FY13, FG22, GM18b, Gar08, GC17, Ham16, Ham18, Har18, HSW14, Hum14, Ivi16, JLSW15, JKK16, JM10, KW16, KMS21, Kar19, KT15, Kim21a, KTZ16, KTZ18, Kua15, LLMA16, Li15, lLW21, Li22, LLZ16, MCW19, MR15, MP15, Mer11a, Mor16, MP21, NH20, Ngu19, Nic06, O’S16, OS17, Ono21, Pal14, Pat19b, Pum20, RS11]. function [Ros17, SS19, Shi16, Suz17, Tou09, TT18b, Tsu18, Vil18a, Wan17, Wit07, Xio16, You12, ZH16c, ZMS15, Zue18]. Functional [EH08, KMT11, Ngu11, Pal14, Suz05, BZ15, Hei18, JP14, Phi11]. Functions [Ada12, AJKM09, AZZ05, And13, And12, BB09b, BST10, Bou11, BM11, Bui12, Bui13, CLN05, Cho13a, CZ10, Dja13, Dum09, El 11, FW10, Gil13, GML12, HRL11, Har07, Has10, HN10, HN11, Hub11a, KP14, Kad12, Kat10, Kim10a, KM05, Koc08, KMT11, LZ11, LW13, MN13, Min12, MY13, Mui12, MS09b, OT05, Ono08, łPS13, Pat10, Pil09, Pol14b, Sch10, SZZ13, Suz05, Tam14, Tan12, Tem10, TT09, Veg11, ZW14b, Zor13, AB19, Agg21, AD16a, Alo19, Ass21, Aym22, BG15a, ES20, BG15b, BB18b, BB20b, BD22, BB14a, BK15a, BT18b, Bor18, Bro10b, BS19, BS17, Bur21, CGG15, Cha20, Che18a, CY18, Che18c, CK17, CP18b, CG13, CGH18, Dah18, Dau14, Def15, DE21, DM19, DK21, Dja11, – DLV20, Ell15, Eve22, Fol21, FT18, FJ20, GY16]. functions [Goo17, Gug21, Hah21, Haj20, HM16, Hau18, HLN19, HK14, Hir22, HL11, Hou21, HSZ20, HHM21, Hun18, Jia20, wJShY21, JL17b, JQ21, yK21, Kas19, KP16, KS16, KE18, Lan19, Lei12, LM15a, LS19, LW20, LY09, LZ14, Lou15, MNZ19, Mad12, MU18, MP22a, MSV18, Mem17, Mem20a, Mem20b, Mér20, Miy15, MV14b, Mur15, NMZJ22, OS11, OO17, Par18a, Par20, Pat19a, PsR21, Pen08, Pey20, Pi20, PT15, Rei21, RGHK20, SS14, Sah16, San09, Sch18a, SS21, Sch15, Sed15, SR19, SST19, SY17, SSS22, Str22, Sun16c, SY19b, Sun19b, TX16, Toh08, Vää15, Van21a, Van12b, Van16a, Yan19, You16b, You17, ZL18b, dS18a]. Fundamental [AM17, LL17, AK15c, Rau16, Sar22]. Further [MZ22, SY19b, Zha22a, Guo21, Sun17c, ZS20].

Galois [Amo21, Amo18, AD16b, Bar14, Ber18, Bru05, Buc11, Car11a, CKW13, hC21, Cop20, CR18, DR16, DD09, Dor20, EJ11, EJ19, Fra21, GNS06, GM18a, GG22, Gir22, GP12, Hin18, IIO20, JM11, Klü12, KRS18, Lim16, Nţ21, Pan11, PRVS08, Ray21, Ray22, Ric13, Sal09, SS15b, SS10, Sha17, SST19, Tay17, Van18]. Galois-invariant [EJ19]. galoisienne [Ric13, Sod21]. galoisiennes [Tay17]. Game [NQ08]. Gamma [Kat10, Kas19, Zha17c]. Gap [PSZ16, HR21]. Gaps [Alk07, GL11, AB19, Ben15, FGT15, KK18]. Garvan [Pat15]. Garza [Höh11]. Gauss [ASV10, Aka14, ASD16, BBCZ05, CD11, Gur11, Has10, LS17a, Liu12b, Moh19, NMZJ22, Van07, ZCX22]. Gaussian [BK12, BK13, Kal18, KP16, KL09, McC12, MM11, MV14a, Sad16, SZZ13]. gcd [CG22, HLT20, Zhu22, Zhu22]. gcd-closed [Zhu22]. General [ACH05, BW21, CH09, KTT06, MS05, Hou21, Ros18]. generalised [Bal19]. Generalization [BTW06, Fu11, Kuo09, LS09a, RD13, Rol11, Yan13, BDTT16, BO19, Isa21, Kob16, LHK18, Mur15, Qua19, ST18, Tan14, YF18, ZC17a]. Generalizations [Bao15, OEL13, Sch14, Tót13, Zha22b, Mis17, SY19a, Zha22a]. Generalized [BCDY15, Che17, Cho18a, CCRT14, Dil20, ELY05, FG12, HRL11, Kat10, KK15, KL12, KLR09, KM09, KR10, KM12c, KL17, KY18, Lin14, Mat12, Ono08, Raj09, Raj11, Sch10, Shp14, Toh08, YX12, AMMS17, Bha20, CC17, CG19b, DFG^{+}14, DL20, GS18, HS22, HKLP09, Hsu20, JOS19, JO20, KO20, Kat17, KwK21, Koy22, KR16, MNZ19, NMZJ22, OS11, OS17, Şia15, Tak21, WW18, Xu19a]. Generalizing [Fel12, GM18b]. generate [KT15, Yas16]. generated [Amo07, Peh16, Pot21, Yas15]. Generating [AJKM09, BB20b, BST10, Pat10, BB18b, MP22a, Mon14, O’S16, RC17]. Generators [CR18, EMG08, Jah10, Pap11, Yas13]. Generic [HM10, Mis16, Kel17]. genre [Tay17]. Genus [BARCVS13, BGW11, Leh08, OPY08, RSW14, AAW16, BCF21, Bro10a, Col12, Gar18a, LW19, Mau17, RPR17, Tay17, Zan16]. Geodesic [Mer22]. Geodesics [SS07a, SS07b]. Geometric [AG12, Mah12, Pon06, Che15, DWW17, Has13b, KM12b, N18, San14, Sar22]. geometrically [Fri16, Liu22]. geometry [Edi05, LM15b]. George [And21c, Che18b]. Germain [Leo18]. Ginzburg [FZ16, ZW14a]. given [Bud20, JLSW15, KT15, Kom17, PS17, Zoe19]. Glaisher [Van12a]. Glaisher-Type [Van12a]. Global [LZ11, Rüh10, Vig12, DP22, Kru16, Leb10, PS22, Sal09, Vää18, Wan17, ZH16c]. Goldbach [BB06, CGZ21, Li16, LZ18d, LZ18c, Liu12d, Liu16, Pol11, ZL18a]. golden [ABCM14, PS15, LS09b]. Göllnitz [XY11, CH05]. gonal [JO20]. Goncharov [CDc20]. Good [IFD08, Vol10, RT17, Tak15, Zor11, Rob20]. Gordon [CH05, MMO08, XY11]. Gowers [Liu11]. Graham [US13]. Graph [FG14, AAR16]. Graphs [LS12, Gro18]. great [Ska21]. greatest [FF15, FZ18]. Greenberg [Do17, Fin14, Kat17, Tak21, Tho06]. GRH [DGM19]. Grid [HM12]. Grimm [LS06]. Groshev [Gho11]. Gross [BKY19, BKY20, Par11]. Grothendieck [Col12]. Group [Car11b, LM14, PRVS08, Roy10, AM09, Amo07, AITA19, AMO17, ATT^{+}16, Baj09, CM20, Col12, Dor20, FN14, GK18a, KRLT20, May12, Mil13, NT14, SS15b, SS10, Tsa17, Van16b, Won18, YL16]. groupe [Amo07, AMO17, SS10]. Groups [Bal08, Bis15, CLM08, CF09, Chi09, Del05, Dum09, EK07, GP12, Hoe10, Jah10, KRS10, Klü12, Lev06, LMP10, LP13, Mat12, Mau12, Pap11, QY11, Res09, Sai10, Sak11, Şen12, Vig12, Vul09, WS10, ZY08, Ari13, AZTM15, BZ15, Bro10b, CeM21, Col12, DM21, EGP21, FZ16, FRcT20, Fri16, Fuj20, GL19, GG22, Gir22, Ga20, Hin18, Ito18, KPW17, Kit13, KRS18, LŞ20, LL22, Lim15a, Liu18, Mat18, Mil18, Nat11, OLG19, PQSW14, PHLS19, Sak14, Sal09, Sar22, Sch21b, SST19, SY14, WZ12, Wit07, Zem17, ZH16c]. growing [MS15]. Growth [LM14, San18b, LŞ20]. guess [PsR21]. Guillera [CC21a]. Gun [Cha12]. Guo [NP18, Vse14]. Guy [CGPY15].

H.2 [Guo21]. H.3 [pFG22]. Habiro [LM15b]. Hadamard [Dja13]. Halász [IW16]. Halberstam [Smi11, Ano15b]. Half [Ala16, CK13, HKKL12, Koh10, Che19, Dai16a, DM20, JM16, MM14, Moo19, PR21, Wal17b, Xio17]. Half-factorial [Ala16]. Half-Integral [CK13, HKKL12, Koh10, Dai16a, JM16, MM14, Moo19, PR21, Wal17b]. half-plane [DM20]. hall [ACS09]. Halton [HKLP09]. Hamme [Guo19b, Guo21, WY20]. Hankel [CK21, DJ22a, HW15, ZS18]. Hankel-type [ZS18]. Harborth [MORS16]. Hardy [CK14, LZ08, Lan09]. Harmonic [CZ09, CWR16, Li13, Meš12, Ros13, XC10, Zha08, Bat17a, CSJ17, Che17, Che22b, FJ20, GS18, Jar22, KP19, MS16a, MW19, MTWZ17, Tau18, WW18, Xu19b, YC17]. Harmonious [KLPP15]. Hashimoto [KRY09]. Hasse [Cla08, Rom19, Tow13]. Hauptmodul [HY18]. Hauptmoduln [BL15]. Hausdorff [BLW21, NSS15, TLZ22]. hauteur [Hua17, Mig15]. Having [IFD08, CGG15]. Hecke [AJKM09, Amo21, And13, Bao19, BB14b, Bro10a, CEP12, CW07, CL22, Dau14, DD09, FW21, GK18a, GJS14, Ghi11, GS17, Hei18, Hyo15, IKS07, JV21, Kim22a, OO17, Pat20, Pen19, Pey20, Res09, Sah11, STW10, TT18a, Ter13, Val14, Wal06, Wal08, Wal17a, Wal17b, WC19, Wie09, ZS20, Zha22a]. Hecke-Type [And13, GJS14, ZS20, Zha22a]. Heegner [AAD11, Liu12a, Mat18]. Height [LW13, Pon06, SS07a, SS07b, Thu08, Via10, CF22, ELO16, Hol19, Hua17, KT15, Mig15, MS18b, VY13, Wan15b]. Heights [Abo08, BM12, Nat09, San18b]. Heilbronn [Won17]. Heini [Ano15b]. Heisenberg [Hyo15, Van16b]. Hensley [Wan12]. Hermite [Bac08, Mue12]. Hermitian [HK14, Hir22, Vul10b, Vul10c]. heuristic [MP16a, MP16b]. heuristics [RS14b]. hexad [EJ19]. Hida [JP14, LV12, Ray22]. Hida-families [Ray22]. high [BJ19, Bae19, MT17, Shp19, Ska17b, Ska19]. Higher [Cha05, HR21, Ula10, Wal13, Ade18, Ade21, Cas12, Cho18b, Cob21, CWR16, Dic21, Hol19, Pum20, Qua19, RGHK20, Zan16, Zho18]. Higher-dimensional [HR21, Hol19]. higher-order [Cas12, Zho18]. Highest [Gou18]. Hilbert [Van16a, Agn22, Aza09, ARTZ19, Bir19, Bor15, BP17, CW07, Chi19, Dem20, Ehl10, FW21, GH14, Hun18, JM11, LM15c, Su16, Tan12, Van12b]. Hirschhorn [CLN05]. Hodge [Anc17, MR06]. Hoffman [HS19b, Wak12b, Wak17]. Holomorphic [Gri11, HJ14, JSS14, KN09a, MP10, Han17, Ste16, Zha16]. holomorphy [MS16b]. Holonomic [PsR21]. Homogeneous [Bor15, Rey13]. Homology [Şen12, AD16b, LŞ20]. horocycles [Aka14]. Hua [LZ18a]. Humbert [Cou08]. Hurwitz [Sch14, Bao14, Bao15, DK21, Kri10, Lau13, Ono21, Rob20, Shp14]. Hyper [BU11b, Gur08]. Hyper- [BU11b]. Hyper-Kloosterman [Gur08]. Hyperbolas [GM12]. Hyperbolic [Şen12, BPZ14, Che18c]. Hyperelliptic [Gla09, Hub11b, MP22b, Sai10, BJ19, Cre18, EGP21, GR11, Gar18a, Jed22, Jun14, Sad16]. Hypergeometric [BK12, Goo17, Has10, HN08, HN10, HN11, KM12b, Kom13, McC10, MR06, Veg11, BG15a, BK13, BS15, BJMV14, Kal18, McC12, MP15, Miy15, Sad16, TSB20, ZL18b]. hyperplanes [Gho11, Pha22]. hypersurface [Le15]. Hypersurfaces [MR06, Pon06, Shp14, Goo17, HK21, Mig15, Miy15, Qua18]. hypothèse [BD10]. hypotheses [DP22]. Hypothesis [Olo09, CQh22, Dud15, WY21, BD10].

I.2 [Guo19b]. Ideal

[Buc11, FP12b, FHL^{+}13, Gra11, Lez12, Pál07, Vig12, Fuj20, Hsu16, Kom17, San09, Zam16]. Ideals [Bos09, Rüh10, Bat17b, BS19, Deb19, PZ16, SD20]. idéaux [Sod21]. idempotent [Wan21]. idempotent-sum-free [Wan21]. identical [Yan19]. identification [Mon14]. Identities [AALW08, BS12, Ber09, CLN05, CH09, GY07, Hua14, JL18a, Kim10a, Kim17b, Kop08, Li13, Liu12c, Sil07, Sun12, Sun17b, Tót13, Xio11, Yee09, AD16a, ACS09, And21c, Bao19, BN14, Bat17a, BCSX20, cC21b, Gug21, JMV16, hKS21, KP19, LW19, LW20, Mac16, MS22, Toh08, Tót18, Tsu18, Wan16, WC19, You16b, ZS20, Zha22a]. Identity [AAW10b, Coo06, Kim10b, Row10, All09, All17, cC21b, CHJL21, CBJ22, HS19b, JKK16, KL14, LHK18, Mér20, Tsu18, ZH16a, ZC17a, Cha10a]. Igusa [dS18a]. Ihara [Klo13]. II [CW19, ABK19, Bor20, BV09, Cha05, CP18a, DM19, Do17, Dum09, EGL21, FHL^{+}13, GY07, Jed22, Kei21, KR10, KTZ18, LP13, LM15a, LX19, MAM06, MO20a, Mun10, PT14, SS18, SX14, TZ12b, Wal08, WS17, ZLZ13]. II.16.12 [BI21]. III [Che22c, FK11, Jia20, MS12b, NV10, Ray21, Sun13]. Ikeda [Hay14]. Ikehara [RD13]. Illusory [FI05]. image [Pea22, Sha17]. Images [Amo21, Tay17, AB19, Tay17]. Imaginary [Klo13, Vul10a, CeM21, Dai16a, GJR19, GZ22, Kom17, LM15c, Mor16, Oga14, Sze19, Tak21]. impaires [Nic06]. impairs [JM10]. implies [cC21b]. Improved [Cre18, Moh19, Zhe15, Pac19]. improvement [Cho20c, SY19b]. Incomplete [BH13, Aka14, KL17]. incongruent [Eld19]. indecomposable [Sha17, Wan17]. Indefinite [Has13a, Vul10c]. Independence [AAD11, Bun12, CO09, GMR11, Ter13, TT18a, Vää16]. Index [Ano05, Ano06, Ano07a, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13, DD10, MS13, SX14, SW06, Xia13, ZLZ13, Ano14, Ano15a, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, Ge18, JKS17, KY20, KW18, Liu18, Mac16, SZZ18]. index- [Liu18]. Index-Conjecture [SX14, Xia13]. Index-Dependent [MS13]. Indices [AK11, ABK19, TY13, CWR16, KM14, San18a, Shi16]. Indivisibility [Chi09, Dai16a]. Induced [BCH08, ZJ19]. induction [DS21]. Inductive [El 11]. Inequalities [AsMS20, Har11b, He20, LRS14, Liu08, Min12, Cio20, hKS21, Vil18a]. Inequality [Bac08, BZ05, Hal12, BS20, Cai18, Deb19, Dil20, EMS21, Haj15, Mu17, Som22, Tre15b, ZL19]. inert [He20, Kra15]. inertes [Kra15]. inertia [Cop20, KRS18]. infinitary [BO19]. Infinite [AAAW17, BN14, BD22, CZ09, JM11, Mas08, MW11, RB18, ACX18, Dau14, Kom17, Leb10, MZ22, Xu19b, Zub20]. Infinitely [ABP09, BCU14, FT15, Gir18b]. Infinitude [MW16]. Infinity [Gek11]. inhomogeneous [Ram17]. Inner [YZ22, AMPS17]. inspired [CBJ22]. Integer [AALW09, HKN10, JQ21, KPT08, Liu08, LY05, LMS^{+}22, Mas08, Wil10, YX12, ZLZ13, APW14, Bad17, Bat17b, Dra11, HKN11, JL20, LR19a, Mac17, RS14a, SZZ18, Ula19]. Integer-valued [JQ21, Mac17]. Integers [GZ11, Gur11, Hu13b, JV08, Kan10, KMT11, LS06, MM11, Ngu11, PV12, Sha09, SZZ13, TZ12a, TZ12b, XY13, AD18, BK22, Ben15, CGG15, Cha18a, Cha08, Cho15, DJ22b, Eld19, Fla19, GK18b, Gou18, HLT20, Hu13a, JM10, KwK21, Kit13, KE17, Liu14, LMO^{+}19, Mos15, MP21, Rib11, Sin09, Šiu16, Ska17b, Ska19, Som22, Suz22, Szc15, Szc16, Vep17, WJ20, XMT16, Yan19]. Integral [ABP09, Ben13, Cen16, CK13, EK20, HSW14, HKKL12, Koh10, LS12, MS12a, Pic10, TX16, ANH14, AMSV21, BB14b, CGG15, Dai16a, GR11, JM16, Jon21, Kim22b, Liu22, MM14, Moo19, PR21, Sha14, Wal17b]. Integrality [Pil09, Sch21a]. Integrals [FP12a, AKMR12, Dau14, DA18, DA19, Li20a, Zor11]. interpolated [Wak17]. Interpolation [Mor11, HL12, Mor22, OS11, Riv09]. Interpretation [KM12a]. interpreting [SR19]. intersecting [ZW14a]. Intersection [BM12, HM10, HB17]. intersections [Zha21]. intersective [Li22]. interval [Bor18]. Intervals [AZ05, Bai13, Baz11, BH13, Cha05, Cha06, Cro07, Dub18, Hu13b, LZ08, Lan09, Mat10, RSW14, Smi13, CTZ16, DGM16, DGM19, Kim22a, Mat16, Sed15, TZ09, Tol06, Yao18]. Invariance [DD10]. Invariant [EJ11, Kab10, Sch18b, BV18, CM20, EJ19, Kim16a, Kum21, Sha14, ZW14a]. Invariants [IN22, Mat08, MP10, RS14b, Ari13, BS10a, Bar13a, GLPW18, INST14, Kle17, KK12, Leb10, PS17, Paz19, XLD22]. Inverse [Chu13, Liu12c, MORS16, Van18, BB20c, Liu19a, QH16, Tsa17]. Inverses [Bai13, BH13]. inversion [CH15]. inversions [CDHS15]. inversive [Lin22, NW05]. investigation [CCH^{+}19]. involutions [HK20]. Involving [GML12, GZ11, GZ12, Har11b, Kim10a, Kum09, WC11, Xio11, YX12, Alo19, Apa18, BD22, CSJ17, Cai18, CG18, Che22b, GN19, Hah21, JV21, KP16, LZ22, LZ15, MCW19, Mac16, MS16a, MW19, MTWZ17, NR16, Sun16a, Sun16b, Sun18, Tau18, WW18, lWfS22, Xu19b, ZH16a]. Irrational [LS05a]. Irrationalité [JM10]. irrationalities [EW15, Oye16]. Irrationality [LT14, Les16, ZZ21, JM10]. irrationals [Mur15]. Irreducibility [BBCM13, CL11a, Nţ21, BB20a, Bil11]. irreducible [BSK17, Din09]. Isogenies [MM13a, DG17, Orr17, Paz19]. isogeny [Ric13]. isomorphisms [SG17]. Isoperimetric [Bal08]. Iterated [FHS11, HHP09, Li20a, MMR11, AKMR12, BS17, PS15]. iterates [Hin18, IJO^{+}21, LTZ20]. iterative [XY11]. IV [AS09, Nat11]. Iwahori [CR18]. Iwasawa [AH07, AMO17, BS10a, Bar13a, Bar13b, INST14, IMO13, Kle17, Kum21, LZ14, Mat08, Pit11, Sch14, Val14, Wit20].

Jackson [Vil18b]. Jacobi [ASD16, Álv14, ES20, BR11, CWR16, FY13, Gir11, Gir12b, GML12, GN19, JS19, Kra14, Oli14, PPT12, RS13c, Sch18a, Toh08, Tsu18]. Jacobian [Sai10]. Jacobians [Álv14, Bis15, Cre18, FS08, Jed22]. Jacquet [Bir19, Fre12]. Jentzsch [CL11a]. jointe [AST22]. juggling [Tou19]. Julia [GR19]. Jumping [GL11, Hol19].

Kac [KL12]. Kaneko [You16b]. Keith [XY14]. Kemperman [Lev06]. kernel [San09]. Kernels [YZ22, Pit11]. Keys [Nit09]. Khintchine [Alv21, DK06, Gho11, Oli22]. Khintchine-type [Alv21]. Kim [CDc20]. Kind [HZZ12, KO10, Mer11b, Ade21, QH19]. kinds [CFTZ14, HW18]. Kirch [Szc15]. Klein [AAW10b]. Kloosterman [BEH10, BH13, CT18, CE07, Gur08, Gur11, Kel10, Shp18, ZH16a]. Koblitz [Zyw11]. Kohnen [IW16, Su16]. Koike [Gug21].

Kontsevich [Vs21]. Korselt [Alr14, BEP10, EG12, Wan18]. Krattenthaler [Vse14]. Kummer [Gra18, KRY09, PS19, PST20, PST21, She14, Yas13, Yas15]. Kurokawa [Mat17b]. Kuznetsov [Mag13].

l.c.m [San21]. lacunary [Kre17, Len17a, Len18]. Lagarias [WY21]. Lagrange [Cai10, Chu13, PMM13]. Lagrangianp [Riv09]. lagrangienne [Riv09]. Laguerre [BB20a, SS15b]. lambda [INST14]. Lambert [Agn22, BW17, CJKM18, LT14, Ye15b]. Lang [Pan11, San18b, VY13]. Langlands [Bir19, FRcT20, Fre12, Won18]. Laplacian [HS19a]. Lara [Har18]. Larcombe [Sun17b, Zha14]. Large [BZ05, CLMR09,

– DLV20, Hal12, Liu14, Liu19a, Pey20, BB18a, BS20, GZ22, GL19, GG22, KY20, KW18, MR12, Mos15, Pri09, Shl12]. Largest [SXJ13, BL13, CW19, Gir17, LLL22, NY21, SYZ14, Zel19]. last [KO20]. Lattice [IFD08, Kel17, Now16, BPZ14, Bor20, Cha15, HL20, JL20, PZ16, ZH16b]. Lattices [FP12b, FHL^{+}13, GNS06, Har11a, HKN12, IFD08, KM06, Kom09, MS12b, Mue12, NV10, Neb13, Sua09, Hou17, KN19, Leo15, Liu19a, Mou17, Wan17, Zem17]. Laurent [MV16, SXJ13]. Law [Li11, CHL19, JTY16]. Laws [KL09, Kra14, Ram17]. lcm [CG22, HLT20, Zhu22]. lcm-sums [CG22]. Least [LW08, CPS18, Che22a, MY21, NJ22, RC17, Tre15b, YF15, Zam16]. Lebesgue [Mol05, Row10]. lecture [ACS09]. Left [Zub20]. Legendre [EG14, Mar14, Sun13]. Lehmer [Bon08, BŻ19, DP08, Liu09, McN13, MW16]. Lemma [CZ10, Klo13, Mer11b]. Lemmermeyer [ATIA20]. Length [LP13, Pon09, SX14, Xia13, HZ18, Nat11, TYZ16]. lengths [GZZ15]. Lenstra [RS14b]. Leopoldt [Gra18]. Lerch [CM12, Che19, FWX21, Xio17]. less [Gou18]. Lev [Hui18]. Level [Boy10, HJ14, Sak11, SFM17, Bro10a, CWW15, Dic15, FM20, Gil17, Gui21, Han17, Mat19, Miz08, PT14, PSY17, PSY18, Sun15a, Tak08, VX18, Wal17a]. level-lowering [Tak08]. levels [Cho20a, Sak14]. LeVeque [Som22]. LeVeque-type [Som22]. Li [OS11, OO17]. lie [Hyo15]. Lift [Ehl10, Mat17b]. Lifting [CF09, SV15, Fin20]. Liftings [Fin12]. Lifts [Hay14, PR21, Sun15a]. Like [BST10, CCS10, PPT12, Gui20, HW15, LV17, Rei21]. likely [BSK17]. Limit [Rau16, Rod14, CbL21, CK17, Fol21, LLZ18, MR15, Miz08, Zha21]. Lindelöf [DM19]. Line [Baj14, Far08, HM12, Nat09, CK20]. Linear [BHPR17, CO09, DH13, EG07, HHP09, HK13, KP14, Kan13, Kaz08, KMW10, Mas07, Smi21, WG11, ABCM14, Bat17b, Bel22, BKS16, BK15b, CK21, Dub18, KTZ16, KTZ18, Kug22, Nat15, Pol14a, Ros17, RL19, Sch18b, Sun19a, WS16, WS17, ZC17b]. Lines [HM12]. Linnik [CPS18]. Liouville [AD16a, El 12]. Lipschitz [Mem20a]. lisses [Mig15]. list [Che22c, Kur21]. little [LS05a]. Littlewood [BHV11, LZ08, Lan09]. Local [BD08, BM12, Čes16, Dic15, Gam14, Jon21, Ked05, Kle17, LS17b, MY13, Olo09, Pic10, Pon16, Rüh10, Tu11, CHL19, DS21, DP22, FW20, Hir22, JK21, KRLT20, Kru16, Las17, LP14, PS22, Smi21, Zha17b, Hua17, Mai11]. local-global [DP22]. locale [Hua17]. Locally [Sha17, BCF16, BCF21, HK21, Kid16]. Location [XZ21]. locaux [Mai11]. loci [KM18a]. Log [LY12, SSU21, Zha14, Dau14]. Log-Behavior [Zha14]. Log-Concavity [LY12, SSU21]. log-polynomial [Dau14]. Logarithmic [BK15a, LZ11, MMR11, Abo08]. logarithms [MU18]. Lommel [DK21]. Long [Har11a, YL16, Wan21]. longest [LY19]. look [Bar13a, GLPW18, Par11]. lost [KKL21, BKW13]. Low [NV10, Sak11, SS07a, SS07b]. Low-Dimensional [NV10]. Low-Level [Sak11]. Lower [AV19, KW16, Kim17a, Kna08, Mue12, Pon06, Via10, Amo07, BLW21, GY16, wJShY21, Pup15]. lower-order [GY16]. lowering [Gol07, Tak08]. LP [Mos07]. LP-Sequences [Mos07]. Lucas [ABCM14, BDGL15, BLMS05, KK15, LLS21, LS05b, Meš16, San17, Şia15, SW08, Sun14, ZC17b]. Lucas-type [SW08]. Lüroth [KM22, LLL22, SYZ14, SFM17, TZ21, TLZ22].

Maass [BT18b, Cho13a, Gue15, JV21, JL17a, LLZ16, Pey20, CWR16]. Macdonaldcg [Toh08]. Machin [LS09b]. Macwilliams [SZ07]. magnetic [LN19]. magnitude [LR19a]. Mahler [AV19, Amo18, DS17, DA19, LR17, Mem20a, Mem20b, Sam10, Sas15, TT18a, Vää15, Vää16]. main [Wit20]. major [Li16]. manifolds [BLW21]. Manin [Bal19, PS22]. Many [ABP09, BGm22, BCU14, BC18, CS21, FT15, Gir18b, MP22b, Sug15, Tro17]. map [PS15]. Mapping [MM11]. Mappings [Koh08, MP16a, MP16b]. Maps [DFG^{+}14, LW13, BT18a, CV17, JKM09, Sch18b]. mark [GM18b]. marked [BK11]. Markoff [AAR16, AA20, Bao14, Bao15, Shp14]. Markov [Vul09, Vul10b]. matching [LY19]. matrices [HP11, HK14, HL11, Mok20, Zhu22]. Maximal [Ngu11, Wan15b, Yu11, Dud14, KT20, Lin18, MS18b, SFM17, TYZ16]. maximally [Poë20]. maximum [Baj09, HR21, KS16]. McIntosh [Mos07]. Mean [And12, CG22, GK13, Har08, Rei21, Wat08, ZLZ13, ABCM14, CFTZ14, CTZ16, Fer22, GL17a, Jun14, KM12b, NMZJ22, Suz17]. means [BK15a, Has13b]. Measure [Sam10, AV19, Amo18, Bud20, Fla18, LR17, Mem20a]. measure-preserving [Mem20a].

Measures [Liu11, BS10a, Bar13a, DS17, GS16, Kas19, LLMA16, Sas15, Sch18b, ZZ21]. Meinardus [Par15b]. Melham [Toh13]. Mellin [Ivi05, Pat19b]. members [LLS21]. Memory [Hub11b]. Menon [Tót18, LHK18, ZC17a]. Menon-type [Tót18]. Merca [CH16, MS22]. meromorphic [Cha20, Oli14]. Mertens [Alk20, ApKK22, Hum14, Lam16, SS19]. metacyclic [May19]. metaplectic [Aka14, HL12, Lu20, Pet16]. Method [Bal08, SS08, AM15, BDTT16, BM21b, Cho18b, MS18a, SA09, XY11]. Metric [BHV11, DH13, KMW10, NRS20, HR21, WLWY22]. Metrical [HK13, ZMS15]. Mild [Sal09, Van21a]. Milne [LY05]. Milnor [Cha12]. Minima [BFCC09, Vul10c, CL19]. Minimal [EK07, LP13, MS12b, mPP22, SX14, Xia13, AAR16, BM21a, Fri16, KW18, KV19, ST19]. Minkowski [GM18b]. minoration [Amo07]. minus [Fuj20, MOS14a]. Mircea [CH16]. miscellany [BCDY15]. Mixed [BHV11, DH13, CT18, GL18, GL17a, JO20, LL22, Wu18, XMT16]. mixed-reduction [LL22]. Miyawaki [Hay14]. Möbius [CH15, NH20, SH08]. mock [BB20b, Bur21, CHJL21, Che18a, CGH18, Fol21, Gar08, HSZ20, yK21, LW20, Sch18a, SR19, Boy10, MOR21, ZL18b]. models [Bis15, BF19, Shl12]. modifications [BB20a]. modulaires [Tay17]. Modular [Alk07, Aza09, BG11, Boy10, Bra14, CW07, CKW13, Chi09, CK13, CWW08, DR10, DM20, Ehl10, EG09, Fol09, Gek11, GH14, Gim13, Gri11, Gun06, GMR11, GY13, HJ14, Har07, KTT06, Kaz08, KL12, KLR09, KM09, KR10, KM12c, Koh10, Kum13, Li11, LV12, Mah12, Mas07, MP10, Mor11, MS13, Moy13, Mui11, MM07, OT05, Pál07, Paz19, Raj09, Raj11, Sua09, Tan12, Ver10, Wie09, Zum11, Agn22, Anc17, BG15a, Bir19, BM21b, BB14b, BF19, Bru18, CF17, Chi19, Col12, Cow22, Dah11, Dai16a, Dic15, Fol21, FM20, Got20, Gra20, Gug21, Han17, Hat16, Hei18, Hou17, Hun18, yK21, Kan14, KPW17, KM18b, LS14, LN19, LR06, Mah19, Mal20, Mat19, MOR21, Mr22, Mor22, PR21, PsR21, Pea22, PP18, SS14, SG17, Sch15, SV15, Sha14, Su16, Tak08]. modular [Tay17, Van12b, Van16a, Won20, Zha16]. Module [Vol10, DP08, ZJ19]. Modules [Bar14, DWW09, KL09, Ber11, Cen16, hC21, EG14, IMO13, Kat21, Pit11]. Moduli [BZ05, Baj14, Hal12, KK10, KM12a, LZ12, BB18a, BS20, Edi05, Fow20, GZ22, Gir19b, JsKL21, Jen05, LMS10, Rif19, Sch21a]. Modulo [Ber07, CKW13, GJS14, Jam12, Kel10, Lin14, Vig12, XY14, AjW21, BD22, CSJ17, CM16, CHJL21, CDHS15, Che22a, Che22c, CGM15, CG19b, pCGyS20, Gir16a, JZ17, Kim16b, Liu19a, Mem20b, Peh16, RS11, RS13a, RSY18, She17, Str15, wS20a, Xio16]. modulus [NW05]. Mohanty [GF18]. Moivre [Gir22]. Moment [Bet10, Hea14, Kei13, Tam14, CQh22, Li17, LZ18b, Pat15, TZ09, ZZ11]. Moments [Cha05, Del05, Dja13, Wal13, AjW21, AsMS20, CT18, Che22c, CK21, EMS21, JS15, LRS14, LLZ16, Mao14, Str22, TX16]. monodromy [KRLT20, Ked05, Las17]. Monogeneity [Rob10]. monogenity [BE21]. monoids [SH08]. Monomial [LW13, Miy15, Hah21, Nom14]. monomials [AV19, DS17]. monotonicity [NSS15]. Monsky [CG19a, Mok20]. Moody [KL12]. Mordell [AM09, BDTT16, KRS10, LL22, Mat17a, Ono21, San18b]. Morita [QY11]. Morphisms [Paz13]. Morse [HW15, HWh20, Vää15]. Morse-like [HW15]. Most [Cha10a]. mother [Mer22]. Motives [Bar13b, Fra21, Las17, LM15b]. Motivic [Sca17, Sou10]. Motzkin [Len14]. moving [CWW15, Le15, Qua18]. MSTD [AMMS17]. Muir [BKB13]. multi [All17, Cas12, CG18, San09]. multi-dimensional [All17]. multi-quadratic [Cas12, San09]. multi-variable [CG18]. multicolor [DE21]. multidimensional [DFG^{+}14, GM18b]. Multifractal [CWW15]. multigrade [Cho20c]. Multilinear [Per12]. Multiple [BH14, Dub14, KO10, Li13, Mac12, Mar14, OEL13, Ono08, Rib11, Ros13, TY13, Wak12b, Zha08, AMSV21, BT18a, BP18, BC10, ELO16, ELO17, EL18, FJ20, GLZ15, Hof17, Jar22, JVW20, Kas19, KH11, KP19, Li20a, LP20, LQ21, Mac16, MTWZ17, Mur17, MS18b, NJ22, Ono21, Ono17, RC17, Sas15, SY19a, She22, Wak17, WV21, Yam15, Yas16]. multiple-zeta [BC10]. Multiplication [CCS13, Ngu11, WZ09, Gam14, Gil17, Kan14, LY21, Tre15a]. Multiplicative [ASV10, Bai13, BH13, CD11, LM14, OSW11, Par18a, Van07, CM20, DM21, Ell15, Liu19a, Nţ21, Pac19, Par20, Peh16, Sed15, WZWQ16, WHZ19]. multiplicatives [LR22]. Multiplicities [BM12, SZ21]. multiplicity [MY21, RPR17, RL19, Wal21]. multiplied [Gil17]. Multiranks [FT18]. multisections [Tür11]. multivariate [LM18, Ros17]. munie [Gil17]. Murata [BC11]. Murty [Cha12]. mutually [Ska21]. Mysterious [Hir06].

Nagell [Mol05]. Narkiewicz [GGW11]. narrow [Mil18]. Natural [Das13, Pon09, dS18a, CGG15]. nature [Che18a]. Near [Far08, Dub18]. Nebentypus [MOR21]. Negative [IKS07, Raj09, Bro10a]. neighborhood [Tou09]. neighborhoods [Oli22]. Néron [Bis15, Ga20]. Nesterenko [BO12, BV11]. Nets [Nat11]. Nevanlinna [Cha20]. newer [BP17]. newform [SVY20]. Newforms [BGW11, MY13, CG15, JM16, JTY16, SV15]. Newly [IJO^{+}21]. Newton [Bla11, CGG15, Ren19]. Nicomachean [hKS21]. nine [Vep17]. niveau [Gil17]. no [RLT22]. Noether [KM18a, Les16]. nombre [Amo07]. nombres [Kra07, Kra15, Mai11]. Non [Bal08, Bou11, Bru18, Bui12, CY18, CWW08, Cow22, DD10, Dem20, Fra21, GGW11, Gra11, Jar22, KN09a, KMT11, LW08, Lei12, MP10, MR12, Mui11, Mui12, Oye16, Şen12, Tan12, Tu11, Via10, Vol10, Wit20, Bro10b, CG19a, Cop20, Dah18, DFV13, Gar18a, Hsu16, KM18b, Liu22, Mr22, Nom14, Pat19a, Pha22, RSY18, SS10, Tre15b]. Non-Abelian [Bal08, SS10, Cop20, Nom14]. Non-Arithmetic [Şen12]. Non-communtative [Wit20]. Non-commutative [Lei12]. non-congruent [CG19a, RSY18]. Non-critical [Bru18]. non-cyclotomic [DFV13]. Non-Dense [Via10]. non-geometrically [Liu22]. Non-Holomorphic [KN09a, MP10]. non-hyperelliptic [Gar18a]. Non-Invariance [DD10]. Non-openness [Fra21]. Non-periodic [Oye16]. Non-Positive [CWW08, KMT11]. Non-Principal [Gra11, Hsu16]. Non-random [Cow22]. Non-rational [Dem20]. Non-Residue [LW08, Tre15b]. Non-Semisimple [Vol10]. non-split [Mr22]. non-subdegenerate [Pha22]. Non-Unique [GGW11]. Non-Vanishing [Bui12, Mui11, Mui12, Tan12, CY18, Jar22, Bro10b, Dah18, KM18b, Pat19a]. Nonanalytic [Tay09]. nonbases [Lin18]. noncongruence [LŞ20]. Nonconsecutive [Hay10]. nondegeneracy [Sug15]. Nonexistence [Bru05, Fin12, GM18a]. Nonlinear [EMW06, EMG08, KPT08, OSW11, Yao18]. Nonmaximal [AAD11]. nonnegative [Yan19]. nonresidue [CPS18]. Nonresidues [TK10, BG17]. Nonsolvable [Bru05, Rob11]. Nontrivial [GK13, Wid11]. Nonvanishing [AZ05, DG13, Dja11, Mer11b]. Nonzero [MN13, KK18, KLkO21]. Norm [Bel22, Cau20, HKN10, Har11b, Lez12, Liu11, McG12, PS11, Gro20, Gro18, HKN11, Mah20, Sze19]. Norm-compatible [Cau20]. Norm-Euclidean [Lez12, McG12]. Normal [Pic10, Čes16, Zha21]. normality [FF15]. Normalized [Gra18, Pon06, Via10]. Normalizers [Zem17]. norms [Ste16]. Note [And12, BE13, Cha06, JSS14, Koh10, Liu12a, LV12, MP12, Mat10, Ska17a, US13, WZWQ16, Zho21, Aym22, Bha20, CL22, Dai16b, hF22, FP15, Gro18, HM16, HYZ17, IW16, wJShY21, Jun14, Kan22, Kri16, KP19, LS18, Liu09, Mat16, MQ18, MP21, OO17, Sah16, SYZ14, Tür18, Wu17, Yan17, You16a]. notebook [KKL21, BKW13]. Nottingham [Ell15]. nullity [Lim15a]. Number [Akh12, AALW09, AK11, Bao10, Ben13, Ber09, Bos09, Bru05, BO12, CCL13, EMG08, FHS11, FG12, GML12, Has13a, Kan13, KTT06, KW18, Klü12, Kol15, Lee10, LO12, LM15a, Mag13, Mam10, MS12a, MMO08, PS11, Rol11, Sha09, Wak12a, Wid11, YX12, ACX18, AV19, APW14, Amo07, ABK19, Aym22, ARTZ19, Bal19, BP15a, Bao08, BB18b, Baš12, BE21, CFTZ14, CFM18, Dai14, DS17, Dra11, Eld19, FT15, FK11, GZ22, GL19, GG22, Gir18a, HS21a, Has13b, HP11, Hsu16, HK20, JK16, KO20, Kei17, Kha19, KKK16, KwK21, Kim22b, KN19, Kra07, KP19, Kug22, Kum21, LDSM19, LH16, LX21, Mai11, May12, Nţ21, Nat11, NT14, Oga14, Par11, PR17, PS19, PST21, PZ16, Pol14a, Pon16, PS15, RS14a, RSS18, Rey16, Rob18, RPR17, San09]. number [San15, Sch14, SZ21, Shl12, SD20, Sto21, Sun19a, Sur20, Suz22, TT14, Tra17, Vep17, Viñ19, WZ12, WS16, WS17, WJ20, XMT16, XLD22, Zel19, ZW14a, Zha15a, Zha21, Zho21]. number-theoretic [CFTZ14]. Numbers [AK12a, ACH05, AB09, BCH08, BEP10, CCS10, CT13, Cha13, CW12, CZ09, Cro07, EMW06, GZ12, Har08, HM12, Höh11, HZZ12, Kan13, KL11, Kob14, Kök13a, Kök13b, Kök13c, Kom13, Lan09, Li12, LY12, Mat10, MM13a, Meš12, Nob12, Olo09, PS11, Pon09, Roy05, Toh13, XC10, YX12, ABCM14, Ade18, Ade21, AM17, AAW16, Alk15, ACB19, BB14a, BCSX20, BL13, Bud20, BŻ19, CCZ15, CMS20, CS21, CF22, Che17, CG19a, Che22b, Cho18a, CGPY15, Dai16a, DL20, Dra17, DA18, EF16, hF22, FGT15, Gib14, GHK^{+}15a, GHK^{+}15b, GR19, Gir14a, HS21b, Has13b, Hic14, HK18, JL22, JW17, JOS19, JO20, Kam08, Kei13, KK15, KLkO21, Klu16, Kob16, KY18, Kra15, KV19, LLS21, LL18, LY21, Len14, Li19, Lin22, LV17, LLW18, MS16a, MW19, MS15, Mat16, MW16, Mok20]. numbers [Nat17, NW05, OS21, mPP22, PS19, PST20, Pot18, QH19, Rau16, RSY18, RLT22, Ros18, RL18, SS15a, San17, San21, ST18, Str15, Sun17b, Sun18, Sun19a, Sun19b, TT18a, Tau18, Tol06, Vää16, WS16, WS17, WS19, WW18, Wei20, Wu18, XMT16, XY19, Xu19b, ZY20, ZLL16]. Numerators [Hay10]. Numerical [Leh08, GT17, RPR17].

Observations [HKN10, HKN11]. Obstruction [Hat16, Bal19, PS22]. obtained [DJ22b, HLN19]. Octahedral [Gan10]. Octonary [AAW10a, AK14, Kök13a, Kök13b, Kök13c, Ala14, AK15b, AK17, RSS18]. Odd [AK12a, And16, CF09, MM07, Rey13, Zha12, BD22, CGM15, pCG21, JS15, Jon21, JM10, LR19b, MP21, RSY18, Rey16, SZ21, Sun17c, Wan15a, Zel19, Nic06]. Oddness [Ber18]. off [DFV13]. Ogg [Dic21]. Omega [MM13b]. One [Bar13b, Boy10, Fu11, Klo13, Liu13b, Mu17, PS11, Roy05, TY13, BCF21, Col12, Dud14, ELO16, FM20, Gui21, LZ18c, Li20a, Per17, Ram17, VX18, Viñ19, Wal21]. one-dimensional [Per17]. One-Parameter [Bar13b]. One-Two-Three [TY13]. one’s [MV14a, Vep17]. only [BD22]. open [Sar22]. openness [Fra21]. Operations [HHP09]. operator [BV18, Mau17]. Operators [AJKM09, CEP12, CW07, Mas07, Mau12, Wal06, Wal08, BB14b, FW21, Gol07, Wal17b]. Optimal [Dum05, Has10, SS15a]. Optimally [Bal08, MR12]. orbifold [CF17]. orbit [Hin19]. Orbits [Kab10, HSW14]. Order [ASV10, Buc11, Lin13a, WG11, Zor13, Ade18, Ade21, Bud20, Bur21, Cas12, EF16, GY16, HSZ20, Kim16b, Kom17, KRS18, LW20, LR19b, LP14, NH20, Par22, Peh16, PQSW14, RL19, SS10, ST19, Toh08, WZ12, Zho18]. ordering [LTZ20]. Orders [AC13, AAD11, Chi09, CD11, Del05, Tro17, Tu11, Yu11, Ala16, ACB19, Kit13, KT20, Len14, LS17b, Lou16, LL17, PZ16, Sch21b]. ordinarity [FT15]. Ordinary [Mat08, Van18]. Origin [FHS11, DS21]. Orthogonal [BG11, DWW09, Sak11, CK21, GK18a, Sak14, Ska21]. orthogonality [Gue15]. Orthorecursive [KK20]. Oscillations [MT21]. Other [BBCZ05]. Overconvergence [Koc08]. Overconvergent [SG17, Bir19, Hsu20]. overorders [HS20]. Overpartition [BL08, Xio16, NS18]. Overpartitions [Lov05, AB15a, AsMS20, And15, CHS15, pCGyS20, HS21a, JL18a, KKL21, LRS14, Mao14, RB18, She16, She17, Zha22b]. overview [Ked05].

packets [AMPS17, FRcT20, Mis16]. Padé [Mer11b, Vää18]. Pair [BPZ14, EJ11, Smi13, Che14, Mac16, Sun16c]. pairing [FN14, Kat21]. Pairs [BG06, BL08, El 12, Lev06, Liu13b, Ula10, AMMS17, BD21, Cha18a, Cho15, Kom17, KLPP15, LZ22, LP16, Mal21, sMT21, NS18, Par15a, Var22, Zub20]. palindromes [Baš12]. Paper [TB06]. parabola [HL20]. Parabolic [KM09, KM12c]. Parameter [Bar13b, ZMS15]. Parameters [Mac12, FRcT20, Smi21]. parametric [Xu19b]. Parametrization [VSF10]. Paramodular [PSY18, GK18a, JLR14, JLR17, PSY17, RT11]. parities [Wei19].

Parity [AB15b, BD15a, Kaa11, Kar19, LZ12, AD16b, DW16, Dai16b, LMW16, NY21, SZ21]. part [DFV13, Kei17, Kum21, Sca17]. Partial [DL22, Kim10a, SXJ13, Zha12, ZT13, All09, Fol21, Gir16a, HYZ17, JKK16, KL14, O’S16, SA09, SFM17, TZ21]. particular [Hic14]. Partition [BB09a, Ber07, Cha10b, Fu11, GY07, Gar10, Kaa11, LZ12, Lin13b, Mor08, Pol14b, Xio11, Yee09, ZW14b, BB20b, BD22, CbL21, CHJL21, Che14, Che22c, CG13, DE21, HHM21, HK20, Kar19, lLW21, MS22, NY21, Nic06, O’S16, Pat15, Pen08, SS21, Tou09, Wan15a, Wan16]. Partitions [BCH08, BS12, Kaa11, Kei17, LW20, Lin14, RS06, XY14, Yan19, ACX18, AB15b, ACS09, BD15a, BB18b, CC17, CCH^{+}19, CM16, Cio20, CGM15, CGH18, CG19b, Dai14, DW16, Dai16b, Gol16, JS15, JZ17, Kei21, LMW16, xMgC22, MY21, Pen19, RS13a, SZ21, Vau15, Yao15, Zho21]. Parts [And13, Dub09, Gar10, RS06, ACX18, BS16a, BB18b, BD22, CGM15, Dai14, Fre19, Gir15, JL18a, JS15, JLl21, MY21, Min22, NY21, Wan15a, Wit07, Zho21]. Patterns [VW17, Rib11, Tou19]. Paul [Ano15c]. Pell [BDGL15, EGL21, FY21, HS22, RLT22, YF19]. Pellian [JZ06]. Pentagonal [Ber09, Kol15, YX12]. Perfect [AK12a, BLMS05, CCZ15, KL11, KW14, KP18, NV10, Gou18, Klu16, Zel19]. Perimeter [BG06]. perimeters [Cho20b]. Period [Ma17, Wit10, CV17, Dau14, HWh20, Tre15a]. period-doubling [HWh20]. Periodic [Gir12b, LT14, Tow13, CV17, Che18c, DJ22b, Mor16, Mur15, Oye16, Pan22]. Periods [ASV10, BS17, Kop08, McC10, Fli11, Smi21, Tak08, Vs21]. Permutation [AAW08a, Zie08, KKK16, QQH15]. Permutations [AK12b]. Perron [RR20]. Perturbed [HJW14, Min22]. Petersson [CbL21, SPY18, YZ22]. Pezzo [Car11b, Hua17]. Pfaffians [Ros08].

Phenomenon [Bou05, Gor19, MS15]. Piatetski [LZ18a, QGX22, WC11]. Picard [Anc17]. piecewise [Sch18b]. Pillai [CPZ17, DL20, EGL21]. Pintz [Li22]. Pisot [HS21b]. places [Oli22]. Plancherel [GS16]. Plane [EJ19, BB21a, BCF16, DM20]. plongement [SS10]. Plongements [Mai11]. plus [MOS14a, MV14a, Su16]. Pochhammer [TT09]. Poincaré [Bla11, DG13, De 07, Ehr09]. Point [BB09a, Bui12, Bui13, Hub11b, PMM13, BCF16, BCF21, CV17, Dja11, HH21, LR19b]. Pointed [Fre12]. Points [ABP09, AAD11, Ben13, BU11b, Cho13b, CCS13, DW09, FG12, GM12, IT10, Liu12a, Mig15, Shp14, Thu08, Tow13, VZ14, Vul10b, Yas13, Zum11, Aka14, AjW21, BB21a, Bad17, BGm22, BLW21, BCU14, CS21, Cha15, hC21, Dic21, Dra11, Fri16, FY17, GR11, HB17, Hin15, Hin18, Hin19, Hol19, Hua17, HL20, JL20, Kel17, KT15, LTZ20, Mat18, Mei18, Mr22, Mor16, NR16, Now16, Pan22, Sha14, Shl12, Ska21, Yas15]. Poisson [Gra07a, Gra07b, Kur09]. polarizations [Orr17]. Polarized [Paz13]. Poles [Hah21]. poly [KL17, KY18, OS21, You16b]. poly-Bernoulli [KL17, KY18, OS21, You16b]. poly-Cauchy [KL17, KY18]. Pólya [May19]. Polygamma [MS09b]. polygonal [JOS19, Wu18]. Polygons [Bla11, GP12]. polylog [CDc20]. polylogarithm [Sca17]. Polylogarithmic [You17]. polylogarithms [Li20a]. Polynomial [Bao19, BO12, BM10, DS07, DE21, DFL08, DJ10, Gün12, HH21, Jon09, LS19, Pol11, VSF10, AM15, BK13, BSK17, CS21, Dau14, Dor20, IJO^{+}21, Klu16, KV19, Ma17, Mad12, MV16, NJ22, Par15b, RR20, Sch15, lWfS22, You16b]. Polynomials [AAW08a, AK11, BG11, BBCM13, BE13, BST10, Cao11, CCRT14, DWW09, Dub11, Dub14, GP12, Gur08, HJW14, HN08, HN21, HM10, IT10, KRY09, KRS10, Kur09, Our09, PPT12, Rob11, Rob15, Sak11, Sun13, Ula12, Zie08, Ade21, AT22, Bad17, BB20a, BGm22, BGW12, Bzd17, CG18, CP18b, CK21, DR22, DJ22a, DFV13, DS17, DS18b, Dud14, DA18, EG14, FZ18, Gir22, HH21, HLN19, Hin19, Jam16, JMV16, KKK16, Kim17a, hKS21, KL17, Kre17, Kru16, LM18, LS17a, Li22, LZ15, Mar14, MP16a, MP16b, MO20a, Min22, MV14b, Mon14, MR12, MP14, N18, Ozd21, PZ18, Pas15, QQH15, Qua19, RZ22, Ros17, Sak14, SA09, SS15b, SST19, Sun15c, Sun16b, WS19, Yam16, Zha17a, Zho21]. Poncelet [LMT18]. positifs [JM10]. position [Qua18, Qua19]. Positive [AALW09, AT20, BZ13, CWW08, GT10, HKN10, HHP09, Koc08, KMT11, Mun10, Oh11, Ot20, APW14, AD18, BK22, Fla19, HKN11, HLT20, Hu13a, JM10, JKO18, Koy22, LRS14, MP21, Rom19, Sar22, Szc15, Szc16, Vep17, Vie10]. Positive-definite [Ot20]. Possess [Far08]. pour [Amo07, AST22, BD10, Bil11, DP08]. Power [ANH14, Bya09, CS10, CL11a, CEO06, Dja13, Hal12, LLZ16, Mor11, Mor22, Moy13, WZ09, And16, BB18a, BGW12, GZ22, Gou18, Hyo15, JK16, KM14, Li17, LZ18b, MU18, MSV18, Mol12, NW05, Sin09, TZ09, Tre15b, Ula19, Wan18, ZZ11, Zhu22]. Powerful [BZ05]. Powers [BDGL15, BLMS05, CEP12, CKW13, CE07, Dub09, DFL08, GZ11, HRL11, Jam12, Kel10, Kum09, LL11, Liu12d, Liu13b, LS09b, MMR11, PV12, BJ19, Bae19, BD22, BHPR17, CG18, Cho18a, CG19b, DL20, EGL21, FY13, GF18, GL18, GL17b, KH11, KP18, Len17b, Len17a, Len18, LZ18d, Liu16, Liu17b, Liu19b, MV16, Mér20, MV14b, Mol12, Mu17, MQ18, QGX22, Rey16, Rif19, Ska17b, Ska19, Str15, Ula19, Van21b, ZL18a, ZG20, Zhu18, Zue18]. practical [ST18, Wei20]. precise [HP11]. Preface [And21a, And21b, Ano07b, BD15b]. Preimages [FHS11]. premiers [Kra15]. Preperiodic [IT10]. Prescribed [MS09a, DM21, HZ18, KMS21, Pey20, Rob15, SS17]. preserve [DS18b]. preserving [Mem20a]. primality [Meš16]. primary [DR22, PZ16]. Prime [AK12a, BBCM13, Bou05, CKW13, Cla08, DW09, El 12, GT10, Gra07a, Gra07b, Kan13, Kel10, KL11, Lee10, LO12, Li12, Liu13b, LS05b, McG12, Pol14a, Pol14b, PV12, Rho09, Sah11, ZLZ10, Alr14, Aym22, Ben15, BHPR17, BL13, CW19, Cho20a, Cho18a, Dud14, hF22, GL20, Has13b, Hu13a, II16, JL20, KMPS20, KM14, Kra15, LL18, LZ18a, Li19, LX21, LP16, MT17, MV14a, Miz08, Mu17, NJ22, Pas15, Pon16, RSY18, RC17, RL18, SD20, Str15, Tro17, Wan18, WY20, Zam16, Zel19, Zha15a, ZL19]. prime-additive [hF22]. prime-generating [RC17]. Prime-Perfect [KL11]. Primes [AK12b, BZ09, Baz11, Bro12, Cha05, Cha06, CLMR09, CP10b, DGM16, EV07, GL11, Kna08, Kou15, Lai10, LL11, MS10, MS09a, Tan09, Van12a, WC11, Wit10, Yas13, Zho18, AC15, AjW21, BB21b, CCZ15, Cai18, Cen16, CW19, Dub18, Dud15, DGM19, ETT17, GL18, Guo21, HKL^{+}21, He20, JKS17, Jon21, Lei12, Leo18, LS18, LZ18a, Lin22, Liu17b, Liu19b, LWC19, MCW19, ME17, MQ18, NH20, Pas15, Peh16, PT15, RS13a, RR20, RSS20, Rob15, Ska17a, wS20a, Suz17, VW17, Yao18, ZG20, Zoe19, Par20]. Primitive [Coh06, Che22a, CK20, JL22, JW17, MT18, mPP22, Tou19, Zoe19]. Primitives [MMR11]. primitivity [DJ15]. Principal [Gra11, BS19, Hsu16, Kid16, May19]. Principalization [AZTM15, Bos09]. Principle [Cla08, DK06, Tow13, Kru16, PS22, Rom19]. priority [GHK^{+}15a, HKN11]. pro [CR18, Mil13, Sal09]. pro- [CR18, Sal09]. pro-dihedral [Mil13]. probability [Hu13a, SD20]. Problem [AG12, Bao05, Bon08, BO12, DFL08, DJ10, Fel12, FM12, HM10, Jah10, LZ08, Liu08, Liu12d, Per12, Pol11, Rol11, WC11, Wan12, AMM17, Ary17, Boc08, Bor20, Bre19, CH16, CPS18, CPZ17, Cho17, DP08, DL20, EGL21, FK11, GT17, Ivi16, Jam16, Les16, Li16, LZ18b, LZ18d, LZ18c, Lu20, xMgC22, Mér20, PS22, SS10, Van18, WLWY22, ZL18a, Zhu18]. problème [DP08, SS10]. Problems [LM11, Nat11, US13, Hat16, HL20, Liu16]. Process [Sch13]. Product [Bou05, CLN05, CH09, Coo06, FW10, Kim10b, Li09, MPY13, RSS20, Suz05, ES20, Cil16, Fow20, FG22, Gou18, Hah21, Lam16, Ono17, Sch18a, TZ21, Zhe15, YZ22]. Products [BKW13, Bzd17, Dja13, Ehl10, HJW14, Hof13, Klü12, KN09b, LM16, ME17, Per12, Ska17b, Ska19, Van12a, AAAW17, BJ19, BD21, BDGL15, BC18, CCH^{+}19, DL22, ELO16, LLS21, Li20a, MZ22, SPY18, dS18a]. Progression [MS09a, BS16b, KP18, San14, SS15b, Van21b, Zho21, Zoe19]. Progressions [AG13, BZ09, CT13, Has13a, Kna08, KN09b, MN13, MC13, Oh11, SSD11, Adi15, AAAW17, BF10, CGZ21, DGM19, EK20, ETT17, Gib14, HKL^{+}21, Kou15, Pas15, RSS20, Rau16, RGHK20, San14, Suz17, Szc15, Vil18a, Zhe15, Zho18]. projectifs [Phi11]. Projective [Nar22, Nat09, Gro18, HS19a, Phi11]. Proof [CGPY15, FWX21, Fu11, Gui21, GL16, GL17b, Guo19a, HR11, Kim10b, LV17, LY05, Mao17, Meš12, VS10, XY14, ZY20, Apa18, BM21a, Gol16, Hui18, JKK16, Kim21a, Pat20, SY19a, Vil18b]. Proofs [Kim10a, XY19, Yee09, LW19, ZZ21]. Properties [Boy10, Gri11, KK10, MMO08, MPY13, Mor11, OT05, Szc16, Ula12, Wea22, BO19, CL11b, CL13, Che14, CHS15, CK20, DWW17, DLZ19, GR14, Haj20, Her16, HKLP09, HHM21, Jen05, yK21, LS14, LMS10, LP20, LR06, MNZ19, Mis17, NH17, RS11, Sch21a, She16, Sun17a, Sun17c, TZ17, Van16b, WV21, Wan15a, Yao15, Zha16]. Property [Dem20, IT10, BŻ19, DS18b, May19, Wei15b, ZMS15, ZC17b]. proportion [BCF16, BCF21, HK21, Rom19]. proportional [ME17]. Prouhet [Cho20c]. prove [PsR21]. Pseudo [Fin12, DR22, Lim15a]. Pseudo-Canonical [Fin12]. pseudo-nullity [Lim15a]. pseudo-polynomials [DR22]. Pseudorandom [CW12, EMW06, EMG08, Liu11, Lin22, Liu14, Liu19a, NW05]. puissances [Sod21]. Pure [DFL08, ANH14, BE21, May19]. Pythagorean [GHK^{+}15a, GHK^{+}15b, VSF10].

q [pFG22]. Quadratic [AALW08, AALW09, AAW10a, AAW10b, AK14, BZ09, BG17, BD08, BP15b, CV17, CCL13, CLM08, CEIK07, Coh06, DS13, EV07, FHS11, Fuk08, GF18, GT10, HS10, Has10, Has13a, Hea14, Hoe10, JR13, Kab10, Kan10, Klo13, Kök13a, Kök13b, Kök13c, Kow06, LW08, Li14, May14, Mr22, Oh11, PS11, Res09, Rüh10, Sha09, wS20a, Tem10, Vul10a, XY13, Zie11, Zum11, APW14, Ala14, AK15b, AAW16, AK17, Ala16, Ary17, ARTZ19, BB21a, BS20, BDTT16, Cas12, CPS18, CK20, Dai16a, EF16, EK20, Eld19, EW15, GJR19, GM18a, GZ22, Hin19, INST14, Jam16, Jon21, JKO18, KT15, KW18, KwK21, Kim22b, KN19, KE17, Kom17, Kru16, Li22, LH16, LM15c, Mal21, Mor16, Oga14, Ot20, Oye16, Pat20, PZ16, RS14a, RSS18, Rez21, RZ22, San09, Sch18b]. quadratic [Sze19, Vep17, Viñ19, Wan17, Xio16, XYZ17, Ye15b]. quadrics [HB17]. quadrilaterals [Cho20b, LZ22]. quadruples [RLT22, SZ15]. Qualified [Bya06, Bya09]. Quantitative [GGW11, Pet05]. Quantum [Ngu11, Mal20]. Quartic [Akh12, BTW06, EW15, Naj12, Wak12a, AK15c, BCF21, BCU14, Cho18b, GJR19, IIO20, LM15c, wS20a, Tak21, Wei19]. quartics [EJ19]. Quasi [MV14a]. Quasi-uniqueness [MV14a]. Quasimodular [Gra20, Roy07, DM15]. Quasimodularity [MS21]. Quaternary [AALW08, AALW09, CC07, GNS06, KN19, AW12, AAW16, Kim22b, LH16, Ye16]. Quaternion [CCL13, Jah10, AB18, CL19, Hir22, Hou17]. Quaternionic [BG15a, Bra14, LV12, Ter13, SG17, Van12b, Van16a]. Quaternionically [Gil17]. quaternionique [Gil17]. Quaternions [Kri10]. question [BB20a, GM18b, VB19]. questions [Har15, Vau15]. Quintic [HM10, KRY09, KRS10, PR17, SA09]. Quintuple [CLN05, Coo06, Kim10b]. Quotient [MM11, CDc20, Oga14, WZWQ16]. Quotients [AW13, CW12, SXJ13, Sin10, Wil12, YXJ13, Zha12, ZT13, AAA15, Fuj20, Gir16a, Guo15, HYZ17, TZ21, Ye15b].

races [Aym22]. Rademacher [BH14].

radial [Fol21]. Radical [Klu16]. Radically [McN13]. radix [Wei15b]. raised [VB19]. Ramanujan [Sil07, AZ05, Alk12, And21c, AH19, BO11, Ber07, BKW13, BI21, BCSX20, Cha10a, Cha10b, cC21b, Chi19, DR10, Dix11, Gug21, Gui20, Gui21, Jam12, JZ17, yK21, KL14, KKL21, KK12, KR16, Lai10, LS12, Lin14, Liu12b, Luc10, Mah19, Par22, Sah16, Sin10, Sun19b, Tót13, UW14, VB19, Vil18b, WJ20, XY11, You16a, Zha05, ZW14b]. Ramanujan-like [Gui20]. Ramanujan-Type [ZW14b, Gui21, JZ17, Liu12b]. Ramification [BZ13, GP12, LRL10, WS10, Yas15, KMS21]. Ramified [DD10, IMO13, Mon14]. Ramsey [ZW14a]. random [Cow22, Hu13a, KM22, MP16a, MP16b]. ranges [Def15]. Rank [BL08, DWW09, FP10, Gar10, Gol07, HR11, JS15, Kow06, Mun10, Álv14, CM16, CHJL21, Che22c, CWR16, Cre18, EMS21, FM20, KL14, LRS14, Li14, MT17, Mao14, Mil18, OLG19, Pot21, Pri09, Wei19, Zha22b].

Rankin [BM21b, JS19, KS13, Lan19, Meh12, Mor22, OS17, Pi20, Vie10, Zha17c]. Ranks [Bri09, pCGyS20, Gla09, Jed22, Wal13, pCG21, DG17, IMO13, LL22, Mat17a]. Rapid [BT18b]. rate [CC21a, Len17b, San18b]. Rath [Cha12]. Ratio [BB09b, CP10b, Has13b]. Rational [Adi15, BU11b, CCL13, Dub09, Dum05, ES12, EV07, FHS11, FG12, GM12, Gün12, HB17, Hin15, LMT18, Pan22, Ska21, Tam14, VZ14, Aka14, BARCVS13, BGm22, BCU14, BS17, Dem20, Fow20, Gir14a, Gui21, Hin18, Hit18, Hua17, JS20, JL17b, KP16, LR19b, Mem17, Mér20, NR16, PST20, Poë20, Rey16, Rob18, SS15a, Vää15]. Rationality [YZ22]. Rationally [EOY05]. rationals [Cil16, Mat17a, Oli22, Peh16]. rationnels [Hua17]. ratios [CGG15, Tre15a]. Rauzy [NSS15]. Ray [Hoe10]. Real [BO12, CLM08, Höh11, JV08, Lou15, Gam14, HS19a, KV19, Kug22, LL17, Oza17, Ste16, Yas16]. realizable [Tsa17]. realization [Ma17, Sca17]. Realizations [PRVS08, KRS18]. reciprocal [DFV13]. Reciprocals [CEO06, And16, Fre19, HN21]. Reciprocity [HS10, Gir19a]. Recognizing [AK12b]. reconstruction [Sar22]. recoverability [BGP15]. Rectangle [BG06]. Recurrence [HLN19, OSW11, BHPR17, Dub18, HS22, Kim21a, RC17, RL19, Xu19a]. Recurrences [GT10, KPT08]. recurrent [ABCM14, ZC17b]. Recursion [BJMV14, OS21]. Rédei [Bri11, Mur15]. Reducibility [HJW14]. reducible [AD16b, Ber18, IJO^{+}21, Ray21, Ray22]. Reduction [AM09, Cao11, LM14, Vol10, Yas13, Álv14, Her16, LL22, Nţ21, Roz18, Tak15, Vs21]. Reductions [Per17, Kim16a, LR19b, Liu22, PS19, Tro17]. refined [Dah11]. Refinement [Jam12, MS22, Zyw11, Wal21]. Regarding [Höh11, Mos07]. Region [HN11]. Regions [Kad12, Aym22, LTZ20, Mer11a]. Regular [CEIK07, Pen19, XY14, BD15a, HS21a, JZ17, NH17, Pen08, RB18, She16]. regularity [ETT17]. Regularized [Ehl10, LQ21, Mac16]. Regulator [Bar14, LW19, Gra18]. Related [Ber09, Bun12, Coo10, EH08, FM12, Has10, JZ06, Koz09, Sch12, Ula12, Ver10, Yee09, All09, ES20, BW17, BB20b, Bre19, CHJL21, CGH18, DJ22a, pFG22, Guo15, KW16, KKL21, Li15, LLL22, Pat15, SY19b, Tau18, WS19, Xu19a]. relates [Bat17b]. Relation [ACH05, KTT06, Suz05, Wak12b, BT18a, BO19, Gue15, Kim21a]. Relations [BB14a, Chu13, EG07, Kaz08, Liu12c, Van07, Veg11, BCSX20, Goo17, Gro20, HLN19, Kug22, LQ21, Ma17, Mur17, PsR21, Vää18, Wak17, Wal17a, XY19]. Relative [JZ06, MMW11, TLZ22, Sod21]. Relatively [BBCM13, El 12, Hu13a, Pon16, SD20]. Relaxation [AD16b]. Remainder [CC16]. Remark [Pol14b, Sug15, TB06, Lim15b, Val14]. Remarks [FG22, MOS14a, Meh12, Zha17a, KMV20]. Rémond [Dil20]. Rényi [LLZ18, Zha21]. Répartition [AST22, Ric13, Tou09]. repeated [JS15]. representable [Mos15]. Representation [AAW16, AC13, FZ06, FZ12, Kök13a, Kök13b, Kök13c, Mol12, Ngu19, PV12, BG15b, CGG15, EF16, Fra21, IIO20, Jia20, wJShY21, KS16, KE17, SV15, Suz17, Wu17, Yan19]. Representations [AALW08, AALW09, AK14, Ala14, AK15b, AK17, Cha08, CKW13, CC07, DD09, DS13, EV07, JM11, Kan10, Oh11, Pan11, Rey16, Sha09, Toh13, XY13, YX12, Ye15b, APW14, Amo21, AD16b, Ber18, BD21, Bil11, CGZ21, Cop20, CR18, GM18a, Hah21, II16, JVW20, JKO18, KwK21, Kim22b, Lim16, Mis16, Nţ21, Nat11, Ot20, Pea22, RS14a, RSS18, Ray21, Ray22, Roz18, Sha17, Sun19a, Tay17, Vep17, WS16, WS17, XMT16, AST22, Bil11, Tay17]. representative [Cob21].

Representing [Hu22, Alk15]. repunit [BŻ19]. residuacity [EW15]. residual [AsMS20, MS21, Pea22, Sha17]. residually [Ber18]. Residue [Bao14, Bya06, Bya09, CS10, DS13, EMW06, EMG08, Gur11, LW08, MS10, Bao15, LS18, LS19, Tre15b]. Residues [WZ09, wS20a]. residuosity [Hit18]. résidus [AST22]. Resonance [BGP15]. respect [Cho18a, mPP22]. Restricted [Kei21, Lin13b, Sun19c, TW19, BKS16, Bor22, DE21, GM19, xMgC22, O’S16]. Restrictions [Aza09, DWW17]. Result [AH07, Lan09, Ter13, Cho20c, RyW22, SZZ18, SW14, Val14]. Resultants [HS10]. Results [Baz11, CEIK07, HK13, PS15, Sah11, Ula12, AB15b, BD15a, Bor18, Bun08, BV09, DW16, Liu16, MORS16, MZ22, QQH15, SSU21, WS19]. Revisions [KM12c]. Revisiting [AD18, cC21b]. RH [DGM16]. Rice [CCS13]. Riemann [Sch14, AD18, BD10, Bet10, CQh22, Dud15, Far08, Fre12, Ivi05, JM10, KTZ16, KTZ18, Li15, MMW11, Mer11a, Olo09, Pat19b, WY21]. right [Zub20]. rightmost [FG22]. Rigid [KRLT20]. Rikuna [CCRT14]. Ring [OPY08, BGW12, Hyo15, LM15b, May14, Mil13, Mor16, SD20, Som22, WZWQ16, WHZ19, WJ20]. Rings [Bao14, DU10, EMW06, EMG08, GNS06, Gur11, JV08, MM11, Bao15, Bat17b, Kit13, LS19, MS16b, SSTW14, Wea22]. Robba [Wea22]. Robin [WY21]. Robinson [GR19]. Roch [Fre12]. Rodríguez [Har18]. Rogers [All09, cC21b, Gug21, LZ15, Sil07]. Romanoff [Kua15]. Root [BO12, BM10, Che22a, LY21, lLW21, Tsa17, Zoe19]. Rooted [BT18a, Zub20]. Roots [CO09, Coh06, Dub11, Dub14, DFV13, FY17, JL22, Mor22, MT18, lWfS22]. rotation [Now16]. Roth [LS09a]. Rounded [FP12b, FHL^{+}13, Küh12]. RSA [Nit09]. Ru [RyW22]. rules [BJMV14]. Runge [SS08]. Ruzsa [DR22].

Saito [Mat17b]. Salem

[Amo07, BK22, Hic14, Sto21]. Same

[ZT13, NY21, Ot20]. Satisfied

[CCS10, Mat19]. satisfying [LS05a, Yam16]. Sato [EGP21, HKL^{+}21, Kuo09, Pan11, Shp19]. scalar [Mau17]. schemes [Sca17]. Schenker [Mis17]. Schinzel [US13]. Schmidt [CG18, Le15, Pha22, Qua18, Yan17]. Scholz [ZY08]. Schreier [Ren19, WS10]. Schröder [LV17, LLW18]. Schröder-like [LV17]. Schröter [LY09]. Schubert [Thu08]. Schur [CC17]. Search [CLMR09, BB21b]. Second [Bet10, Hea14, HZZ12, Mer11b, Ade21, Bur21, CQh22, May12, Pat15, Zel19]. second-order [Bur21]. Section [LS09b]. Sections [AG13]. Segments [HM12]. Selberg [BM21b, CTZ16, KS13, Mor22, MS07, OS11, OS17, łPS13, Pi20, SSS22, Zha05, Zha17c]. selected [EL18]. selectivity [LS17b]. Self [BS12, DW09, Pic10]. Self-Conjugate [BS12]. Self-Dual [Pic10]. Self-Points [DW09]. Selmer [Ari13, BZ15, Bro10b, Chi09, FN14, KPW17, Li14, LM14, Lim15a, Mat12, Mat18]. Semi [Mag13, Her16, Kra07, LR19a, Vs21, Kra07]. semi-abelian [Her16]. Semi-Adelic [Mag13]. semi-canonical [Vs21]. semi-diagonal [LR19a]. Semi-stable [Kra07]. semi-stables [Kra07]. Semigroup [Sch13, WZWQ16, WHZ19]. Semigroups [Leh08, GT17, RPR17, Wan21]. Semisimple [Vol10]. semismooth [Suz22]. separated [Kre17]. Separation [BM10]. Septic [TK10, HHM21]. Sequence [Bun12, Coo10, DS07, LT14, MS09a, Sch12, Sun12, Ula12, Zha14, HW15, KL14, Lin22, LP14, San17, Vää15]. Sequences [BB09b, BDGL15, BLMS05, HN08, LP13, Liu11, LS05b, Mos07, OSW11, SX14, Sun14, TZ12a, TZ12b, WG11, Xia13, ABCM14, Bel22, BHPR17, BGP15, CHL19, DJ22a, Dub18, DJ22b, GF18, HS22, Hic14, HKLP09, HWh20, LLS21, Liu18, Mad12, MS15, NJ22, PQSW14, QGX22, Sch18a, SK13, Şia15, Sun17c, Wan21, YL16, ZC17b, ZS18]. Series [Alk15, BK12, Bla11, Bun12, CL11a, CH09, CZ09, Chu13, Coo10, CEO06, DG13, De 07, Ehr09, FP12a, KN09a, KM12a, Kob14, KM06, Kri10, LT14, McC10, MMO08, Mor11, MR06, Moy13, MW11, SXJ13, Tay09, Wal06, Wal08, Wil12, Agn22, All09, And16, ACS09, BW17, BK13, BS15, BK15a, BGW12, Bun08, BV09, BV11, BP17, CJKM18, CH05, Che18a, CC21a, Che22b, Coo09, Dah18, Eve22, FW21, Gui20, Gui21, HM06, HN21, Hyo15, JM16, JS19, Kal18, Kan14, Kid16, KK19, KKL21, KM12b, Kob16, Kra14, Kri16, KP19, LM15a, LLM18, Lim15b, Lim16, Liu12b, Liu13a, MAM06, Mat17b, McC12, MOR21, Miz08, Moo19, Mor22, Ono17, Oza17, PT14, Pat19a, Pet16, RZ22, Sad16, SVY20, Su16, TT18a, Tan19, TSB20, Tsu18, Ula19, Vää18, Wal17a]. series [Wal17b, WLWY22, WW18, Wil19, Xu19b, Ye15b, HM06]. Set [Lan09, Pol11, RS06, AC15, Alr14, BLW21, Bud20, EG12, Liu14, Liu19b, MV14a, Mem17, MR12, RPR17, Szc15, Szc16, Wan18, Yan19]. Sets [BB09a, BEP10, Bya06, Bya09, DJ10, FS11, HHP09, Kur09, MC13, NQ08, ZT13, ApKK22, AMMS17, BG15a, BK15b, CWW15, Cil16, DWW17, Din21, HS22, Hin19, LL18, LZ18a, LLL22, LLZ18, Mac16, Pac19, PHLS19, Poë20, Pon16, RR20, Ros17, SFM17, VW17, Zha15b, Zha21, Zhe15, Zhu22]. setting [BS15, McC12]. Seven [Lov05]. Seven-Colored [Lov05]. Sevens [Hir06]. Seventh [CS10, Li17]. Several [AZZ05, ELO17, FS11, ZW14b, JS19]. Sextenary [AAW10b]. sextic [ANH14, BS10b, BCU14, GJR19, TT14, Wei19]. Sextuples [Cho15]. Shadows [Har11a]. Shafarevich [Del05, Dum09, LL22]. Shapiro [LZ18a, QGX22, WC11]. Sharing [Pil09]. sharp [Lin22, VY13]. Sheaf [BEH10]. shears [Kel17]. shift [Wei15b].

Shifted [Bro12, GY07, CW19, GC17, JV21, LWC19, PT15, San21]. Shiftless [GY07]. Shifts [Bet10]. Shimura [Cho13b, PT14, PR21, Pur13, SG17]. Short [Aka14, AZ05, Bai13, Baz11, Bor18, BH13, Cha05, Cha06, Cro07, Koh10, LZ08, Lan09, LY05, Mat10, RSW14, Smi13, Bha20, CTZ16, DGM16, DGM19, IW16, Kim22a, Kou15, Mat16, Sed15, TZ09, Tol06, Yao18]. shortest [Nat11]. shrinking [WLWY22]. Shuffle [Sou10, LQ21, Wak17]. Sidon [Din21, Pac19]. Siegel [BM21b, Bro10a, CW07, CL22, Das13, DR10, Dic15, FW21, GS17, JLR17, Mat17b, MP15, Miz08, OPY08, PSY17, RSW14, Sah11, Tay17, Wal06, Wal08, Wal17a, Wal17b, Zam16]. Sierpiński [Jon09]. Sieve [BZ05, FI05, Hal12, BB18a, BS20, GZ22]. sieving [Gor19]. Sign [GS17, HY18, HKKL12, KM14, MM14, RSW14, HP11, IW16, KM18b, Zha16]. Sign-change [HY18]. signature [RZ22]. similar [LTZ20]. Similarity [Kab10]. Simple [Cho13a, Wak12a, LS17b, Sug15, Viñ19]. Simplest [GJR19]. Simplices [EOY05]. Simultaneous [KM18b, Lag10a, Lag10b, Roy05, SSD11, STW10, FY21, WLWY22, YF19]. simultaneously [BLW21]. Sine [Ono08]. single [AAW16]. Singular [And15, Baj14, Ge18, KK10, LZ12, AB15a, BL15, CHS15, Edi05, Fow20, Jen05, LMS10, NS18, Poë20, Rif19, She17]. Sinnott [KT20]. six [CG19b]. sixteen [Par22, RS14a]. Sixth [Mah20]. Size [Alk07, DJ10, EK07, Baj09, NY21, Pac19, SS17, TT18b]. sizes [Kei17, SA09]. Skałba [LS05b]. Slater [Kur21, Sil07]. slope [Vie10]. slopes [Ren19]. Small [Bal08, Bru05, BD08, FZ06, FZ12, FS08, Fuk08, GL20, Gir12a, HK13, Pot18, Roy10, Xio17, Ben15, CF22, Hin18, Kei17, Mér20, Min22, NR16, Pea22, Wil19]. Smallest [And13, Din09, Gar10, CM20, CP10a, HL11, JL18a]. Smooth [Cro07, Mat10, BB21a, LWC19, Mat16, Mig15]. soluble [HK21]. solution [HS22, RS13d]. Solutions [Akh12, Bao10, Gün12, JR13, Kan13, Liu08, Mor16, Ngu11, SS08, Wak12a, Bao08, Eld19, GL20, Gra20, LR19a, MT17, Sze19, Tan18, YF15, YF19]. solvability [FY21]. Solvable [HM10]. Solving [FLS12, KP14, Sch15, Cho18b]. Some [AAW08a, ES20, BO11, BU11b, DWW17, El 11, pFG22, Guo15, Har11a, KTT06, KMV20, Li13, LQ21, Liu12c, LLW18, Meh12, Ono08, Pat19b, ST11, Sun17c, TZ17, TY13, VZ14, WC19, WG11, Xu19b, Yas13, ZS20, Zie08, ACS16, Ari13, ARTZ19, BP15a, BB20b, BD22, Bat17a, CP17, CP18a, CGH18, GS18, GR14, GL16, Hic14, Ivi16, JMV16, KM18a, KKL21, KE17, Len17a, Len18, MS22, Mao17, MW19, Sed15, SR19, SST19, Sun17a, Tan19, Vää18, XY19, Xio16, Zha22a, ZS18, ZLL16]. somme [AST22]. Sommes [LR22]. Somos [Gug21, Xu19a]. Sophie [Leo18]. Space [Gan10, HK14, Hir22, Ma17, Mig15, Roe14, Sch21a, Su16, Yas16]. Spaces [Cha12, EG07, Kab10, Mah12, Bor15, HS19a, May14, Phi11, SPY18, Tra17]. Spacing [Kur09]. sparse [Ros17]. Special [Aga10, Bou11, Bro10b, Cai10, Cao11, FP12a, MS09b, TT09, Bad17, JS19, Kug22, LZ22, Mor22, Pat19a]. Specializations [CCRT14]. Specified [Kan13]. Spectra [Pon09, Vul09, Vul10b]. spectral [Dic15, Haj20, Wu17]. Spectrum [PMM13]. speed [TLZ22]. Spherical [HK14, Hir22, RZ22]. Spiegelungssatz [HR11]. Spinor [Hay14, JKO18, RSW14]. Split [LM14, Mr22]. Splitting [GP12, Her16, Cen16, Pol14a]. Square [DU10, Dja13, Dum09, GK13, HM12, Ivi05, LM11, MM11, MY13, MT18, Ros17, Alr14, Bai16, BS20, CTZ16, FP15, JLSW15, JL22, LZ18c, Mor22, Şia15, Sun15a, Tsa17]. Square-Free [DU10, Ros17, Bai16, JLSW15, JL22]. Square-full [MT18]. squared [Sof18]. squarefree [MP14, Pas15, PSY17, Tol06]. Squarefull [CT13, AM17]. Squareness [Aga10]. Squares [ACH05, BCH08, BS16b, Cai10, CKO05, LL11, Liu13b, LY05, Mas08, Ros08, SS15b, Smi13, Vau15, BCSX20, Cha15, Cha18b, Cho15, Cio20, Hu22, Kim22b, KP18, Mos15, Par18a, RS13a, RY16, Ska17a, SC16, Sun19c, UW14, Wu18, XY19]. Stability [JK21, Zha17c]. Stabilizers [Cul12]. stable [JL17b, Kra07]. stables [Kra07]. Star [DK06, KO10, TY13, Wak12b]. Stark [AMO17, DR16, Nom14, Par11]. Starting [FGT15]. Statistics [BJ19, Bae19, Gil13, Kur09, Rho09]. Steiger [Li22]. Steiner [EJ11, EJ19]. Steinhaus [Mah20]. Steinitz [SS10]. Step [BB09b]. Stern [Bun12, Coo10, DS07, Len19, SW13, Ula12]. Stickelberger [Cas12, Sod21]. Stickelberger-type [Cas12]. Stieltjes [HWh20, OS17]. Stirling [Ade18, Ade21, HZZ12, KY18, KP19, QH19]. stochastic [TZ17]. strange [AK15a, Pat15]. Straus [US13]. strict [AD16b]. Strings [Tan09]. Strip [Far08, Pat19b]. Strong [Cha12, Wal21]. Strongly [BF19, NV10]. Structural [WV21]. Structure [Car11a, HKN12, Hua14, Lev06, Wan21, Zie11, Gil17, PQSW14, Sod21]. structures [Anc17]. study [AM15, All09]. Sturm [Mau17]. subconvex [Agg21]. Subconvexity [Dah18, Ass21]. subdegenerate [Pha22]. Subfields [Wid11, KW18]. subgeneral [Qua18, Qua19]. subgroup [Peh16, SS17]. Subgroups [KK10, AM09, DM21, GS17, LŞ20, LM18, Mér20, NT14, Pot21, Shk18]. Sublattices [Küh12]. subrings [Shl12]. subschemes [Sar22]. subsequence [JL18b, LY19]. subsequences [GZZ15, HZ18, HKLP09, LDSM19]. subset [Baj09, PHLS19]. Subsets [Via10, Ben15, Liu14, NSS15]. subspace [Le15, Pha22, Qua18, Qua19, Yan17]. subspaces [Alv21, BB14b]. Substitutive [Sur20]. Subvarieties [Via10, RyW22]. Such [BB09a]. Sum [BBS08, Bou05, Din21, EL18, KW12, Lee10, LP13, MS05, OEL13, SX14, TZ12a, TZ12b, Wil05, Xia13, YX12, ABCM14, AM15, APW14, Apa18, AH19, Baj09, Che19, CP17, CP18a, Dau14, ELO17, Eve22, FW20, GZZ15, GLPW18, Ge18, Gir16a, Gir19b, GC17, GLZ15, HZ18, HLS11, Isa21, Kim21a, KLkO21, KE18, KMPS20, KR16, LDSM19, LX21, MY21, Mol12, PQSW14, PHLS19, Rey16, RL18, Suz17, Van21b, WJ20, Wan21, XTY14, Xio17, Yam15, Ye15b, YL16, ZW14a]. sum-of-digits [Eve22]. summands [ACX18, CGZ21]. summation [Ass21, BP18, Liu12b, Vil18b, WW18]. Sums [ACH05, Alk12, BBS08, BCH08, BLS07, BBCZ05, BH10, Bla11, BBCM13, BEH10, BH13, CK14, CM12, CKO05, CE07, EOY05, ELY05, El 12, EH08, Fre19, Gir12a, GML12, GZ11, GZ12, Gur08, Gur11, HRL11, Ham13, Has13a, JRW11, Kel10, KE18, KN09b, LS12, LM18, LY05, Mac12, MS10, Mas08, OEL08, OSW11, RR11, Ros13, Ros08, Roy07, Sau15, Sha09, Smi13, Van07, Wat08, XC10, Zha08, AS09, Aka14, AK16, Alk20, Alo19, Álv14, ES20, BP15a, BP18, BS16a, Bat17a, BH14, BCSX20, Bor22, BC18, CSJ17, CG22, CT18, CDHS15, Che17, CG18, CP18b, Cho20a, Cho15, CY14, Cow22, DS17, FWX21, Gir14a, Gir14b, Gir15, Gir16b, Gir17, Gir18a, Gir18b, Gir19a, GL17a, GS18, GL16, GL17b, Ham16, HLT20, Hou21, Jar22, JV21, JL17a, JL18b, JOS19]. sums [JO20, Kei13, KP16, KH11, Kim17b, KE17, KP18, LP17, Len17b, Len17a, Len18, Liu19b, Mac17, sMT21, MTWZ17, Mis17, Moh19, Mos15, MT21, MS18a, NW05, NMZJ22, Par18b, Par18a, PHLS19, RS13b, Rez21, RGHK20, Sad16, SY17, Shp18, Sin09, Sof18, SVY20, Sun16a, Sun17a, Sun18, Sun19c, Sun19b, Tsu15, Wei15a, Wu18, WS20b, XMT16, XY19, XYZ17, YC17, YX14, Yao18, Ye15a, ZH16a, ZL18b, ZCX22]. Sumset [EK07, SY14]. Sumsets [Bal08, MC13, TW19]. Sun [GL16, LV17, Mao17, Meš12, RS13d, XY19]. sup [Ste16]. sup-norms [Ste16]. Super [Sun15c, YC17, MTWZ17]. Supercharacters [CT18]. supercongruence [Guo19b, Guo21, WY20]. Supercongruences [Len18, Sun16b, Tau18, pFG22, Gor19]. Superelliptic [BU11b]. Supersingular [BG11, DWW09, Vol10, XLD22, EG14, Lei12]. Support [Per12]. supported [KM14]. Surface [Gil17, Li11, MMW11, HH21, Hua17]. Surfaces [Car11b, Ehl10, EJ11, MM13a, Vol10, VZ14, Wie09, Anc17, Bal19, BCU14, Gam14, GL19, Ito18, RT17, Rom19, Tre15a]. Survey [KM05, Luc10]. Swisher [pFG22]. Sylvester [All17, Zha15b]. Symbol [CS10, GN19]. symbolic [JMV16, JVW20]. Symbols [Gir11, Gir12b, Wie09, BK11, Cow22, pCG21]. Symmetric [AS06, Dja13, Dum09, HRL11, Rey13, Hin15, II16, MU18, MSV18, Sun15a]. symmetric-square [Sun15a]. Symmetry [AA20, BK11, FS11]. Symplectic [Cul12, Mau12, Mue12, QY11]. System [JZ06, PS15]. Systems [BG11, FHS11, HK13, KMW10, Lag10a, Lag10b, AK15c, BK15b, Cau20, Har15, KRLT20, MT17, Ram17, SSTW14, Sur20, Wei15b]. Szego [CL11a, LZ15]. Szemerédi [Li22]. Szösz [Boc08].

T [Ano15c]. tagged [ACX18]. tails [BC10]. take [Gir18b]. taken [CW19]. Tamagawa [Tür18]. tame [San09]. tamely [IMO13]. target [WLWY22]. targets [Le15]. Tarry [Cho17]. Tate [Kuo09, Cen16, Del05, Dum09, EGP21, FN14, GH14, HKL^{+}21, LL22, LM15b, Pan11, PP18, Shp19, Vol10]. Tau [AZ05]. Tauraso [Guo19a]. Taylor [LS14, MP22a, SST19]. Teichmüller [Col12]. Telescoping [MS18a]. tensor [BD21, You12]. Tenth [Bya09]. Terai [Miy11]. Term [Dub11, EH08, KPT08, Bor20, CTZ16, HW18, KP18, LZ18b, MP14, San17]. Terms [HS22, BDGL15, CFTZ14, CP18b, GY16, MV16, Oza17, Sof18, ZL18b]. Ternary [CEIK07, JOS19, Kan10, BDTT16, Cai18, Dud14, Jon21, JKO18, KwK21, LMO^{+}19, MR12, Ot20, Zha17a]. test [CS21, Zor11]. testing [Mér20]. Tetrahedral [Şen12]. th [AsMS20, FY17, lLW21, MS21, San17, Toh08, Tre15b]. Their [Gil13, Hua14, KM09, KM12c, Mac12, Mor11, OT05, Sak11, Veg11, AjW21, BSM16, Gui20, He20, HKLP09, INST14, Kal18, Kam08, OS21, San18a, Sar22, You17]. Theme [GHHP06, Jon09]. Theorem [BTW06, Ber09, Cai10, CL11a, DK06, Fre12, Gim13, Har08, Höh11, KK08, Kum13, Lev06, LS09a, Lov05, Pet05, PSZ16, Raj11, RD13, Rod14, Ros13, Yan13, Yee09, ZY08, Zha08, Alv21, ApKK22, BM21a, CP08, Che15, CPS18, CK17, Dau14, Deb19, Dic21, ELO16, Fer22, Gho11, Gol16, Ham18, IW16, KO20, Kua15, Le15, LZ18a, LS05a, LLZ18, MR15, Oli22, Par15b, Pat15, Pet16, Pha22, QH16, Qua18, Qua19, Rob20, Ska17a, Tak08, Tan14, Tür18, Yan17, Zha15a, Zha17b, Zha21, ZJ19, CC16, Kol15, Smi11]. Theorems [AT20, BL09, LV12, MM13b, Pan11, AM15, ASD16, Cas12, HR21, Pea22, Rei21, Tou19, Yao18]. Theoretic [KTT06, MMO08, CFTZ14, LM15a]. Théorie [AMO17]. Theory [Bar13b, Car11a, CDc20, DH13, JM16, Li12, QY11, AMO17, BP15a, Cha20, Che15, Che18a, Ge18, LZ14, Nat11, PS19, PST20, TZ21, WLWY22, Won18, Wu17]. There [RLT22]. Theta [AALW08, Car11a, CZ10, Ehl10, Hub11a, Kim10a, KM06, Kop08, Kri10, Liu12c, RZ22, Wal06, Wal08, All09, ES20, BB20b, BCSX20, Bur21, CHJL21, Che18a, CGH18, Fol21, FW21, FT18, FY13, Gar08, HSZ20, JKK16, Kan14, KL14, Kra14, LW20, LY09, Mat17b, Sch18a, SR19, Sun19b, Toh08, Tsu18, ZL18b]. thin [Gue15, Shp19]. Third [HSZ20]. Third-order [HSZ20]. Thirty [BB06]. Thirty-two [BB06]. Three [CF09, CKO05, KPT08, Kök13b, sMT21, PSZ16, RY16, Sch10, TY13, Vin14, HB17, KLkO21, KP18, Ska21, ZY20]. Three-Variable [Vin14]. Thue [Akh12, GJR19, HW15, HWh20, JZ06, Vää15, Wak12a]. time [CS21]. times [Gir18b]. Titchmarsh [AG12, Deb19, Fel12]. Tits [ACS16, ACB19]. topographs [AA20]. Topological [LP20, DS18b]. topology [Szc15, Szc16]. Tori [Per12, Per17]. toric [Nar22]. Tornheim [Ono21]. Torsion [CCS13, Dum05, Hol19, JS20, LŞ20, LRL10, Şen12, Yas13, AjW21, Cen16, hC21, Col12, FY17, Gam14, MP22a, Pri09, Yas15]. torus [Lei21]. Totally [CCL13, JV08, CL19, CF22, Fla19, Hou17, Kug22, LL17, Mon14, Oza17, Pot18]. totients [FP15]. Towers [Has10, JM11, Mat12, CeM21, Ren19]. Trace [MS12a, MMW11, Fla19]. Traces [BL15, KK10, LZ12, MP10, BJ19, Bae19, Edi05, Jen05, LMS10]. Transcendence [Bun12, Coo10, JSS14, MW11, Roy05, TT09]. Transcendental [Gun06, Alk15]. Transference [DK06]. Transform [BKB13, Ivi05, Fin14]. Transformation [Dix11, Bat17b, BSM16, CC16, Ham16, Kra14, WC19]. transformations [BS15, BW21, Nat15]. transforms [BS10a, Bar13a, Pat19b]. translate [CK20]. translations [NR16]. tree [ACS16, ACB19, BT18a, Len19]. Trees [AC13, SS15a, Zub20]. Tri [Ber09]. Tri-Pentagonal [Ber09]. Triangle [BG06, Wie09]. Triangle-Rectangle [BG06]. triangles [LZ22]. triangulable [BD21]. Triangular [ACH05, BCH08, BCSX20, KLkO21, Sun19a, Sun19b, WS16, WS17, XY19]. trianguline [BD21]. Trigonometric [BBCZ05, BH10, ES20, Kim17b]. Trihedra [EJ11]. Trilinear [Shp18]. trinomial [EGP21, JKS17]. Trinomials [PR17, Wil10, BS10b, Har12]. TRIP [DFG^{+}14]. Triple [FW10, OEL08, AMSV21, CFM18, FG22, Hah21, LX19, Pet16, Sch18a]. Triples [Fow20, GHK^{+}15a, GHK^{+}15b, Wan15a, Wan16]. triprojectif [Mig15]. triprojective [Mig15]. Trotter [Pan11]. truncated [DJ22b, RR20, Kol15]. Truncations [CL11a, Far08]. Tsfasman [Leb10]. Tsunogai [KRY09]. tuple [Dai14]. Tuples [VSF10, Liu09, MS16b, Rib11, Tol06]. Turán [Boc08, Yan13]. Twelve [Lin13a]. Twin [Cha13, hKS21, AC15]. Twisted [BEH10, Ehl10, FZ06, FZ12, Hea14, KW12, Kel10, Str22, AH19, Isa21, KW16, LMS10, Ren19, Wei15a, You12]. Twisting [JLR14, JLR17]. Twists [Aga10, Bui13, Cla08, Gar18a, Kow06, Ula10, Cob21, INST14, Jed14, Li14, Wei19]. Two [AAR16, And13, AS06, Baz11, BBCM13, CM12, DWW09, Har15, Hsu16, Hu13b, JRW11, Klo13, Kök13c, LL11, Liu13b, LZ15, Liu16, Mac12, MS16a, RR11, Smi13, Suz05, TY13, Tót13, Xio11, Zha05, APW14, AD16a, BB06, Bur21, CFTZ14, CC21a, CL22, Dic15, FW20, FZ18, GS17, HW18, HL20, KRS18, Kur21, LW20, Mal20, MU18, Miy15, ME17, NR16, NW05, Rez21, RL18, Ska17a, SC16, Suz17]. Two-color [AAR16]. Two-Dimensional [AS06]. twos [Vep17]. Type [And13, Cai10, Cha10b, DH13, El 12, Gan10, GJS14, Koh08, Lin14, Sil07, Van12a, ZW14b, Zha08, Alv21, AITA19, AZTM15, ATT^{+}16, Bor15, Cas12, Che22c, CH15, GS18, Gui21, HL12, JZ17, Kim16a, LL22, LW20, Liu12b, Liu16, Mah19, Ono21, RT11, Som22, SW08, Tót18, Vää18, Val14, WC19, WW18, Yao18, ZS20, Zha22a, ZS18]. types [JS20, KY20].

Ultrametric [Sam10]. unary [KW14].

Unbounded [Bet10, Baš12].

Undecidability [JV08]. Uniform [Ste16, Toh13]. Uniformity [Liu11]. Uniformly [Zoe19, BK15b, Mem20b]. unifying [GLPW18]. unimodal [KL14, Sch18a]. Unimodular [Har11a, HKN12, Kom09, Neb13, Wil10, Kel17]. Uninhibited [De 07]. Unique [GGW11, LL18, SH08]. uniqueness [MV14a, Par20]. Unit [LM11, Shi16, Zie11, Bzd17, DFV13, Sch21b]. Unitary [CF09, Cho13b, Kid16, FRcT20, HK14, Tsa17]. unités [AMO17]. Units [Bel09, Fol09, Vig12, AM17, AK15c, AMO17, LL17, Mem17, MO20b, Ngu19, Par11, SC16]. unity [KK20, lWfS22]. Universal [JO20, Won20, EG14, JOS19, Wu18]. Universality [Lau13]. unlike [LZ18d, Liu17b, Liu19b, Mu17, MQ18, ZL18a]. unramified [AP08]. unreduced [LM16]. unsplittable [YL16]. Upper [Akh12, HHP09, KM18a, Zel19, BŻ19, Che22a, Edd16, Ivi16, Pac19, YF15]. Urbana [Ell15]. Use [Min12, RC17]. Using [AK12b, JVW20, Liu19a, SS15a, Zor11].

Valeurs [Nic06]. valuation [HP11].

Valuations [HZZ12, QH19, Ula19]. Value [Aga10, And12, Har08, Kur09, Roy10, Wat08, ZLZ13, AMSV21, CFTZ14, CG22, Fer22, FZ18, FG22, Gir18b, GL17a, Jun14, MU18, NR16, NMZJ22, Rei21, Suz17]. value-distribution [MU18]. Valued [Gim13, Mas07, Tay09, BM21b, CF17, Čes16, FM20, Got20, JQ21, Mac17, Mal20, SS14, SV15, Wil19]. Values [Aga10, AK11, BK12, Bou11, BM11, FW10, FP12a, GMR11, Kal18, KO10, KS13, KMT11, Li13, Mac12, MN13, Mér20, MM07, MS09b, OEL13, łPS13, Pat10, San09, TY13, Tem10, TT09, Wak12b, BT18a, BFLT10, Bad17, BL15, BC10, Bro10b, Bru18, Cha20, CW19, – DLV20, ELO16, ELO17, EL18, FJ20, Gir16b, Gir17, GC17, GLZ15, HH21, Har18, Hof17, Jar22, JS19, Jia20, JVW20, Kan22, Kas19, KS16, Klu16, KP19, Kug22, Lan19, LP20, LQ21, LWC19, Mac16, MAM06, Mor22, Mur17, MS18b, MP14, MP21, Nic06, PT14, Pas15, Pat19a, Pey20, Pi20, Ros17, Sas15, SY19a, She22, SSS22, TT18a, VR20, Wak17, WV21, Xio17, Yam15, Yas16]. vanish [AAAW17]. Vanishing [BGW11, Buc11, Bui12, Che19, MZ22, Mui11, Mui12, Tan12, Tan19, Bro10b, CY18, Dah18, Jar22, KM18b, Pat19a]. Variable [Vin14, CG18, Li20a]. Variables [AZZ05, Cai10, CF09, Sch10, JS19, KE17, Kre17, Mal21, Mu17, RS14a, Vep17, ZL18a, ZL19]. Variance [Kna08, Liu12a, RGHK20]. Variant [AG12, DFL08, DJ10, RR20, Smi11, CPZ17, Won17]. variants [CPS18, Def15]. variation [Par15b]. Variations [GHHP06, Jon09, Ula10, BB06]. Varieties [Cho13b, Cul12, FP10, GH14, Paz13, Per12, Sai10, Thu08, Via10, BGm22, Bru18, Dem20, Her16, Hol19, Kim16a, Orr17, Pal14, Sug15]. Vector [CF17, Gim13, Mas07, NQ08, OSW11, Tay09, Wil19, BM21b, FM20, Got20, Mal20, SS14, SV15]. Vector-Valued [Gim13, Mas07, Tay09, CF17, Wil19, BM21b, FM20, Got20, Mal20, SS14, SV15]. Vectorial [PP18]. Vectors [CW12, MS12b, Fin14, JL20, JLR14, Zha16, Zor11]. versa [San14]. Version [Pet05]. Vertical [Far08, MN13, SSS22]. Very [BSK17]. VI [Cha18b]. Via [BH10, HS10, KMW10, MM13a, Ade18, Ade21, ACB19, CC16, CG19a, Hin18, YF18]. vice [San14]. Viète [MO20b]. view [KH11]. visible [LTZ20]. Vladut [Leb10]. voisinage [Tou09]. Vojta [Dil20, RyW22]. Volkenborn [DA18]. Volume [Ano05, Ano06, Ano07a, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13, Ano14, Ano15a, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21]. Voronoï [Ass21].

W. [GL16, Mao17, Meš12]. Walum [Bor22]. Waring [Li16, LZ18d, LZ18c, Liu12d, Liu16, ZL18a]. Warnaar [JKK16]. Watt [Har08]. Weak [KRS10, Do17, Tho06]. Weakening [McN13]. Weakly [Gri11, HJ14, JSS14, hF22, Han17, Zha16]. Weber [FK11, Har07]. Weierstrass [Dic21, Fin20, LR17]. Weight [Bra14, CK13, CWW08, ELY05, Gri11, HKKL12, Koh10, Raj09, Ter13, AAA15, BP11, BB14b, Dai16a, Dic21, Dic15, JM16, KP16, Mau17, MM14, Moo19, NMZJ22, PR21, PSY17, Roe14, Ste16, Wal17b, Wil19]. Weighted [Che22c, Li11, Mac12, Nob12, OLG19, OEL13, Alk20, BLW21, CP17, CP18a, ELO17, GLZ15, KKL21, KE18, LDSM19, MOS14a, MORS16, MT21, SC16]. weights [PSY18]. Weil [AM09, AB09, BGm22, Kat21, KRS10, LL22, Mat17a, SV15]. Weil-pairing [Kat21]. Well [FP12b, FHL^{+}13, Küh12]. Well-Rounded [FP12b, FHL^{+}13, Küh12]. Weyl [Li11, Yao18]. Which [Min12, Suz05, AAAW17, BK22, LLS21, MW16, Mos15]. Whipple [WW18]. Whipple-type [WW18]. whole [Yas16]. Whose [AK14, AITA19, Cho15, DS18b]. Wieferich [BB21b]. Wiener [RD13]. Wild [Cop20]. Wildly [DD10]. Wilf [HMST16]. Wilson [CP08]. Winquist [CLN05]. wise [Hu13a, SD20]. without [Dub18, JS15]. Witt [Fin14, Mil13, Ren19]. Wolstenholme [Ros13, Zha08]. WP [ZH16b]. Wreath [Klü12]. Writing [ABCM14, Mur15]. Wronskians [MMO08].

Yee [MS22, TB06]. Yokoi [IN22].

Z [AC15, Hyo15]. Z. [GL16, Mao17, Meš12]. Z.-W. [GL16, Mao17, Meš12]. zêta [JM10]. Zagier [Vs21, BKY19, BKY20, Fla18, Jam16]. Zaharescu [TB06]. Zannier [PZ18]. Zeilberger [Liu17a]. Zeng [NP18]. Zero [BGW11, GZZ15, Gek11, Gil13, GY16, HN11, Kad12, LP13, Mer11a, MS07, SX14, She22, Xia13, Zor13, AM15, Aym22, CP17, CP18a, GLPW18, Ge18, HZ18, LDSM19, MR15, PQSW14, PHLS19, Ram17, RL19, Sca17, TYZ16, YL16, ZW14a]. Zero-Free [HN11, Kad12, Mer11a]. zero-one [Ram17]. Zero-Sum [LP13, SX14, Xia13, GZZ15, AM15, CP17, CP18a, GLPW18, Ge18, HZ18, LDSM19, PHLS19, YL16, ZW14a]. Zeroes [Rod14, Moo19]. Zeros [AT22, Cho13a, Far08, Fuk08, GK13, Gun06, HJ14, Han17, JSS14, Suz05, DS18b, KTZ16, KTZ18, Lou15, MR15, SA09, XZ21, Zam16]. Zeta [Alk12, Bet10, Cou08, Far08, FP12a, HS19a, HN07, HN10, HN11, Hea14, Ivi05, Kad12, KO10, Lau13, LZ11, Li09, Li13, Mac12, OEL13, Rod14, RSW14, Sas15, TY13, Wak12b, AD18, AMSV21, BT18a, BB14a, BC10, BJMV14, DK21, ELO16, ELO17, EL18, GLZ15, Haj20, Har18, Hof17, Ivi16, Jar22, JVW20, JM10, Kan22, Kas19, KTZ16, KTZ18, KP19, Li15, LP20, LQ21, Lou15, Ma17, Mac16, MR15, Mer11a, MO20a, Mur17, MS18b, MP21, Ono21, Pat19b, SS21, SY19a, She22, Sun16c, Wak17, WV21, Yas16, You16b, You17, dS18a]. Zeta-Function [Ivi05, Lau13, Ivi16, Li15, MP21]. Zeta-Star [KO10, TY13, Wak12b]. Zhang [Fla18]. Ziv [FZ16, ZW14a]. Zsigmondy [ZJ19].

References

Abe:2020:STM

[AA20] Ryuji Abe and Iain R. Aitchison. Symmetry of topographs of Markoff forms. International Journal of Number Theory (IJNT), 16(7):1573–1598, August 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                                    tronic). URL https://www.
                                                    worldscientific.com/doi/
                                                    10.1142/S1793042120500839.

Alaca:2015:FCC

[AAA15] Ayşe Alaca, Şaban Alaca, and Zafer Selçuk Aygin. Fourier coefficients of a class of eta quotients of weight 2. International Journal of Number Theory (IJNT), 11(8):2381–2392, December 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                                    tronic). URL https://www.
                                                    worldscientific.com/doi/
                                                    10.1142/S1793042115501109.

Alaca:2017:IPC

[AAAW17] Ayşe Alaca, Şaban Alaca, Zafer Selcuk Aygin, and Kenneth S. Williams. Infinite products with coefficients which vanish on certain arithmetic progressions. International Journal of Number Theory (IJNT), 13(5):1095– 1117, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                                    tronic). URL https://www.
                                                    worldscientific.com/doi/
                                                    10.1142/S1793042117500592.

Astaneh-Asl:2011:IHP

[AAD11] Ali Astaneh-Asl and Hassan Daghigh. Independence of Heegner points for nonmaximal orders. International Journal of Number Theory (IJNT), 7(3): 663–669, May 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                                    tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042111004241.

Alaca:2008:TFI

[AALW08] Ayşe Alaca, Şaban Alaca, Mathieu F. Lemire, and Kenneth S. Williams. Theta function identities and representations by certain quaternary quadratic forms. International Journal of Number Theory (IJNT), 4(2):219– 239, April 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001304.

Alaca:2009:NRP

[AALW09] Ayşe Alaca, Şaban Alaca, Mathieu F. Lemire, and Kenneth S. Williams. The number of representations of a positive integer by certain quaternary quadratic forms. International Journal of Number Theory (IJNT), 5(1):13–40, February 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109001943.

Abe:2016:TCM

[AAR16] Ryuji Abe, Iain R. Aitchison, and Benoît Rittaud. Two-color Markoff graph and minimal forms. International Journal of Number Theory (IJNT), 12(4):1093–1122, June 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042116500676.

Akbary:2008:SCP

[AAW08a] Amir Akbary, Sean Alaric, and Qiang Wang. On some classes of permutation polynomials. International Journal of Number Theory (IJNT), 4(1):121–133, February 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042108001249.

Alaca:2008:BCF

[AAW08b] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. Berndt’s curious formula. International Journal of Number Theory (IJNT), 4(4):677– 689, August 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042108001614.

Alaca:2010:FOQ

[AAW10a] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. Fourteen octonary quadratic forms. International Journal of Number Theory (IJNT), 6(1):37–50, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211000279X.

Alaca:2010:SQF

[AAW10b] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. Sextenary quadratic forms and an identity of Klein and Fricke. International Journal of Number Theory (IJNT), 6(1):169– 183, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110002880.

Alaca:2016:RNC

[AAW16] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. Representation numbers of certain quaternary quadratic forms in a genus consisting of a single class. International Journal of Number Theory (IJNT), 12(6):1529–1573, September 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500950.

Angles:2009:WNC

[AB09] Bruno Anglès and Tatiana Beliaeva. On Weil numbers in cyclotomic fields. International Journal of Number Theory (IJNT), 5(5):871–884, August 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002432.

Ahmed:2015:NCA

[AB15a] Zakir Ahmed and Nayandeep Deka Baruah. New congruences for Andrews’ singular overpartitions. International Journal of Number Theory (IJNT), 11(7):2247–2264, November 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115501018.

Ahmed:2015:PRB

[AB15b] Zakir Ahmed and Nayandeep Deka Baruah. Parity results for broken 5diamond, 7-diamond and 11diamond partitions. International Journal of Number Theory (IJNT), 11(2):527–542, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211550027X.

Aryasomayajula:2018:EAC

[AB18] Anilatmaja Aryasomayajula and Baskar Balasubramanyam. Estimates of automorphic cusp forms over quaternion algebras. International Journal of Number Theory (IJNT), 14(4):1143– 1170, May 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500719.

Achenjang:2019:GCI

[AB19] Niven Achenjang and Aaron Berger. On gaps in the closures of images of divisor functions. International Journal of Number Theory (IJNT), 15(5):1023– 1036, June 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500556.

Abrate:2014:WSA

[ABCM14] Marco Abrate, Stefano Barbero, Umberto Cerruti, and Nadir Murru. Writing π as sum of arctangents with linear recurrent sequences, golden mean and Lucas numbers. International Journal of Number Theory (IJNT), 10(5):1309– 1319, August 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500286.

Ayad:2019:INF

[ABK19] Mohamed Ayad, Rachid Bouchenna, and Omar Kihel. Indices in a number field II. International Journal of Number Theory (IJNT), 15(1):89–103, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501750.

Abouzaid:2008:HLA

[Abo08] Mourad Abouzaid. Heights and logarithmic gcd on algebraic curves. International Journal of Number Theory (IJNT), 4(2):177–197, April 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001298.

Alvanos:2009:CAC

[ABP09] Paraskevas Alvanos, Yuri Bilu, and Dimitrios Poulakis. Characterizing algebraic curves with infinitely many integral points. International Journal of Number Theory (IJNT), 5(4):585– 590, June 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002274.

Arenas-Carmona:2013:EOT

[AC13] Luis Arenas-Carmona. Eichler orders, trees and representation fields. International Journal of Number Theory (IJNT), 9(7):1725–1741, November 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211350053X.

Adam:2015:BFS

[AC15] David Adam and Jean-Luc Chabert. On Bhargava’s factorials of the set of twin primes in z and in F_{q}[T]. International Journal of Number Theory (IJNT), 11(6):1941–1959, September 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500840.

Arenas-Carmona:2019:CEN

[ACB19] Luis Arenas-Carmona and Claudio Bravo. Computing embedding numbers and branches of orders via extensions of the Bruhat–Tits tree. International Journal of Number Theory (IJNT), 15(10):2067–2088, November 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119501136.

Adiga:2005:GRB

[ACH05] Chandrashekar Adiga, Shaun Cooper, and Jung Hun Han. A general relation between sums of squares and sums of triangular numbers. International Journal of Number Theory (IJNT), 1(2):175– 182, June 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000078.

Andrews:2009:SIA

[ACS09] George E. Andrews, Sylvie Corteel, and Carla D. Savage. On q-series identities arising from lecture hall partitions. International Journal of Number Theory (IJNT), 5(2):327– 337, March 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002134.

Arenas-Carmona:2016:SBB

[ACS16] Luis Arenas-Carmona and Ignacio Saavedra. On some branches of the Bruhat–Tits tree. International Journal of Number Theory (IJNT), 12(3): 813–831, May 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500524.

Adansie:2018:NIF

[ACX18] Prince Adansie, Shane Chern, and Ernest X. W. Xia. New infinite families of congruences for the number of tagged parts over partitions with designated summands. International Journal of Number Theory (IJNT), 14(7):1935–1942, August 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118501154.

Alaca:2016:LIT

[AD16a]  Şaban Alaca and Greg Doyle.

Liouville identities with two functions. International Journal of Number Theory (IJNT), 12(5):1345–1363, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042116500822.

Ash:2016:RSP

[AD16b] Avner Ash and Darrin Doud. Relaxation of strict parity for reducible Galois representations attached to the homology of GL(3, Z). International Journal of Number Theory (IJNT), 12(2):361–381, March 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500226.

Alladi:2018:RRZ

[AD18] Krishnaswami Alladi and Colin Defant. Revisiting the Riemann zeta function at positive even integers. International Journal of Number Theory (IJNT), 14(7):1849–1856, August 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501105.

Adam:2012:AAF

[Ada12] David Adam. Abel arithmetic functions in finite characteristic. International Journal of Number Theory (IJNT), 8 (6):1557–1568, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042112500959.

Adelberg:2018:AAS

[Ade18] Arnold Adelberg. The p-adic analysis of Stirling numbers via higher order Bernoulli numbers. International Journal of Number Theory (IJNT), 14(10): 2767–2779, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042118501671.

Adelberg:2021:AAS

[Ade21]  Arnold Adelberg.  The 2-

adic analysis of Stirling numbers of the second kind via higher order Bernoulli numbers and polynomials. International Journal of Number Theory (IJNT), 17(01):55–70, February 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042121500044.

Adiceam:2015:RAA

[Adi15]  Faustin Adiceam.  Ratio-

nal approximation and arithmetic progressions. International Journal of Number Theory (IJNT), 11(2):451–486, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115500232.

Akbary:2012:GVT

[AG12] Amir Akbary and Dragos Ghioca. A geometric variant of Titchmarsh divisor problem. International Journal of Number Theory (IJNT), 8(1):53–69, February 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500030.

Alvarado:2013:APC

[AG13] Alejandra Alvarado and Edray Herber Goins. Arithmetic progressions on conic sections. International Journal of Number Theory (IJNT), 9(6):1379–1393, September 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500322.

Agashe:2010:SSV

[Aga10] Amod Agashe. Squareness in the special L-value and special L-values of twists. International Journal of Number Theory (IJNT), 6(5):1091– 1111, August 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003393.

Aggarwal:2021:NSB

[Agg21] Keshav Aggarwal. A new subconvex bound for GL(3) Lfunctions in the t-aspect. International Journal of Number Theory (IJNT), 17(05):1111– 1138, June 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500275.

Alkabouss:2018:E

[AGL18] Sakha A. Alkabouss, Tarek Garici, and Jesse Larone. On the equation Res_{x}(P(x), x^{2} + sx + t) = a. International Journal of Number Theory (IJNT), 14(4):1073–1079, May 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500653.

Agnihotri:2022:LSA

[Agn22] Rishabh Agnihotri. Lambert series associated to Hilbert modular form. International Journal of Number Theory (IJNT), 18(06):1335–1349, July 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500671.

Angles:2007:RI

[AH07] Bruno Anglès and Thomas Herreng. On a result of Iwasawa. International Journal of Number Theory (IJNT), 3(1):19– 41, March 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042107000791.

Aygin:2019:RCS

[AH19] Zafer Selcuk Aygin and Nankun Hong. Ramanujan’s convolution sum twisted by Dirichlet characters. International Journal of Number Theory (IJNT), 15(1):137–152, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042119500027.

Aouissi:2019:FWC

[AITA19] Siham Aouissi, Moulay Chrif Ismaili, Mohamed Talbi, and Abdelmalek Azizi. Fields Q( ^{3} \sqrt{d}, ζ_{3}) whose 3-class group is of type (9, 3). International Journal of Number Theory (IJNT), 15(7):1437–1447, August 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042119500817.

AlHajjShehadeh:2009:GFH

[AJKM09] Hala Al Hajj Shehadeh, Samar Jaafar, and Kamal Khuri-Makdisi. Generating functions for Hecke operators. International Journal of Number Theory (IJNT), 5(1):125–140, February 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109001979.

Akbary:2021:MTP

[AjW21] Amir Akbary and Peng jie Wong. On the moments of torsion points modulo primes and their applications. International Journal of Number Theory (IJNT), 17(04):1029– 1045, May 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500226.

Ayad:2011:CDV

[AK11] Mohamed Ayad and Omar Kihel. Common divisors of values of polynomials and common factors of indices in a number field. International Journal of Number Theory (IJNT), 7(5):1173–1194, August 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004526.

Acquaah:2012:PFO

[AK12a] Peter Acquaah and Sergei Konyagin. On prime factors of odd perfect numbers. International Journal of Number Theory (IJNT), 8(6):1537–1540, September 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500935.

Ayad:2012:RPU

[AK12b] Mohamed Ayad and Omar Kihel. Recognizing the primes using permutations. International Journal of Number Theory (IJNT), 8(8):2045–2057, December 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112501163.

Alaca:2014:RCOa

[AK14]  Şaban Alaca and Yavuz Kesi-

cioğlu. Representations by certain octonary quadratic forms whose coefficients are 1, 2, 3 and 6. International Journal of Number Theory (IJNT), 10(1):133–150, February 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500851.

Ahlgren:2015:DSF

[AK15a] Scott Ahlgren and Byungchan Kim. Dissections of a “strange” function. International Journal of Number Theory (IJNT), 11(5):1557–1562, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115400072.

Alaca:2015:ROQ

[AK15b] Ayşe Alaca and M. Nesibe Kesicioğlu. Representations by octonary quadratic forms with coefficients 1, 3 or 9. International Journal of Number Theory (IJNT), 11(8):2353–2368, December 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115501080.

Aoki:2015:SFU

[AK15c] Miho Aoki and Yasuhiro Kishi.

On systems of fundamental units of certain quartic fields. International Journal of Number Theory (IJNT), 11(7):2019– 2035, November 2015. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500864.

Alaca:2016:ECS

[AK16]  Şaban Alaca and Yavuz Kesi-

cioğlu. Evaluation of the convolution sums ∑_{l+27m=n} σ(l)σ(m) and ∑_{l+32m=n} σ(l)σ(m). International Journal of Number Theory (IJNT), 12(1):1– 13, February 2016. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500019.

Alaca:2017:ROQ

[AK17] Ayşe Alaca and M. Nesibe Kesicioğlu. Representations by octonary quadratic forms with coefficients 1, 2, 3 or 6. International Journal of Number Theory (IJNT), 13(3):735– 749, April 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500385.

Akarsu:2014:SIG

[Aka14] Emek Demirci Akarsu. Short incomplete Gauss sums and rational points on metaplectic horocycles. International Journal of Number Theory (IJNT), 10(6):1553–1576, September 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500444.

Akhtari:2012:UBN

[Akh12]  Shabnam Akhtari.  Upper

bounds for the number of solutions to quartic Thue equations. International Journal of Number Theory (IJNT), 8(2):335– 360, March 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500200.

Amdeberhan:2012:II

[AKMR12] Tewodros Amdeberhan, Christoph Koutschan, Victor H. Moll, and Eric S. Rowland. The iterated integrals of ln(1 + x^{n}). International Journal of Number Theory (IJNT), 8(1):71–94, February 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500042.

Alaca:2014:RCOb

[Ala14]  Ayşe Alaca.  Representa-

tions by certain octonary quadratic forms with coefficients 1, 3 or 9. International Journal of Number Theory (IJNT), 10(6):1365–1384, September 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211450033X.

Alan:2016:HFD

[Ala16] Murat Alan. Half-factorial domains and quadratic orders. International Journal of Number Theory (IJNT), 12(2):465– 472, March 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500275.

Alkan:2007:ASG

[Alk07] Emre Alkan. Average size of gaps in the Fourier expansion of modular forms. International Journal of Number Theory (IJNT), 3(2):207– 215, June 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107000870.

Alkan:2012:RSB

[Alk12] Emre Alkan. Ramanujan sums and the Burgess zeta function. International Journal of Number Theory (IJNT), 8 (8):2069–2092, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042112501187.

Alkan:2015:SRT

[Alk15] Emre Alkan. Series representing transcendental numbers that are not U-numbers. International Journal of Number Theory (IJNT), 11(3):869– 892, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500487.

Alkan:2020:BBW

[Alk20] Emre Alkan. Biased behavior of weighted Mertens sums. International Journal of Number Theory (IJNT), 16(3):547– 577, April 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500281.

Alladi:2009:NCS

[All09] Krishnaswami Alladi. A new combinatorial study of the Rogers–Fine identity and a related partial theta series. International Journal of Number Theory (IJNT), 5(7):1311– 1320, November 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042109002675.

Alladi:2017:MDE

[All17]  Krishnaswami Alladi.  A

multi-dimensional extension of Sylvester’s identity. International Journal of Number Theory (IJNT), 13(10):2487– 2504, November 2017. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211750138X.

Aloui:2019:ESI

[Alo19] Karam Aloui. Exponential sums involving q-digital functions. International Journal of Number Theory (IJNT), 15(6): 1143–1172, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042119500635.

Alrasasi:2014:KSS

[Alr14] Ibrahim Alrasasi. The Korselt set of the square of a prime. International Journal of Number Theory (IJNT), 10(4):875– 884, June 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500043.

Alvarez:2014:RRJ

[Álv14] A. Álvarez. The p-rank of the reduction modp of Jacobians and Jacobi sums. International Journal of Number Theory (IJNT), 10(8):2097–2114, December 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042114500705.

Alvey:2021:KTT

[Alv21] Daniel C. Alvey. A Khintchinetype theorem for affine subspaces. International Journal of Number Theory (IJNT), 17(06):1323–1341, July 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S179304212150038X.

Akbary:2009:RSM

[AM09] Amir Akbary and V. Kumar Murty. Reduction modp of subgroups of the Mordell–Weil group of an elliptic curve. International Journal of Number Theory (IJNT), 5(3):465– 487, May 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042109002225.

Adhikari:2015:PMS

[AM15] Sukumar Das Adhikari and Eshita Mazumdar. The polynomial method in the study of zero-sum theorems. International Journal of Number Theory (IJNT), 11(5):1451– 1461, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115400114.

Aktas:2017:FUC

[AM17] Kevser Aktaş and M. Ram Murty. Fundamental units and consecutive squarefull numbers. International Journal of Number Theory (IJNT), 13(1):243– 252, February 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500142.

Adhikari:2017:AED

[AMM17] S. D. Adhikari, Anirban Mukhopadhyay, and M. Ram Murty. The analog of the Erdős distance problem in finite fields. International Journal of Number Theory (IJNT), 13(9):2319– 2333, October 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117501275.

Asada:2017:FPG

[AMMS17] Megumi Asada, Sarah Manski, Steven J. Miller, and Hong Suh. Fringe pairs in generalized MSTD sets. International Journal of Number Theory (IJNT), 13(10):2653–2675, November 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117501470.

Amoroso:2007:MPE

[Amo07] Francesco Amoroso. Une minoration pour l’exposant du groupe des classes d’un corps engendré par un nombre de Salem. (French) [A lower bound for the exponent of the group of classes of a field generated by a number of Salem]. International Journal of Number Theory (IJNT), 3(2):217– 229, June 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107000882.

Assim:2017:TIU

[AMO17] Jilali Assim, Youness Mazigh, and Hassan Oukhaba. Théorie d’Iwasawa des unités de Stark et groupe de classes. (French) [Iwasawa theory of Stark units and class group]. International Journal of Number Theory (IJNT), 13(5):1165– 1190, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500634.

Amoroso:2018:MMG

[Amo18] Francesco Amoroso. Mahler measure on Galois extensions. International Journal of Number Theory (IJNT), 14(6):1605– 1617, July 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500963.

Amoros:2021:IGR

[Amo21] Laia Amorós. Images of Galois representations in mod p Hecke algebras. International Journal of Number Theory (IJNT), 17(05):1265– 1285, June 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042121500354.

Aubert:2017:PDI

[AMPS17] Anne-Marie Aubert, Sergio Mendes, Roger Plymen, and Maarten Solleveld. On Lpackets and depth for SL_{2}(K) and its inner form. International Journal of Number Theory (IJNT), 13(10):2545–2568, November 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117501421.

Amdeberhan:2021:TIA

[AMSV21] Tewodros Amdeberhan, Victor H. Moll, Armin Straub, and Christophe Vignat. A triple integral analog of a multiple zeta value. International Journal of Number Theory (IJNT), 17(02):223–237, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400102.

Ancona:2017:DHS

[Anc17] Giuseppe Ancona. Degeneration of Hodge structures over Picard modular surfaces. International Journal of Number Theory (IJNT), 13(5):1145– 1164, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500622.

Andrews:2007:FD

[And07]  George E. Andrews.  A

Fine dream. International Journal of Number Theory (IJNT), 3(3):325–334, September 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042107000948.

Andrade:2012:NMV

[And12] Julio Andrade. A note on the mean value of L-functions in function fields. International Journal of Number Theory (IJNT), 8(7):1725–1740, November 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500996.

Andersen:2013:HTC

[And13]  Nickolas Andersen.  Hecke-

type congruences for two smallest parts functions. International Journal of Number Theory (IJNT), 9(3):713– 728, May 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112501564.

Andrews:2015:SO

[And15] George E. Andrews. Singular overpartitions. International Journal of Number Theory (IJNT), 11(5):1523–1533, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042115400059.

Anders:2016:OBC

[And16] Katie Anders. Odd behavior in the coefficients of reciprocals of binary power series. International Journal of Number Theory (IJNT), 12(3):635– 648, May 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211650041X.

Andrews:2021:Pa

[And21a] George E. Andrews. Preface. International Journal of Number Theory (IJNT), 17 (02):v–vii, March 2021. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120020017.

Andrews:2021:Pb

[And21b] George E. Andrews. Preface. International Journal of Number Theory (IJNT), 17 (03):v–vii, April 2021. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120020029.

Andrews:2021:RDI

[And21c]  George E. Andrews.  The

Ramanujan—Dyson identities and George Beck’s congruence conjectures. International Journal of Number Theory (IJNT), 17(02):239–249, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120400060.

Ahmad:2014:PIB

[ANH14] Shahzad Ahmad, Toru Nakahara, and Syed Muhammad Husnine. Power integral bases for certain pure sextic fields. International Journal of Number Theory (IJNT), 10(8):2257–2265, December 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042114500778.

Anonymous:2005:AIV

[Ano05]  Anonymous.  Author in-

dex (volume 1). International Journal of Number Theory (IJNT), 1(4):603–605, December 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042105000376.

Anonymous:2006:AIV

[Ano06]  Anonymous.  Author in-

dex (volume 2). International Journal of Number Theory (IJNT), 2(4):617–619, December 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042106000760.

Anonymous:2007:AIV

[Ano07a]  Anonymous.  Author in-

dex (volume 3). International Journal of Number Theory (IJNT), 3(4):653–654, December 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107001164.

Anonymous:2007:P

[Ano07b] Anonymous. Preface. International Journal of Number Theory (IJNT), 3(3):v– vi, September 2007. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107001061.

Anonymous:2008:AIV

[Ano08] Anonymous. Author index (volume 4). International Journal of Number Theory (IJNT), 4(6):1069–1072, December 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001900.

Anonymous:2009:AIV

[Ano09] Anonymous. Author index (volume 5). International Journal of Number Theory (IJNT), 5(8):1489–1493, December 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002766.

Anonymous:2010:AIV

[Ano10] Anonymous. Author index (volume 6). International Journal of Number Theory (IJNT), 6(8):1945–1950, December 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003885.

Anonymous:2011:AIV

[Ano11] Anonymous. Author index (volume 7). International Journal of Number Theory (IJNT), 7(8):2289–2295, December 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111005118.

Anonymous:2012:AIV

[Ano12] Anonymous. Author index (volume 8). International Journal of Number Theory (IJNT), 8(8):2093–2098, December 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112990011.

Anonymous:2013:AIV

[Ano13] Anonymous. Author index (volume 9). International Journal of Number Theory (IJNT), 9(8):2129–2135, December 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113990017.

Anonymous:2014:AIV

[Ano14] Anonymous. Author index (volume 10). International Journal of Number Theory (IJNT), 10(8):2267–2273, December 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114990012.

Anonymous:2015:AIV

[Ano15a] Anonymous. Author index (volume 11). International Journal of Number Theory (IJNT), 11(8):2513–2520, December 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115990018.

Anonymous:2015:HHB

[Ano15b] Anonymous. Heini Halberstam — biography. International Journal of Number Theory (IJNT), 11(5):xix–xxii, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115760027.

Anonymous:2015:PBB

[Ano15c] Anonymous. Paul T. Bateman — biography. International Journal of Number Theory (IJNT), 11(5):xv–xviii, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115760015.

Anonymous:2016:AIV

[Ano16] Anonymous. Author index (volume 12). International Journal of Number Theory (IJNT), 12(8):2343–2350, December 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116990013.

Anonymous:2017:AIV

[Ano17]  Anonymous.  Author in-

dex (volume 13). International Journal of Number Theory (IJNT), 13(10):2791–2798, November 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117990019.

Anonymous:2018:AIV

[Ano18]  Anonymous.  Author in-

dex (volume 14). International Journal of Number Theory (IJNT), 14(10):2781–2789, November 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118990014.

Anonymous:2019:AIV

[Ano19]  Anonymous.  Author in-

dex (volume 15). International Journal of Number Theory (IJNT), 15(10):2241–2248, November 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211999001X.

Anonymous:2020:AIV

[Ano20]  Anonymous.  Author in-

dex (volume 16). International Journal of Number Theory (IJNT), 16(10):2379–2385, November 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120990018.

Anonymous:2021:AIV

[Ano21]  Anonymous.  Author in-

dex (volume 17). International Journal of Number Theory (IJNT), 17(10):2395–2402, November 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121990013.

Ash:2008:EUA

[AP08] Avner Ash and David Pollack. Everywhere unramified automorphic cohomology for SL_{3}(Z). International Journal of Number Theory (IJNT), 4(4):663–675, August 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001602.

Ash:2017:TEE

[AP17] Avner Ash and David Pollack. Toward an explicit eigencurve for GL(3). International Journal of Number Theory (IJNT), 13(5):1213– 1231, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117500658.

Apagodu:2018:EPC

[Apa18] Moa Apagodu. Elementary proof of congruences involving sum of binomial coefficients. International Journal of Number Theory (IJNT), 14(6):1547– 1557, July 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500938.

Arango-pineros:2022:MTC

[ApKK22] Santiago Arango-piñeros, Daniel Keliher, and Christopher Keyes. Mertens’ theorem for Chebotarev sets. International Journal of Number Theory (IJNT), 18(08):1823–1842, September 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S1793042122500932.

Anavi:2013:CF

[APP13] Aria Anavi, Paul Pollack, and Carl Pomerance. On congruences of the form σ(n) ≡ a(modn). International Journal of Number Theory (IJNT), 9(1):115–124, February 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042112501266.

Alaca:2014:NRP

[APW14]  Şaban Alaca, Lerna Pehli-

van, and Kenneth S. Williams.

On the number of representations of a positive integer as a sum of two binary quadratic forms. International Journal of Number Theory (IJNT), 10(6):1395–1420, September 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500353.

Aribam:2013:ISG

[Ari13] Chandrakant S. Aribam. On μinvariants of Selmer groups of some CM elliptic curves. International Journal of Number Theory (IJNT), 9(5):1199– 1214, August 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500206.

Azizi:2019:HCF

[ARTZ19] Abdelmalek Azizi, Mohammed Rezzougui, Mohammed Taous, and Abdelkader Zekhnini. On the Hilbert 2-class field of some quadratic number fields. International Journal of Number Theory (IJNT), 15(4):807– 824, May 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211950043X.

Aryan:2017:QDP

[Ary17]  Farzad Aryan.  On the

quadratic divisor problem. International Journal of Number Theory (IJNT), 13(6):1457– 1471, July 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117500816.

Ayuso:2006:NST

[AS06] Pedro Fortuny Ayuso and Fritz Schweiger. A new symmetric two-dimensional algorithm. International Journal of Number Theory (IJNT), 2(4):489–498, December 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042106000681.

Adolphson:2009:ESI

[AS09] Alan Adolphson and Steven Sperber. Exponential sums on A^{n}. IV. International Journal of Number Theory (IJNT), 5 (5):747–764, August 2009. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042109002353.

Al-Shaghay:2016:ABC

[ASD16] Abdullah Al-Shaghay and Karl Dilcher. Analogues of the binomial coefficient theorems of Gauss and Jacobi. International Journal of Number Theory (IJNT), 12(8):2125–2145, December 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211650127X.

Al-saedi:2020:IDT

[AsMS20] Ali Al-saedi, Thomas Morrill, and Holly Swisher. Inequalities for the d-th residual crank moments of overpartitions. International Journal of Number Theory (IJNT), 16(7):1599– 1606, August 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500840.

Assing:2021:AVS

[Ass21] Edgar Assing. Adelic Voronoï summation and subconvexity for GL(2) L-functions in the depth aspect. International Journal of Number Theory (IJNT), 17(07):1537–1568, August 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500470.

Amri:2022:RJC

[AST22] Myriam Amri, Lukas Spiegelhofer, and Jörg Thuswaldner. Répartition jointe dans les classes de résidus de la somme des chiffres pour deux représentations d’Ostrowski. International Journal of Number Theory (IJNT), 18(05): 955–976, June 2022. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500506.

Ahmadi:2010:MOG

[ASV10] Omran Ahmadi, Igor E. Shparlinski, and José Felipe Voloch. Multiplicative order of Gauss periods. International Journal of Number Theory (IJNT), 6(4):877– 882, June 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003290.

Anbar:2020:BTP

[AT20] Nurdagül Anbar and Seher Tutdere. Belyi’s theorems in positive characteristic. International Journal of Number Theory (IJNT), 16(6):1355– 1368, July 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500712.

AlHamdani:2022:ZBC

[AT22] Summer Al Hamdani and Khang Tran. Zeros of a binomial combination of Chebyshev polynomials. International Journal of Number Theory (IJNT), 18(04):799– 811, May 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500427.

Aouissi:2020:CL

[ATIA20] Siham Aouissi, Mohamed Talbi, Moulay Chrif Ismaili, and Abdelmalek Azizi. On a conjecture of Lemmermeyer. International Journal of Number Theory (IJNT), 16(7):1407– 1424, August 2020. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500748.

Azizi:2016:GT

[ATT^{+}16] Abdelmalek Azizi, Mohamed Talbi, Mohammed Talbi, Aïssa Derhem, and Daniel C. Mayer. The group Gal(k_{3}^{(2)}|k) for k = Q(\sqrt{−3}, \sqrt{d}) of type (3, 3). International Journal of Number Theory (IJNT), 12(7):1951– 1986, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042116501207.

Akhtari:2019:LBM

[AV19] Shabnam Akhtari and Jeffrey D. Vaaler. Lower bounds for Mahler measure that depend on the number of monomials. International Journal of Number Theory (IJNT), 15(7):1425– 1436, August 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500805.

Alaca:2012:QF

[AW12] Ayşe Alaca and Kenneth S.

Williams. On the quaternary forms x^{2} + y^{2} + 2z^{2} + 3t^{2}, x^{2} + 2y^{2} + 2z^{2} + 6t^{2}, x^{2} + 3y^{2} + 3z^{2} + 6t^{2} and 2x^{2} + 3y^{2} + 6z^{2} + 6t^{2}. International Journal of Number Theory (IJNT), 8(7):1661–1686, November 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500972.

Achter:2013:QEC

[AW13] Jeffrey D. Achter and Siman Wong. Quotients of elliptic curves over finite fields. International Journal of Number Theory (IJNT), 9(6):1395– 1412, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042113500334.

Awtrey:2012:DAF

[Awt12] Chad Awtrey. Dodecic 3-adic fields. International Journal of Number Theory (IJNT), 8(4): 933–944, June 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500558.

Aymone:2022:NPN

[Aym22] Marco Aymone. A note on prime number races and zero free regions for L functions. International Journal of Number Theory (IJNT), 18(01): 1–8, February 2022. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500014.

Alkan:2005:NRT

[AZ05] Emre Alkan and Alexandru Zaharescu. Nonvanishing of the Ramanujan Tau function in short intervals. International Journal of Number Theory (IJNT), 1(1):45– 51, March 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000029.

Azaiez:2009:RHM

[Aza09] Najib Ouled Azaiez. Restrictions of Hilbert modular forms. International Journal of Number Theory (IJNT), 5(1):67–80, February 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109001931.

Azizi:2015:PCG

[AZTM15] Abdelmalek Azizi, Abdelkader Zekhnini, Mohammed Taous, and Daniel C. Mayer. Principalization of 2-class groups of type (2, 2, 2) of biquadratic fields. International Journal of Number Theory (IJNT), 11(4):1177– 1215, June 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500645.

Alkan:2005:AFS

[AZZ05] Emre Alkan, Alexandru Zaharescu, and Mohammad Zaki. Arithmetical functions in several variables. International Journal of Number Theory (IJNT), 1(3):383–399, September 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304210500025X.

Bacher:2008:NIH

[Bac08] Roland Bacher. A new inequality for the Hermite constants. International Journal of Number Theory (IJNT), 4(3):363– 386, June 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001390.

Badziahin:2017:FSF

[Bad17] Dmitry Badziahin. Finding special factors of values of polynomials at integer points. International Journal of Number Theory (IJNT), 13(1):209–228, February 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500129.

Bae:2019:STH

[Bae19] Sunghan Bae. Statistics of traces of high powers of the Frobenius for biquadratic function fields over finite fields. International Journal of Number Theory (IJNT), 15(8):1675– 1691, September 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042119500933.

Baier:2013:MIS

[Bai13] S. Baier. Multiplicative inverses in short intervals. International Journal of Number Theory (IJNT), 9(4):877– 884, June 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500036.

Baier:2016:ECS

[Bai16] Stephan Baier. Elliptic curves with square-free Δ. International Journal of Number Theory (IJNT), 12(3):737– 764, May 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500482.

Bajnok:2009:MSS

[Baj09] Béla Bajnok. On the maximum size of a (k, l)-sum-free subset of an abelian group. International Journal of Number Theory (IJNT), 5(6):953–971, September 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002481.

Bajolet:2014:FSM

[Baj14] Aurelien Bajolet. Finding singular moduli on a complex line. International Journal of Number Theory (IJNT), 10(3):763– 777, May 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042113501212.

Balandraud:2008:IMN

[Bal08] Éric Balandraud. The isoperimetric method in non-Abelian groups with an application to optimally small sumsets. International Journal of Number Theory (IJNT), 4(6):927–958, December 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001821.

Balestrieri:2019:BMO

[Bal19] Francesca Balestrieri. Brauer– Manin obstruction and families of generalised Châtelet surfaces over number fields. International Journal of Number Theory (IJNT), 15(2):289– 308, March 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211950012X.

Baoulina:2005:PC

[Bao05]  Ioulia Baoulina.  On a

problem of Carlitz. International Journal of Number Theory (IJNT), 1(2):293– 307, June 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000194.

Baoulina:2006:EFF

[Bao06] Ioulia Baoulina. On the equation over a finite field. International Journal of Number Theory (IJNT), 2(3):351–363, September 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000607.

Baoulina:2008:NSE

[Bao08]  Ioulia Baoulina.  On the

number of solutions to the equation (x_{1} + ⋯ + x_{n})^{2} = ax_{1} ⋯ x_{n} in a finite field. International Journal of Number Theory (IJNT), 4(5):797– 817, October 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001675.

Baoulina:2010:NSC

[Bao10] Ioulia Baoulina. On the number of solutions to certain diagonal equations over finite fields. International Journal of Number Theory (IJNT), 6 (1):1–14, February 2010. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110002776.

Baoulina:2014:MHE

[Bao14]  Ioulia N. Baoulina.  On

Markoff–Hurwitz equations over residue class rings. International Journal of Number Theory (IJNT), 10(2):421–454, March 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113501029.

Baoulina:2015:GMH

[Bao15] Ioulia N. Baoulina. Generalizations of the Markoff–Hurwitz equations over residue class rings. International Journal of Number Theory (IJNT), 11(3): 909–930, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115500505.

Bao:2019:PIB

[Bao19] Dianbin Bao. Polynomial identities between Hecke eigenforms. International Journal of Number Theory (IJNT), 15(10): 2135–2150, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042119501173.

Barman:2013:ALI

[Bar13a]  Rupam Barman.  Another

look at Iwasawa λ-invariants of p-adic measures on Z_{p}^{n} and Γ-transforms. International Journal of Number Theory (IJNT), 9(5):1289–1299, August 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500279.

Barth:2013:ITO

[Bar13b] Peter Barth. Iwasawa theory for one-parameter families of motives. International Journal of Number Theory (IJNT), 9 (2):257–319, March 2013. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501357.

Bartel:2014:FEG

[Bar14] Alex Bartel. Factor equivalence of Galois modules and regulator constants. International Journal of Number Theory (IJNT), 10(1):1–12, February 2014. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500772.

Bautista-Ancona:2013:GFC

[BARCVS13] Víctor Bautista-Ancona, Martha Rzedowski-Calderón, and Gabriel Villa-Salvador. Genus fields of cyclic l-extensions of rational function fields. International Journal of Number Theory (IJNT), 9(5):1249–1262, August 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500243.

Basic:2012:DPD

[Baš12] Bojan Bašić. On d-digit palindromes in different bases: the number of bases is unbounded. International Journal of Number Theory (IJNT), 8(6):1387– 1390, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042112500819.

Batir:2017:SCI

[Bat17a]  Necdet Batir.  On some

combinatorial identities and harmonic sums. International Journal of Number Theory (IJNT), 13(7):1695–1709, August 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211750097X.

Batson:2017:LTR

[Bat17b] Scott C. Batson. The linear transformation that relates the canonical and coefficient embeddings of ideals in cyclotomic integer rings. International Journal of Number Theory (IJNT), 13(9):2277–2297, October 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117501251.

Bazzanella:2011:TCR

[Baz11] Danilo Bazzanella. Two conditional results about primes in short intervals. International Journal of Number Theory (IJNT), 7(7):1753–1759, November 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004563.

Borwein:2006:TTG

[BB06] Jonathan M. Borwein and David M. Bradley. Thirtytwo Goldbach variations. International Journal of Number Theory (IJNT), 2(1):65– 103, March 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL http://

docserver.carma.newcastle.
edu.au/301/;   https://www.
worldscientific.com/doi/
10.1142/S1793042106000383.

Baccar:2009:SSP

[BB09a] N. Baccar and F. Ben Saïd.

On sets such that the partition function is even from a certain point on. International Journal of Number Theory (IJNT), 5(3):407– 428, May 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002195.

Bell:2009:BSF

[BB09b] Jason P. Bell and Jonathan W.

Bober. Bounded step functions and factorial ratio sequences. International Journal of Number Theory (IJNT), 5 (8):1419–1431, December 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042109002742.

Bayad:2014:RBB

[BB14a] Abdelmejid Bayad and Matthias Beck. Relations for Bernoulli– Barnes numbers and Barnes zeta functions. International Journal of Number Theory (IJNT), 10(5):1321–1335, August 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042114500298.

Boylan:2014:HOC

[BB14b] Matthew Boylan and Kenny Brown. Hecke operators on certain subspaces of integral weight modular forms. International Journal of Number Theory (IJNT), 10(7):1909–1919, November 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042114500614.

Baier:2018:LSP

[BB18a] Stephan Baier and Arpit Bansal. The large sieve with power moduli for Z[i]. International Journal of Number Theory (IJNT), 14(10):2737–2756, November 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118501658.

Baruah:2018:EGF

[BB18b] Nayandeep Deka Baruah and Nilufar Mana Begum. Exact generating functions for the number of partitions into distinct parts. International Journal of Number Theory (IJNT), 14(7):1995–2011, August 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118501191.

Banerjee:2020:IQC

[BB20a] Pradipto Banerjee and Ranjan Bera. An irreducibility question concerning modifications of Laguerre polynomials. International Journal of Number Theory (IJNT), 16(5):1031– 1051, June 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500530.

Baruah:2020:GFC

[BB20b] Nayandeep Deka Baruah and Nilufar Mana Begum. Generating functions and congruences for some partition functions related to mock theta functions. International Journal of Number Theory (IJNT), 16(2):423– 446, March 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500220.

Baustian:2020:ABD

[BB20c] Falko Baustian and Vladimir Bobkov. On asymptotic behavior of Dirichlet inverse. International Journal of Number Theory (IJNT), 16(6):1337– 1354, July 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500700.

Badr:2021:BSP

[BB21a] Eslam Badr and Francesc Bars.

Bielliptic smooth plane curves and quadratic points. International Journal of Number Theory (IJNT), 17(04):1047– 1066, May 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500238.

Bamunoba:2021:SCW

[BB21b] Alex Samuel Bamunoba and Jonas Bergström. A search for c-Wieferich primes. International Journal of Number Theory (IJNT), 17(07):1599– 1616, August 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500500.

Bonciocat:2013:ICS

[BBCM13] Nicolae Ciprian Bonciocat, Yann Bugeaud, Mihai Cipu, and Maurice Mignotte. Irreducibility criteria for sums of two relatively prime polynomials. International Journal of Number Theory (IJNT), 9 (6):1529–1539, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042113500413.

Beck:2005:DAG

[BBCZ05] Matthias Beck, Bruce C. Berndt, O-Yeat Chan, and Alexandru Zaharescu. Determinations of analogues of Gauss sums and other trigonometric sums. International Journal of Number Theory (IJNT), 1(3):333–356, September 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000200.

Balasuriya:2008:CES

[BBS08] Sanka Balasuriya, William D.

Banks, and Igor E. Shparlinski. Congruences and exponential sums with the sum of aliquot divisors function. International Journal of Number Theory (IJNT), 4(6):903–909, December 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304210800178X.

Borwein:2010:DTM

[BC10] Jonathan M. Borwein and O-Yeat Chan. Duality in tails of multiple-zeta values. International Journal of Number Theory (IJNT), 6(3): 501–514, May 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL http://

docserver.carma.newcastle.
edu.au/1218/;   https://
www.worldscientific.com/
doi/10.1142/S1793042110003058.

Ballot:2011:MD

[BC11] Christian Ballot and Mireille Car. On Murata densities. International Journal of Number Theory (IJNT), 7(7):1717– 1736, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S179304211100440X.

Bush:2018:FPM

[BC18] Albert Bush and Ernie Croot. Few products, many h-fold sums. International Journal of Number Theory (IJNT), 14 (8):2107–2128, September 2018. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118501270.

Bennett:2015:GFE

[BCDY15] Michael A. Bennett, Imin Chen, Sander R. Dahmen, and Soroosh Yazdani. Generalized Fermat equations: a miscellany. International Journal of Number Theory (IJNT), 11(1): 1–28, February 2015. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211530001X.

Bhargava:2016:PPC

[BCF16] Manjul Bhargava, John Cremona, and Tom Fisher. The proportion of plane cubic curves over Q that everywhere locally have a point. International Journal of Number Theory (IJNT), 12(4):1077– 1092, June 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500664.

Bhargava:2021:PGO

[BCF21] Manjul Bhargava, John Cremona, and Tom Fisher. The proportion of genus one curves over Q defined by a binary quartic that everywhere locally have a point. International Journal of Number Theory (IJNT), 17(04):903– 923, May 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500147.

Baruah:2008:SSS

[BCH08] Nayandeep Deka Baruah, Shaun Cooper, and Michael Hirschhorn. Sums of squares and sums of triangular numbers induced by partitions of 8. International Journal of Number Theory (IJNT), 4(4):525–538, August 2008. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304210800150X.

Bian:2020:RTF

[BCSX20] Min Bian, Shane Chern, Doris D. M. Sang, and Ernest X. W. Xia. Ramanujan’s theta function identities and the relations between sums of squares and sums of triangular numbers. International Journal of Number Theory (IJNT), 16(6):1275– 1294, July 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500669.

Bremner:2014:CDQ

[BCU14] Andrew Bremner, Ajai Choudhry, and Maciej Ulas. Constructions of diagonal quartic and sextic surfaces with infinitely many rational points. International Journal of Number Theory (IJNT), 10(7):1675–1698, November 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211450050X.

Brueggeman:2008:LCD

[BD08] Sharon Brueggeman and Darrin Doud. Local corrections of discriminant bounds and small degree extensions of quadratic base fields. International Journal of Number Theory (IJNT), 4(3):349– 361, June 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042108001389.

Balazard:2010:CBD

[BD10] Michel Balazard and Anne De Roton. Sur un critère de Báez–Duarte pour l’hypothèse de Riemann. (French) [On a criterion of Báez–Duarte for the Riemann Hypothesis]. International Journal of Number Theory (IJNT), 6(4):883– 903, June 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003307.

Baruah:2015:PRR

[BD15a] Nayandeep Deka Baruah and Kuwali Das. Parity results for 7-regular and 23regular partitions. International Journal of Number Theory (IJNT), 11(7):2221–2238, November 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500992.

Berndt:2015:P

[BD15b] Bruce C. Berndt and Harold G.

Diamond. Preface. International Journal of Number Theory (IJNT), 11(5): v–xiii, August 2015. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115020017.

Berger:2021:TTP

[BD21] Laurent Berger and Giovanni Di Matteo. On triangulable tensor products of B-pairs and trianguline representations. International Journal of Number Theory (IJNT), 17(10):2221–2233, November 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500846.

Baruah:2022:IFC

[BD22] Nayandeep Deka Baruah and Hirakjyoti Das. Infinite families of congruences modulo powers of 2 for some partition functions involving only odd parts. International Journal of Number Theory (IJNT), 18(08):1843–1862, September 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500944.

Bravo:2015:PPT

[BDGL15] Jhon J. Bravo, Pranabesh Das, Sergio Guzmán, and Shanta Laishram. Powers in products of terms of Pell’s and Pell–Lucas sequences. International Journal of Number Theory (IJNT), 11(4):1259– 1274, June 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500682.

Blackwell:2016:GMM

[BDTT16] Sarah Blackwell, Gabriel Durham, Katherine Thompson, and Tiffany Treece. A generalization of a method of Mordell to ternary quadratic forms. International Journal of Number Theory (IJNT), 12(8):2081– 2105, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042116501256.

Borwein:2013:NBP

[BE13] Peter Borwein and Tamás Erdélyi. A note on Barker polynomials. International Journal of Number Theory (IJNT), 9 (3):759–767, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S179304211250159X.

BenYakkou:2021:MCP

[BE21] Hamid Ben Yakkou and Lhoussain El Fadil. On monogenity of certain pure number fields defined by x^{p^{r}} − m. International Journal of Number Theory (IJNT), 17(10):2235–2242, November 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500858.

Booher:2010:ECT

[BEH10] Jeremy Booher, Anastassia Etropolski, and Amanda Hittson. Evaluations of cubic twisted Kloosterman sheaf sums. International Journal of Number Theory (IJNT), 6 (6):1349–1365, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042110003538.

Belliard:2009:ACC

[Bel09] Jean-Robert Belliard. Asymptotic cohomology of circular units. International Journal of Number Theory (IJNT), 5 (7):1205–1219, November 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S179304210900264X.

Bellah:2022:NFE

[Bel22]  Elisa Bellah.  Norm form

equations and linear divisibility sequences. International Journal of Number Theory (IJNT), 18(06):1239–1259, July 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500634.

Bennett:2010:DE

[BEN10] Michael A. Bennett, Jordan S. Ellenberg, and Nathan C. Ng. The Diophantine equation A^{4} + 2^{δ}B^{2} = C^{n}. International Journal of Number Theory (IJNT), 6(2):311– 338, March 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110002971.

Bennett:2013:IPC

[Ben13]  Michael A. Bennett.  In-

tegral points on congruent number curves. International Journal of Number Theory (IJNT), 9(6):1619–1640, September 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500474.

Benatar:2015:ESP

[Ben15] Jacques Benatar. The existence of small prime gaps in subsets of the integers. International Journal of Number Theory (IJNT), 11(3):801– 833, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211550044X.

Bouallegue:2010:KNS

[BEP10] Kais Bouallègue, Othman Echi, and Richard G. E. Pinch. Korselt numbers and sets. International Journal of Number Theory (IJNT), 6(2):257– 269, March 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110002922.

Berndt:2007:RCP

[Ber07]  Bruce C. Berndt.  Ra-

manujan’s congruences for the partition function modulo 5, 7, and 11. International Journal of Number Theory (IJNT), 3(3):349–354, September 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107000961.

Berkovich:2009:TPN

[Ber09] Alexander Berkovich. The tripentagonal number theorem and related identities. International Journal of Number Theory (IJNT), 5(8):1385–1399, December 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002705.

Berger:2011:AFD

[Ber11] Laurent Berger. A p-adic family of dihedral (ϕ, Γ)-modules. International Journal of Number Theory (IJNT), 7(7):1825– 1834, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042111004770.

Berger:2018:ORR

[Ber18]  Tobias Berger.  Oddness

of residually reducible Galois representations. International Journal of Number Theory (IJNT), 14(5):1329– 1345, June 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500835.

Bettin:2010:SMR

[Bet10]  Sandro Bettin.  The sec-

ond moment of the Riemann zeta function with unbounded shifts. International Journal of Number Theory (IJNT), 6(8):1933–1944, December 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003861.

Belabas:2010:DCP

[BF10] Karim Belabas and Étienne Fouvry. Discriminants cubiques et progressions arithmétiques. (French) [Cubic discriminants and arithmetic progressions].

International Journal of Number Theory (IJNT), 6(7):1491– 1529, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042110003605.

Bruin:2019:SMM

[BF19] Peter Bruin and Andrea Ferraguti. Strongly modular models of Q -curves. International Journal of Number Theory (IJNT), 15(3):505– 526, April 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211950026X.

Bayer-Fluckiger:2009:EMC

[BFCC09] Eva Bayer-Fluckiger, Jean-Paul Cerri, and Jérôme Chaubert. Euclidean minima and central division algebras. International Journal of Number Theory (IJNT), 5(7):1155–1168, November 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002614.

Baczkowski:2010:V

[BFLT10] Daniel Baczkowski, Michael Filaseta, Florian Luca, and Ognian Trifonov. On values of d(n!)/m!, φ(n!)/m! and σ(n!)/m!. International Journal of Number Theory (IJNT), 6(6):1199–1214, September 2010. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042110003435.

Bremner:2006:TRP

[BG06] Andrew Bremner and Richard K. Guy. Triangle-rectangle pairs with a common area and a common perimeter. International Journal of Number Theory (IJNT), 2(2):217– 223, June 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042106000504.

Baba:2011:OSM

[BG11] Srinath Baba and Håkan Granath. Orthogonal systems of modular forms and supersingular polynomials. International Journal of Number Theory (IJNT), 7(1):249–259, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004009.

Baba:2015:QMF

[BG15a] Srinath Baba and Håkan Granath. Quaternionic modular forms and exceptional sets of hypergeometric functions. International Journal of Number Theory (IJNT), 11(2):631– 643, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115500347.

Balasubramanian:2015:ARF

[BG15b] R. Balasubramanian and Sumit Giri. On additive representation functions. International Journal of Number Theory (IJNT), 11(4):1165– 1176, June 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500633.

Banks:2017:QNB

[BG17] William D. Banks and Victor Z.

Guo. Quadratic nonresidues below the Burgess bound. International Journal of Number Theory (IJNT), 13(3):751– 759, April 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500397.

Berardini:2022:WPA

[BGm22] Elena Berardini and Alejandro J. Giangreco-maidana. Weil polynomials of abelian varieties over finite fields with many rational points. International Journal of Number Theory (IJNT), 18(07):1591– 1603, August 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500804.

Birbrair:2015:RSR

[BGP15] Lev Birbrair, Marlon Gomes, and Wanderley Pereira. Resonance sequences and recoverability. International Journal of Number Theory (IJNT), 11 (2):495–506, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042115500256.

Boylan:2011:VFC

[BGW11] Matthew Boylan, Sharon Anne Garthwaite, and John Webb. On the vanishing of Fourier coefficients of certain genus zero newforms. International Journal of Number Theory (IJNT), 7(5):1229–1245, August 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004290.

Birmajer:2012:FPR

[BGW12] Daniel Birmajer, Juan B. Gil, and Michael Weiner. Factoring polynomials in the ring of formal power series over Z. International Journal of Number Theory (IJNT), 8(7):1763– 1776, November 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042112501011.

Beck:2010:FTC

[BH10] Matthias Beck and Mary Halloran. Finite trigonometric character sums via discrete Fourier analysis. International Journal of Number Theory (IJNT), 6(1):51–67, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042110002806.

Browning:2013:IKS

[BH13] T. D. Browning and A. Haynes.

Incomplete Kloosterman sums and multiplicative inverses in short intervals. International Journal of Number Theory (IJNT), 9(2):481–486, March 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501448.

Bayad:2014:MDR

[BH14] Abdelmejid Bayad and Yoshinori Hamahata. Multiple Dedekind–Rademacher sums in function fields. International Journal of Number Theory (IJNT), 10(5):1291–1307, August 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500274.

Bharadwaj:2020:SNG

[Bha20] Abhishek T. Bharadwaj. A short note on generalized Euler–Briggs constants. International Journal of Number Theory (IJNT), 16(4):823– 839, May 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500426.

Bertok:2017:LCP

[BHPR17] Csanád Bertók, Lajos Hajdu, István Pink, and Zsolt Rábai. Linear combinations of prime powers in binary recurrence sequences. International Journal of Number Theory (IJNT), 13(2):261–271, March 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500166.

Bugeaud:2011:MCC

[BHV11] Yann Bugeaud, Alan Haynes, and Sanju Velani. Metric considerations concerning the mixed Littlewood conjecture. International Journal of Number Theory (IJNT), 7(3):593– 609, May 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004289.

Bhatnagar:2021:EIC

[BI21] Gaurav Bhatnagar and Mourad E. H. Ismail. On Entry II.16.12: a continued fraction of Ramanujan. International Journal of Number Theory (IJNT), 17(02):251– 266, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400072.

Billerey:2011:CIP

[Bil11]  Nicolas Billerey.  Critères

d’irréductibilité pour les représentations des courbes elliptiques. (French) [Irreducibility criteria for the representations of elliptic curves].

International Journal of Number Theory (IJNT), 7(4):1001– 1032, June 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004538.

Birkbeck:2019:JLC

[Bir19] Christopher Birkbeck. The Jacquet–Langlands correspondence for overconvergent Hilbert modular forms. International Journal of Number Theory (IJNT), 15(3):479–504, April 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500258.

Biswas:2015:GCN

[Bis15] Saikat Biswas. Groups of components of Néron models of Jacobians and Brauer groups. International Journal of Number Theory (IJNT), 11(2):621– 629, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500335.

Bae:2019:SPT

[BJ19] Sunghan Bae and Hwanyup Jung. Statistics for products of traces of high powers of the Frobenius class of hyperelliptic curves in even characteristic. International Journal of Number Theory (IJNT), 15(7):1519– 1530, August 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042119500878.

Byrnes:2014:RRH

[BJMV14] Alyssa Byrnes, Lin Jiu, Victor H. Moll, and Christophe Vignat. Recursion rules for the hypergeometric zeta function. International Journal of Number Theory (IJNT), 10 (7):1761–1782, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042114500547.

Boulet:2011:SMD

[BK11] Cilanne Boulet and Kağan Kurşungöz. Symmetry of kmarked Durfee symbols. International Journal of Number Theory (IJNT), 7(1):215–230, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111003971.

Barman:2012:CVG

[BK12] Rupam Barman and Gautam Kalita. Certain values of Gaussian hypergeometric series and a family of algebraic curves. International Journal of Number Theory (IJNT), 8(4):945– 961, June 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211250056X.

Barman:2013:PGH

[BK13] Rupam Barman and Gautam Kalita. On the polynomial x^{d}+ax+b over F_{q} and Gaussian hypergeometric series. International Journal of Number Theory (IJNT), 9(7):1753–1763, November 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500553.

Berndt:2015:LMD

[BK15a] Bruce C. Berndt and Sun Kim.

Logarithmic means and double series of Bessel functions. International Journal of Number Theory (IJNT), 11(5):1535– 1556, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115400060.

Broderick:2015:DES

[BK15b] Ryan Broderick and Dmitry Kleinbock. Dimension estimates for sets of uniformly badly approximable systems of linear forms. International Journal of Number Theory (IJNT), 11(7):2037–2054, November 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500876.

Bayless:2016:CCE

[BK16] Jonathan Bayless and Paul Kinlaw. Consecutive coincidences of Euler’s function. International Journal of Number Theory (IJNT), 12(4):1011– 1026, June 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500639.

BenNasr:2022:AIW

[BK22] Mabrouk Ben Nasr and Hassen Kthiri. On algebraic integers which are 2-Salem elements in positive characteristic. International Journal of Number Theory (IJNT), 18(07):1637– 1650, August 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304212250083X.

Belaghi:2013:BMT

[BKB13] Mahmoud Jafari Shah Belaghi, Sergey Khrushchev, and Agamirza E. Bashirov. On Bauer–Muir transform of continued fractions. International Journal of Number Theory (IJNT), 9(2):321–332, March 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112501369.

Bringmann:2006:BBA

[BKK06] Kathrin Bringmann, Benjamin Kane, and Winfried Kohnen. On the boundary behavior of automorphic forms. International Journal of Number Theory (IJNT), 2(2):187– 194, June 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042106000565.

Bibak:2016:RLC

[BKS16] Khodakhast Bibak, Bruce M.

Kapron, and Venkatesh Srinivasan. On a restricted linear congruence. International Journal of Number Theory (IJNT), 12(8):2167–2171, December 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211650130X.

Berndt:2013:EPR

[BKW13] Bruce C. Berndt, Byungchan Kim, and Kenneth S. Williams. Euler products in Ramanujan’s Lost Notebook. International Journal of Number Theory (IJNT), 9(5):1313–1349, August 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500292.

Byeon:2019:CGZ

[BKY19] Dongho Byeon, Taekyung Kim, and Donggeon Yhee. A conjecture of Gross and Zagier: Case E(Q)_{tor} ≅ Z/3Z. International Journal of Number Theory (IJNT), 15(9):1793– 1800, October 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119501008.

Byeon:2020:CGZ

[BKY20] Dongho Byeon, Taekyung Kim, and Donggeon Yhee. A conjecture of Gross and Zagier: Case EQ_{tor} ≅ Z/2Z ⊕ Z/2Z, Z/2Z ⊕ Z/4Z or Z/2Z ⊕ Z/6Z. International Journal of Number Theory (IJNT), 16(7):1567– 1572, August 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500827.

Bringmann:2008:RCO

[BL08] Kathrin Bringmann and Jeremy Lovejoy. Rank and congruences for overpartition pairs. International Journal of Number Theory (IJNT), 4(2):303– 322, April 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001353.

Bandini:2009:CTE

[BL09] A. Bandini and I. Longhi. Control theorems for elliptic curves over function fields. International Journal of Number Theory (IJNT), 5(2):229– 256, March 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002067.

Bravo:2013:LPF

[BL13] Jhon J. Bravo and Florian Luca. On the largest prime factor of the k-Fibonacci numbers. International Journal of Number Theory (IJNT), 9(5):1351– 1366, August 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042113500309.

Beneish:2015:TSV

[BL15] Lea Beneish and Hannah Larson. Traces of singular values of Hauptmoduln. International Journal of Number Theory (IJNT), 11(3):1027– 1048, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500542.

Blache:2011:NPC

[Bla11] Régis Blache. Newton polygons for character sums and Poincaré series. International Journal of Number Theory (IJNT), 7(6):1519–1542, September 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004368.

Bugeaud:2005:PPL

[BLMS05] Yann Bugeaud, Florian Luca, Maurice Mignotte, and Samir Siksek. On perfect powers in Lucas sequences. International Journal of Number Theory (IJNT), 1(3):309– 332, September 2005. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000236.

Becheanu:2007:SCD

[BLS07] Mircea Becheanu, Florian Luca, and Igor E. Shparlinski. On the sums of complementary divisors. International Journal of Number Theory (IJNT), 3(4):635–648, December 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107001152.

Beresnevich:2021:LBH

[BLW21] Victor Beresnevich, Jason Levesley, and Ben Ward. A lower bound for the Hausdorff dimension of the set of weighted simultaneously approximable points over manifolds. International Journal of Number Theory (IJNT), 17(08):1795–1814, September 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500639.

Bugeaud:2010:PRS

[BM10] Yann Bugeaud and Maurice Mignotte. Polynomial root separation. International Journal of Number Theory (IJNT), 6(3):587–602, May 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003083.

Bui:2011:CVD

[BM11] H. M. Bui and Micah B. Milinovich. Central values of derivatives of Dirichlet L-functions. International Journal of Number Theory (IJNT), 7(2):371– 388, March 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004125.

Busch:2012:LHE

[BM12] Vincenz Busch and Jan Steffen Müller. Local heights on elliptic curves and intersection multiplicities. International Journal of Number Theory (IJNT), 8(6):1477–1484, September 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211250087X.

Bartolome:2017:E

[BM17] B. Bartolomé and P. Mihăilescu.

On equation X^{n} − 1 = BZ^{n}. International Journal of Number Theory (IJNT), 13(3): 549–570, April 2017. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500282.

Ballantine:2021:CPM

[BM21a] Cristina Ballantine and Mircea Merca. Combinatorial proof of the minimal excludant theorem. International Journal of Number Theory (IJNT), 17(08): 1765–1779, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042121500615.

Bouganis:2021:RSM

[BM21b] Thanasis Bouganis and Salvatore Mercuri. On the Rankin– Selberg method for vectorvalued Siegel modular forms. International Journal of Number Theory (IJNT), 17(05): 1207–1242, June 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042121500330.

Bowman:2006:CF

[BMW06] D. Bowman, J. McLaughlin, and N. J. Wyshinski. A qcontinued fraction. International Journal of Number Theory (IJNT), 2(4):523–547, December 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304210600070X.

Baruah:2014:IFA

[BN14] Nayandeep Deka Baruah and Kallol Nath. Infinite families of arithmetic identities for doubled distinct t-cores for t = 3, 4, …, 10. International Journal of Number Theory (IJNT), 10(1):85–113, February 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500826.

Baruah:2011:SCD

[BO11] Nayandeep Deka Baruah and Kanan Kumari Ojah. Some congruences deducible from Ramanujan’s cubic continued fraction. International Journal of Number Theory (IJNT), 7(5):1331–1343, August 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004745.

Budarina:2012:PNW

[BO12] Natalia Budarina and Hugh O’Donnell. On a problem of Nesterenko: When is the closest root of a polynomial a real number? International Journal of Number Theory (IJNT), 8(3): 801–811, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500455.

Burnett:2019:GID

[BO19] Joseph Vade Burnett and Otto Vaughn Osterman. A generalization of the infinitary divisibility relation: Algebraic and analytic properties. International Journal of Number Theory (IJNT), 15(9):1771– 1792, October 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500994.

Boca:2008:PES

[Boc08]  Florin P. Boca.  A prob-

lem of Erdős, Szösz and Turán concerning Diophantine approximations. International Journal of Number Theory (IJNT), 4(4):691–708, August 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042108001626.

Bonciocat:2008:CLP

[Bon08] Nicolae Ciprian Bonciocat. Congruences and Lehmer’s problem. International Journal of Number Theory (IJNT), 4 (4):587–596, August 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042108001547.

Borovoi:2015:HSH

[Bor15] Mikhail Borovoi. Homogeneous spaces of Hilbert type. International Journal of Number Theory (IJNT), 11(2):397– 405, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115500207.

Bordelles:2018:SIR

[Bor18] Olivier Bordellès. Short interval results for certain arithmetic functions. International Journal of Number Theory (IJNT), 14(2):535–548, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500331.

Bordelles:2020:ETL

[Bor20] Olivier Bordellès. On the error term of a lattice counting problem, II. International Journal of Number Theory (IJNT), 16(5):1153–1160, June 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304212050058X.

Bordelles:2022:RDS

[Bor22] Olivier Bordellès. On restricted divisor sums associated to the Chowla–Walum conjecture. International Journal of Number Theory (IJNT), 18(01):19–25, February 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500038.

Bosca:2009:PIA

[Bos09] Sébastien Bosca. Principalization of ideals in Abelian extensions of number fields. International Journal of Number Theory (IJNT), 5(3):527– 539, May 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002213.

Bourgain:2005:MSP

[Bou05] J. Bourgain. More on the sum–product phenomenon in prime fields and its applications. International Journal of Number Theory (IJNT), 1 (1):1–32, March 2005. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042105000108.

Bouganis:2011:NAC

[Bou11]  Thanasis Bouganis.  Non–

Abelian congruences between special values of L-functions of elliptic curves: the CM case. International Journal of Number Theory (IJNT), 7(7):1883– 1934, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S179304211100468X.

Boylan:2010:APC

[Boy10] Matthew Boylan. Arithmetic properties of certain level one Mock modular forms. International Journal of Number Theory (IJNT), 6(1):185–202, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110002855.

Barcau:2011:CCF

[BP11] Mugurel Barcau and Vicenţiu Paşol. mod p congruences for cusp forms of weight four for Γ_{0}(pN). International Journal of Number Theory (IJNT), 7(2):341–350, March 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211100406X.

Banescu:2015:AES

[BP15a] Magdalena Bănescu and Dumitru Popa. Asymptotic evaluations for some double sums in number theory. International Journal of Number Theory (IJNT), 11(7):2073–2085, November 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500906.

Bugeaud:2015:QA

[BP15b] Yann Bugeaud and Tomislav Pejković. Quadratic approximation in Q_{p}. International Journal of Number Theory (IJNT), 11(1):193–209, February 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500128.

Butzer:2017:FAC

[BP17] Paul L. Butzer and Tibor K.

Pogány. A fresh approach to classical Eisenstein series and the newer Hilbert–Eisenstein series. International Journal of Number Theory (IJNT), 13(4): 885–911, May 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500464.

Banescu:2018:MAS

[BP18] Magdalena Bănescu and Dumitru Popa. A multiple Abel summation formula and asymptotic evaluations for multiple sums. International Journal of Number Theory (IJNT), 14(4):1197– 1210, May 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500732.

Boca:2014:PCH

[BPZ14] Florin P. Boca, Alexandru A. Popa, and Alexandru Zaharescu. Pair correlation of hyperbolic lattice angles. International Journal of Number Theory (IJNT), 10(8):1955–1989, December 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042114500651.

Bringmann:2011:EFC

[BR11] Kathrin Bringmann and Olav K. Richter. Exact formulas for coefficients of Jacobi forms. International Journal of Number Theory (IJNT), 7(3):825– 833, May 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004617.

Brasca:2014:QMF

[Bra14] Riccardo Brasca. Quaternionic modular forms of any weight. International Journal of Number Theory (IJNT), 10(1):31– 53, February 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500796.

Bremner:2019:PER

[Bre19] Andrew Bremner. On a problem of Erdős related to common factor differences. International Journal of Number Theory (IJNT), 15(5):1059– 1068, June 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500581.

Bringmann:2009:CDR

[Bri09] Kathrin Bringmann. Congruences for Dyson’s ranks. International Journal of Number Theory (IJNT), 5(4):573– 584, June 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002262.

Brink:2011:RFD

[Bri11]  David Brink.  Rédei fields

and dyadic extensions. International Journal of Number Theory (IJNT), 7(1):231–240, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111003983.

Brown:2010:FNH

[Bro10a] Jim Brown. The first negative Hecke eigenvalue of genus 2 Siegel cuspforms with level n ≥ 1. International Journal of Number Theory (IJNT), 6(4): 857–867, June 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003277.

Brown:2010:SVF

[Bro10b] Jim Brown. Special values of L-functions on GSp_{4} × GL_{2} and the non-vanishing of Selmer groups. International Journal of Number Theory (IJNT), 6(8):1901–1926, December 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003769.

Broughan:2012:SPB

[Bro12] Kevin A. Broughan. On shifted primes and balanced primes. International Journal of Number Theory (IJNT), 8(8):2017– 2033, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S179304211250114X.

Brueggeman:2005:NCN

[Bru05]  Sharon Brueggeman.  The

nonexistence of certain nonsolvable Galois extensions of number fields of small degree. International Journal of Number Theory (IJNT), 1(1):155– 160, March 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042105000121.

Brunault:2018:NCE

[Bru18] François Brunault. Non-critical equivariant L-values of modular abelian varieties. International Journal of Number Theory (IJNT), 14(9):2517–2542, October 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501506.

Barman:2010:IIT

[BS10a] Rupam Barman and Anupam Saikia. Iwasawa λ-invariants and Γ-transforms of p-adic measures on Z_{p}^{n}. International Journal of Number Theory (IJNT), 6(8):1819–1829, December 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003824.

Bremner:2010:CST

[BS10b] Andrew Bremner and Blair K.

Spearman. Cyclic sextic trinomials x^{6} + Ax + B. International Journal of Number Theory (IJNT), 6(1):161–167, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110002843.

Baruah:2012:ISC

[BS12] Nayandeep Deka Baruah and Bipul Kumar Sarmah. Identities for self-conjugate 7and 9-core partitions. International Journal of Number Theory (IJNT), 8(3):653– 667, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500376.

Barman:2015:CTH

[BS15] Rupam Barman and Neelam Saikia. Certain transformations for hypergeometric series in the p-adic setting. International Journal of Number Theory (IJNT), 11(2):645– 660, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115500359.

Banks:2016:FPD

[BS16a] William D. Banks and Igor E. Shparlinski. Fractional parts of Dedekind sums. International Journal of Number Theory (IJNT), 12(5):1137– 1147, August 2016. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211650069X.

Bremner:2016:SAP

[BS16b] Andrew Bremner and Samir Siksek. Squares in arithmetic progression over cubic fields. International Journal of Number Theory (IJNT), 12(5):1409– 1414, August 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211650086X.

Burnette:2017:PIR

[BS17] Charles Burnette and Eric Schmutz. Periods of iterated rational functions. International Journal of Number Theory (IJNT), 13(5):1301– 1315, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500713.

Browning:2019:AAF

[BS19] T. D. Browning and E. Sofos. Averages of arithmetic functions over principal ideals. International Journal of Number Theory (IJNT), 15(3):547– 567, April 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500283.

Baier:2020:LSI

[BS20] Stephan Baier and Rajneesh Kumar Singh. The large sieve inequality with square moduli for quadratic extensions of function fields. International Journal of Number Theory (IJNT), 16 (09):1907–1922, October 2020. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500979.

Bary-Soroker:2017:BPC

[BSK17] Lior Bary-Soroker and Gady Kozma. Is a bivariate polynomial with ±1 coefficients irreducible? Very likely! International Journal of Number Theory (IJNT), 13(4):933– 936, May 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042117500488.

Bayad:2016:FFB

[BSM16] A. Bayad, D. D. Somashekara, and K. Narasimha Murthy. On finite forms of Bailey’s _{2}ψ_{2} transformation formulae and their applications. International Journal of Number Theory (IJNT), 12(8):2189–2200, December 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116501323.

Bostan:2010:GFC

[BST10] Alin Bostan, Bruno Salvy, and Khang Tran. Generating functions of Chebyshevlike polynomials. International Journal of Number Theory (IJNT), 6(7):1659–1667, November 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003691.

Bachmann:2018:RTM

[BT18a] Henrik Bachmann and Tatsushi Tanaka. Rooted tree maps and the derivation relation for multiple zeta values. International Journal of Number Theory (IJNT), 14(10):2657–2662, November 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118501592.

Booker:2018:RCF

[BT18b] Andrew R. Booker and Holger Then. Rapid computation of L-functions attached to Maass forms. International Journal of Number Theory (IJNT), 14(5):1459–1485, June 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500896.

Bennett:2006:GTB

[BTW06] Michael A. Bennett, Alain Togbé, and P. G. Walsh. A generalization of a theorem of Bumby on quartic Diophantine equations. International Journal of Number Theory (IJNT), 2(2):195– 206, June 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000474.

Bremner:2011:X

[BU11a] Andrew Bremner and Maciej Ulas. On x^{a} ± y^{b} ± z^{c} ± w^{d} = 0, 1/a + 1/b + 1/c + 1/d = 1. International Journal of Number Theory (IJNT), 7(8):2081–2090, December 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111005076.

Bremner:2011:RPS

[BU11b] Andrew Bremner and Maciej Ulas. Rational points on some hyper- and superelliptic curves. International Journal of Number Theory (IJNT), 7(1):115– 132, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111003946.

Buckingham:2011:FGI

[Buc11] Paul Buckingham. The fractional Galois ideal for arbitrary order of vanishing. International Journal of Number Theory (IJNT), 7(1):87–99, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004010.

Budarina:2020:EEM

[Bud20] Natalia Budarina. An effective estimate for the measure of the set of p-adic numbers with a given order of approximation. International Journal of Number Theory (IJNT), 16(3):651– 672, April 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500335.

Bui:2012:NVD

[Bui12] H. M. Bui. Non-vanishing of Dirichlet L-functions at the central point. International Journal of Number Theory (IJNT), 8(8):1855–1881, December 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042112501060.

Bui:2013:TAF

[Bui13]  H. M. Bui.  Twists of

automorphic L-functions at the central point. International Journal of Number Theory (IJNT), 9(4):1015–1053, June 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500115.

Bundschuh:2008:ARC

[Bun08] Peter Bundschuh. Arithmetical results on certain q-series, I. International Journal of Number Theory (IJNT), 4(1):25–43, February 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001201.

Bundschuh:2012:TAI

[Bun12]  Peter Bundschuh.  Tran-

scendence and algebraic independence of series related to Stern’s sequence. International Journal of Number Theory (IJNT), 8(2):361–376, March 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500212.

Burson:2021:CTS

[Bur21]  Hannah Burson.  Combi-

natorics of two second-order mock theta functions. International Journal of Number Theory (IJNT), 17(02):285– 295, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400229.

Bundschuh:2009:ARC

[BV09] Peter Bundschuh and Keijo Väänänen. Arithmetical results on certain q-series, II. International Journal of Number Theory (IJNT), 5(7):1231–1245, November 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002663.

Bundschuh:2011:AND

[BV11] Peter Bundschuh and Keijo Väänänen. An application of Nesterenko’s dimension estimate to p-adic q-series. International Journal of Number Theory (IJNT), 7(2):431– 447, March 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004071.

Bandini:2018:AOI

[BV18] Andrea Bandini and Maria Valentino. On the Atkin U_{t}-operator for Γ_{1}(t)-invariant Drinfeld cusp forms. International Journal of Number Theory (IJNT), 14(10):2599–2616, November 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042118501555.

Banerjee:2017:AEL

[BW17] Shubho Banerjee and Blake Wilkerson. Asymptotic expansions of Lambert series and related q-series. International Journal of Number Theory (IJNT), 13(8):2097–2113, September 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117501135.

Bowman:2021:GFT

[BW21] Douglas Bowman and Sarah Wesley. General Fine transformations I. International Journal of Number Theory (IJNT), 17(02):267–283, March 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042120400163.

Bremner:2018:E

[BX18] Andrew Bremner and Tho Nguyen Xuan. The equation (w+x+y+ z)(1/w + 1/x + 1/y + 1/z) = n. International Journal of Number Theory (IJNT), 14(5):1229– 1246, June 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500768.

Byard:2006:QRD

[Bya06] Kevin Byard. On qualified residue difference sets. International Journal of Number Theory (IJNT), 2(4):591–597, December 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000747.

Byard:2009:TPQ

[Bya09] Kevin Byard. Tenth power qualified residue difference sets. International Journal of Number Theory (IJNT), 5(5):797– 803, August 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002389.

Baier:2005:LSI

[BZ05] Stephan Baier and Liangyi Zhao. Large sieve inequality with characters for powerful moduli. International Journal of Number Theory (IJNT), 1 (2):265–279, June 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042105000170.

Baier:2009:PQP

[BZ09] Stephan Baier and Liangyi Zhao. On primes in quadratic progressions. International Journal of Number Theory (IJNT), 5(6):1017–1035, September 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002523.

Bouw:2013:ECF

[BZ13] Irene I. Bouw and Leonardo Zapponi. Existence of covers with fixed ramification in positive characteristic. International Journal of Number Theory (IJNT), 9(6):1475–1489, September 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500383.

Backhausz:2015:AFE

[BZ15] Tibor Backhausz and Gergely Zábrádi. Algebraic functional equations and completely faithful Selmer groups. International Journal of Number Theory (IJNT), 11(4):1233– 1257, June 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500670.

Burek:2019:NUB

[BŻ19] Dominik Burek and Błażej Żmija. A new upper bound for numbers with the Lehmer property and its application to repunit numbers. International Journal of Number Theory (IJNT), 15(7):1463–1468, August 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500830.

Bzdga:2017:PCP

[Bzd17]  Bartłomiej Bzdega.  Prod-

ucts of cyclotomic polynomials on unit circle. International Journal of Number Theory (IJNT), 13(10):2515–2530, November 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117501408.

Cai:2010:LFS

[Cai10]  Yingchun Cai.  Lagrange’s

four squares theorem with variables of special type. International Journal of Number Theory (IJNT), 6(8):1801–1817, December 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003812.

Cai:2018:TDI

[Cai18]  Yingchun Cai.  A ternary

Diophantine inequality involving primes. International Journal of Number Theory (IJNT), 14(8):2257–2268, September 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501361.

Cao:2011:SDR

[Cao11] Wei Cao. A special degree reduction of polynomials over finite fields with applications. International Journal of Number Theory (IJNT), 7(4):1093– 1102, June 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004277.

Carls:2011:GTC

[Car11a] Robert Carls. Galois theory of the canonical theta structure. International Journal of Number Theory (IJNT), 7(1):173– 202, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111003934.

Carter:2011:BGP

[Car11b]  Andrea C. Carter.  The

Brauer group of Del Pezzo surfaces. International Journal of Number Theory (IJNT), 7(2): 261–287, March 2011. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111003910.

Castillo:2012:HOS

[Cas12] Daniel Macias Castillo. On higher-order Stickelberger-type theorems for multi-quadratic extensions. International Journal of Number Theory (IJNT), 8(1):95–110, February 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500054.

Cauchi:2020:NCS

[Cau20]  Antonio Cauchi.  Norm-

compatible systems of cohomology classes for GU(2, 2). International Journal of Number Theory (IJNT), 16(3):461– 510, April 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500244.

Choudhry:2022:DEI

[CBJ22] Ajai Choudhry, Iliya Bluskov, and Alexander James. A diophantine equation inspired by Brahmagupta’s identity. International Journal of Number Theory (IJNT), 18(04):905– 911, May 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500476.

Castano-bernard:2021:PPL

[CbL21] Carlos Castaño-bernard and Florian Luca. On Petersson’s partition limit formula. International Journal of Number Theory (IJNT), 17(06):1365– 1378, July 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500408.

Choie:2007:RBF

[CC07] Y. Choie and Y. Chung. Representations of binary forms by quaternary forms. International Journal of Number Theory (IJNT), 3(3):471–474, September 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042107000997.

Chan:2016:ATF

[CC16] Heng Huat Chan and Teoh Guan Chua. An alternative transformation formula for the Dedekind η-function via the Chinese Remainder Theorem. International Journal of Number Theory (IJNT), 12(2):513– 526, March 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500329.

Cao:2017:GSP

[CC17] Zhu Cao and Shi-Chao Chen.

On generalized Schur’s partitions. International Journal of Number Theory (IJNT), 13(6): 1381–1391, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042117500774.

Chen:2021:AGT

[CC21a] Xiaojing Chen and Wenchang Chu. q-analogues of Guillera’s two series for π^{±2} with convergence rate 27/64. International Journal of Number Theory (IJNT), 17(01):71–90, February 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500056.

Chan:2021:CAC

[cC21b] Hei chi Chan. Chasing after cancellations: Revisiting a classic identity that implies the Rogers–Ramanujan identities.

International Journal of Number Theory (IJNT), 17(02):297– 310, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400266.

Cha:2019:IPE

[CCH^{+}19] Byungchul Cha, Adam Claman, Joshua Harrington, Ziyu Liu, Barbara Maldonado, Alexander Miller, Ann Palma, Tony W. H. Wong, and Hongkwon Yi. An investigation on partitions with equal products. International Journal of Number Theory (IJNT), 15(8):1731–1744, September 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500969.

Cerri:2013:ETD

[CCL13] Jean-Paul Cerri, Jérôme Chaubert, and Pierre Lezowski. Euclidean totally definite quaternion fields over the rational field and over quadratic number fields. International Journal of Number Theory (IJNT), 9(3):653– 673, May 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501540.

Cullinan:2014:SGR

[CCRT14] John Cullinan, Celeste Cass, Alexander Rasmussen, and Darko Trifunovski. Specializations of generalized Rikuna polynomials. International Journal of Number Theory (IJNT), 10(3):585–600, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL

https://www.worldscientific.
com/doi/10.1142/S1793042113501108.

Chan:2010:CSA

[CCS10] Heng Huat Chan, Shaun Cooper, and Francesco Sica. Congruences satisfied by Apérylike numbers. International Journal of Number Theory (IJNT), 6(1):89–97, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110002879.

Clark:2013:TPE

[CCS13] Pete L. Clark, Brian Cook, and James Stankewicz. Torsion points on elliptic curves with complex multiplication (with an appendix by Alex Rice). International Journal of Number Theory (IJNT), 9(2):447– 479, March 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501436.

Cai:2015:PNF

[CCZ15] Tianxin Cai, Deyi Chen, and Yong Zhang. Perfect numbers and Fibonacci primes (I). International Journal of Number Theory (IJNT), 11(1):159–169, February 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042115500098.

Cosgrave:2011:MOC

[CD11] John B. Cosgrave and Karl Dilcher. The multiplicative orders of certain Gauss factorials. International Journal of Number Theory (IJNT), 7(1):145– 171, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211100396X.

Corwin:2020:PQG

[CDc20] David Corwin and Ishai Dancohen. The polylog quotient and the goncharov quotient in computational Chabauty– Kim Theory I. International Journal of Number Theory (IJNT), 16(08):1859–1905, September 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500967.

Chen:2015:DSI

[CDHS15] Yiwang Chen, Nicholas Dunn, Campbell Hewett, and Shashwat Silas. Dedekind sums s(a, b) and inversions modulo b. International Journal of Number Theory (IJNT), 11(8):2325– 2339, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042115501067.

Choi:2007:CSP

[CE07] H. Timothy Choi and Ronald Evans. Congruences for sums of powers of Kloosterman sums. International Journal of Number Theory (IJNT), 3(1):105– 117, March 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107000821.

Chan:2007:FRR

[CEIK07] Wai Kiu Chan, A. G. Earnest, Maria Ines Icaza, and Ji Young Kim. Finiteness results for regular definite ternary quadratic forms over. International Journal of Number Theory (IJNT), 3(4):541–556, December 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107001103.

Chems-eddin:2021:CGC

[CeM21] Mohamed Mahmoud Chemseddin and Katharina Müller. 2class groups of cyclotomic towers of imaginary biquadratic fields and applications. International Journal of Number Theory (IJNT), 17(08):1781–1793, September 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500627.

Centeleghe:2016:ITM

[Cen16] Tommaso Giorgio Centeleghe.

Integral Tate modules and splitting of primes in torsion fields of elliptic curves. International Journal of Number Theory (IJNT), 12(1):237–248, February 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500147.

Cooper:2006:RBP

[CEO06] Joshua N. Cooper, Dennis Eichhorn, and Kevin O’Bryant. Reciprocals of binary power series. International Journal of Number Theory (IJNT), 2(4):499–522, December 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000693.

Carney:2012:PEF

[CEP12] Alexander Carney, Anastassia Etropolski, and Sarah Pitman. Powers of the etafunction and Hecke operators. International Journal of Number Theory (IJNT), 8(3):599– 611, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500339.

Cesnavicius:2016:LFV

[Čes16] Kestutis Česnavičius. Local factors valued in normal domains. International Journal of Number Theory (IJNT), 12(1):249– 272, February 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500159.

Chan:2009:COD

[CF09] Ping-Shun Chan and Yuval Z.

Flicker. Cyclic odd degree base change lifting for unitary groups in three variables. International Journal of Number Theory (IJNT), 5(7):1247–1309, November 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002687.

Candelori:2017:VVM

[CF17] Luca Candelori and Cameron Franc. Vector-valued modular forms and the modular orbifold of elliptic curves. International Journal of Number Theory (IJNT), 13(1):39– 63, February 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211750004X.

Checcoli:2022:CTP

[CF22] S. Checcoli and A. Fehm. Constructing totally p-adic numbers of small height. International Journal of Number Theory (IJNT), 18(03):501– 514, April 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500294.

Cipu:2018:NED

[CFM18] Mihai Cipu, Yasutsugu Fujita, and Takafumi Miyazaki. On the number of extensions of a Diophantine triple. International Journal of Number Theory (IJNT), 14(3):899– 917, April 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500549.

Cao:2014:DBT

[CFTZ14] Xiaodong Cao, Jun Furuya, Yoshio Tanigawa, and Wenguang Zhai. On the differences between two kinds of mean value formulas of numbertheoretic error terms. International Journal of Number Theory (IJNT), 10(5):1143–1170, August 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042114500195.

Cui:2013:CDB

[CG13] Su-Ping Cui and Nancy Shan Shan Gu. Congruences for k dots bracelet partition functions. International Journal of Number Theory (IJNT), 9(8):1885– 1894, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042113500644.

Chow:2015:DN

[CG15] Sam Chow and Alexandru Ghitza. Distinguishing newforms. International Journal of Number Theory (IJNT), 11(3): 893–908, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500499.

Chen:2018:DSI

[CG18] Qi-Fei Chen and Victor J. W.

Guo. On the divisibility of sums involving powers of multivariable Schmidt polynomials. International Journal of Number Theory (IJNT), 14(2):365– 370, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500240.

Cheng:2019:NCN

[CG19a] Weidong Cheng and Xuejun Guo. The non-congruent numbers via Monsky’s formula. International Journal of Number Theory (IJNT), 15(4):677– 711, May 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500362.

Cui:2019:CMP

[CG19b] Su-Ping Cui and Nancy S. S.

Gu. Congruences modulo powers of 2 for generalized Frobenius partitions with six colors. International Journal of Number Theory (IJNT), 15(6):1173– 1181, July 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042119500647.

Chaubey:2022:MVE

[CG22] Sneha Chaubey and Shivani Goel. Mean value estimates of gcd and lcm-sums. International Journal of Number Theory (IJNT), 18(03):629– 654, April 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500348.

Cegielski:2015:NRF

[CGG15] Patrick Cégielski, Serge Grigorieff, and Irène Guessarian. Newton representation of functions over natural integers having integral difference ratios. International Journal of Number Theory (IJNT), 11(7):2109–2139, November 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211550092X.

Cui:2018:CSP

[CGH18] Su-Ping Cui, Nancy S. S. Gu, and Li-Jun Hao. Congruences for some partitions related to mock theta functions. International Journal of Number Theory (IJNT), 14(4):1055– 1071, May 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500641.

Cui:2015:CPO

[CGM15] Su-Ping Cui, Wen Xiang Gu, and Zhen Sheng Ma. Congruences for partitions with odd parts distinct modulo 5. International Journal of Number Theory (IJNT), 11(7):2151– 2159, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042115500943.

Chua:2015:PCG

[CGPY15] Lynn Chua, Benjamin Gunby, Soohyun Park, and Allen Yuan. Proof of a conjecture of Guy on class numbers. International Journal of Number Theory (IJNT), 11(4):1345– 1355, June 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500724.

Cantarini:2021:CAG

[CGZ21] Marco Cantarini, Alessandro Gambini, and Alessandro Zaccagnini. Cesàro averages for Goldbach representations with summands in arithmetic progressions. International Journal of Number Theory (IJNT), 17(10):2379–2393, November 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500937.

Chen:2005:SEG

[CH05] Sin-Da Chen and Sen-Shan Huang. On the series expansion of the Göllnitz–Gordon continued fraction. International Journal of Number Theory (IJNT), 1(1):53–63, March 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000030.

Chen:2009:GSP

[CH09] Sin-Da Chen and Sen-Shan Huang. On general series– product identities. International Journal of Number Theory (IJNT), 5(6):1129–1148, September 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002596.

Chou:2015:EFT

[CH15] Wun-Seng Chou and Leetsch C. Hsu. An extension of Flecktype Möbius function and inversion. International Journal of Number Theory (IJNT), 11 (6):1807–1819, September 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042115500785.

Chen:2016:PMM

[CH16] Yong-Gao Chen and Da-Wei He. On a problem of Mircea Merca. International Journal of Number Theory (IJNT), 12(7):2017–2024, November 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042116501220.

Chan:2005:HMP

[Cha05] Tsz Ho Chan. Higher moments of primes in short intervals II. International Journal of Number Theory (IJNT), 1(2):207– 214, June 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000169.

Chan:2006:NPS

[Cha06] Tsz Ho Chan. A note on primes in short intervals. International Journal of Number Theory (IJNT), 2(1):105– 110, March 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000437.

Chapman:2008:RIF

[Cha08] Robin Chapman. Representations of integers by the form x^{2} + xy + y^{2} + z^{2} + zt + t^{2}. International Journal of Number Theory (IJNT), 4(5): 709–714, October 2008. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001638.

Chan:2010:RCCa

[Cha10a]  Hei-Chi Chan.  Ramanu-

jan’s cubic continued fraction and an analog of his “Most Beautiful Identity”. International Journal of Number Theory (IJNT), 6(3):673– 680, May 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003150.

Chan:2010:RCCb

[Cha10b] Hei-Chi Chan. Ramanujan’s cubic continued fraction and Ramanujan type congruences for a certain partition function. International Journal of Number Theory (IJNT), 6(4):819– 834, June 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003241.

Chatterjee:2012:SCM

[Cha12] Tapas Chatterjee. The strong Chowla–Milnor spaces and a conjecture of Gun, Murty and Rath. International Journal of Number Theory (IJNT), 8(5):1301–1314, August 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500765.

Chan:2013:TCN

[Cha13] Tsz Ho Chan. Twin cubefull numbers. International Journal of Number Theory (IJNT), 9(1):17–26, February 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501229.

Chan:2015:FAS

[Cha15] Tsz Ho Chan. Factors of almost squares and lattice points on circles. International Journal of Number Theory (IJNT), 11(5):1701–1708, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115400205.

Chan:2018:CFA

[Cha18a]  Tsz Ho Chan.  Common

factors among pairs of consecutive integers. International Journal of Number Theory (IJNT), 14(3):871–880, April 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500525.

Chan:2018:FAS

[Cha18b] Tsz Ho Chan. Finding almost squares VI. International Journal of Number Theory (IJNT), 14(5):1347– 1356, June 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500847.

Chalebgwa:2020:NTA

[Cha20] Taboka Prince Chalebgwa.

Nevanlinna theory and algebraic values of certain meromorphic functions. International Journal of Number Theory (IJNT), 16(7):1607–1636, August 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500852.

Chan:2021:CEU

[Cha21] Tsz Ho Chan. The congruence equation ā + b̄ ≡ c̄ mod p. International Journal of Number Theory (IJNT), 17(02):329– 340, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400011.

Chen:2014:APP

[Che14]  Shi-Chao Chen.  Arith-

metic properties of a partition pair function. International Journal of Number Theory (IJNT), 10(6):1583–1594, September 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500468.

Chen:2015:DTG

[Che15]  W. W. L. Chen.  Daven-

port’s theorem in geometric discrepancy theory. International Journal of Number Theory (IJNT), 11(5):1437–1449, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115400047.

Chen:2017:GHN

[Che17]  Kwang-Wu Chen.  Gen-

eralized harmonic numbers and Euler sums. International Journal of Number Theory (IJNT), 13(2):513–528, March 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500883.

Chen:2018:DNT

[Che18a] Bin Chen. On the dual nature theory of bilateral series associated to mock theta functions. International Journal of Number Theory (IJNT), 14(1):63– 94, February 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500069.

Chern:2018:CGB

[Che18b] Shane Chern. On a conjecture of George Beck. International Journal of Number Theory (IJNT), 14(3):647– 651, April 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500392.

Cherubini:2018:APF

[Che18c] Giacomo Cherubini. Almost periodic functions and hyperbolic counting. International Journal of Number Theory (IJNT), 14(9):2343–2368, October 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501439.

Chen:2019:VCH

[Che19]  Shi-Chao Chen.  Vanish-

ing of the coefficients of a half Lerch sum. International Journal of Number Theory (IJNT), 15(2):251–263, March 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042119500106.

Chen:2022:EUB

[Che22a]  Bo Chen.  Explicit upper

bound on the least primitive root modulo p^{2}. International Journal of Number Theory (IJNT), 18(02):389– 395, March 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500233.

Chern:2022:CIH

[Che22b] Shane Chern. On a congruence involving harmonic series and Bernoulli numbers. International Journal of Number Theory (IJNT), 18(08):1691–1712, September 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500865.

Chern:2022:WPR

[Che22c] Shane Chern. Weighted partition rank and crank moments. III. A list of Andrews– Beck type congruences modulo 5, 7, 11 and 13. International Journal of Number Theory (IJNT), 18(01):141– 163, February 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500117.

Chida:2009:IOS

[Chi09] Masataka Chida. Indivisibility of orders of Selmer groups for modular forms. International Journal of Number Theory (IJNT), 5(2):271–280, March 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002080.

Chiriac:2019:ECR

[Chi19] Liubomir Chiriac. On the equality case of the Ramanujan conjecture for Hilbert modular forms. International Journal of Number Theory (IJNT), 15(10):2107–2114, November 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211950115X.

Chan:2021:MTF

[CHJL21] Song Heng Chan, Nankun Hong, Jerry, and Jeremy Lovejoy. A mock theta function identity related to the partition rank modulo 3 and 9. International Journal of Number Theory (IJNT), 17(02):311– 327, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042120400254.

Cai:2019:LBL

[CHL19] Zhaodong Cai, A. J. Hildebrand, and Junxian Li. A local Benford Law for a class of arithmetic sequences. International Journal of Number Theory (IJNT), 15(3):613– 638, April 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042119500325.

Cho:2013:SZM

[Cho13a] Peter Jaehyun Cho. Simple zeros of Maass L-functions. International Journal of Number Theory (IJNT), 9(1):167–178, February 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211250131X.

Chojecki:2013:DCP

[Cho13b] Przemyslaw Chojecki. Density of crystalline points on unitary Shimura varieties. International Journal of Number Theory (IJNT), 9(3):729– 744, May 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112501576.

Choudhry:2015:SIW

[Cho15]  Ajai Choudhry.  Sextuples

of integers whose sums in pairs are squares. International Journal of Number Theory (IJNT), 11(2):543–555, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500281.

Choudhry:2017:NAT

[Cho17] Ajai Choudhry. A new approach to the Tarry–Escott problem. International Journal of Number Theory (IJNT), 13 (2):393–417, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042117500233.

Choe:2018:GCN

[Cho18a] Chunghyok Choe. Generalized cyclotomic numbers with respect to prime powers. International Journal of Number Theory (IJNT), 14(7):2083– 2093, August 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501257.

Choudhry:2018:NMS

[Cho18b]  Ajai Choudhry.  A new

method of solving certain quartic and higher degree diophantine equations. International Journal of Number Theory (IJNT), 14(8):2129–2154, September 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501282.

Cho:2020:CSD

[Cho20a]  Bumkyu Cho.  Convolu-

tion sums of a divisor function for prime levels. International Journal of Number Theory (IJNT), 16(3):537– 546, April 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304212050027X.

Choudhry:2020:BQE

[Cho20b] Ajai Choudhry. Brahmagupta quadrilaterals with equal perimeters and equal areas. International Journal of Number Theory (IJNT), 16(3):523– 535, April 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500268.

Choudhry:2020:IPR

[Cho20c] Ajai Choudhry. An improvement of Prouhet’s 1851 result on multigrade chains. International Journal of Number Theory (IJNT), 16(7):1425– 1432, August 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304212050075X.

Chen:2015:APA

[CHS15] Shi-Chao Chen, Michael D. Hirschhorn, and James A. Sellers. Arithmetic properties of Andrews’ singular overpartitions. International Journal of Number Theory (IJNT), 11(5):1463–1476, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115400011.

Chu:2013:DIS

[Chu13] Wenchang Chu. Derivative inverse series relations and Lagrange expansion formula. International Journal of Number Theory (IJNT), 9(4):1001– 1013, June 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500103.

Cilleruelo:2016:PSR

[Cil16] Javier Cilleruelo. On product sets of rationals. International Journal of Number Theory (IJNT), 12(5):1415–1420, August 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500871.

Ciolan:2020:AIP

[Cio20] Alexandru Ciolan. Asymptotics and inequalities for partitions into squares. International Journal of Number Theory (IJNT), 16(1):121–143, February 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500074.

Chakraborty:2018:AEL

[CJKM18] Kalyan Chakraborty, Abhishek Juyal, Shiv Datt Kumar, and Bibekananda Maji. An asymptotic expansion of a Lambert series associated to cusp forms. International Journal of Number Theory (IJNT), 14(1):289– 299, February 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500173.

Choie:2013:FCM

[CK13] Young Ju Choie and Winfried Kohnen. On the Fourier coefficients of modular forms of half-integral weight. International Journal of Number Theory (IJNT), 9(8):1879– 1883, December 2013. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500632.

Can:2014:CAC

[CK14] Mümün Can and Veli Kurt. Character analogues of certain Hardy–Berndt sums. International Journal of Number Theory (IJNT), 10(3):737– 762, May 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113501200.

Cho:2017:CLT

[CK17] Peter J. Cho and Henry H. Kim. Central limit theorem for Artin L-functions. International Journal of Number Theory (IJNT), 13(1):1–14, February 2017. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500014.

Cohen:2020:TLP

[CK20] Stephen D. Cohen and Giorgos Kapetanakis. The translate and line properties for 2primitive elements in quadratic extensions. International Journal of Number Theory (IJNT), 16(09):2027–2040, October 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120501055.

Cigler:2021:HDL

[CK21] J. Cigler and C. Krattenthaler.

Hankel determinants of linear combinations of moments of orthogonal polynomials. International Journal of Number Theory (IJNT), 17(02):341– 369, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400321.

Choi:2005:STS

[CKO05] S. K. K. Choi, A. V. Kumchev, and R. Osburn. On sums of three squares. International Journal of Number Theory (IJNT), 1(2):161–173, June 2005. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042105000054.

Chen:2013:MGR

[CKW13] Imin Chen, Ian Kiming, and Gabor Wiese. On modular Galois representations modulo prime powers. International Journal of Number Theory (IJNT), 9(1):91–113, February 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112501254.

Chambert-Loir:2011:TJS

[CL11a] Antoine Chambert-Loir. The theorem of Jentzsch–Szegő on an analytic curve: Application to the irreducibility of truncations of power series. International Journal of Number Theory (IJNT), 7(7):1807–1823, November 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004691.

Chen:2011:CP

[CL11b] Yong-Gao Chen and Jing-Rui Lou. The congruent properties for r_{s}(n). International Journal of Number Theory (IJNT), 7(6):1595–1602, September 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004885.

See erratum [CL13].

Chen:2013:ECP

[CL13] Yong-Gao Chen and Jing-Rui Lou. Erratum: “The congruent properties for r_{s}(n)”. International Journal of Number Theory (IJNT), 9(3):811, May 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113920012.

See [CL11b].

Cerri:2019:CEM

[CL19] Jean-Paul Cerri and Pierre Lezowski. Computation of Euclidean minima in totally definite quaternion fields. International Journal of Number Theory (IJNT), 15(1):43– 66, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501725.

Chen:2022:NAB

[CL22] Guohua Chen and Weiping Li.

A note on arithmetic behavior of Hecke eigenvalues of Siegel cusp forms of degree two. International Journal of Number Theory (IJNT), 18(01):9– 18, February 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500026.

Clark:2008:AHP

[Cla08] Pete L. Clark. An “anti-Hasse principle” for prime twists. International Journal of Number Theory (IJNT), 4(4):627– 637, August 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001572.

Chakraborty:2008:ECG

[CLM08] Kalyan Chakraborty, Florian Luca, and Anirban Mukhopadhyay. Exponents of class groups of real quadratic fields. International Journal of Number Theory (IJNT), 4(4):597– 611, August 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001559.

Chaumont:2009:CSL

[CLMR09] Alain Chaumont, Johannes Leicht, Tom Müller, and Andreas Reinhart. The continuing search for large elite primes. International Journal of Number Theory (IJNT), 5(2):209– 218, March 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002031.

Chan:2005:EFQ

[CLN05] Heng Huat Chan, Zhi-Guo Liu, and Say Tiong Ng. Elliptic functions and the quintuple, Hirschhorn and Winquist product identities. International Journal of Number Theory (IJNT), 1(1):33– 43, March 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042105000017.

Chan:2012:TCA

[CM12] Song Heng Chan and Renrong Mao. Two congruences for Appell–Lerch sums. International Journal of Number Theory (IJNT), 8(1):111–123, February 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500066.

Chan:2016:RCP

[CM16] Song Heng Chan and Renrong Mao. The rank and crank of partitions modulo 3. International Journal of Number Theory (IJNT), 12(4):1027– 1053, June 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500640.

Chang:2020:SIF

[CM20] Ben Chang and Greg Martin.

The smallest invariant factor of the multiplicative group. International Journal of Number Theory (IJNT), 16(6):1377– 1405, July 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500736.

Cao:2020:CAN

[CMS20] Hui-Qin Cao, Yuri Matiyasevich, and Zhi-Wei Sun. Congruences for Apéry numbers β_{n} = ∑_{k=0}^{n} \binom{n}{k}^{2}\binom{n+k}{k}. International Journal of Number Theory (IJNT), 16(5):981– 1003, June 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500505.

Carr:2009:LIR

[CO09] Richard Carr and Cormac O’Sullivan. On the linear independence of roots. International Journal of Number Theory (IJNT), 5(1):161–171, February 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002018.

Cobbe:2021:RRH

[Cob21] Alessandro Cobbe. A representative of RΓ(N, T) for higher dimensional twists of Z_{p}^{r}(1). International Journal of Number Theory (IJNT), 17(08):1925– 1949, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042121500706.

Cohen:2006:PRQ

[Coh06]  Joseph Cohen.  Primitive

roots in quadratic fields. International Journal of Number Theory (IJNT), 2(1):7– 23, March 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000425.

Collas:2012:AGT

[Col12]  Benjamin Collas.  Action

of a Grothendieck–Teichmüller group on torsion elements of full Teichmüller modular groups of genus one. International Journal of Number Theory (IJNT), 8(3):763– 787, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500431.

Cooper:2006:QPI

[Coo06] Shaun Cooper. The quintuple product identity. International Journal of Number Theory (IJNT), 2(1):115– 161, March 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000401.

Cooper:2009:CES

[Coo09] Shaun Cooper. Construction of Eisenstein series for Γ_{0}(p). International Journal of Number Theory (IJNT), 5(5):765– 778, August 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002365.

Coons:2010:TSR

[Coo10] Michael Coons. The transcendence of series related to Stern’s diatomic sequence. International Journal of Number Theory (IJNT), 6(1):211–217, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110002958.

Coppola:2020:WGR

[Cop20] Nirvana Coppola. Wild Galois representations: Elliptic curves over a 2-adic field with non-abelian inertia action. International Journal of Number Theory (IJNT), 16(6):1199– 1208, July 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304212050061X.

Coulangeon:2008:EZF

[Cou08] Renaud Coulangeon. On Epstein’s zeta function of Humbert forms. International Journal of Number Theory (IJNT), 4(3): 387–401, June 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001407.

Cowan:2022:NRB

[Cow22] Alex Cowan. Non-random behavior in sums of modular symbols. International Journal of Number Theory (IJNT), 18(04): 879–903, May 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500464.

Chapman:2008:AWT

[CP08] Robin Chapman and Hao Pan. q-analogues of Wilson’s theorem. International Journal of Number Theory (IJNT), 4(4):539–547, August 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001511.

Chao:2010:NBS

[CP10a] Kuok Fai Chao and Roger Plymen. A new bound for the smallest x with π(x) > li(x). International Journal of Number Theory (IJNT), 6(3):681– 690, May 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003125.

Copil:2010:RCP

[CP10b] Vlad Copil and Laurenţiu Panaitopol. On the ratio of consecutive primes. International Journal of Number Theory (IJNT), 6(1):203–210, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110002934.

Chintamani:2017:SWZ

[CP17] Mohan N. Chintamani and Prabal Paul. On some weighted zero-sum constants. International Journal of Number Theory (IJNT), 13(2):301–308, March 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500191.

Chintamani:2018:SWZ

[CP18a] Mohan N. Chintamani and Prabal Paul. On some weighted zero-sum constants II. International Journal of Number Theory (IJNT), 14(2):383– 397, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500264.

Cho:2018:EBC

[CP18b] Bumkyu Cho and Ho Park. Evaluating binomial convolution sums of divisor functions in terms of Euler and Bernoulli polynomials. International Journal of Number Theory (IJNT), 14(2):509– 525, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500318.

Chen:2018:ECV

[CPS18] Evan Chen, Peter S. Park, and Ashvin A. Swaminathan. Elliptic curve variants of the least quadratic nonresidue problem and Linnik’s theorem. International Journal of Number Theory (IJNT), 14(1):255–288, February 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500161.

Chim:2017:VPP

[CPZ17] Kwok Chi Chim, István Pink, and Volker Ziegler. On a variant of Pillai’s problem. International Journal of Number Theory (IJNT), 13(7):1711– 1727, August 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500981.

Chirre:2022:SMS

[CQh22] Andrés Chirre and Oscar E.

Quesada-herrera. The second moment of S_{n}(t) on the Riemann hypothesis. International Journal of Number Theory (IJNT), 18(06):1203–1226, July 2022. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500610.

Cornut:2018:GPI

[CR18] Christophe Cornut and Jishnu Ray. Generators of the prop Iwahori and Galois representations. International Journal of Number Theory (IJNT), 14(1):37–53, February 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500045.

Creutz:2018:IRB

[Cre18] Brendan Creutz. Improved rank bounds from 2-descent on hyperelliptic Jacobians. International Journal of Number Theory (IJNT), 14(6):1709– 1713, July 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S179304211850104X.

Croot:2007:SNS

[Cro07] Ernie Croot. Smooth numbers in short intervals. International Journal of Number Theory (IJNT), 3(1):159– 169, March 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042107000833.

Caranay:2010:ESP

[CS10] Perlas C. Caranay and Renate Scheidler. An efficient seventh power residue symbol algorithm. International Journal of Number Theory (IJNT), 6(8):1831–1853, December 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003770.

Carpenter:2021:PTT

[CS21] Ryan Carpenter and Charles L. Samuels. A polynomial time test to detect numbers with many exceptional points. International Journal of Number Theory (IJNT), 17(04):973– 990, May 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500184.

Cai:2017:CIH

[CSJ17] Tianxin Cai, Zhongyan Shen, and Lirui Jia. A congruence involving harmonic sums modulo p^{α}q^{β}. International Journal of Number Theory (IJNT), 13(5):1083–1094, June 2017. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500580.

Chan:2013:SNA

[CT13] Tsz Ho Chan and Kai Man Tsang. Squarefull numbers in arithmetic progressions. International Journal of Number Theory (IJNT), 9(4):885– 901, June 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500048.

Chavez:2018:SMM

[CT18] Ángel Chávez and George Todd. Supercharacters and mixed moments of Kloosterman sums. International Journal of Number Theory (IJNT), 14(4):1023–1032, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042118500616.

Cao:2016:MSA

[CTZ16] Xiaodong Cao, Yoshio Tanigawa, and Wenguang Zhai. On the mean square of an arithmetical error term of the Selberg class in short intervals. International Journal of Number Theory (IJNT), 12(6):1675– 1701, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042116501037.

Cullinan:2012:SSA

[Cul12]  John Cullinan.  Symplectic

stabilizers with applications to Abelian varieties. International Journal of Number Theory (IJNT), 8(2):321–334, March 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500194.

Canci:2017:QMP

[CV17] Jung Kyu Canci and Solomon Vishkautsan. Quadratic maps with a periodic critical point of period 2. International Journal of Number Theory (IJNT), 13(6):1393–1417, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042117500786.

Caulk:2007:HOH

[CW07] Suzanne Caulk and Lynne H. Walling. Hecke operators on Hilbert–Siegel modular forms. International Journal of Number Theory (IJNT), 3(3):391– 420, September 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042107001048.

Chen:2012:DPN

[CW12] Zhixiong Chen and Arne Winterhof. On the distribution of pseudorandom numbers and vectors derived from Euler–Fermat quotients. International Journal of Number Theory (IJNT), 8(3):631– 641, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500352.

Chen:2019:VTL

[CW19] Bin Chen and Jie Wu. On values taken by the largest prime factor of shifted primes (II). International Journal of Number Theory (IJNT), 15(5):935– 944, June 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500507.

Conley:2016:HMJ

[CWR16] Charles H. Conley and Martin Westerholt-Raum. Harmonic Maaß–Jacobi forms of degree 1 with higher rank indices. International Journal of Number Theory (IJNT), 12 (7):1871–1897, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042116501165.

Cooper:2008:CMF

[CWW08] Yaim Cooper, Nicholas Wage, and Irena Wang. Congruences for modular forms of nonpositive weight. International Journal of Number Theory (IJNT), 4(1):1–13, February 2008. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001171.

Chen:2015:MAL

[CWW15] Haibo Chen, Xiaohua Wang, and Zhixiong Wen. Multifractal analysis on the level sets described by moving averages. International Journal of Number Theory (IJNT), 11 (7):2175–2189, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042115500967.

Cooper:2014:ECS

[CY14] Shaun Cooper and Dongxi Ye. Evaluation of the convolution sums ∑_{l+20m=n} σ(l)σ(m), ∑_{4l+5m=n} σ(l)σ(m) and ∑_{2l+5m=n} σ(l)σ(m). International Journal of Number Theory (IJNT), 10(6):1385– 1394, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042114500341.

Chen:2018:NVF

[CY18] Guohua Chen and Xiaofei Yan. Non-vanishing of the first derivative of GL(3) × GL(2) L-functions. International Journal of Number Theory (IJNT), 14(3):847–869, April 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500513.

Chu:2009:ISH

[CZ09] Wenchang Chu and Deyin Zheng. Infinite series with harmonic numbers and central binomial coefficients. International Journal of Number Theory (IJNT), 5(3):429– 448, May 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002171.

Chu:2010:BBL

[CZ10] Wenchang Chu and Wenlong Zhang. Bilateral Bailey lemma and false theta functions. International Journal of Number Theory (IJNT), 6(3):515– 577, May 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211000306X.

Duran:2018:ENP

[DA18] Ugur Duran and Mehmet Acikgoz. On ρ, q -Euler numbers and polynomials associated with ρ, q-Volkenborn integrals. International Journal of Number Theory (IJNT), 14(1):241– 253, February 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211850015X.

Duran:2019:AME

[DA19] Ugur Duran and Mehmet Acikgoz. On applications for Mahler expansion associated with p-adic q-integrals. International Journal of Number Theory (IJNT), 15(1):67– 84, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501737.

Dahmen:2011:RMA

[Dah11] Sander R. Dahmen. A refined modular approach to the Diophantine equation x^{2} + y^{2n} = z^{3}. International Journal of Number Theory (IJNT), 7(5):1303–1316, August 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004472.

Dahl:2018:SDD

[Dah18] Alexander Dahl. Subconvexity for a double Dirichlet series and non-vanishing of L-functions. International Journal of Number Theory (IJNT), 14(6):1573– 1604, July 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500951.

Dai:2014:CNT

[Dai14] Haobo Dai. Congruences for the number of 4-tuple partitions with distinct even parts. International Journal of Number Theory (IJNT), 10(8):2037– 2043, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042114500687.

Dai:2016:CHI

[Dai16a]  Haobo Dai.  Coefficients

of half-integral weight modular forms and indivisibility of class numbers of imaginary quadratic fields. International Journal of Number Theory (IJNT), 12(2):561– 565, March 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500354.

Dai:2016:NPB

[Dai16b] Haobo Dai. A note on the parity of broken 7-diamond partitions. International Journal of Number Theory (IJNT), 12 (6):1669–1674, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042116501025.

Das:2013:NDE

[Das13] Soumya Das. On the natural densities of eigenvalues of a Siegel cusp form of degree 2. International Journal of Number Theory (IJNT), 9(1):9–15, February 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501217.

Daughton:2014:HCT

[Dau14] Austin Daughton. A Hecke correspondence theorem for automorphic integrals with infinite log-polynomial sum period functions. International Journal of Number Theory (IJNT), 10(7):1857–1879, November 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500596.

Dewitt:2009:FGR

[DD09] Meghan Dewitt and Darrin Doud. Finding Galois representations corresponding to certain Hecke eigenclasses. International Journal of Number Theory (IJNT), 5(1):1– 11, February 2009. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109001888.

DelCorso:2010:NII

[DD10] Ilaria Del Corso and Roberto Dvornicich. Non-invariance of the index in wildly ramified extensions. International Journal of Number Theory (IJNT), 6(8):1855–1868, December 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003836.

DeAzevedoPribitkin:2007:UPS

[De 07] Wladimir De Azevedo Pribitkin. Uninhibited Poincaré series. International Journal of Number Theory (IJNT), 3(3):335–347, September 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304210700095X.

Dilcher:2021:PAR

[DE21] Karl Dilcher and Larry Ericksen. Polynomial analogues of restricted multicolor b-ary partition functions. International Journal of Number Theory (IJNT), 17(02):371–391, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400096.

Debaene:2019:ECI

[Deb19]  Korneel Debaene.  Explicit

counting of ideals and a Brun– Titchmarsh inequality for the Chebotarev density theorem. International Journal of Number Theory (IJNT), 15(5): 883–905, June 2019. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500477.

Defant:2015:RVD

[Def15] Colin Defant. On ranges of variants of the divisor functions that are dense. International Journal of Number Theory (IJNT), 11(6):1905–1912, September 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500827.

Delaunay:2005:MOT

[Del05]  Christophe Delaunay.  Mo-

ments of the orders of Tate– Shafarevich groups. International Journal of Number Theory (IJNT), 1(2):243– 264, June 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042105000133.

Demeio:2020:NRV

[Dem20] Julian Lawrence Demeio. Nonrational varieties with the Hilbert property. International Journal of Number Theory (IJNT), 16(4):803–822, May 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500414.

Dasaratha:2014:GFM

[DFG^{+}14] Krishna Dasaratha, Laure Flapan, Thomas Garrity, Chansoo Lee, Cornelia Mihaila, Nicholas Neumann-Chun, Sarah Peluse, and Matthew Stoffregen. A generalized family of multidimensional continued fractions: TRIP maps. International Journal of Number Theory (IJNT), 10(8):2151–2186, December 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042114500730.

Dujella:2008:PVP

[DFL08] Andrej Dujella, Clemens Fuchs, and Florian Luca. A polynomial variant of a problem of Diophantus for pure powers. International Journal of Number Theory (IJNT), 4(1):57–71, February 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001225.

Dobrowolski:2013:NCP

[DFV13] E. Dobrowolski, M. Filaseta, and A. F. Vincent. The noncyclotomic part of f(x)x^{n}+g(x) and roots of reciprocal polynomials off the unit circle. International Journal of Number Theory (IJNT), 9(7):1865– 1877, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042113500620.

Das:2013:NPS

[DG13] Soumya Das and Satadal Ganguly. Nonvanishing of Poincaré series on average. International Journal of Number Theory (IJNT), 9(1):1–8, February 2013. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501205.

Daniels:2017:REC

[DG17] Harris B. Daniels and Hannah Goodwillie. On the ranks of elliptic curves with isogenies. International Journal of Number Theory (IJNT), 13 (9):2215–2227, October 2017.

CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117501202.

Dudek:2016:PES

[DGM16] Adrian W. Dudek, Loïc Grenié, and Giuseppe Molteni. Primes in explicit short intervals on RH. International Journal of Number Theory (IJNT), 12(5): 1391–1407, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042116500858.

Dudek:2019:ESI

[DGM19] Adrian W. Dudek, Loïc Grenié, and Giuseppe Molteni. Explicit short intervals for primes in arithmetic progressions on GRH. International Journal of Number Theory (IJNT), 15(4): 825–862, May 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042119500441.

Dickinson:2013:MTM

[DH13] Detta Dickinson and Mumtaz Hussain. The metric theory of mixed type linear forms. International Journal of Number Theory (IJNT), 9(1):77–90, February 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112501278.

Dickson:2015:LSE

[Dic15] Martin Dickson. Local spectral equidistribution for degree two Siegel modular forms in level and weight aspects. International Journal of Number Theory (IJNT), 11(2):341– 396, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500190.

Dicks:2021:HWA

[Dic21] Robert Dicks. A higher weight analogue of Ogg’s theorem on Weierstrass points. International Journal of Number Theory (IJNT), 17(05):1155– 1162, June 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500299.

Dill:2020:GVR

[Dil20] Gabriel A. Dill. Generalized Vojta–Rémond inequality. International Journal of Number Theory (IJNT), 16(1):107–120, February 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500062.

Ding:2009:SIF

[Din09] Shanshan Ding. Smallest irreducible of the form x^{2} − dy^{2}. International Journal of Number Theory (IJNT), 5(3):449– 456, May 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042109002158.

Ding:2021:SEF

[Din21] Yuchen Ding. Sum of elements in finite Sidon sets. International Journal of Number Theory (IJNT), 17(04):991– 1001, May 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500196.

Dixit:2011:ATF

[Dix11]  Atul Dixit.  Analogues of

a transformation formula of Ramanujan. International Journal of Number Theory (IJNT), 7(5):1151–1172, August 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211100454X.

Dujella:2010:SSP

[DJ10] Andrej Dujella and Ana Jurasić. On the size of sets in a polynomial variant of a problem of Diophantus. International Journal of Number Theory (IJNT), 6(7):1449– 1471, November 2010. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003575.

Daileda:2015:PDC

[DJ15] R. Daileda and N. Jones. On primitivity of Dirichlet characters. International Journal of Number Theory (IJNT), 11(6):1913–1939, September 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500839.

Dilcher:2022:HDS

[DJ22a] Karl Dilcher and Lin Jiu.

Hankel determinants of sequences related to Bernoulli and Euler polynomials. International Journal of Number Theory (IJNT), 18(02):331– 359, March 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304212250021X.

Dubickas:2022:DIO

[DJ22b] Art = uras Dubickas and Lukas Jonuška. Divisibility of integers obtained from truncated periodic sequences. International Journal of Number Theory (IJNT), 18(01):165–174, February 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500129.

Djankovic:2011:NFA

[Dja11] Goran Djanković. Nonvanishing of the family of Γ_{1}(q)- automorphic L-functions at the central point. International Journal of Number Theory (IJNT), 7(6):1423–1439, September 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004800.

Djankovic:2013:EHP

[Dja13]  Goran Djanković.  Euler–

Hadamard products and power moments of symmetric square L-functions. International Journal of Number Theory (IJNT), 9(3):621–639, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S1793042112501527.

Dodson:2006:KTT

[DK06] M. M. Dodson and S. Kristensen. Khintchine’s theorem and transference principle for star bodies. International Journal of Number Theory (IJNT), 2(3):431–453, September 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042106000668.

Dixit:2021:HZF

[DK21] Atul Dixit and Rahul Kumar. On Hurwitz zeta function and Lommel functions. International Journal of Number Theory (IJNT), 17(02):393– 404, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120400047.

Ddamulira:2020:PPK

[DL20] Mahadi Ddamulira and Florian Luca. On the problem of Pillai with k-generalized Fibonacci numbers and powers of 3. International Journal of Number Theory (IJNT), 16(7):1643– 1666, August 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500876.

Du:2022:PFP

[DL22] Lara Du and Jeffrey C. Lagarias. Partial factorizations of products of binomial coefficients. International Journal of Number Theory (IJNT), 18(03):691–728, April 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500373.

Du:2019:CP

[DLZ19] Julia Q. D. Du, Edward Y. S.

Liu, and Jack C. D. Zhao. Congruence properties of p_{k}(n). International Journal of Number Theory (IJNT), 15(6):1267– 1290, July 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500714.

Dewar:2013:AFC

[DM13] Michael Dewar and M. Ram Murty. An asymptotic formula for the coefficients of j(z). International Journal of Number Theory (IJNT), 9(3):641– 652, May 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042112501539.

Das:2015:QE

[DM15] Soumya Das and Jaban Meher. On quasimodular eigenforms. International Journal of Number Theory (IJNT), 11(3):835– 842, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115500451.

Dixit:2019:LCF

[DM19] Anup B. Dixit and V. Kumar Murty. The Lindelöf class of L-functions II. International Journal of Number Theory (IJNT), 15(10):2201–2221, November 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042119501215.

Duncan:2020:MFD

[DM20] John F. R. Duncan and David A. McGady. Modular forms on the double half-plane. International Journal of Number Theory (IJNT), 16(09):1989–2003, October 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304212050102X.

Downey:2021:CMG

[DM21] Jenna Downey and Greg Martin. Counting multiplicative groups with prescribed subgroups. International Journal of Number Theory (IJNT), 17(09):2087–2112, October 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500780.

Do:2017:FFC

[Do17] Thong Nguyen Quang Do. Sur une forme faible de la conjecture de Greenberg, II. (French) [On a weak form of the Greenberg conjecture, II]. International Journal of Number Theory (IJNT), 13(4):1061– 1070, May 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500567.

Doris:2020:CGG

[Dor20] Christopher Doris. Computing the Galois group of a polynomial over a p-adic field. International Journal of Number Theory (IJNT), 16(08):1767–1801, September 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304212050092X.

David:2008:PLA

[DP08] Sinnou David and Amílcar Pacheco. Le problème de Lehmer abélien pour un module de Drinfel’d. (French) [The Lehmer abelien problem for a Drinfel’d module]. International Journal of Number Theory (IJNT), 4(6):1043–1067, December 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042108001870.

Dvornicich:2022:DFE

[DP22] Roberto Dvornicich and Laura Paladino. On the division fields of an elliptic curve and an effective bound to the hypotheses of the local-global divisibility. International Journal of Number Theory (IJNT), 18(07):1567– 1590, August 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500798.

Dewar:2010:RCS

[DR10] Michael Dewar and Olav K. Richter. Ramanujan congruences for Siegel modular forms. International Journal of Number Theory (IJNT), 6(7):1677– 1687, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S179304211000371X.

Dejou:2016:GBS

[DR16] Gaelle Dejou and Xavier-François Roblot. The Galois Brumer–Stark conjecture for SL_{2}(F_{3})-extensions. International Journal of Number Theory (IJNT), 12(1):165–188, February 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211650010X.

Delaygue:2022:PPP

[DR22] É. Delaygue and T. Rivoal. On primary pseudo-polynomials (around Ruzsa’s conjecture). International Journal of Number Theory (IJNT), 18(07): 1613–1636, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042122500828.

Draziotis:2011:NIP

[Dra11] Konstantinos A. Draziotis. On the number of integer points on the elliptic curve y^{2} = x^{3} + Ax. International Journal of Number Theory (IJNT), 7(3):611– 621, May 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004149.

Drappeau:2017:ADD

[Dra17] Sary Drappeau. On the average distribution of divisors of friable numbers. International Journal of Number Theory (IJNT), 13(1):153–193, February 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500105.

Dilcher:2007:PAS

[DS07] Karl Dilcher and Kenneth B.

Stolarsky. A polynomial analogue to the Stern sequence. International Journal of Number Theory (IJNT), 3(1):85– 103, March 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304210700081X.

Drmota:2013:EDR

[DS13] Michael Drmota and Mariusz Skałba. Equidistribution of divisors in residue classes and representations by binary quadratic forms. International Journal of Number Theory (IJNT), 9(8):2011–2018, December 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500711.

Dobrowolski:2017:MMP

[DS17] Edward Dobrowolski and Chris Smyth. Mahler measures of polynomials that are sums of a bounded number of monomials. International Journal of Number Theory (IJNT), 13(6):1603– 1610, July 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500907.

duSautoy:2018:NBE

[dS18a]  Marcus du Sautoy.  Natu-

ral boundaries for Euler products of Igusa zeta functions of elliptic curves. International Journal of Number Theory (IJNT), 14(8):2317–2331, September 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501415.

Dubon:2018:EED

[DS18b] Eric Dubon and Juan Matías Sepulcre. On the existence of equivalent Dirichlet polynomials whose zeros preserve a topological property. International Journal of Number Theory (IJNT), 14(3):713– 725, April 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500458.

Dummigan:2021:CLO

[DS21] Neil Dummigan and David Spencer. Congruences of local origin and automorphic induction. International Journal of Number Theory (IJNT), 17(07):1617–1629, August 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500512.

Dofs:2022:E

[DT22] Erik Dofs and Nguyen Xuan Tho. The equation x_{1}x_{2} + x_{2}x_{3} + x_{3}x_{4} + x_{4}x_{1} = n. International Journal of Number Theory (IJNT), 18(01):75–87, February 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500075.

David:2010:SFD

[DU10] Chantal David and Jorge Jiménez Urroz. Square-free discriminants of Frobenius rings. International Journal of Number Theory (IJNT), 6(6):1391– 1412, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042110003599.

Dubickas:2009:FPR

[Dub09] Artūras Dubickas. On the fractional parts of rational powers. International Journal of Number Theory (IJNT), 5(6): 1037–1048, September 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042109002535.

Dubickas:2011:RPD

[Dub11]  Artūras Dubickas.  Roots

of polynomials with dominant term. International Journal of Number Theory (IJNT), 7(5):1217–1228, August 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004575.

Dubickas:2014:PMR

[Dub14] Artūras Dubickas. Polynomials with multiple roots at 1. International Journal of Number Theory (IJNT), 10(2):391– 400, March 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113501005.

Dubickas:2018:IPN

[Dub18] Artūras Dubickas. Intervals without primes near elements of linear recurrence sequences. International Journal of Number Theory (IJNT), 14(2):567– 579, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500355.

Duda:2014:MCT

[Dud14] Dominik Duda. The maximal coefficient of ternary cyclotomic polynomials with one free prime. International Journal of Number Theory (IJNT), 10(4):1067–1080, June 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500158.

Dudek:2015:RHD

[Dud15] Adrian W. Dudek. On the Riemann hypothesis and the difference between primes. International Journal of Number Theory (IJNT), 11(3):771– 778, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500426.

Dummigan:2005:RTO

[Dum05] Neil Dummigan. Rational torsion on optimal curves. International Journal of Number Theory (IJNT), 1(4):513–531, December 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000340.

Dummigan:2009:SSF

[Dum09]  Neil Dummigan.  Symmet-

ric square L-functions and Shafarevich–Tate groups, II. International Journal of Number Theory (IJNT), 5(7):1321– 1345, November 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042109002699.

Delaunay:2009:SPE

[DW09] Christophe Delaunay and Christian Wuthrich. Self-points on elliptic curves of prime conductor. International Journal of Number Theory (IJNT), 5(5):911–932, August 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002456.

Dai:2016:MPR

[DW16] Haobo Dai and Yudong Wang. More parity results for broken 9-diamond partitions. International Journal of Number Theory (IJNT), 12(7):1937–1949, November 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116501190.

Dobi:2009:SRT

[DWW09] Doris Dobi, Nicholas Wage, and Irena Wang. Supersingular rank two Drinfel’d modules and analogs of Atkin’s orthogonal polynomials. International Journal of Number Theory (IJNT), 5(5):885– 895, August 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002444.

Dai:2017:SGP

[DWW17] Yuxia Dai, Chun Wei, and Shengyou Wen. Some geometric properties of sets defined by digit restrictions. International Journal of Number Theory (IJNT), 13(1):65– 75, February 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500051.

Eddin:2016:EUB

[Edd16] Sumaia Saad Eddin. An explicit upper bound for |L(1, χ)| when χ(2) = 1 and χ is even. International Journal of Number Theory (IJNT), 12(8):2299–2315, December 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116501372.

Edixhoven:2005:AGT

[Edi05] Bas Edixhoven. On the padic geometry of traces of singular moduli. International Journal of Number Theory (IJNT), 1(4):495–497, December 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000327.

Earnest:2016:OBQ

[EF16] A. G. Earnest and Robert W. Fitzgerald. The order of binary quadratic forms from representation numbers. International Journal of Number Theory (IJNT), 12(3):679– 690, May 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500457.

El-Guindy:2007:LCR

[EG07]  Ahmad El-Guindy.  Linear

congruences and relations on spaces of cusp forms. International Journal of Number Theory (IJNT), 3(4):529–539, December 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107001097.

El-Guindy:2009:FEM

[EG09] Ahmad El-Guindy. Fourier expansions with modular form coefficients. International Journal of Number Theory (IJNT), 5(8):1433–1446, December 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042109002717.

Echi:2012:KS

[EG12] Othman Echi and Nejib Ghanmi.

The Korselt set of pq. International Journal of Number Theory (IJNT), 8(2):299– 309, March 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500170.

El-Guindy:2014:LDM

[EG14] Ahmad El-Guindy. Legendre Drinfeld modules and universal supersingular polynomials. International Journal of Number Theory (IJNT), 10(5):1277– 1289, August 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500262.

Erazo:2021:PPC

[EGL21] Harold S. Erazo, Carlos A.

Gómez, and Florian Luca. On Pillai’s problem with Xcoordinates of Pell equations and powers of 2 II. International Journal of Number Theory (IJNT), 17(10):2251–2277, November 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500871.

Emory:2021:TST

[EGP21] Melissa Emory, Heidi Goodson, and Alexandre Peyrot. Towards the Sato–Tate groups of trinomial hyperelliptic curves. International Journal of Number Theory (IJNT), 17(10):2175– 2206, November 2021. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500822.

Ernvall-Hytonen:2008:ETA

[EH08] Anne-Maria Ernvall-Hytönen. On the error term in the approximate functional equation for exponential sums related to cusp forms. International Journal of Number Theory (IJNT), 4(5):747–756, October 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001730.

Ehlen:2010:TBP

[Ehl10]  Stephan Ehlen.  Twisted

Borcherds products on Hilbert modular surfaces and the regularized theta lift. International Journal of Number Theory (IJNT), 6(7):1473–1489, November 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003642.

Ehrenpreis:2009:EPS

[Ehr09]  Leon Ehrenpreis.  Eisen-

stein and Poincaré series on SL(3, r). International Journal of Number Theory (IJNT), 5(8):1447–1475, December 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002729.

Elsenhans:2011:CSG

[EJ11] Andreas-Stephan Elsenhans and Jörg Jahnel. Cubic surfaces with a Galois invariant pair of Steiner trihedra. International Journal of Number Theory (IJNT), 7(4):947– 970, June 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004253.

Elsenhans:2019:PQG

[EJ19] Andreas-Stephan Elsenhans and Jörg Jahnel. Plane quartics with a Galois-invariant Steiner hexad. International Journal of Number Theory (IJNT), 15(5):1075–1109, June 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211950060X.

Eliahou:2007:BMS

[EK07] Shalom Eliahou and Michel Kervaire. Bounds on the minimal sumset size function in groups. International Journal of Number Theory (IJNT), 3 (4):503–511, December 2007. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042107001085.

Earnest:2020:IQF

[EK20] A. G. Earnest and Ji Young Kim. Integral quadratic forms avoiding arithmetic progressions. International Journal of Number Theory (IJNT), 16(10): 2141–2148, November 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042120501109.

ElBachraoui:2011:IFS

[El 11] Mohamed El Bachraoui. Inductive formulas for some arithmetic functions. International Journal of Number Theory (IJNT), 7(8):2261–2268, December 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211100509X.

ElBachraoui:2012:ASL

[El 12] Mohamed El Bachraoui. Arithmetic sums of Liouville type over relatively prime pairs. International Journal of Number Theory (IJNT), 8(6):1503– 1512, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042112500893.

Eie:2018:SFM

[EL18] Minking Eie and Tung-Yang Lee. Sum formulas of multiple zeta values with selected arguments. International Journal of Number Theory (IJNT), 14(10):2617–2630, November 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501567.

Eldin:2019:NIS

[Eld19] Ramy F. Taki Eldin. On the number of incongruent solutions to a quadratic congruence over algebraic integers. International Journal of Number Theory (IJNT), 15(1):105–130, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501762.

Elliott:2015:AMF

[Ell15] P. D. T. A. Elliott. Additive and multiplicative functions from Nottingham to Urbana–Champaign. International Journal of Number Theory (IJNT), 11(5):1357–1366, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115400138.

Eie:2016:DTP

[ELO16] Minking Eie, Wen-Chin Liaw, and Yao Lin Ong. The decomposition theorem of products of multiple zeta values of height one. International Journal of Number Theory (IJNT), 12(1):15–25, February 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500020.

Eie:2017:SWS

[ELO17] Minking Eie, Wen-Chin Liaw, and Yao Lin Ong. Several weighted sum formulas of multiple zeta values. International Journal of Number Theory (IJNT), 13(9):2253–2264, October 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117501238.

Eie:2005:EGE

[ELY05] Minking Eie, Wen-Chin Liaw, and Fu-Yao Yang. On evaluation of generalized Euler sums of even weight. International Journal of Number Theory (IJNT), 1(2):225– 242, June 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000182.

El-Mahassni:2008:DCD

[EMG08] Edwin D. El-Mahassni and Domingo Gomez. On the distribution of counter-dependent nonlinear congruential pseudorandom number generators in residue rings. International Journal of Number Theory (IJNT), 4(6):1009–1018, December 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042108001857.

Eyyunni:2021:IBF

[EMS21] Pramod Eyyunni, Bibekananda Maji, and Garima Sood. An inequality between finite analogues of rank and crank moments. International Journal of Number Theory (IJNT), 17(02):405–423, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400217.

El-Mahassni:2006:DNC

[EMW06] Edwin D. El-Mahassni and Arne Winterhof. On the distribution of nonlinear congruential pseudorandom numbers in residue rings. International Journal of Number Theory (IJNT), 2(1):163– 168, March 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000413.

Eie:2005:EDE

[EOY05] Minking Eie, Yao Lin Ong, and Fu Yao Yang. Evaluating double Euler sums over rationally deformed simplices. International Journal of Number Theory (IJNT), 1(3):439–458, September 2005. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042105000273.

ElBasraoui:2012:REF

[ES12] Abdelkrim El Basraoui and Abdellah Sebbar. Rational equivariant forms. International Journal of Number Theory (IJNT), 8(4):963– 981, June 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500571.

Bachraoui:2020:SNS

[ES20] Mohamed El Bachraoui and József Sándor. Some new sums of q-trigonometric and related functions through a theta product of Jacobi. International Journal of Number Theory (IJNT), 16(08):1803–1817, September 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500931.

Elsholtz:2017:RPA

[ETT17] Christian Elsholtz, Niclas Technau, and Robert Tichy. On the regularity of primes in arithmetic progressions. International Journal of Number Theory (IJNT), 13(5):1349– 1361, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117500750.

Evans:2007:RRP

[EV07] Ronald Evans and Mark Van Veen. Rational representations of primes by binary quadratic forms. International Journal of Number Theory (IJNT), 3(4):513–528, December 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107000936.

Everlove:2022:DSA

[Eve22]  Corey Everlove.  Dirich-

let series associated to sumof-digits functions. International Journal of Number Theory (IJNT), 18(04):777– 798, May 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500415.

Evans:2015:QRQ

[EW15] Ronald J. Evans and Kenneth S. Williams. Quartic residuacity and the quadratic character of certain quadratic irrationalities. International Journal of Number Theory (IJNT), 11(8):2487–2503, December 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115501158.

Farag:2008:DTR

[Far08]  Hany M. Farag.  Dirichlet

truncations of the Riemann zeta function in the critical strip possess zeros near every vertical line. International Journal of Number Theory (IJNT), 4(4):653–662, August 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042108001596.

Felix:2012:GTD

[Fel12] Adam Tyler Felix. Generalizing the Titchmarsh divisor problem. International Journal of Number Theory (IJNT), 8(3): 613–629, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500340.

Ferrari:2022:BMV

[Fer22] Matteo Ferrari. On a basic mean value theorem with explicit exponents. International Journal of Number Theory (IJNT), 18(03):673– 690, April 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500361.

Fernandez:2015:ANG

[FF15] José L. Fernández and Pablo Fernández. Asymptotic normality and greatest common divisors. International Journal of Number Theory (IJNT), 11(1):89–126, February 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115500062.

Fanali:2012:NRP

[FG12] Stefania Fanali and Massimo Giulietti. On the number of rational points of generalized Fermat curves over finite fields. International Journal of Number Theory (IJNT), 8(4):1087– 1097, June 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500650.

Flynn:2014:GCD

[FG14] Ryan Flynn and Derek Garton. Graph components and dynamics over finite fields. International Journal of Number Theory (IJNT), 10(3):779– 792, May 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113501224.

Fukunaga:2022:RRC

[FG22] Kengo Fukunaga and Kohta Gejima. Remarks on the rightmost critical value of the triple product L-function. International Journal of Number Theory (IJNT), 18(06):1351– 1366, July 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500683.

Filaseta:2015:SGB

[FGT15] M. Filaseta, S. Graham, and O. Trifonov. Starting with gaps between k-free numbers. International Journal of Number Theory (IJNT), 11(5):1411– 1435, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115400199.

Fukshansky:2013:WRI

[FHL^{+}13] Lenny Fukshansky, Glenn Henshaw, Philip Liao, Matthew Prince, Xun Sun, and Samuel Whitehead. On well-rounded ideal lattices II. International Journal of Number Theory (IJNT), 9(1):139–154, February 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501291.

Faber:2011:NRI

[FHS11] Xander Faber, Benjamin Hutz, and Michael Stoll. On the number of rational iterated preimages of the origin under quadratic dynamical systems. International Journal of Number Theory (IJNT), 7 (7):1781–1806, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042111004162.

Friedlander:2005:IS

[FI05] J. B. Friedlander and H. Iwaniec. The illusory sieve. International Journal of Number Theory (IJNT), 1(4):459– 494, December 2005. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042105000303.

Finotti:2012:NPC

[Fin12] Luís R. A. Finotti. Nonexistence of pseudo-canonical liftings. International Journal of Number Theory (IJNT), 8(1): 31–51, February 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042112500029.

Finotti:2014:CWV

[Fin14] Luís R. A. Finotti. Computations with Witt vectors and the Greenberg transform. International Journal of Number Theory (IJNT), 10(6):1431–1458, September 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500377.

Finotti:2020:WCC

[Fin20] Luís R. A. Finotti. Weierstrass coefficients of the canonical lifting. International Journal of Number Theory (IJNT), 16(2): 397–422, March 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042120500219.

Furusho:2020:AMF

[FJ20] Hidekazu Furusho and David Jarossay. p-adic multiple Lfunctions and cyclotomic multiple harmonic values. International Journal of Number Theory (IJNT), 16(2):361– 375, March 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500190.

Fukuda:2011:WCN

[FK11] Takashi Fukuda and Keiichi Komatsu. Weber’s class number problem in the cyclotomic Z_{2}- extension of Q, III. International Journal of Number Theory (IJNT), 7(6):1627–1635, September 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004782.

Faye:2015:E

[FL15] Bernadette Faye and Florian Luca. On the equation ϕ(X^{m} − 1) = X^{n} − 1. International Journal of Number Theory (IJNT), 11(5):1691–1700, August 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115400187.

Flammang:2018:ZZM

[Fla18] V. Flammang. On the Zhang– Zagier measure. International Journal of Number Theory (IJNT), 14(10):2663–2671, November 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501609.

Flammang:2019:ATT

[Fla19]  V. Flammang.  The abso-

lute trace of totally positive algebraic integers. International Journal of Number Theory (IJNT), 15(1):173–181, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500064.

Flicker:2011:CFP

[Fli11]  Yuval Z. Flicker.  Cusp

forms on GSp(4) with SO(4)- periods. International Journal of Number Theory (IJNT), 7 (4):855–919, June 2011. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004186.

Filaseta:2012:S

[FLS12] M. Filaseta, S. Laishram, and N. Saradha. Solving n(n + d) ⋯ (n + (k − 1)d) = by^{2} with P (b) ≤ Ck. International Journal of Number Theory (IJNT), 8(1):161–173, February 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500091.

Felix:2012:PFR

[FM12] Adam Tyler Felix and M. Ram Murty. A problem of Fomenko’s related to Artin’s conjecture. International Journal of Number Theory (IJNT), 8(7):1687– 1723, November 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042112500984.

Franc:2020:CVV

[FM20] Cameron Franc and Geoffrey Mason. Constructions of vectorvalued modular forms of rank four and level one. International Journal of Number Theory (IJNT), 16(5):1111– 1152, June 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500578.

Fisher:2014:CCT

[FN14] Tom Fisher and Rachel Newton. Computing the Cassels– Tate pairing on the 3-Selmer group of an elliptic curve. International Journal of Number Theory (IJNT), 10(7):1881– 1907, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042114500602.

Folsom:2009:CMU

[Fol09] Amanda Folsom. A characterization of the modular units. International Journal of Number Theory (IJNT), 5(2):303– 310, March 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002110.

Folsom:2021:AEP

[Fol21]  Amanda Folsom.  Asymp-

totic expansions, partial theta functions, and radial limit differences of mock modular and modular forms. International Journal of Number Theory (IJNT), 17(02):425– 434, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400126.

Fowler:2020:TSM

[Fow20]  Guy Fowler.  Triples of

singular moduli with rational product. International Journal of Number Theory (IJNT), 16(10):2149–2166, November 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120501110.

Fehm:2010:RAV

[FP10] Arno Fehm and Sebastian Petersen. On the rank of Abelian varieties over ample fields. International Journal of Number Theory (IJNT), 6(3):579– 586, May 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003071.

Friedman:2012:SVD

[FP12a] Eduardo Friedman and Aldo Pereira. Special values of Dirichlet series and zeta integrals. International Journal of Number Theory (IJNT), 8(3): 697–714, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500406.

Fukshansky:2012:WRI

[FP12b] Lenny Fukshansky and Kathleen Petersen. On wellrounded ideal lattices. International Journal of Number Theory (IJNT), 8(1):189–206, February 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211250011X.

Freiberg:2015:NST

[FP15] Tristan Freiberg and Carl Pomerance. A note on square totients. International Journal of Number Theory (IJNT), 11 (8):2265–2276, December 2015. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211550102X.

Frantzen:2021:NOV

[Fra21] Maike Ella Elisabeth Frantzen. Non-openness of v-adic Galois representation for a-motives. International Journal of Number Theory (IJNT), 17(01):33– 53, February 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500032.

Feng:2020:ELP

[FRcT20] Tony Feng, Niccolò Ronchetti, and Cheng chiang Tsai. Epipelagic Langlands parameters and Lpackets for unitary groups. International Journal of Number Theory (IJNT), 16(7):1449– 1491, August 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500773.

FreixasIMontplet:2012:JLC

[Fre12] Gerard Freixas I.Montplet.

The Jacquet–Langlands correspondence and the arithmetic Riemann–Roch theorem for pointed curves. International Journal of Number Theory (IJNT), 8(1):1–29, February 2012. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500017.

Fregoli:2019:SRF

[Fre19] Reynold Fregoli. Sums of reciprocals of fractional parts. International Journal of Number Theory (IJNT), 15(4):789– 797, May 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500416.

Friedman:2016:MDB

[Fri16] Joshua S. Friedman. On the minimal distance between elliptic fixed points for geometrically finite Fuchsian groups.

International Journal of Number Theory (IJNT), 12(6):1663– 1668, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042116501013.

Ford:2008:CFF

[FS08] Kevin Ford and Igor Shparlinski. On curves over finite fields with Jacobians of small exponent. International Journal of Number Theory (IJNT), 4(5):819–826, October 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001687.

Freiman:2011:SSC

[FS11] Gregory A. Freiman and Yonutz V. Stanchescu. Sets with several centers of symmetry. International Journal of Number Theory (IJNT), 7 (5):1115–1135, August 2011. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004174.

Freitas:2015:COF

[FT15] Nuno Freitas and Panagiotis Tsaknias. Criteria for pordinarity of families of elliptic curves over infinitely many number fields. International Journal of Number Theory (IJNT), 11(1):81–87, February 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500050.

Fu:2018:MCT

[FT18] Shishuo Fu and Dazhao Tang.

Multiranks and classical theta functions. International Journal of Number Theory (IJNT), 14(2):549–566, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042118500343.

Fu:2011:CPO

[Fu11] Shishuo Fu. Combinatorial proof of one congruence for the broken 1-diamond partition and a generalization. International Journal of Number Theory (IJNT), 7(1):133–144, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004022.

Fujii:2020:MQI

[Fuj20] Satoshi Fujii. On minus quotients of ideal class groups of cyclotomic fields. International Journal of Number Theory (IJNT), 16(09):2013–2026, October 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120501043.

Fukshansky:2008:SZQ

[Fuk08] Lenny Fukshansky. Small zeros of quadratic forms over.

International Journal of Number Theory (IJNT), 4(3):503– 523, June 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001481.

Feigon:2010:EAC

[FW10] Brooke Feigon and David Whitehouse. Exact averages of central values of triple product L-functions. International Journal of Number Theory (IJNT), 6(7):1609–1624, November 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211000368X.

Fraser:2020:LSC

[FW20] Robert Fraser and James Wright. The local sum conjecture in two dimensions. International Journal of Number Theory (IJNT), 16(08): 1667–1699, September 2020. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500888.

Fretwell:2021:HOH

[FW21] Dan Fretwell and Lynne Walling. Hecke operators on Hilbert– Siegel theta series. International Journal of Number Theory (IJNT), 17(09):1965–1996, October 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S179304212150072X.

Fan:2021:PCC

[FWX21] Yan Fan, Liuquan Wang, and Ernest X. W. Xia. Proof of a conjectural congruence of Chan for Appell–Lerch sums. International Journal of Number Theory (IJNT), 17(04):1003– 1011, May 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500202.

Fukuhara:2013:BFE

[FY13] Shinji Fukuhara and Yifan Yang. Bases for S_{k}(Γ_{1}(4)) and formulas for even powers of the Jacobi theta function. International Journal of Number Theory (IJNT), 9(8):1973–1993, December 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500693.

Fukuda:2017:RDE

[FY17] Kohei Fukuda and Sho Yoshikawa.

The 12-th roots of the discriminant of an elliptic curve and the torsion points. International Journal of Number Theory (IJNT), 13(4):1037– 1060, May 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500555.

Fu:2021:SSP

[FY21] Ruiqin Fu and Hai Yang. On the solvability of the simultaneous Pell equations x^{2} − ay^{2} = 1 and y^{2} − bz^{2} = v_{1}^{2}. International Journal of Number Theory (IJNT), 17(09):1997– 2008, October 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500731.

Flicker:2006:TCS

[FZ06] Yuval Z. Flicker and Dmitrii Zinoviev. Twisted character of a small representation of Gl(4). International Journal of Number Theory (IJNT), 2 (3):329–350, September 2006. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042106000589.

Flicker:2012:CTC

[FZ12] Yuval Z. Flicker and Dmitrii Zinoviev. Computation of a twisted character of a small representation of Gl(3, E). International Journal of Number Theory (IJNT), 8(5):1153– 1230, August 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500704.

Fan:2016:EGZ

[FZ16] Yushuang Fan and Qinghai Zhong. On the Erdős– Ginzburg–Ziv constant of groups of the form C_{2}^{r} ⊕ C_{n}. International Journal of Number Theory (IJNT), 12(4):913– 943, June 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500573.

Frenkel:2018:EGC

[FZ18] Péter E. Frenkel and Gergely Zábrádi. Estimating the greatest common divisor of the value of two polynomials. International Journal of Number Theory (IJNT), 14(9):2543–2554, October 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501518.

Gonzalez-aviles:2020:BGN

[Ga20] Cristian D. González-avilés.

Brauer groups and Néron class groups. International Journal of Number Theory (IJNT), 16(10):2275–2292, November 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL

https://www.worldscientific.
com/doi/10.1142/S179304212050116X.

Gamzon:2014:LTA

[Gam14] Adam Gamzon. Local torsion on abelian surfaces with real multiplication by. International Journal of Number Theory (IJNT), 10(7):1807–1827, November 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500572.

Ganguly:2010:DSC

[Gan10] Satadal Ganguly. On the dimension of the space of cusp forms of octahedral type. International Journal of Number Theory (IJNT), 6(4):767– 783, June 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042110003198.

Garthwaite:2008:CMT

[Gar08] Sharon Anne Garthwaite. The coefficients of the ω(q) mock theta function. International Journal of Number Theory (IJNT), 4(6):1027–1042, December 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042108001869.

Garvan:2010:CAS

[Gar10] F. G. Garvan. Congruences for Andrews’ smallest parts partition function and new congruences for Dyson’s rank. International Journal of Number Theory (IJNT), 6(2):281– 309, March 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211000296X.

Garcia:2018:TNH

[Gar18a] Elisa Lorenzo García. Twists of non-hyperelliptic curves of genus 3. International Journal of Number Theory (IJNT), 14(6):1785–1812, July 2018.

ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042118501075.

Garnek:2018:DEC

[Gar18b]  Jedrzej Garnek.  On p-

degree of elliptic curves. International Journal of Number Theory (IJNT), 14(3):693– 704, April 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500434.

Gonzalez:2015:F

[GBL15] Juan José Alba González, Pedro Berrizbeitia, and Florian Luca. On the formula F_{p} = u^{2} + pv^{2}. International Journal of Number Theory (IJNT), 11(1):185– 191, February 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500116.

Gong:2017:SSF

[GC17] Ming-Liang Gong and Yong-Gao Chen. The shifted sum of the first n values of Euler’s function. International Journal of Number Theory (IJNT), 13(5):1245– 1251, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500671.

Ge:2018:ICZ

[Ge18]  Fan Ge.  On the index

conjecture in zero-sum theory: Singular case. International Journal of Number Theory (IJNT), 14(2):355–364, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500239.

Gekeler:2011:ZDD

[Gek11] Ernst-Ulrich Gekeler. Zero distribution and decay at infinity of Drinfeld modular coefficient forms. International Journal of Number Theory (IJNT), 7(3): 671–693, May 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004307.

Garcia-Fritz:2018:QSP

[GF18] Natalia Garcia-Fritz. Quadratic sequences of powers and Mohanty’s conjecture. International Journal of Number Theory (IJNT), 14(2):479–507, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500306.

Gillibert:2022:GCU

[GG22] Jean Gillibert and Pierre Gillibert. Galois covers of P^{1} and number fields with large class groups. International Journal of Number Theory (IJNT), 18(06):1261– 1288, July 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042122500646.

Gao:2011:QAN

[GGW11] Weidong Gao, Alfred Geroldinger, and Qinghong Wang. A quantitative aspect of nonunique factorizations: the Narkiewicz constants. International Journal of Number Theory (IJNT), 7(6):1463–1502, September 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004721.

Getz:2014:ACT

[GH14] Jayce R. Getz and Heekyoung Hahn. Algebraic cycles and Tate classes on Hilbert modular varieties. International Journal of Number Theory (IJNT), 10(1):161–176, February 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500875.

Grekos:2006:VTC

[GHHP06] G. Grekos, L. Haddad, C. Helou, and J. Pihko. Variations on a theme of Cassels for additive bases. International Journal of Number Theory (IJNT), 2(2): 249–265, June 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000553.

Ghitza:2011:DHE

[Ghi11] Alexandru Ghitza. Distinguishing Hecke eigenforms. International Journal of Number Theory (IJNT), 7(5):1247– 1253, August 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004708.

Gil:2015:APF

[GHK^{+}15a] Byung Keon Gil, Ji-Woo Han, Tae Hyun Kim, Ryun Han Koo, Bon Woo Lee, Jaehoon Lee, Kyeong Sik Nam, Hyeon Woo Park, and Poo-Sung Park. Acknowledgment of priority: “Frobenius numbers of Pythagorean triples”. International Journal of Number Theory (IJNT), 11(3):1061– 1062, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115920013.

Gil:2015:FNP

[GHK^{+}15b] Byung Keon Gil, Ji-Woo Han, Tae Hyun Kim, Ryun Han Koo, Bon Woo Lee, Jaehoon Lee, Kyeong Sik Nam, Hyeon Woo Park, and Poo-Sung Park. Frobenius numbers of Pythagorean triples. International Journal of Number Theory (IJNT), 11(2):613– 619, March 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500323.

Ghosh:2011:KGT

[Gho11] Anish Ghosh. A Khintchine– Groshev theorem for affine hyperplanes. International Journal of Number Theory (IJNT), 7(4):1045–1064, June 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042111004228.

Gibson:2014:DFN

[Gib14] Jason Gibson. The distribution of r-free numbers in arithmetic progressions. International Journal of Number Theory (IJNT), 10(3):559–563, May 2014. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113501078.

Gillespie:2013:FAF

[Gil13] Tim Gillespie. Factorization of automorphic L-functions and their zero statistics. International Journal of Number Theory (IJNT), 9(6):1367– 1378, September 2013. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500310.

Gillibert:2017:SAM

[Gil17]  Florence Gillibert.  Surface

abéliennes à multiplication quaternionique munie d’une structure de niveau Γ_{0}(N). (French) [Quaternionically multiplied abelian surface with a level structure Γ_{0}(N)]. International Journal of Number Theory (IJNT), 13(5):1317– 1333, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500725.

Gimenez:2013:ECT

[Gim13] Jose Gimenez. Eichler cohomology theorem for vectorvalued modular forms. International Journal of Number Theory (IJNT), 9(7):1765–1788, November 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500565.

Girstmair:2011:CFJ

[Gir11] Kurt Girstmair. Continued fractions and Jacobi symbols. International Journal of Number Theory (IJNT), 7(6):1543– 1555, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042111004848.

Girstmair:2012:DSS

[Gir12a] Kurt Girstmair. Dedekind sums with small denominators. International Journal of Number Theory (IJNT), 8(8):1965– 1970, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042112501102.

Girstmair:2012:PCF

[Gir12b] Kurt Girstmair. Periodic continued fractions and Jacobi symbols. International Journal of Number Theory (IJNT), 8 (6):1519–1525, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042112500911.

Girstmair:2014:ARN

[Gir14a]  Kurt Girstmair.  Approxi-

mation of rational numbers by Dedekind sums. International Journal of Number Theory (IJNT), 10(5):1241–1244, August 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500237.

Girstmair:2014:CED

[Gir14b]  Kurt Girstmair.  A cri-

terion for the equality of Dedekind sums modZ. International Journal of Number Theory (IJNT), 10(3):565– 568, May 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211350108X.

Girstmair:2015:FPD

[Gir15] Kurt Girstmair. On the fractional parts of Dedekind sums. International Journal of Number Theory (IJNT), 11(1):29– 38, February 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042115500013.

Girstmair:2016:CMA

[Gir16a] Kurt Girstmair. Congruences modulo 4 for the alternating sum of the partial quotients. International Journal of Number Theory (IJNT), 12(1):93– 96, February 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500056.

Girstmair:2016:DSE

[Gir16b] Kurt Girstmair. On Dedekind sums with equal values. International Journal of Number Theory (IJNT), 12(2):473– 481, March 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500287.

Girstmair:2017:LVD

[Gir17] Kurt Girstmair. The largest values of Dedekind sums. International Journal of Number Theory (IJNT), 13(6):1579– 1583, July 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117500889.

Girstmair:2018:DSA

[Gir18a] Kurt Girstmair. Dedekind sums in the p-adic number field. International Journal of Number Theory (IJNT), 14(2):527– 533, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211850032X.

Girstmair:2018:DST

[Gir18b]  Kurt Girstmair.  Dedekind

sums take each value infinitely many times. International Journal of Number Theory (IJNT), 14(4):1009– 1012, May 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500598.

Girstmair:2019:RRF

[Gir19a] Kurt Girstmair. On a recent reciprocity formula for Dedekind sums. International Journal of Number Theory (IJNT), 15(7):1469–1472, August 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500842.

Girstmair:2019:MDS

[Gir19b] Kurt Girstmair. On the moduli of a Dedekind sum. International Journal of Number Theory (IJNT), 15(4):863– 867, May 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500453.

Girstmair:2022:CPG

[Gir22] Kurt Girstmair. Chebyshev polynomials and Galois groups of De Moivre polynomials. International Journal of Number Theory (IJNT), 18(08):1735– 1748, September 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042122500889.

Gaal:2019:SQS

[GJR19] István Gaál, Borka Jadrijević, and László Remete. Simplest quartic and simplest sextic Thue equations over imaginary quadratic fields. International Journal of Number Theory (IJNT), 15(1):11– 27, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501695.

Garvan:2014:HTC

[GJS14] Frank G. Garvan and Chris Jennings-Shaffer. Hecke-type congruences for Andrews’ sptfunction modulo 16 and 32. International Journal of Number Theory (IJNT), 10(2):375– 390, March 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500991.

Garunkstis:2013:DMS

[GK13] Ramūnas Garunkštis and Justas Kalpokas. The discrete mean square of the Dirichlet L-function at nontrivial zeros of another Dirichlet L-function. International Journal of Number Theory (IJNT), 9(4):945– 963, June 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500085.

Gallenkamper:2018:HAO

[GK18a] Jonas Gallenkämper and Aloys Krieg. The Hecke algebras for the orthogonal group SO(2, 3) and the paramodular group of degree 2. International Journal of Number Theory (IJNT), 14(9):2409–2423, October 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501464.

Garunkstis:2018:ADB

[GK18b] Ramūnas Garunkštis and Laima Kaziulytė. Asymptotic distribution of Beurling integers. International Journal of Number Theory (IJNT), 14(10):2555– 2569, November 2018. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211850152X.

Goldston:2011:JCG

[GL11] D. A. Goldston and A. H.

Ledoan. Jumping champions and gaps between consecutive primes. International Journal of Number Theory (IJNT), 7 (6):1413–1421, September 2011. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S179304211100471X.

Guo:2016:PSC

[GL16] Victor J. W. Guo and Ji-Cai Liu. Proof of some conjectures of Z.-W. Sun on the divisibility of certain double sums. International Journal of Number Theory (IJNT), 12(3):615– 623, May 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500391.

Gong:2017:MVM

[GL17a] Ming-Liang Gong and Ya-Li Li. On the mean value of mixed exponential sums. International Journal of Number Theory (IJNT), 13(6):1515– 1530, July 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117500841.

Guo:2017:PCS

[GL17b] Victor J. W. Guo and Ji-Cai Liu. Proof of a congruence on sums of powers of q-binomial coefficients. International Journal of Number Theory (IJNT), 13(6):1571– 1577, July 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117500877.

Ge:2018:DAM

[GL18] Wenxu Ge and Huake Liu. Diophantine approximation with mixed powers of primes. International Journal of Number Theory (IJNT), 14(7):1903– 1918, August 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501130.

Gillibert:2019:ESL

[GL19] Jean Gillibert and Aaron Levin. Elliptic surfaces over P^{1} and large class groups of number fields. International Journal of Number Theory (IJNT), 15(10):2151–2162, November 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119501185.

Gao:2020:SPS

[GL20] Gaiyun Gao and Zhixin Liu.

Small prime solutions to diagonal Diophantine equations. International Journal of Number Theory (IJNT), 16(3): 447–460, April 2020. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500232.

Glass:2009:RHC

[Gla09] Darren Glass. The 2-ranks of hyperelliptic curves with extra automorphisms. International Journal of Number Theory (IJNT), 5(5):897–910, August 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042109002468.

Gao:2018:ULZ

[GLPW18] Weidong Gao, Yuanlin Li, Jiangtao Peng, and Guoqing Wang. A unifying look at zero-sum invariants. International Journal of Number Theory (IJNT), 14(3):705–711, April 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500446.

Guo:2015:FWS

[GLZ15] Li Guo, Peng Lei, and Jianqiang Zhao. Families of weighted sum formulas for multiple zeta values. International Journal of Number Theory (IJNT), 11(3):997– 1025, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115500530.

Goins:2012:PHR

[GM12] Edray Herber Goins and Kevin Mugo. Points on hyperbolas at rational distance. International Journal of Number Theory (IJNT), 8(4):911– 922, June 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112500534.

Gamzon:2018:NCG

[GM18a] Adam Gamzon and Lance Edward Miller. Nonexistence of certain Galois representations for quadratic fields. International Journal of Number Theory (IJNT), 14(5):1505– 1524, June 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500914.

Garrity:2018:GMQ

[GM18b] Thomas Garrity and Peter McDonald. Generalizing the Minkowski question mark function to a family of multidimensional continued fractions. International Journal of Number Theory (IJNT), 14(9):2473– 2516, October 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211850149X.

Gluncic:2019:CRC

[GM19] Matko Glunčić and Ivica Martinjak. A class of Srestricted compositions. International Journal of Number Theory (IJNT), 15(2):361– 371, March 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500180.

Gomez-Molleda:2012:CNF

[GML12] M. A. Gómez-Molleda and Joan-C. Lario. Class number formulas for certain bicyclic biquadratic number fields as finite sums involving Jacobi elliptic functions. International Journal of Number Theory (IJNT), 8(5):1257– 1270, August 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211250073X.

Gun:2011:AIV

[GMR11] Sanoli Gun, M. Ram Murty, and Purusottam Rath. Algebraic independence of values of modular forms. International Journal of Number Theory (IJNT), 7(4):1065– 1074, June 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004769.

Guo:2019:CIJ

[GN19] Victor J. W. Guo and He-Xia Ni. A q-congruence involving the Jacobi symbol. International Journal of Number Theory (IJNT), 15(10):1977–1981, November 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042119501082.

Godinho:2021:DEX

[GN21] Hemar Godinho and Victor G. L. Neumann. The diophantine equation x^{2}+p^{a}q^{b} = y^{q}. International Journal of Number Theory (IJNT), 17(09):2113– 2130, October 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500792.

Gaborit:2006:ELG

[GNS06] Philippe Gaborit, Ann Marie Natividad, and Patrick Solé. Eisenstein lattices, Galois rings and quaternary codes. International Journal of Number Theory (IJNT), 2(2):289– 303, June 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000577.

Goldfeld:2007:RLO

[Gol07] Dorian Goldfeld. Rank lowering operators on GL(n, R). International Journal of Number Theory (IJNT), 3(3):365–375, September 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107000985.

Golze:2016:BPT

[Gol16]  Hiram Golze.  A bijective

proof of a theorem on 3core partitions. International Journal of Number Theory (IJNT), 12(6):1483–1496, September 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500913.

Goodson:2017:HFR

[Goo17] Heidi Goodson. Hypergeometric functions and relations to Dwork hypersurfaces. International Journal of Number Theory (IJNT), 13(2):439– 485, March 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117500269.

Gorodetsky:2019:QCA

[Gor19] Ofir Gorodetsky. q-congruences, with applications to supercongruences and the cyclic sieving phenomenon. International Journal of Number Theory (IJNT), 15(9):1919–1968, October 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042119501069.

Gottesman:2020:AVV

[Got20] Richard Gottesman. The arithmetic of vector-valued modular forms on Γ_{0}(2). International Journal of Number Theory (IJNT), 16(2):241– 289, March 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500141.

Goudout:2018:HPP

[Gou18] Élie Goudout. Highest perfect power of a product of integers less than x. International Journal of Number Theory (IJNT), 14(7):2043–2044, August 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118501233.

Greve:2012:RPS

[GP12] Christian Greve and Sebastian Pauli. Ramification polygons, splitting fields, and Galois groups of Eisenstein polynomials. International Journal of Number Theory (IJNT), 8 (6):1401–1424, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042112500832.

Gallegos-Ruiz:2011:IPH

[GR11] Homero R. Gallegos-Ruiz. Sintegral points on hyperelliptic curves. International Journal of Number Theory (IJNT), 7(3): 803–824, May 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004435.

Grobner:2014:SAP

[GR14] Harald Grobner and A. Raghuram. On some arithmetic properties of automorphic forms of GL_{m} over a division algebra. International Journal of Number Theory (IJNT), 10(4):963– 1013, June 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500110.

Gillibert:2019:JRN

[GR19] Pierre Gillibert and Gabriele Ranieri. Julia Robinson numbers. International Journal of Number Theory (IJNT), 15 (8):1565–1599, September 2019.

ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042119500908.

Granville:2007:EPD

[Gra07a] Andrew Granville. Erratum: “Prime Divisors Are Poisson Distributed”. International Journal of Number Theory (IJNT), 3(4):649–651, December 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042107001073.

See [Gra07b].

Granville:2007:PDP

[Gra07b] Andrew Granville. Prime divisors are Poisson distributed. International Journal of Number Theory (IJNT), 3(1): 1–18, March 2007. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042107000778.

See erratum [Gra07a].

Graves:2011:NPE

[Gra11] Hester Graves. Has a nonprincipal Euclidean ideal. International Journal of Number Theory (IJNT), 7(8):2269– 2271, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042111004988.

Gras:2018:AKL

[Gra18] Georges Gras. The p-adic Kummer–Leopoldt constant: Normalized p-adic regulator. International Journal of Number Theory (IJNT), 14(2):329– 337, March 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500203.

Grabner:2020:QFS

[Gra20] Peter J. Grabner. Quasimodular forms as solutions of modular differential equations. International Journal of Number Theory (IJNT), 16(10):2233– 2274, November 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042120501158.

Griffin:2011:DPC

[Gri11]  Michael Griffin.  Divisibil-

ity properties of coefficients of weight 0 weakly holomorphic modular forms. International Journal of Number Theory (IJNT), 7(4):933– 941, June 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004599.

Grosu:2018:NPN

[Gro18] Codruţ Grosu. A note on projective norm graphs. International Journal of Number Theory (IJNT), 14(1):55– 62, February 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500057.

Grossi:2020:NRA

[Gro20] Giada Grossi. On norm relations for Asai–Flach classes. International Journal of Number Theory (IJNT), 16(10):2311– 2377, November 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042120501183.

Goldberg:2016:PMC

[GS16] David Goldberg and Dani Szpruch. Plancherel measures for coverings of p-adic SL_{2}(F). International Journal of Number Theory (IJNT), 12(7):1907– 1936, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042116501189.

Gun:2017:SCF

[GS17] S. Gun and J. Sengupta. Sign changes of Fourier coefficients of Siegel cusp forms of degree two on Hecke congruence subgroups. International Journal of Number Theory (IJNT), 13(10):2597–2625, November 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117501457.

Goral:2018:CSG

[GS18] Haydar Göral and Doğa Can Sertbaş. A congruence for some generalized harmonic type sums. International Journal of Number Theory (IJNT), 14(4): 1033–1046, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042118500628.

Gottesman:2010:QRP

[GT10] Richard Gottesman and Kwokfung Tang. Quadratic recurrences with a positive density of prime divisors. International Journal of Number Theory (IJNT), 6(5):1027– 1045, August 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003368.

Gu:2017:FPC

[GT17] Ze Gu and Xilin Tang. The Frobenius problem for a class of numerical semigroups. International Journal of Number Theory (IJNT), 13(5):1335– 1347, June 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500749.

Guerreiro:2015:ORT

[Gue15] João Guerreiro. An orthogonality relation for a thin family of GL(3) Maass forms. International Journal of Number Theory (IJNT), 11(8):2277–2294, December 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115501031.

Gugg:2021:CKS

[Gug21] Chadwick Gugg. On conjectures of Koike and Somos for modular identities for the Rogers–Ramanujan functions. International Journal of Number Theory (IJNT), 17(02):435– 472, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400138.

Guillera:2020:BRL

[Gui20]  Jesús Guillera.  Bilateral

Ramanujan-like series for 1/π^{k} and their congruences. International Journal of Number Theory (IJNT), 16(09):1969–1988, October 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120501018.

Guillera:2021:PRR

[Gui21] Jesús Guillera. Proof of a rational Ramanujan-type series for 1/π. The fastest one in level 3. International Journal of Number Theory (IJNT), 17(02):473– 477, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400242.

Gun:2006:TZC

[Gun06]  Sanoli Gun.  Transcenden-

tal zeros of certain modular forms. International Journal of Number Theory (IJNT), 2(4):549–553, December 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000711.

Gunaydin:2012:RSP

[Gün12]  Ayhan Günaydin.  Ratio-

nal solutions of polynomial– exponential equations. International Journal of Number Theory (IJNT), 8(6):1391–1399, September 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500820.

Guo:2015:SCR

[Guo15] Victor J. W. Guo. Some congruences related to the q-Fermat quotients. International Journal of Number Theory (IJNT), 11(3):1049– 1060, May 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500554.

Guo:2019:PCC

[Guo19a] Victor J. W. Guo. Proof of a q-congruence conjectured by Tauraso. International Journal of Number Theory (IJNT), 15(1):37–41, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501713.

Guo:2019:ASV

[Guo19b] Victor J. W. Guo. A qanalogue of the (I.2) supercongruence of Van Hamme. International Journal of Number Theory (IJNT), 15(1):29– 36, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501701.

Guo:2021:FQA

[Guo21] Victor J. W. Guo. A further qanalogue of Van Hamme’s (h.2) supercongruence for primes p ≡ 3 (mod 4). International Journal of Number Theory (IJNT), 17(05):1201– 1206, June 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500329.

Gurak:2008:PHK

[Gur08] S. Gurak. Polynomials for hyper-Kloosterman sums. International Journal of Number Theory (IJNT), 4(6):959–972, December 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001808.

Gurak:2011:GKS

[Gur11] S. Gurak. Gauss and Kloosterman sums over residue rings of algebraic integers. International Journal of Number Theory (IJNT), 7(1):101–114, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111003958.

Gross:2013:F

[GV13] Samuel S. Gross and Andrew F.

Vincent. On the factorization of f(n) for f(x) in Z[x]. International Journal of Number Theory (IJNT), 9(5):1225– 1236, August 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211350022X.

Garvan:2007:SSP

[GY07] Frank G. Garvan and Hamza Yesilyurt. Shifted and shiftless partition identities II. International Journal of Number Theory (IJNT), 3(1):43– 84, March 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042107000808.

Gunnells:2013:MFE

[GY13] Paul E. Gunnells and Dan Yasaki. Modular forms and elliptic curves over the cubic field of discriminant −23. International Journal of Number Theory (IJNT), 9(1):53–76, February 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501242.

Gillespie:2016:ZCL

[GY16] Timothy L. Gillespie and Yangbo Ye. Zero correlation with lower-order terms for automorphic L-functions. International Journal of Number Theory (IJNT), 12(1):27– 55, February 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500032.

Guo:2011:FSA

[GZ11] Victor J. W. Guo and Jiang Zeng. Factors of sums and alternating sums involving binomial coefficients and powers of integers. International Journal of Number Theory (IJNT), 7(7):1959–1976, November 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004812.

Guo:2012:NCS

[GZ12] Victor J. W. Guo and Jiang Zeng. New congruences for sums involving Apéry numbers or central Delannoy numbers. International Journal of Number Theory (IJNT), 8(8):2003–2016, December 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042112501138.

Gao:2022:LSP

[GZ22] Peng Gao and Liangyi Zhao.

The large sieve with power moduli in imaginary quadratic number fields. International Journal of Number Theory (IJNT), 18(08):1713–1733, September 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500877.

Gao:2015:ZSS

[GZZ15] Weidong Gao, Pingping Zhao, and Jujuan Zhuang. Zerosum subsequences of distinct lengths. International Journal of Number Theory (IJNT), 11 (7):2141–2150, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042115500931.

Hahn:2021:PTP

[Hah21] Heekyoung Hahn. Poles of triple product L-functions involving monomial representations. International Journal of Number Theory (IJNT), 17(02): 479–486, March 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042120400291.

Hajli:2015:AI

[Haj15] Mounir Hajli. On an arithmetic inequality on P_{Q}^{1}. International Journal of Number Theory (IJNT), 11(4):1227– 1231, June 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500669.

Hajli:2020:SPC

[Haj20] Mounir Hajli. The spectral properties of a continuous family of zeta functions. International Journal of Number Theory (IJNT), 16(4):693– 717, May 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500359.

Halupczok:2012:NBL

[Hal12] Karin Halupczok. A new bound for the large sieve inequality with power moduli. International Journal of Number Theory (IJNT), 8(3):689– 695, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211250039X.

Hamahata:2013:DDS

[Ham13]  Yoshinori Hamahata.  De-

nominators of Dedekind sums in function fields. International Journal of Number Theory (IJNT), 9(6):1423–1430, September 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500358.

Hamahata:2016:DST

[Ham16] Yoshinori Hamahata. Dedekind sums and the transformation formula in function fields. International Journal of Number Theory (IJNT), 12(8):2061– 2072, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042116501232.

Hamahata:2018:CTF

[Ham18] Yoshinori Hamahata. Chowla’s theorem over function fields. International Journal of Number Theory (IJNT), 14(6):1689– 1698, July 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501026.

Hanamoto:2017:ZWH

[Han17] Seiichi Hanamoto. Zeros of weakly holomorphic modular forms of level 5. International Journal of Number Theory (IJNT), 13(1):195–207, February 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500117.

Hart:2007:NCM

[Har07] William B. Hart. A new class of modular equation for Weber functions. International Journal of Number Theory (IJNT), 3(1):141– 157, March 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042107000845.

Harman:2008:WMV

[Har08] Glyn Harman. Watt’s mean value theorem and Carmichael numbers. International Journal of Number Theory (IJNT), 4(2):241–248, April 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042108001316.

Harada:2011:CSU

[Har11a] Masaaki Harada. Construction of some unimodular lattices with long shadows. International Journal of Number Theory (IJNT), 7(5):1345– 1358, August 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004794.

Harvey:2011:CDI

[Har11b] M. P. Harvey. Cubic Diophantine inequalities involving a norm form. International Journal of Number Theory (IJNT), 7(8):2219–2235, December 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111005052.

Harrington:2012:FT

[Har12] Joshua Harrington. On the factorization of the trinomials x^{n} + cx^{n−1} + d. International Journal of Number Theory (IJNT), 8(6):1513–1518, September 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211250090X.

Harrington:2015:TQC

[Har15] Joshua Harrington. Two questions concerning covering systems. International Journal of Number Theory (IJNT), 11 (6):1739–1750, September 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S179304211550075X.

Harada:2018:LRF

[Har18] Ryotaro Harada. On Lara Rodríguez’ full conjecture for double zeta values in function fields. International Journal of Number Theory (IJNT), 14(4): 1081–1092, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042118500677.

Hasegawa:2010:AOT

[Has10]  Takehiro Hasegawa.  On

asymptotically optimal towers over quadratic fields related to Gauss hypergeometric functions. International Journal of Number Theory (IJNT), 6(5):989–1009, August 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003344.

Hashimoto:2013:AFC

[Has13a] Yasufumi Hashimoto. Asymptotic formulas for class number sums of indefinite binary quadratic forms on arithmetic progressions. International Journal of Number Theory (IJNT), 9(1):27–51, February 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042112501230.

Hassani:2013:RAG

[Has13b] Mehdi Hassani. On the ratio of the arithmetic and geometric means of the prime numbers and the number e. International Journal of Number Theory (IJNT), 9(6):1593–1603, September 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500450.

Hatley:2016:OCM

[Hat16] Jeffrey Hatley. Obstruction criteria for modular deformation problems. International Journal of Number Theory (IJNT), 12(1):273–285, February 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500160.

Haukkanen:2018:DAF

[Hau18] Pentti Haukkanen. Derivation of arithmetical functions under the Dirichlet convolution.

International Journal of Number Theory (IJNT), 14(5):1257– 1264, June 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500781.

Haynes:2010:NDN

[Hay10] Alan K. Haynes. Numerators of differences of nonconsecutive Farey fractions. International Journal of Number Theory (IJNT), 6(3):655– 666, May 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003113.

Hayashida:2014:SFM

[Hay14] Shuichi Hayashida. On the spinor L-function of Miyawaki– Ikeda lifts. International Journal of Number Theory (IJNT), 10(2):297–307, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042113500930.

Heath-Brown:2017:RPI

[HB17] D. R. Heath-Brown. Rational points on the intersection of three quadrics. International Journal of Number Theory (IJNT), 13(2):273– 289, March 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117500178.

Chen:2021:GCT

[hC21] Chien hua Chen. Galois criterion for torsion points of Drinfeld modules. International Journal of Number Theory (IJNT), 17(10):2315–2326, November 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500901.

He:2020:IIP

[He20] Zilong He. Inequalities for inert primes and their applications. International Journal of Number Theory (IJNT), 16(08): 1819–1832, September 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042120500943.

Heap:2014:TSM

[Hea14] Winston Heap. The twisted second moment of the Dedekind zeta function of a quadratic field. International Journal of Number Theory (IJNT), 10 (1):235–281, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S1793042113500917.

Heim:2018:HMF

[Hei18] Bernhard Heim. Hecke’s modular forms by functional equations. International Journal of Number Theory (IJNT), 14(5): 1247–1256, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S179304211850077X.

Hertgen:2016:SPR

[Her16] Alan Hertgen. Splitting properties of the reduction of semiabelian varieties. International Journal of Number Theory (IJNT), 12(8):2241–2264, December 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116501359.

Fang:2022:NWP

[hF22] Jin hui Fang. A note on weakly prime-additive numbers. International Journal of Number Theory (IJNT), 18(01):175– 178, February 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500130.

Hajdu:2021:PVS

[HH21] L. Hajdu and O. Herendi. Polynomial values of surface point counting polynomials. International Journal of Number Theory (IJNT), 17(01):15–32, February 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500020.

Huber:2021:APS

[HHM21] Timothy Huber, Mayra Huerta, and Nathaniel Mayes. Arithmetic properties of septic partition functions. International Journal of Number Theory (IJNT), 17(02):499–512, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304212040031X.

Hegyvari:2009:ICL

[HHP09] Norbert Hegyvári, François Hennecart, and Alain Plagne. Iterated compositions of linear operations on sets of positive upper density. International Journal of Number Theory (IJNT), 5(6):981–997, September 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304210900250X.

Hichri:2014:BES

[Hic14] Hachem Hichri. Beta expansion for some particular sequences of Salem numbers. International Journal of Number Theory (IJNT), 10(8):2135–2149, December 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500729.

Hindes:2015:RPC

[Hin15] Wade Hindes. Rational points on certain families of symmetric equations. International Journal of Number Theory (IJNT), 11(6):1821–1838, September 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042115500797.

Hindes:2018:CGG

[Hin18] Wade Hindes. Classifying Galois groups of small iterates via rational points. International Journal of Number Theory (IJNT), 14(5):1403– 1426, June 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500872.

Hindes:2019:FOP

[Hin19] Wade Hindes. Finite orbit points for sets of quadratic polynomials. International Journal of Number Theory (IJNT), 15(8):1693–1719, September 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500945.

Hirschhorn:2006:CMS

[Hir06] Michael D. Hirschhorn. The case of the mysterious sevens. International Journal of Number Theory (IJNT), 2(2):213– 216, June 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042106000486.

Hironaka:2022:SFL

[Hir22] Yumiko Hironaka. Spherical functions and local densities on the space of p-adic quaternion Hermitian forms. International Journal of Number Theory (IJNT), 18(03):575– 612, April 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042122500324.

Hittmeir:2018:CAR

[Hit18]  Markus Hittmeir.  Com-

putational aspects of rational residuosity. International Journal of Number Theory (IJNT), 14(3):631–645, April 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500380.

Haddock:2014:ZWH

[HJ14] Andrew Haddock and Paul Jenkins. Zeros of weakly holomorphic modular forms of level 4. International Journal of Number Theory (IJNT), 10(2): 455–470, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042113501030.

Harrington:2014:RCP

[HJW14] Joshua Harrington, Lenny Jones, and Daniel White. The reducibility of constantperturbed products of cyclotomic polynomials. International Journal of Number Theory (IJNT), 10(1):13–29, February 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

worldscientific.com/doi/
10.1142/S1793042113500784.

Hussain:2013:MRS

[HK13] M. Hussain and S. Kristensen.

Metrical results on systems of small linear forms. International Journal of Number Theory (IJNT), 9(3):769– 782, May 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112501606.

Hironaka:2014:SFS

[HK14] Yumiko Hironaka and Yasushi Komori. Spherical functions on the space of p-adic unitary Hermitian matrices. International Journal of Number Theory (IJNT), 10(2):513– 558, March 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113501066.

Huang:2018:FN

[HK18] Hui Huang and Manuel Kauers.

D-finite numbers. International Journal of Number Theory (IJNT), 14(7):1827– 1848, August 2018. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501099.

Huh:2020:NEC

[HK20] Jisun Huh and Byungchan Kim. The number of equivalence classes arising from partition involutions. International Journal of Number Theory (IJNT), 16(5):925– 939, June 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500475.

Hirakawa:2021:HCP

[HK21] Yoshinosuke Hirakawa and Yoshinori Kanamura. How to calculate the proportion of everywhere locally soluble diagonal hypersurfaces. International Journal of Number Theory (IJNT), 17(10):2361–2377, November 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500925.

Hulse:2012:SFC

[HKKL12] Thomas A. Hulse, E. Mehmet Kiral, Chan Ieong Kuan, and Li-Mei Lim. The sign of Fourier coefficients of halfintegral weight cusp forms. International Journal of Number Theory (IJNT), 8(3):749– 762, May 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211250042X.

Hammonds:2021:EST

[HKL^{+}21] Trajan Hammonds, Casimir Kothari, Noah Luntzlara, Steven J. Miller, Jesse Thorner, and Hunter Wieman. The explicit Sato–Tate conjecture for primes in arithmetic progressions. International Journal of Number Theory (IJNT), 17(08):1905–1923, September 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304212150069X.

Hofer:2009:DPG

[HKLP09] Roswitha Hofer, Peter Kritzer, Gerhard Larcher, and Friedrich Pillichshammer. Distribution properties of generalized Van Der Corput–Halton sequences and their subsequences. International Journal of Number Theory (IJNT), 5(4):719– 746, June 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042109002328.

Han:2010:FNP

[HKN10] Jeong Soon Han, Hee Sik Kim, and J. Neggers. The Fibonaccinorm of a positive integer: Observations and conjectures. International Journal of Number Theory (IJNT), 6(2):371– 385, March 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003009.

See acknowledgement of priority [HKN11].

Han:2011:APF

[HKN11] Jeong Soon Han, Hee Sik Kim, and J. Neggers. Acknowledgment of priority: “The Fibonacci-norm of a positive integer: Observations and conjectures”. International Journal of Number Theory (IJNT), 7(3): 853–854, May 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004927.

See [HKN10].

Hentschel:2012:CEU

[HKN12] Michael Hentschel, Aloys Krieg, and Gabriele Nebe. On the classification of even unimodular lattices with a complex structure. International Journal of Number Theory (IJNT), 8(4): 983–992, June 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500583.

Kim:2021:TNQ

[hKS21] Seon hong Kim and Kenneth B. Stolarsky. Twin Nicomachean q-identities and conjectures for the associated discriminants, polynomials, and inequalities. International Journal of Number Theory (IJNT), 17(03):621– 645, April 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400175.

Hong:2011:ABS

[HL11] Shaofang Hong and Raphael Loewy. Asymptotic behavior of the smallest eigenvalue of matrices associated with completely even functions (modr).

International Journal of Number Theory (IJNT), 7(6):1681– 1704, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S179304211100437X.

Hill:2012:AIM

[HL12] Richard Hill and David Loeffler. p-adic interpolation of metaplectic forms of cohomological type. International Journal of Number Theory (IJNT), 8(7):1613–1660, November 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500960.

Huang:2020:TLP

[HL20] Jing-Jing Huang and Huixi Li.

On two lattice points problems about the parabola. International Journal of Number Theory (IJNT), 16(4):719– 729, May 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500360.

Heim:2019:RRP

[HLN19] Bernhard Heim, Florian Luca, and Markus Neuhauser. Recurrence relations for polynomials obtained by arithmetic functions. International Journal of Number Theory (IJNT), 15(6):1291–1303, July 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042119500726.

Hare:2011:SD

[HLS11] Kevin G. Hare, Shanta Laishram, and Thomas Stoll. The sum of digits of n and n^{2}. International Journal of Number Theory (IJNT), 7(7):1737–1752, November 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004319.

Hilberdink:2020:CSC

[HLT20] Titus Hilberdink, Florian Luca, and László Tóth. On certain sums concerning the gcd’s and lcm’s of k positive integers. International Journal of Number Theory (IJNT), 16(1):77– 90, February 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500049.

Hbaib:2006:BDC

[HM06] M. Hbaib and M. Mkaouar. Sur le bêta-développement de 1 dans le corps des séries formelles. (French) [On the beta-development of 1 in the body of formal series]. International Journal of Number Theory (IJNT), 2(3):365–378, September 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042106000619.

Hoshi:2010:FIP

[HM10] Akinari Hoshi and Katsuya Miyake. On the field intersection problem of solvable quintic generic polynomials. International Journal of Number Theory (IJNT), 6(5):1047– 1081, August 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211000337X.

Haukkanen:2012:ANL

[HM12] Pentti Haukkanen and Jorma K.

Merikoski. Asymptotics for numbers of line segments and lines in a square grid. International Journal of Number Theory (IJNT), 8(5):1145– 1152, August 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500698.

Hamieh:2016:NAD

[HM16] Alia Hamieh and M. Ram Murty. A note on q-analogues of Dirichlet L-functions. International Journal of Number Theory (IJNT), 12(3):765– 773, May 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500494.

Han:2016:CWF

[HMST16] Sandie Han, Ariane M. Masuda, Satyanand Singh, and Johann Thiel. The u, v-Calkin–Wilf forest. International Journal of Number Theory (IJNT), 12(5):1311–1328, August 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042116500809.

Hassen:2007:EZF

[HN07] Abdul Hassen and Hieu D. Nguyen. The error zeta function. International Journal of Number Theory (IJNT), 3 (3):439–453, September 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/
doi/10.1142/S1793042107001000.

Hassen:2008:HBP

[HN08] Abdul Hassen and Hieu D. Nguyen. Hypergeometric Bernoulli polynomials and Appell sequences. International Journal of Number Theory (IJNT), 4(5):767–774, October 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001754.

Hassen:2010:HZF

[HN10] Abdul Hassen and Hieu D. Nguyen. Hypergeometric zeta functions. International Journal of Number Theory (IJNT), 6(1):99–126, February 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304211000282X.

Hassen:2011:ZFR

[HN11] Abdul Hassen and Hieu D.

Nguyen. A zero-free region for hypergeometric zeta functions. International Journal of Number Theory (IJNT), 7(4):1033– 1043, June 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004678.

Heim:2021:PRE

[HN21] Bernhard Heim and Markus Neuhauser. Polynomials and reciprocals of Eisenstein series. International Journal of Number Theory (IJNT), 17(02):487– 497, March 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120400199.

Hoelscher:2010:RCG

[Hoe10] Jing Long Hoelscher. Ray class groups of quadratic and cyclotomic fields. International Journal of Number Theory (IJNT), 6(5):1169–1182, August 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042110003447.

Hofmann:2013:BP

[Hof13] Eric Hofmann. Borcherds products for U(1, 1). International Journal of Number Theory (IJNT), 9(7):1801–1820, November 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113500589.

Hoffman:2017:MZV

[Hof17] Michael E. Hoffman. On multiple zeta values of even arguments. International Journal of Number Theory (IJNT), 13 (3):705–716, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S179304211750035X.

Hohn:2011:TGR

[Höh11] Gerald Höhn. On a theorem of Garza regarding algebraic numbers with real conjugates. International Journal of Number Theory (IJNT), 7(4):943– 945, June 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111004320.

Holmes:2019:TPH

[Hol19]  David Holmes.  Torsion

points and height jumping in higher-dimensional families of abelian varieties. International Journal of Number Theory (IJNT), 15(9):1801–1826, October 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211950101X.

Hou:2017:CAM

[Hou17]  Xiaolu Hou.  Construction

of Arakelov-modular lattices over totally definite quaternion algebras. International Journal of Number Theory (IJNT), 13(7):1855–1880, August 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117501020.

Hou:2021:ESC

[Hou21]  Fei Hou.  On exponential

sums for coefficients of general L-functions. International Journal of Number Theory (IJNT), 17(08):1851–1872, September 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042121500664.

Heuberger:2011:PDA

[HP11] Clemens Heuberger and Helmut Prodinger. A precise description of the p-adic valuation of the number of alternating sign matrices. International Journal of Number Theory (IJNT), 7(1):57–69, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111003892.

Habsieger:2011:SCP

[HR11] Laurent Habsieger and Emmanuel Royer. Spiegelungssatz: a combinatorial proof for the 4-rank. International Journal of Number Theory (IJNT), 7(8):2157–2170, December 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111005106.

Haynes:2021:HDG

[HR21] Alan Haynes and Juan J. Ramirez. Higher-dimensional gap theorems for the maximum metric. International Journal of Number Theory (IJNT), 17(07):1665–1670, August 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500548.

Haessig:2011:FSP

[HRL11] C. Douglas Haessig and Antonio Rojas-León. L-functions of symmetric powers of the generalized Airy family of exponential sums. International Journal of Number Theory (IJNT), 7(8):2019–2064, December 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042111005040.

Hambleton:2010:QRR

[HS10] S. Hambleton and V. Scharaschkin. Quadratic reciprocity via resultants. International Journal of Number Theory (IJNT), 6(6):1413–1417, September 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211000354X.

Hartmann:2019:ZDL

[HS19a] L. Hartmann and M. Spreafico.

Zeta determinant of the Laplacian on the real projective spaces. International Journal of Number Theory (IJNT), 15 (2):373–388, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042119500192.

Hirose:2019:HCI

[HS19b] Minoru Hirose and Nobuo Sato.

Hoffman’s conjectural identity. International Journal of Number Theory (IJNT), 15(1):167– 171, February 2019. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042119500052.

Hofmann:2020:CO

[HS20] Tommy Hofmann and Carlo Sircana. On the computation of overorders. International Journal of Number Theory (IJNT), 16(4):857– 879, May 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S179304212050044X.

Hao:2021:NRO

[HS21a] Robert X. J. Hao and Erin Y. Y. Shen. On the number of l-regular overpartitions. International Journal of Number Theory (IJNT), 17(09):2153– 2173, October 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500810.

Hare:2021:CPN

[HS21b] Kevin G. Hare and Nikita Sidorov. Conjugates of Pisot numbers. International Journal of Number Theory (IJNT), 17(06):1307–1321, July 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500378.

Hajdu:2022:TRS

[HS22] L. Hajdu and P. Sebestyén. Terms of recurrence sequences in the solution sets of generalized Pell equations. International Journal of Number Theory (IJNT), 18(07):1605– 1612, August 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500816.

Hsu:2016:TCN

[Hsu16] Catherine Hsu. Two classes of number fields with a nonprincipal Euclidean ideal. International Journal of Number Theory (IJNT), 12(4):1123– 1136, June 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500688.

Hsu:2020:FCO

[Hsu20] Chi-Yun Hsu. Fourier coefficients of the overconvergent generalized eigenform associated to a CM form. International Journal of Number Theory (IJNT), 16(6):1185– 1197, July 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500608.

Huang:2014:IOF

[HSW14] Hsiu-Lien Huang, Chia-Liang Sun, and Julie Tzu-Yueh Wang. Integral orbits over function fields. International Journal of Number Theory (IJNT), 10(8):2187–2204, December 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042114500742.

Hu:2020:TOM

[HSZ20] Qiuxia Hu, Hanfei Song, and Zhizheng Zhang. Third-order mock theta functions. International Journal of Number Theory (IJNT), 16(1):91–106, February 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042120500050.

Hu:2013:PRP

[Hu13a] Jerry Hu. The probability that random positive integers are k-wise relatively prime. International Journal of Number Theory (IJNT), 9(5):1263– 1271, August 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042113500255.

Hu:2013:DID

[Hu13b] Yong Hu. On the distribution of integers with divisors in two consecutive intervals. International Journal of Number Theory (IJNT), 9(4):903– 915, June 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304211350005X.

Hu:2022:RSC

[Hu22]  Liqun Hu.  Representing

squares by cubic forms. International Journal of Number Theory (IJNT), 18(04):839– 848, May 2022. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042122500440.

Huang:2014:CIT

[Hua14] I-Chiau Huang. Convolution identities and their structure. International Journal of Number Theory (IJNT), 10(2):471– 482, March 2014. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042113501042.

Huang:2017:DLP

[Hua17] Zhizhong Huang. Distribution locale des points rationnels de hauteur bornée sur une surface de del Pezzo de degré 6. (French) [Local distribution of rational points of bounded height on a surface of del Pezzo of degree 6]. International Journal of Number Theory (IJNT), 13(7):1895–1930, August 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042117501044.

Huber:2011:DEC

[Hub11a] Tim Huber. Differential equations for cubic theta functions. International Journal of Number Theory (IJNT), 7(7):1945– 1957, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042111004873.

Hubrechts:2011:MEH

[Hub11b]  Hendrik Hubrechts.  Mem-

ory efficient hyperelliptic curve point counting. International Journal of Number Theory (IJNT), 7(1):203–214, February 2011. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042111004034.

Huicochea:2018:PCL

[Hui18] Mario Huicochea. A proof of a conjecture of Lev. International Journal of Number Theory (IJNT), 14(10):2583–2597, November 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501543.

Humphries:2014:MCF

[Hum14] Peter Humphries. On the Mertens conjecture for function fields. International Journal of Number Theory (IJNT), 10(2): 341–361, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/
                                        doi/10.1142/S1793042113500978.

Hung:2018:DAA

[Hun18] Pin-Chi Hung. On the derivative of anticyclotomic p-adic Lfunctions for Hilbert modular forms. International Journal of Number Theory (IJNT), 14 (3):615–630, April 2018. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118500379.

Huong:2018:DE

[Huo18] Pha.m Lan Huong. On the Diophantine equation x^{3} + y^{3} + z^{3} = q. International Journal of Number Theory (IJNT), 14(8):2205–2217, September 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042118501324.

Han:2015:EHD

[HW15] Guo-Niu Han and Wen Wu. Evaluations of the Hankel determinants of a Thue– Morse-like sequence. International Journal of Number Theory (IJNT), 11(6):1887–1904, September 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042115500815.

Hou:2018:CTE

[HW18] Qing-Hu Hou and Yu-Shan Wang. Constant term evaluation and two kinds of congruence. International Journal of Number Theory (IJNT), 14(7):2013–2022, August 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118501208.

Hu:2020:ASD

[HWh20] Yining Hu and Guoniu Weihan. On the automaticity of sequences defined by the Thue– Morse and period-doubling Stieltjes continued fractions. International Journal of Number Theory (IJNT), 16(10):2187– 2212, November 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/
doi/10.1142/S1793042120501134.

Hu:2018:SCF

[HY18] Bingyang Hu and Dongxi Ye. Sign-change of the Fourier coefficients of a Hauptmodul for Γ_{0}(2) . International Journal of Number Theory (IJNT), 14(8):2269–2276, September 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S1793042118501373.

Yan:2020:DEM

[hY20]  Xiao hui Yan.  The Dio-

phantine equation m^{2} + n^{2}x + 2mn^{y} = m + n^{2z}. International Journal of Number Theory (IJNT), 16(08):1701–1708, September 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S179304212050089X.

Hyodo:2015:FPS

[Hyo15] Fumitake Hyodo. A formal power series of a Hecke ring associated with the Heisenberg lie algebra over z_{p}. International Journal of Number Theory (IJNT), 11(8):2305–2323, December 2015. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042115501055.

Hu:2017:NAE

[HYZ17] Hui Hu, Yueli Yu, and Yanfen Zhao. A note on approximation efficiency and partial quotients of Engel continued fractions. International Journal of Number Theory (IJNT), 13(9):2433– 2443, October 2017. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042117501329.

Han:2018:ZSS

[HZ18] Dongchun Han and Hanbin Zhang. On zero-sum subsequences of prescribed length. International Journal of Number Theory (IJNT), 14(1):167– 191, February 2018. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042118500112.

Hong:2012:AVS

[HZZ12] Shaofang Hong, Jianrong Zhao, and Wei Zhao. The 2-adic valuations of Stirling numbers of the second kind. International Journal of Number Theory (IJNT), 8(4):1057– 1066, June 2012. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042112500625.

Interlando:2008:FAG

[IFD08] J. Carmelo Interlando, André Luiz Flores, and Trajano Pires Da Nóbrega Neto. A family of asymptotically good lattices having a lattice in each dimension. International Journal of Number Theory (IJNT), 4(1):147–154, February 2008. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042108001262.

Ishitsuka:2016:SDR

[II16] Yasuhiro Ishitsuka and Tetsushi Ito. On the symmetric determinantal representations of the Fermat curves of prime degree. International Journal of Number Theory (IJNT), 12(4):955– 967, June 2016. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042116500597.

Ishitsuka:2020:ECM

[IIO20] Yasuhiro Ishitsuka, Tetsushi Ito, and Tatsuya Ohshita. Explicit calculation of the mod4 Galois representation associated with the Fermat quartic. International Journal of Number Theory (IJNT), 16(4):881– 905, May 2020. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042120500451.

Illig:2021:NRP

[IJO^{+}21] Peter Illig, Rafe Jones, Eli Orvis, Yukihiko Segawa, and Nick Spinale. Newly reducible polynomial iterates. International Journal of Number Theory (IJNT), 17(06):1405– 1427, July 2021. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042121500433.

Iwaniec:2007:FNH

[IKS07] H. Iwaniec, W. Kohnen, and J. Sengupta. The first negative Hecke eigenvalue. International Journal of Number Theory (IJNT), 3(3):355–363, September 2007. ISSN 1793-0421 (print), 1793-7310 (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S1793042107001024.

Itoh:2013:RTR

[IMO13] Tsuyoshi Itoh, Yasushi Mizusawa, and Manabu Ozaki. On the Z_{p}-ranks of tamely ramified Iwasawa modules. International Journal of Number Theory (IJNT), 9(6):1491–1503, September 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S1793042113500395.

Isikay:2022:YIA

[IN22] Sevcan Işıkay and Ayten PekiṄ.OnY okoi′sinvariantsandtheAnkeny − −Artin − −Chowlaconjecture.InternationalJournalofNumberTheory(IJNT ), 18 (03) : 473−−484, April2022.ISSN1793−

0421(print), 1793−7310(electronic).URLhttpswwwworldscientificcomdoi101142S1793042122500270.

Ichimura:2014:AIL

[INST14] Humio Ichimura, Shoichi Nakajima, and Hiroki Sumida-Takahashi. On the 2-adic Iwasawa lambda invariants of the p-cyclotomic fields and their quadratic twists. International Journal of Number Theory (IJNT), 10 (2):283–296, March 2014. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042113500929.

Isaacson:2021:TGC

[Isa21] Brad Isaacson. A twisted generalization of the classical Dedekind sum. International Journal of Number Theory (IJNT), 17 (02):513–528, March 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120400059.

Ih:2010:FPP

[IT10]  Su-Ion Ih and Thomas J. Tucker.  A

finiteness property for preperiodic points of Chebyshev polynomials. International Journal of Number Theory (IJNT), 6(5):

1011–1025, August 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003356.

Ito:2018:FBG

[Ito18] Kazuhiro Ito. Finiteness of Brauer groups of K3 surfaces in characteristic 2. International Journal of Number Theory (IJNT), 14 (6):1813–1825, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501087.

Ivic:2005:MTS

[Ivi05] Aleksandar Ivić. The Mellin transform of the square of Riemann’s zeta-function. International Journal of Number Theory (IJNT), 1(1):65–73, March 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042105000042.

Ivic:2016:SUB

[Ivi16] Aleksandar Ivić. On some upper bounds for the zeta-function and the Dirichlet divisor problem. International Journal of Number Theory (IJNT), 12(8):2231–2239, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042116501347.

Inam:2016:SNB

[IW16] Ilker Inam and Gabor Wiese. A short note on the Bruiner–Kohnen sign equidistribution conjecture and Halász’ theorem. International Journal of Number Theory (IJNT), 12 (2):357–360, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500214.

Jahangiri:2010:GAQ

[Jah10]  Majid Jahangiri.  Generators of arith-

metic quaternion groups and a Diophantine problem. International Journal of Number Theory (IJNT), 6(6):1311–1328, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042110003551.

Jameson:2012:RRC

[Jam12] Marie Jameson. A refinement of Ramanujan’s congruences modulo powers of 7 and 11. International Journal of Number Theory (IJNT), 8(4):865–879, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500510.

Jameson:2016:PZQ

[Jam16] Marie Jameson. A problem of Zagier on quadratic polynomials and continued fractions. International Journal of Number Theory (IJNT), 12(1):121–141, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500081.

Jarossay:2022:NVC

[Jar22] David Jarossay. Non-vanishing of certain cyclotomic multiple harmonic sums and application to the non-vanishing of certain padic cyclotomic multiple zeta values. International Journal of Number Theory (IJNT), 18(04):849–877, May 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500452.

Jedrzejak:2014:TFC

[Jed14] Tomasz Jedrzejak. On twists of the Fermat cubic x^{3} + y^{3} = 2. International Journal of Number Theory (IJNT), 10(1):

55–72, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500802.

Judrzejak:2022:RFH

[Jed22] Tomasz Jedrzejak. Ranks in the family of hyperelliptic Jacobians of y^{2} = x^{5} + ax II. International Journal of Number Theory (IJNT), 18(04):813–837, May 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500439.

Jenkins:2005:APT

[Jen05] Paul Jenkins. p-adic properties for traces of singular moduli. International Journal of Number Theory (IJNT), 1(1):103– 107, March 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S179304210500011X.

Jiang:2020:VRF

[Jia20] Xing-Wang Jiang. On the values of representation functions III. International Journal of Number Theory (IJNT), 16(3): 511–522, April 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500256.

Jhorar:2016:PBC

[JK16] Bablesh Jhorar and Sudesh K. Khanduja. On power basis of a class of algebraic number fields. International Journal of Number Theory (IJNT), 12(8):2317– 2321, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116501384.

Jo:2021:SAL

[JK21] Yeongseong Jo and M. Krishnamurthy. Stability of Asai local factors for GL(2). International Journal of Number Theory (IJNT), 17(03):529–546, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120400035.

Ji:2016:CPP

[JKK16] Kathy Q. Ji, Byungchan Kim, and Jang Soo Kim. Combinatorial proof of a partial theta function identity of Warnaar. International Journal of Number Theory (IJNT), 12(6): 1475–1482, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500901.

Jaafar:2009:M

[JKM09] Samar Jaafar and Kamal Khuri-Makdisi. On the maps from X(4p) to X(4). International Journal of Number Theory (IJNT), 5 (5):831–844, August 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002390.

Ju:2018:SRP

[JKO18] Jangwon Ju, Kyoungmin Kim, and Byeong-Kweon Oh. Spinor representations of positive definite ternary quadratic forms. International Journal of Number Theory (IJNT), 14(2):581–594, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500367.

Jakhar:2017:CPD

[JKS17] Anuj Jakhar, Sudesh K. Khanduja, and Neeraj Sangwan. Characterization of primes dividing the index of a trinomial. International Journal of Number Theory (IJNT), 13(10): 2505–2514, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501391.

Jiang:2017:SFC

[JL17a] Yujiao Jiang and Guangshi Lü. On sums of Fourier coefficients of Maass cusp forms. International Journal of Number Theory (IJNT), 13(5):1233–1243, June 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211750066X.

Jones:2017:ESR

[JL17b] Rafe Jones and Alon Levy. Eventually stable rational functions. International Journal of Number Theory (IJNT), 13(9): 2299–2318, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117501263.

Jang:2018:IOE

[JL18a] Min-Joo Jang and Jeremy Lovejoy. Identities for overpartitions with even smallest parts. International Journal of Number Theory (IJNT), 14(7):2023–2033, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211850121X.

Jiang:2018:CSS

[JL18b] Xing-Wang Jiang and Ya-Li Li. On the cardinality of subsequence sums. International Journal of Number Theory (IJNT), 14(3):661–668, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500410.

Jiang:2020:APV

[JL20] Yujiao Jiang and Guangshi Lü. On an analogue of prime vectors among integer lattice points in ellipsoids for automorphic forms. International Journal of Number Theory (IJNT), 16(1):145–160, February 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500086.

Jing:2022:CSF

[JL22] Mengyao Jing and Huaning Liu. Consecutive square-free numbers and squarefree primitive roots. International Journal of Number Theory (IJNT), 18(01): 205–226, February 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500154.

Just:2021:FCP

[JLl21] Matthew Just and Noah Lebowitz-lockard. On factorizations into coprime parts. International Journal of Number Theory (IJNT), 17(08):1825–1850, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042121500652.

Johnson-Leung:2014:TPV

[JLR14] Jennifer Johnson-Leung and Brooks Roberts. Twisting of paramodular vectors. International Journal of Number Theory (IJNT), 10 (4):1043–1065, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500146.

Johnson-Leung:2017:TSP

[JLR17] Jennifer Johnson-Leung and Brooks Roberts. Twisting of Siegel paramodular forms. International Journal of Number Theory (IJNT), 13(7):1755–1854, August 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117501019.

Jacobson:2015:CAC

[JLSW15] M. J. Jacobson, Jr., Y. Lee, R. Scheidler, and H. C. Williams. Construction of all cubic function fields of a given square-free discriminant. International Journal of Number Theory (IJNT), 11(6):1839– 1885, September 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500803.

Jouhet:2010:IEI

[JM10] Frédéric Jouhet and Elie Mosaki. Irrationalité aux entiers impairs positifs d’un q-analogue de la fonction zêta de Riemann. (French) [Irrationality to positive odd integers of a qanalogue of the Riemann zeta function]. International Journal of Number Theory (IJNT), 6 (5):959–988, August 2010. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003332.

Joshi:2011:IHC

[JM11] Kirti Joshi and Cameron Mcleman. Infinite Hilbert class field towers from Galois representations. International Journal of Number Theory (IJNT), 7(1):1–8, February 2011. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111003879.

Jagathesan:2016:TNE

[JM16] T. Jagathesan and M. Manickam. Theory of newforms for Eisenstein series of half-integral weight. International Journal of Number Theory (IJNT), 12(3):725–735, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500470.

Jiu:2016:SAS

[JMV16] Lin Jiu, Victor H. Moll, and Christophe Vignat. A symbolic approach to some identities for Bernoulli–Barnes polynomials. International Journal of Number Theory (IJNT), 12(3):649–662, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500421.

Ju:2020:UMS

[JO20] Jangwon Ju and Byeong-Kweon Oh. Universal mixed sums of generalized 4- and 8-gonal numbers. International Journal of Number Theory (IJNT), 16(3):603–627, April 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304212050030X.

Jones:2009:PVT

[Jon09] Lenny Jones. Polynomial variations on a theme of Sierpiński. International Journal of Number Theory (IJNT), 5(6):999– 1015, September 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002511.

Jones:2021:LDD

[Jon21] Edna Jones. Local densities of diagonal integral ternary quadratic forms at odd primes. International Journal of Number Theory (IJNT), 17(03):547–575, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120400357.

Ju:2019:TUS

[JOS19] Jangwon Ju, Byeong-Kweon Oh, and Bangnam Seo. Ternary universal sums of generalized polygonal numbers. International Journal of Number Theory (IJNT), 15(4):655–675, May 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500350.

Jha:2014:AFE

[JP14] Somnath Jha and Aprameyo Pal. Algebraic functional equation for Hida family. International Journal of Number Theory (IJNT), 10(7):1649–1674, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042114500493.

Jones:2021:IVD

[JQ21] Gareth Jones and Shi Qiu. Integer-valued definable functions in R_{an,exp}. International Journal of Number Theory (IJNT), 17(08): 1739–1752, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500597.

Jones:2013:SCF

[JR13] Marvin Jones and Jeremy Rouse. Solutions of the cubic Fermat equation in quadratic fields. International Journal of Number Theory (IJNT), 9(6):1579–1591, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042113500449.

Jabuka:2011:WTD

[JRW11] Stanislav Jabuka, Sinai Robins, and Xinli Wang. When are two Dedekind sums equal? International Journal of Number Theory (IJNT), 7(8):2197–2202, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042111005088.

Jennings-Shaffer:2015:RCM

[JS15] Chris Jennings-Shaffer. Rank and crank moments for partitions without repeated odd parts. International Journal of Number Theory (IJNT), 11(3):683–703, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042115500372.

Jha:2019:RCB

[JS19] Abhash Kumar Jha and Brundaban Sahu. Rankin–Cohen brackets on Jacobi forms of several variables and special values of certain Dirichlet series. International Journal of Number Theory (IJNT), 15(5):925– 933, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119500490.

Jeon:2020:TRE

[JS20] Daeyeol Jeon and Andreas Schweizer. Torsion of rational elliptic curves over different types of cubic fields. International Journal of Number Theory (IJNT), 16(6): 1307–1323, July 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500682.

Jain-sharma:2021:EBB

[JsKL21] Niraek Jain-sharma, Tanmay Khale, and Mengzhen Liu. Explicit Burgess bound for composite moduli. International Journal of Number Theory (IJNT), 17(10):2207– 2219, November 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500834.

Jennings-Shaffer:2014:NTZ

[JSS14] Chris Jennings-Shaffer and Holly Swisher. A note on the transcendence of zeros of a certain family of weakly holomorphic forms. International Journal of Number Theory (IJNT), 10(2):309–317, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500942.

Jameson:2016:BLC

[JTY16] Marie Jameson, Jesse Thorner, and Lynnelle Ye. Benford’s Law for coefficients of newforms. International Journal of Number Theory (IJNT), 12(2):483–494, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042116500299.

Jung:2014:NMV

[Jun14] Hwanyup Jung. A note on the mean value of L(1, χ) in the hyperelliptic ensemble. International Journal of Number Theory (IJNT), 10(4):859–874, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500031.

Jarden:2008:UFR

[JV08] Moshe Jarden and Carlos R. Videla. Undecidability of families of rings of totally real integers. International Journal of Number Theory (IJNT), 4(5):835–850, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042108001705.

Jha:2021:ESC

[JV21] Abhash Kumar Jha and Lalit Vaishya. Estimates of shifted convolution sums involving Fourier coefficients of Hecke–Maass eigenform. International Journal of Number Theory (IJNT), 17(07):1631–1643, August 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042121500524.

Jiu:2020:ACM

[JVW20] Lin Jiu, Christophe Vignat, and Tanay Wakhare. Analytic continuation for multiple zeta values using symbolic representations. International Journal of Number Theory (IJNT), 16(3):579–602, April 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042120500293.

Jones:2017:PCN

[JW17]  Lenny Jones and Daniel White.  On

primitive covering numbers. International Journal of Number Theory (IJNT), 13(1): 27–37, February 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500038.

Jadrijevic:2006:SRP

[JZ06] Borka Jadrijević and Volker Ziegler. A system of relative Pellian equations and a related family of relative Thue equations. International Journal of Number Theory (IJNT), 2 (4):569–590, December 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042106000735.

Jin:2017:RTC

[JZ17] Hai-Tao Jin and Li Zhang. Ramanujantype congruences for ℓ-regular partitions modulo 3, 5, 11 and 13. International Journal of Number Theory (IJNT), 13(8):1995– 2006, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211750107X.

Kaavya:2011:CPP

[Kaa11] S. J. Kaavya. Crank 0 partitions and the parity of the partition function. International Journal of Number Theory (IJNT), 7(3):793–801, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004381.

Kable:2010:AIO

[Kab10] Anthony C. Kable. An arithmetical invariant of orbits of affine actions and its application to similarity classes of quadratic spaces. International Journal of Number Theory (IJNT), 6(6):1215–1253, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042110003460.

Kadiri:2012:EZF

[Kad12] Habiba Kadiri. Explicit zero-free regions for Dedekind zeta functions. International Journal of Number Theory (IJNT), 8(1): 125–147, February 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500078.

Kalita:2018:VGH

[Kal18] Gautam Kalita. Values of Gaussian hypergeometric series and their connections to algebraic curves. International Journal of Number Theory (IJNT), 14(1):1–18, February 2018. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211850001X.

Kamano:2008:ABN

[Kam08]  Ken Kamano.  p-adic q-Bernoulli num-

bers and their denominators. International Journal of Number Theory (IJNT), 4(6): 911–925, December 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304210800181X.

Kane:2010:RIT

[Kan10]  Ben Kane.  Representations of integers

by ternary quadratic forms. International Journal of Number Theory (IJNT), 6(1): 127–158, February 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110002831.

Kane:2013:ANS

[Kan13] Daniel M. Kane. An asymptotic for the number of solutions to linear equations in prime numbers from specified Chebotarev classes. International Journal of Number Theory (IJNT), 9(4):1073–1111, June 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042113500139.

Kani:2014:BTS

[Kan14] Ernst Kani. Binary theta series and modular forms with complex multiplication. International Journal of Number Theory (IJNT), 10 (4):1025–1042, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500134.

Kandhil:2022:NDZ

[Kan22] Neelam Kandhil. A note on Dedekind zeta values at 1/2. International Journal of Number Theory (IJNT), 18(06): 1289–1299, July 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500658.

Karhadkar:2019:PPF

[Kar19] Kedar Karhadkar. Parity of the partition function p(n, k). International Journal of Number Theory (IJNT), 15(4):799– 805, May 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119500428.

Kashio:2019:AMA

[Kas19] Tomokazu Kashio. p-adic measures associated with zeta values and p-adic log multiple gamma functions. International Journal of Number Theory (IJNT), 15(5): 991–1007, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119500532.

Katayama:2010:GGF

[Kat10] Koji Katayama. Generalized gamma functions with characters. International Journal of Number Theory (IJNT), 6(7):1625– 1657, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003629.

Kataoka:2017:GGC

[Kat17] Takenori Kataoka. On Greenberg’s generalized conjecture for complex cubic fields. International Journal of Number Theory (IJNT), 13(3):619–631, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500312.

Katen:2021:EWP

[Kat21] Jeffrey Katen. Explicit Weil-pairing for Drinfeld modules. International Journal of Number Theory (IJNT), 17(09):2131– 2152, October 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500809.

Kazalicki:2008:LRC

[Kaz08] Matija Kazalicki. Linear relations for coefficients of Drinfeld modular forms. International Journal of Number Theory (IJNT), 4(2):171–176, April 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042108001274.

Kokluce:2017:ESC

[KE17] Bülent Köklüce and Hasan Eser. Evaluation of some convolution sums and representation of integers by certain quadratic forms in 12 variables. International Journal of Number Theory (IJNT), 13(3):775–799, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117500415.

Kiuchi:2018:SWA

[KE18] Isao Kiuchi and Sumaia Saad Eddin. Sums of weighted averages of gcd-sum functions. International Journal of Number Theory (IJNT), 14(10):2699–2728, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118501634.

Kedlaya:2005:LMA

[Ked05] Kiran S. Kedlaya. Local monodromy of p-adic differential equations: an overview. International Journal of Number Theory (IJNT), 1(1):109–154, March 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304210500008X.

Keil:2013:MEE

[Kei13] Eugen Keil. Moment estimates for exponential sums over k-free numbers. International Journal of Number Theory (IJNT), 9(3):607–619, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501515.

Keith:2017:PSN

[Kei17] William J. Keith. Partitions into a small number of part sizes. International Journal of Number Theory (IJNT), 13(1):229– 241, February 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500130.

Keith:2021:RKC

[Kei21]  William J. Keith.  Restricted k-color

partitions, II. International Journal of Number Theory (IJNT), 17(03):591–601, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042120400151.

Kelmer:2010:DTK

[Kel10] Dubi Kelmer. Distribution of twisted Kloosterman sums modulo prime powers. International Journal of Number Theory (IJNT), 6 (2):271–280, March 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110002910.

Kelmer:2017:LPC

[Kel17] Dubi Kelmer. Lattice points in a circle for generic unimodular shears. International Journal of Number Theory (IJNT), 13 (2):291–300, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211750018X.

Kim:2011:AVM

[KH11] Min-Soo Kim and Su Hu. A p-adic view of multiple sums of powers. International Journal of Number Theory (IJNT), 7(8): 2273–2288, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111005027.

Khanduja:2019:DCA

[Kha19] Sudesh K. Khanduja. The discriminant of compositum of algebraic number fields. International Journal of Number Theory (IJNT), 15(2):353–360, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119500167.

Kidwell:2016:UCL

[Kid16] Keenan Kidwell. Unitary completions of locally algebraic and locally analytic principal series of p-adic GL_{2}. International Journal of Number Theory (IJNT), 12(7): 1765–1781, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116501086.

Kim:2010:CPC

[Kim10a] Byungchan Kim. Combinatorial proofs of certain identities involving partial theta functions. International Journal of Number Theory (IJNT), 6(2):449–460, March 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003046.

Kim:2010:BPQ

[Kim10b] Sun Kim. A bijective proof of the quintuple product identity. International Journal of Number Theory (IJNT), 6(2):247– 256, March 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042110002909.

Kim:2016:AFI

[Kim16a] Sungjin Kim. Average of the first invariant factor of the reductions of abelian varieties of CM type. International Journal of Number Theory (IJNT), 12(2):445–463, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500263.

Kim:2016:OMA

[Kim16b] Sungjin Kim. On the order of a modulo n, on average. International Journal of Number Theory (IJNT), 12(8):2073– 2080, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116501244.

Kim:2017:LBD

[Kim17a] Kwang-Seob Kim. Lower bounds for discriminants of polynomials. International Journal of Number Theory (IJNT), 13(10): 2769–2776, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117501524.

Kim:2017:ITD

[Kim17b] Sun Kim. Identities for trigonometric divisor sums. International Journal of Number Theory (IJNT), 13(4):957–975, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117500506.

Kim:2021:CPR

[Kim21a] Sun Kim. A combinatorial proof of a recurrence relation for the sum of divisors function. International Journal of Number Theory (IJNT), 17(03):647–653, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042120400084.

Kim:2021:EVV

[Kim21b] Sungjin Kim. On the equations φ(n) = φ(n+k) and φ(p−1) = φ(q−1). International Journal of Number Theory (IJNT), 17(06): 1287–1305, July 2021. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042121500366.

Kim:2022:HEC

[Kim22a] Jiseong Kim. On Hecke eigenvalues of cusp forms in almost all short intervals. International Journal of Number Theory (IJNT), 18 (04):729–742, May 2022. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042122500385.

Kim:2022:NRS

[Kim22b] Kyoungmin Kim. The number of representations of squares by integral quaternary quadratic forms. International Journal of Number Theory (IJNT), 18(04):757– 775, May 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042122500403.

Kitajima:2013:OEG

[Kit13] Takahiro Kitajima. On the orders of even K-groups of rings of integers in cyclotomic Z_{p}-extensions of Q. International Journal of Number Theory (IJNT), 9(7):1713– 1724, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500528.

Khanduja:2008:TD

[KK08] Sudesh K. Khanduja and Munish Kumar. On a theorem of Dedekind. International Journal of Number Theory (IJNT), 4(6): 1019–1025, December 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001833.

Kang:2010:APT

[KK10] Soon-Yi Kang and Chang Heon Kim. Arithmetic properties of traces of singular moduli on congruence subgroups. International Journal of Number Theory (IJNT), 6(8): 1755–1768, December 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003757.

Konstantinou:2012:RID

[KK12] Elisavet Konstantinou and Aristides Kontogeorgis. Ramanujan invariants for discriminants congruent to 5(mod24). International Journal of Number Theory (IJNT), 8 (1):265–287, February 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500157.

Keskin:2015:GFL

[KK15] Refık Keskin and Olcay Karaatli. Generalized Fibonacci and Lucas numbers of the form 5x^{2}. International Journal of Number Theory (IJNT), 11(3):931–944, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115500517.

Kaushik:2018:GBN

[KK18] Surjeet Kaushik and Narasimha Kumar. On the gaps between nonzero Fourier coefficients of eigenforms with CM. International Journal of Number Theory (IJNT), 14 (1):95–101, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500070.

Kim:2019:ACD

[KK19] Byungchan Kim and Eunmi Kim. Asymptotics of certain double series in Eulerian form. International Journal of Number Theory (IJNT), 15(5):1009–1021, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119500544.

Kalmynin:2020:OEU

[KK20] A. B. Kalmynin and P. R. Kosenko. Orthorecursive expansion of unity. International Journal of Number Theory (IJNT), 16 (6):1209–1226, July 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500621.

Kim:2016:NPP

[KKK16] Kwang Yon Kim, Ryul Kim, and Jin Song Kim. On the number of permutation polynomials over a finite field. International Journal of Number Theory (IJNT), 12(6): 1519–1528, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500949.

Kim:2021:WOR

[KKL21] Byungchan Kim, Eunmi Kim, and Jeremy Lovejoy. On weighted overpartitions related to some q-series in Ramanujan’s lost notebook. International Journal of Number Theory (IJNT), 17(03):603–619, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120400345.

Kuo:2009:GLD

[KL09] Wentang Kuo and Yu-Ru Liu. Gaussian laws on Drinfeld modules. International Journal of Number Theory (IJNT), 5(7): 1179–1203, November 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002638.

Knopfmacher:2011:PPN

[KL11] Arnold Knopfmacher and Florian Luca. On prime-perfect numbers. International Journal of Number Theory (IJNT), 7(7):1705– 1716, November 2011. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042111004447.

Kim:2012:FGK

[KL12] Henry H. Kim and Kyu-Hwan Lee. A family of generalized Kac–Moody algebras and deformation of modular forms. International Journal of Number Theory (IJNT), 8(5):1107– 1131, August 2012. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112500674.

Kim:2014:RUS

[KL14] Byungchan Kim and Jeremy Lovejoy. The rank of a unimodal sequence and a partial theta identity of Ramanujan. International Journal of Number Theory (IJNT), 10 (4):1081–1098, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211450016X.

Komatsu:2017:GIP

[KL17] Takao Komatsu and Florian Luca. Generalized incomplete poly-Bernoulli polynomials and generalized incomplete poly-Cauchy polynomials. International Journal of Number Theory (IJNT), 13(2):371– 391, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042117500221.

Kleine:2017:LBI

[Kle17]  Sören Kleine.  Local behavior of Iwa-

sawa’s invariants. International Journal of Number Theory (IJNT), 13(4):1013– 1036, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042117500543.

Kim:2021:STN

[KLkO21] Daejun Kim, Jeongwon Lee, and Byeong kweon Oh. A sum of three nonzero triangular numbers. International Journal of Number Theory (IJNT), 17(10):2279– 2300, November 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500883.

Klosin:2013:ILD

[Klo13] Krzysztof Klosin. On Ihara’s lemma for degree one and two cohomology over imaginary quadratic fields. International Journal of Number Theory (IJNT), 9(6):1541– 1561, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500425.

Kozek:2015:HP

[KLPP15] Mark Kozek, Florian Luca, Paul Pollack, and Carl Pomerance. Harmonious pairs. International Journal of Number Theory (IJNT), 11 (5):1633–1651, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115400151.

Knopp:2009:ECG

[KLR09] Marvin Knopp, Joseph Lehner, and Wissam Raji. Eichler cohomology for generalized modular forms. International Journal of Number Theory (IJNT), 5(6):1049– 1059, September 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002547.

Kluners:2012:DNF

[Klü12] Jürgen Klüners. The distribution of number fields with wreath products as Galois groups. International Journal of Number Theory (IJNT), 8(3):845–858, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112500492.

Klurman:2016:RPN

[Klu16] Oleksiy Klurman. Radical of perfect numbers and perfect numbers among polynomial values. International Journal of Number Theory (IJNT), 12(3):585–591, May 2016. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500378.

Knopfmacher:2005:SFC

[KM05] A. Knopfmacher and M. E. Mays. A survey of factorization counting functions. International Journal of Number Theory (IJNT), 1 (4):563–581, December 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042105000315.

Kohnen:2006:TSA

[KM06] Winfried Kohnen and Riccardo Salvati Manni. On the theta series attached to D_{m}^{+}-lattices. International Journal of Number Theory (IJNT), 2(1):1–5, March 2006. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042106000449.

Knopp:2009:PGM

[KM09] Marvin Knopp and Geoffrey Mason. Parabolic generalized modular forms and their characters. International Journal of Number Theory (IJNT), 5(5):845–857, August 2009. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042109002419. See revi-

sions [KM12c].

Khuri-Makdisi:2012:MIE

[KM12a] Kamal Khuri-Makdisi. Moduli interpretation of Eisenstein series. International Journal of Number Theory (IJNT), 8(3):

715–748, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500418.

Kinjo:2012:HSA

[KM12b] Kensaku Kinjo and Yuken Miyasaka. Hypergeometric series and arithmetic–geometric mean over 2-adic fields. International Journal of Number Theory (IJNT), 8(3):831– 844, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112500480.

Knopp:2012:RPG

[KM12c] Marvin Knopp and Geoffrey Mason. Revisions to “Parabolic Generalized Modular Forms and Their Characters”. International Journal of Number Theory (IJNT), 8(4):859– 864, June 2012. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500509. See [KM09].

Kohnen:2014:SCF

[KM14] Winfried Kohnen and Yves Martin. Sign changes of Fourier coefficients of cusp forms supported on prime power indices. International Journal of Number Theory (IJNT), 10(8):1921–1927, December 2014. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500626.

Khuri-Makdisi:2018:UBS

[KM18a] Kamal Khuri-Makdisi. Upper bounds for some Brill–Noether loci over a finite field. International Journal of Number Theory (IJNT), 14(3):739–749, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500471.

Kumari:2018:SNV

[KM18b] Moni Kumari and M. Ram Murty. Simultaneous non-vanishing and sign changes of Fourier coefficients of modular forms. International Journal of Number Theory (IJNT), 14(8): 2291–2301, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118501397.

Kalle:2022:ARL

[KM22] Charlene Kalle and Marta Maggioni. On approximation by random Lüroth expansions. International Journal of Number Theory (IJNT), 18(05):1013–1046, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042122500531.

Koh:2020:ESE

[KMPS20] Doowon Koh, Mozhgan Mirzaei, Thang Pham, and Chun-Yen Shen. Exponential sum estimates over prime fields. International Journal of Number Theory (IJNT), 16(2): 291–308, March 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500153.

Karemaker:2021:CFF

[KMS21] Valentijn Karemaker, Sophie Marques, and Jeroen Sijsling. Cubic function fields with prescribed ramification. International Journal of Number Theory (IJNT), 17(09): 2019–2053, October 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500755.

Komori:2011:FED

[KMT11] Yasushi Komori, Kohji Matsumoto, and Hirofumi Tsumura. Functional equations for double L-functions and values at nonpositive integers. International Journal of Number Theory (IJNT), 7(6):1441–1461, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042111004551.

Kumar:2020:SRF

[KMV20] Balesh Kumar, Jay Mehta, and G. K. Viswanadham. Some remarks on the Fourier coefficients of cusp forms. International Journal of Number Theory (IJNT), 16(09): 1935–1943, October 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500992.

Kleinbock:2010:MDA

[KMW10] Dmitry Kleinbock, Gregory Margulis, and Junbo Wang. Metric Diophantine approximation for systems of linear forms via dynamics. International Journal of Number Theory (IJNT), 6(5):1139–1168, August 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003423.

Katsurada:2009:DAA

[KN09a] Masanori Katsurada and Takumi Noda. Differential actions on the asymptotic expansions of non-holomorphic Eisenstein series. International Journal of Number Theory (IJNT), 5(6):1061–1088, September 2009. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042109002559.

Konyagin:2009:SPC

[KN09b] Sergei V. Konyagin and Melvyn B. Nathanson. Sums of products of congruence classes and of arithmetic progressions. International Journal of Number Theory (IJNT), 5(4):625–634, June 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002286.

Kirschmer:2019:QQL

[KN19] Markus Kirschmer and Gabriele Nebe. Quaternary quadratic lattices over number fields. International Journal of Number Theory (IJNT), 15(2):309–325, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119500131.

Knafo:2008:ELB

[Kna08] Emmanuel Knafo. Effective lower bound for the variance of distribution of primes in arithmetic progressions. International Journal of Number Theory (IJNT), 4(1): 45–56, February 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042108001213.

Kaneko:2010:KDM

[KO10] Masanobu Kaneko and Yasuo Ohno. On a kind of duality of multiple zeta-star values. International Journal of Number Theory (IJNT), 6(8):1927–1932, December 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211000385X.

Kara:2020:AGF

[KO20] Yasemin Kara and Ekin Ozman. Asymptotic generalized Fermat’s last theorem over number fields. International Journal of Number Theory (IJNT), 16(5):907–924, June 2020. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500463.

Kobayashi:2014:NSD

[Kob14] Mitsuo Kobayashi. A new series for the density of abundant numbers. International Journal of Number Theory (IJNT), 10 (1):73–84, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500814.

Kobayashi:2016:GSD

[Kob16] Mitsuo Kobayashi. A generalization of a series for the density of abundant numbers. International Journal of Number Theory (IJNT), 12(3):671–677, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500445.

Kochubei:2008:DCE

[Koc08] Anatoly N. Kochubei. Dwork–Carlitz exponential and overconvergence for additive functions in positive characteristic. International Journal of Number Theory (IJNT), 4(3):453–460, June 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001444.

Kohl:2008:CCT

[Koh08] Stefan Kohl. On conjugates of Collatztype mappings. International Journal of Number Theory (IJNT), 4(1):117–120, February 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042108001237.

Kohnen:2010:SNF

[Koh10]  Winfried Kohnen.  A short note on

Fourier coefficients of half-integral weight modular forms. International Journal of Number Theory (IJNT), 6(6):1255–1259, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042110003484.

Kokluce:2013:RNC

[Kök13a] Bülent Köklüce. The representation numbers of certain octonary quadratic forms. International Journal of Number Theory (IJNT), 9 (5):1125–1139, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500164.

Kokluce:2013:RNTa

[Kök13b] Bülent Köklüce. The representation numbers of three octonary quadratic forms. International Journal of Number Theory (IJNT), 9 (2):505–516, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112501461.

Kokluce:2013:RNTc

[Kök13c] Bülent Köklüce. Representation numbers of two octonary quadratic forms. International Journal of Number Theory (IJNT), 9(7):1641– 1648, November 2013. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042113500486.

Kolitsch:2015:AAT

[Kol15] Louis W. Kolitsch. Another approach to the Truncated Pentagonal Number Theorem. International Journal of Number Theory (IJNT), 11(5):1563–1569, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115400084.

Kominers:2009:CEE

[Kom09] Scott Duke Kominers. Configurations of extremal even unimodular lattices. International Journal of Number Theory (IJNT), 5(3):457–464, May 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304210900216X.

Komatsu:2013:HCN

[Kom13] Takao Komatsu. Hypergeometric Cauchy numbers. International Journal of Number Theory (IJNT), 9(3):545–560, May 2013. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501473.

Komatsu:2017:IFP

[Kom17] Toru Komatsu. An infinite family of pairs of imaginary quadratic fields with ideal classes of a given order. International Journal of Number Theory (IJNT), 13(2):253–260, March 2017. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500154.

Konig:2020:DDE

[Kön20] Joachim König. On the mod-p distribution of discriminants of G-extensions. International Journal of Number Theory (IJNT), 16(4):767–785, May 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500396.

Kopeliovich:2008:TCI

[Kop08] Yaacov Kopeliovich. Theta constant identities at periods of coverings of degree 3. International Journal of Number Theory (IJNT), 4 (5):725–733, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001663.

Koukoulopoulos:2015:PSA

[Kou15] Dimitris Koukoulopoulos. Primes in short arithmetic progressions. International Journal of Number Theory (IJNT), 11(5):1499– 1521, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042115400035.

Kowalski:2006:RQT

[Kow06] E. Kowalski. On the rank of quadratic twists of elliptic curves over function fields. International Journal of Number Theory (IJNT), 2(2):267–288, June 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042106000528.

Koymans:2022:GCE

[Koy22] Peter Koymans. The generalized Catalan equation in positive characteristic. International Journal of Number Theory (IJNT), 18 (02):269–276, March 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304212250018X.

Kozuma:2009:ECR

[Koz09] Rintaro Kozuma. Elliptic curves related to cyclic cubic extensions. International Journal of Number Theory (IJNT), 5(4): 591–623, June 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002304.

Kaczorowski:2014:SLE

[KP14] J. Kaczorowski and A. Perelli. Solving linear equations in L-functions. International Journal of Number Theory (IJNT), 10(3):569–584, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113501091.

Kilic:2016:ESI

[KP16] Emrah Kiliç and Helmut Prodinger. Evaluation of sums involving Gaussian q-binomial coefficients with rational weight functions. International Journal of Number Theory (IJNT), 12(2):495–504, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500305.

Koutsianas:2018:PPS

[KP18] Angelos Koutsianas and Vandita Patel. Perfect powers that are sums of squares in a three term arithmetic progression. International Journal of Number Theory (IJNT), 14(10): 2729–2735, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501646.

Kuba:2019:NHN

[KP19] Markus Kuba and Alois Panholzer. A note on harmonic number identities, Stirling series and multiple zeta values. International Journal of Number Theory (IJNT), 15(7):1323–1348, August 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211950074X.

Kirschenhofer:2008:FTT

[KPT08] P. Kirschenhofer, A. Pethő, and J. M. Thuswaldner. On a family of three term nonlinear integer recurrences. International Journal of Number Theory (IJNT), 4(1): 135–146, February 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001250.

Kim:2017:FAS

[KPW17] Chan-Ho Kim, Robert Pollack, and Tom Weston. On the freeness of anticyclotomic Selmer groups of modular forms. International Journal of Number Theory (IJNT), 13 (6):1443–1455, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500804.

Knopp:2010:ECG

[KR10] Marvin Knopp and Wissam Raji. Eichler cohomology for generalized modular forms II. International Journal of Number Theory (IJNT), 6(5):1083–1090, August 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211000340X.

Kuhn:2016:EFG

[KR16] Patrick Kühn and Nicolas Robles. Explicit formulas of a generalized Ramanujan sum. International Journal of Number Theory (IJNT), 12(2):383–408, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500238.

Kraus:2007:CES

[Kra07] Alain Kraus. Courbes elliptiques semi-stables sur les corps de nombres. (French) [Semi-stable elliptical curves on number fields]. International Journal of Number Theory (IJNT), 3 (4):611–633, December 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042107001127.

Krauel:2014:JTS

[Kra14] Matthew Krauel. A Jacobi theta series and its transformation laws. International Journal of Number Theory (IJNT), 10(6):1343– 1354, September 2014. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042114500316.

Kraus:2015:EFN

[Kra15] Alain Kraus. Équation de Fermat et nombres premiers inertes. (French) [Fermat’s equation and inert prime numbers]. International Journal of Number Theory (IJNT), 11(8): 2341–2351, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115501079.

Kreso:2017:DES

[Kre17] Dijana Kreso. Diophantine equations in separated variables and lacunary polynomials. International Journal of Number Theory (IJNT), 13(8):2055–2074, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S179304211750110X.

Krieg:2010:TSH

[Kri10] Aloys Krieg. Theta series over the Hurwitz quaternions. International Journal of Number Theory (IJNT), 6(1):25–36, February 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042110002788.

Krishnamoorthy:2016:NFC

[Kri16] Srilakshmi Krishnamoorthy. A note on the Fourier coefficients of a Cohen–Eisenstein series. International Journal of Number Theory (IJNT), 12(5):1149–1161, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500706.

Katz:2020:RLS

[KRLT20] Nicholas M. Katz, Antonio Rojas-León, and Pham Huu Tiep. Rigid local systems with monodromy group the Conway group Co_{2}. International Journal of Number Theory (IJNT), 16(2):341–360, March 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500189.

Kida:2010:CBQ

[KRS10] Masanari Kida, Yūichi Rikuna, and Atsushi Sato. Classifying Brumer’s quintic polynomials by weak Mordell–Weil groups. International Journal of Number Theory (IJNT), 6(3):691–704, May 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003162.

Konig:2018:GRI

[KRS18] Joachim König, Daniel Rabayev, and Jack Sonn. Galois realizations with inertia groups of order two. International Journal of Number Theory (IJNT), 14(7):1983– 1994, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S179304211850118X.

Krumm:2016:LGP

[Kru16] David Krumm. A local–global principle in the dynamics of quadratic polynomials. International Journal of Number Theory (IJNT), 12(8):2265–2297, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042116501360.

Kida:2009:QPH

[KRY09] Masanari Kida, Guénaël Renault, and Kazuhiro Yokoyama. Quintic polynomials of Hashimoto–Tsunogai, Brumer and Kummer. International Journal of Number Theory (IJNT), 5(4):555–571, June 2009. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002250.

Kasten:2013:CVR

[KS13] Hendrik Kasten and Claus-Günther Schmidt. The critical values of Rankin–Selberg convolutions. International Journal of Number Theory (IJNT), 9(1):205–256, February 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112501345.

Komatsu:2014:BBC

[KS14] Takao Komatsu and László Szalay. Balancing with binomial coefficients. International Journal of Number Theory (IJNT), 10(7): 1729–1742, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500523.

Kiss:2016:MVA

[KS16] Sándor Z. Kiss and Csaba Sándor. On the maximum values of the additive representation functions. International Journal of Number Theory (IJNT), 12(4):1055– 1075, June 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042116500652.

Kettlestrings:2015:CPG

[KT15] David Kettlestrings and Jeffrey Lin Thunder. Counting points of given height that generate a quadratic extension of a function field. International Journal of Number Theory (IJNT), 11(2):569–592, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211550030X.

Kolster:2020:ACS

[KT20] Manfred Kolster and Reza Taleb. The p-adic Coates–Sinnott Conjecture over maximal orders. International Journal of Number Theory (IJNT), 16(6):1227–1246, July 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500633.

Kanemitsu:2006:SNT

[KTT06] Shigeru Kanemitsu, Yoshio Tanigawa, and Haruo Tsukada. Some number theoretic applications of a general modular relation. International Journal of Number Theory (IJNT), 2 (4):599–615, December 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042106000759.

Koutsaki:2016:ZLC

[KTZ16] K. Paolina Koutsaki, Albert Tamazyan, and Alexandru Zaharescu. On the zeros of linear combinations of derivatives of the Riemann zeta function. International Journal of Number Theory (IJNT), 12(6):1703– 1723, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116501049.

Koutsaki:2018:ZLC

[KTZ18] K. Paolina Koutsaki, Albert Tamazyan, and Alexandru Zaharescu. On the zeros of linear combinations of derivatives of the Riemann zeta function, II. International Journal of Number Theory (IJNT), 14(2): 371–382, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500252.

Kuan:2015:FFA

[Kua15] Yen-Liang Kuan. A function field analogue of Romanoff’s theorem. International Journal of Number Theory (IJNT), 11(7): 2161–2173, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500955.

Kuga:2022:LRS

[Kug22] Seiji Kuga. On linear relations for special L-values over certain totally real number fields. International Journal of Number Theory (IJNT), 18(02):227–244, March 2022. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500166.

Kuhnlein:2012:WRS

[Küh12] Stefan Kühnlein. Well-rounded sublattices. International Journal of Number Theory (IJNT), 8(5):1133–1144, August 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112500686.

Kumchev:2009:BAE

[Kum09] Angel V. Kumchev. A binary additive equation involving fractional powers. International Journal of Number Theory (IJNT), 5 (2):281–292, March 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002092.

Kumar:2013:ACT

[Kum13] Narasimha Kumar. A 2-adic control theorem for modular curves. International Journal of Number Theory (IJNT), 9(6):1605– 1618, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500462.

Kumakawa:2021:ILI

[Kum21] Naoki Kumakawa. On the Iwasawa λinvariant of the cyclotomic Z_{2}-extension of Q(\sqrt{pq}) and the 2-part of the class number of Q(\sqrt{pq}, \sqrt{2 + \sqrt{2}}). International Journal of Number Theory (IJNT), 17 (04):931–958, May 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500160.

Kuo:2009:GST

[Kuo09] Wentang Kuo. A generalization of the Sato–Tate Conjecture. International Journal of Number Theory (IJNT), 5(1):173– 184, February 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304210900202X.

Kurlberg:2009:PSS

[Kur09] Pär Kurlberg. Poisson spacing statistics for value sets of polynomials. International Journal of Number Theory (IJNT), 5(3):489–513, May 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002237.

Kursungoz:2021:CCT

[Kur21] Kağan Kurşungöz. A combinatorial construction for two formulas in Slater’s list. International Journal of Number Theory (IJNT), 17(03):655–663, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120400114.

Kudin:2019:CRA

[KV19] Alexey Kudin and Denis Vasilyev. Counting real algebraic numbers with bounded derivative of minimal polynomial. International Journal of Number Theory (IJNT), 15(10): 2223–2239, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119501227.

Kaczorowski:2012:STE

[KW12] Jerzy Kaczorowski and Kazimierz Wiertelak. On the sum of the twisted Euler function. International Journal of Number Theory (IJNT), 8(7):1741–1761, November 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304211250100X.

Komatsu:2014:PUF

[KW14] Hiroyuki Komatsu and Takao Watanabe. Perfect unary forms over certain cubic fields. International Journal of Number Theory (IJNT), 10(5):1337–1342, August 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042114500304.

Kaczorowski:2016:LER

[KW16] Jerzy Kaczorowski and Kazimierz Wiertelak. Lower estimates related to the twisted Dedekind function. International Journal of Number Theory (IJNT), 12(7):1801– 1811, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116501116.

Kim:2018:NFL

[KW18] Henry H. Kim and Zack Wolske. Number fields with large minimal index containing quadratic subfields. International Journal of Number Theory (IJNT), 14(9):2333–2342, October 2018. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501427.

Kim:2021:NRI

[KwK21] Kyoungmin Kim and Yeong wook Kwon. The number of representations of integers by generalized Bell ternary quadratic forms. International Journal of Number Theory (IJNT), 17(04):887–901, May 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500135.

Komatsu:2018:GSN

[KY18] Takao Komatsu and Paul Thomas Young. Generalized Stirling numbers with poly-Bernoulli and poly-Cauchy numbers. International Journal of Number Theory (IJNT), 14 (5):1211–1222, June 2018. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042118500744.

Kida:2020:CTL

[KY20] Masanari Kida and Hiromichi Yanai. CMtypes with large index of degeneracy. International Journal of Number Theory (IJNT), 16(2):325–340, March 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500177.

Lagarias:2010:CSS

[Lag10a] Jeffrey C. Lagarias. Cyclic systems of simultaneous congruences. International Journal of Number Theory (IJNT), 6(2):219–245, March 2010. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/

S1793042110002892. See erratum [Lag10b].

Lagarias:2010:ECS

[Lag10b] Jeffrey C. Lagarias. Erratum: “Cyclic Systems of Simultaneous Congruences”. International Journal of Number Theory (IJNT), 6(4):??, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042110003320. See [Lag10a].

Laishram:2010:CRP

[Lai10] Shanta Laishram. On a conjecture on Ramanujan primes. International Journal of Number Theory (IJNT), 6(8):1869– 1873, December 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003848.

Lalin:2010:CB

[Lal10] Matilde N. Lalín. On a conjecture by Boyd. International Journal of Number Theory (IJNT), 6(3):705–711, May 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003174.

Lamzouri:2016:BMP

[Lam16] Youness Lamzouri. A bias in Mertens’ product formula. International Journal of Number Theory (IJNT), 12(1):97–109, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042116500068.

Languasco:2009:CRE

[Lan09] Alessandro Languasco. A conditional result on the exceptional set for Hardy–Littlewood numbers in short intervals. International Journal of Number Theory (IJNT), 5(6):933– 951, September 2009. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304210900247X.

Lanphier:2019:DCC

[Lan19] Dominic Lanphier. Determining cuspforms from critical values of convolution L-functions and Rankin–Cohen brackets. International Journal of Number Theory (IJNT), 15(7): 1403–1412, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500787.

Laskar:2017:LMF

[Las17] Abhijit Laskar. On local monodromy filtrations attached to motives. International Journal of Number Theory (IJNT), 13(3):801–817, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500427.

Laurincikas:2013:UHZ

[Lau13] Antanas Laurinčikas. On the universality of the Hurwitz zeta-function. International Journal of Number Theory (IJNT), 9(1): 155–165, February 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501308.

Lemos:2019:NFW

[LDSM19] Abílio Lemos, Allan De Oliveira Moura, Anderson T. Silva, and B. K. Moriya. On the number of fully weighted zerosum subsequences. International Journal of Number Theory (IJNT), 15(5):1051– 1057, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304211950057X.

Le:2015:SST

[Le15] Giang Le. Schmidt’s subspace theorem for moving hypersurface targets. International Journal of Number Theory (IJNT), 11(1): 139–158, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500086.

Lebacque:2010:TVI

[Leb10] Philippe Lebacque. On Tsfasman–Vlăduţ invariants of infinite global fields. International Journal of Number Theory (IJNT), 6(6): 1419–1448, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003526.

Lee:2010:SPF

[Lee10] K. S. Enoch Lee. On the sum of a prime and a Fibonacci number. International Journal of Number Theory (IJNT), 6(7):1669– 1676, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003708.

Leher:2008:BGN

[Leh08] Eli Leher. Bounds for the genus of numerical semigroups. International Journal of Number Theory (IJNT), 4(5):827– 834, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042108001699.

Lei:2012:NCA

[Lei12] Antonio Lei. Non-commutative p-adic Lfunctions for supersingular primes. International Journal of Number Theory (IJNT), 8(8):1813–1830, December 2012. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112501047.

Leitner:2021:CCT

[Lei21] Marianne Leitner. Convolutions on the complex torus. International Journal of Number Theory (IJNT), 17(06):1343– 1363, July 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042121500391.

Lengyel:2014:EAO

[Len14] Tamás Lengyel. Exact p-adic orders for differences of Motzkin numbers. International Journal of Number Theory (IJNT), 10(3):653–667, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113501157.

Lengyel:2017:SLS

[Len17a] Tamás Lengyel. On some lacunary sums of even powers of binomial coefficients. International Journal of Number Theory (IJNT), 13 (9):2381–2392, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501305.

Lengyel:2017:RAC

[Len17b] Tamás Lengyel. On the rate of p-adic convergence of sums of powers of binomial coefficients. International Journal of Number Theory (IJNT), 13(8):2075–2091, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042117501111.

Lengyel:2018:SSL

[Len18]  Tamás Lengyel.  Supercongruences for

some lacunary sums of powers of binomial coefficients. International Journal of Number Theory (IJNT), 14(8):2155–2164, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118501294.

Lennerstad:2019:DSB

[Len19]  Håkan Lennerstad.  The n-dimensional

Stern–Brocot tree. International Journal of Number Theory (IJNT), 15(6):1219– 1236, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119500672.

Leonardis:2015:CAL

[Leo15] Antonino Leonardis. Cyclotomic approximation lattices. International Journal of Number Theory (IJNT), 11(2):557–567, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042115500293.

Leonetti:2018:CSG

[Leo18] Paolo Leonetti. A characterization of Sophie Germain primes. International Journal of Number Theory (IJNT), 14(3):653– 660, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118500409.

Leshin:2016:DIN

[Les16] Jonah Leshin. On the degree of irrationality in Noether’s problem. International Journal of Number Theory (IJNT), 12(5): 1209–1218, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500743.

Lev:2006:CPA

[Lev06] Vsevolod F. Lev. Critical pairs in Abelian groups and Kemperman’s structure theorem. International Journal of Number Theory (IJNT), 2(3):379–396, September 2006. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042106000620.

Lezowski:2012:ENE

[Lez12]  Pierre Lezowski.  Examples of norm-

Euclidean ideal classes. International Journal of Number Theory (IJNT), 8(5):1315– 1333, August 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112500777.

Liu:2016:NDQ

[LH16] Huafeng Liu and Liqun Hu. On the number of divisors of a quaternary quadratic form. International Journal of Number Theory (IJNT), 12(5):1219–1235, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042116500755.

Li:2018:GMI

[LHK18] Yan Li, Xiaoyu Hu, and Daeyeoul Kim. A generalization of Menon’s identity with Dirichlet characters. International Journal of Number Theory (IJNT), 14(10):2631– 2639, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118501579.

Li:2009:EPD

[Li09] Xian-Jin Li. On the Euler product of the Dedekind zeta function. International Journal of Number Theory (IJNT), 5(2): 293–301, March 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002109.

Li:2011:WWL

[Li11] Xiaoqing Li. A weighted Weyl law for the modular surface. International Journal of Number Theory (IJNT), 7(1):241– 248, February 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111003995.

Li:2012:EFT

[Li12] Xian-Jin Li. On the explicit formula in the theory of prime numbers. International Journal of Number Theory (IJNT), 8(3):589–597, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500327.

Li:2013:SIH

[Li13] Zhong-Hua Li. Some identities in the harmonic algebra concerned with multiple zeta values. International Journal of Number Theory (IJNT), 9(3):783–798, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112501618.

Li:2014:QTE

[Li14] Zane Kun Li. Quadratic twists of elliptic curves with 3-Selmer rank 1. International Journal of Number Theory (IJNT), 10(5): 1191–1217, August 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500213.

Li:2015:EFR

[Li15] Xian-Jin Li. On the explicit formula related to Riemann’s zeta-function. International Journal of Number Theory (IJNT), 11(8): 2451–2486, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115501146.

Li:2016:EMA

[Li16] Taiyu Li. Enlarged major arcs in the Waring–Goldbach problem. International Journal of Number Theory (IJNT), 12(1): 205–217, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500123.

Li:2017:SPM

[Li17] Jinjiang Li. On the seventh power moment of Δ(x). International Journal of Number Theory (IJNT), 13(3):571–591, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117500294.

Li:2019:DEP

[Li19] Sanhua Li. On a Diophantine equation with prime numbers. International Journal of Number Theory (IJNT), 15(8):1601– 1616, September 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119300011.

Li:2020:IIP

[Li20a] Jiangtao Li. Iterated integrals on products of one variable multiple polylogarithms. International Journal of Number Theory (IJNT), 16(10):2167–2186, November 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042120501122.

Li:2020:BC

[Li20b] Jiyou Li. On the Borwein conjecture. International Journal of Number Theory (IJNT), 16(5):1053–1066, June 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500542.

Li:2022:PSS

[Li22] Guoquan Li. The Pintz–Steiger–Szemerédi estimate for intersective quadratic polynomials in function fields. International Journal of Number Theory (IJNT), 18(02): 417–466, March 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500257.

Lim:2015:PND

[Lim15a] Meng Fai Lim. On the pseudo-nullity of the dual fine Selmer groups. International Journal of Number Theory (IJNT), 11(7):

2055–2063, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500888.

Lim:2015:RCA

[Lim15b] Meng Fai Lim. A remark on the M_{H}(G)- conjecture and Akashi series. International Journal of Number Theory (IJNT), 11(1): 269–297, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500165.

Lim:2016:ASC

[Lim16] Meng Fai Lim. Akashi series, characteristic elements and congruence of Galois representations. International Journal of Number Theory (IJNT), 12(3):593–613, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211650038X.

Lin:2013:ECF

[Lin13a] Bernard L. S. Lin. On the expansion of a continued fraction of order twelve. International Journal of Number Theory (IJNT), 9(8): 2019–2031, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500723.

Lin:2013:RCP

[Lin13b] Bernard L. S. Lin. The restricted 3-colored partition function mod3. International Journal of Number Theory (IJNT), 9(7):1789– 1799, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500577.

Lin:2014:NRT

[Lin14] Bernard L. S. Lin. New Ramanujan type congruence modulo 7 for 4-colored generalized Frobenius partitions. International Journal of Number Theory (IJNT), 10(3):637–639, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113501133.

Ling:2018:CMA

[Lin18] Dengrong Ling. A construction of maximal asymptotic nonbases. International Journal of Number Theory (IJNT), 14(4):919–923, May 2018. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500550.

Lind:2022:SED

[Lin22] Martin Lind. A sharp estimate of the discrepancy of a concatenation sequence of inversive pseudorandom numbers with consecutive primes. International Journal of Number Theory (IJNT), 18(07): 1409–1416, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500713.

Liu:2008:PIS

[Liu08] Yuancheng Liu. On the problem of integer solutions to decomposable form inequalities. International Journal of Number Theory (IJNT), 4(5):859–872, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042108001766.

Liu:2009:NLT

[Liu09]  Huaning Liu.  A note on Lehmer k-

tuples. International Journal of Number Theory (IJNT), 5(7):1169–1178, November 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042109002626.

Liu:2011:GUN

[Liu11] Huaning Liu. Gowers uniformity norm and pseudorandom measures of the pseudorandom binary sequences. International Journal of Number Theory (IJNT), 7(5): 1279–1302, August 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004137.

Liu:2012:NVD

[Liu12a] Sheng-Chi Liu. A note on the variance of the distribution of Heegner points. International Journal of Number Theory (IJNT), 8 (2):525–529, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500273.

Liu:2012:GSR

[Liu12b]  Zhi-Guo Liu.  Gauss summation and

Ramanujan-type series for 1/π. International Journal of Number Theory (IJNT), 8(2):289– 297, March 2012. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112500169.

Liu:2012:SIR

[Liu12c] Zhi-Guo Liu. Some inverse relations and theta function identities. International Journal of Number Theory (IJNT), 8(8):1977– 2002, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501126.

Liu:2012:WGP

[Liu12d] Zhixin Liu. On Waring–Goldbach problem for fifth powers. International Journal of Number Theory (IJNT), 8(5):1247– 1256, August 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042112500728.

Liu:2013:DSE

[Liu13a] Zhi-Guo Liu. On the q-derivative and q-series expansions. International Journal of Number Theory (IJNT), 9(8):2069– 2089, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500759.

Liu:2013:POP

[Liu13b] Zhixin Liu. On pairs of one prime, two prime squares and powers of 2. International Journal of Number Theory (IJNT), 9(6): 1413–1421, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500346.

Liu:2014:LFP

[Liu14] Huaning Liu. Large family of pseudorandom subsets of the set of the integers not exceeding N. International Journal of Number Theory (IJNT), 10(5):1121–1141, August 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042114500183.

Liu:2016:TRP

[Liu16] Huafeng Liu. Two results on powers of two in Waring–Goldbach type problems. International Journal of Number Theory (IJNT), 12(7):1813–1825, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042116501128.

Liu:2017:AAZ

[Liu17a] Jichun Liu. A q-analogue of Apagodu– Zeilberger congruence. International Journal of Number Theory (IJNT), 13(8):2175– 2179, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501172.

Liu:2017:DAU

[Liu17b]  Zhixin Liu.  Diophantine approximation

by unlike powers of primes. International Journal of Number Theory (IJNT), 13(9): 2445–2452, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501330.

Liu:2018:IFS

[Liu18] Chao Liu. On the index-r-free sequences over finite cyclic groups. International Journal of Number Theory (IJNT), 14(6): 1627–1636, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500987.

Liu:2019:LFP

[Liu19a] Huaning Liu. Large families of pseudorandom binary lattices by using the multiplicative inverse modulo P. International Journal of Number Theory (IJNT), 15(3):527–546, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500271.

Liu:2019:ESS

[Liu19b]  Y. Liu.  Exceptional set for sums of

unlike powers of primes. International Journal of Number Theory (IJNT), 15(2): 339–352, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500155.

Liu:2022:CNG

[Liu22] Chunhui Liu. Control of the non-geometrically integral reductions. International Journal of Number Theory (IJNT), 18(01): 47–60, February 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500051.

Liu:2011:DTS

[LL11] Zhixin Liu and Guangshi Lü. Density of two squares of primes and powers of 2. International Journal of Number Theory (IJNT), 7 (5):1317–1329, August 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004605.

Louboutin:2017:FUF

[LL17] Stéphane R. Louboutin and Jun Ho Lee. Fundamental units for a family of totally real cubic orders and the diophantine equation u(u + a)(u + 2a) = v(v + 1). International Journal of Number Theory (IJNT), 13(7): 1729–1746, August 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500993.

Lebowitz-Lockard:2018:AUS

[LL18] Noah Lebowitz-Lockard. Additively unique sets of prime numbers. International Journal of Number Theory (IJNT), 14(10):2757– 2765, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211850166X.

Lei:2022:MWR

[LL22] Antonio Lei and Meng Fai Lim. Mordell– Weil ranks and Tate–Shafarevich groups of elliptic curves with mixed-reduction type over cyclotomic extensions. International Journal of Number Theory (IJNT), 18(02): 303–330, March 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500208.

Lin:2022:ESR

[LLL22] Shuyi Lin, Jinjun Li, and Manli Lou. Exceptional sets related to the largest digits in Lüroth expansions. International Journal of Number Theory (IJNT), 18(07): 1429–1443, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500737.

Li:2018:ASF

[LLM18] Long Li, Xin Li, and Rui Mao. Appell series F_{1} over finite fields. International Journal of Number Theory (IJNT), 14(3):727–738, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211850046X.

Leinonen:2016:AME

[LLMA16] Leena Leinonen, Marko Leinonen, and Tapani Matala-Aho. On approximation measures of q-exponential function. International Journal of Number Theory (IJNT), 12(1): 287–303, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500172.

Laishram:2021:MLS

[LLS21] Shanta Laishram, Florian Luca, and Mark Sias. On members of Lucas sequences which are products of Catalan numbers. International Journal of Number Theory (IJNT), 17 (06):1487–1515, July 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500457.

Liu:2018:SCD

[LLW18] Ji-Cai Liu, Long Li, and Su-Dan Wang. Some congruences on Delannoy numbers and Schröder numbers. International Journal of Number Theory (IJNT), 14(7):2035– 2041, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042118501221.

Li:2021:RPF

[lLW21] Ya li Li and Jie Wu. On the k-th root partition function. International Journal of Number Theory (IJNT), 17(09):2071– 2085, October 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500779.

Liu:2016:PMA

[LLZ16] Huafeng Liu, Shuai Li, and Deyu Zhang. Power moments of automorphic L-function attached to Maass forms. International Journal of Number Theory (IJNT), 12(2): 427–443, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500251.

Liu:2018:ESE

[LLZ18] Jia Liu, Meiying Lü, and Zhenliang Zhang. On the exceptional sets in Erdős–Rényi limit theorem of β-expansion. International Journal of Number Theory (IJNT), 14(7): 1919–1934, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118501142.

Li:2011:DCU

[LM11] Yan Li and Lianrong Ma. Double coverings and unit square problems for cyclotomic fields. International Journal of Number Theory (IJNT), 7(7):1935–1944, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042111004836.

Lim:2014:GSG

[LM14] Meng Fai Lim and V. Kumar Murty. The growth of the Selmer group of an elliptic curve with split multiplicative reduction. International Journal of Number Theory (IJNT), 10(3):675–687, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113501170.

Li:2015:NTA

[LM15a] Brian Li and Bill Mance. Number theoretic applications of a class of Cantor series fractal functions, II. International Journal of Number Theory (IJNT), 11(2): 407–435, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500219.

Lo:2015:GTM

[LM15b] Catharine Wing Kwan Lo and Matilde Marcolli. F_{ζ}-geometry, Tate motives, and the Habiro ring. International Journal of Number Theory (IJNT), 11(2):311–339, March 2015. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500189.

Lynch:2015:QFE

[LM15c] Rodney Lynch and Patrick Morton. The quartic Fermat equation in Hilbert class fields of imaginary quadratic fields. International Journal of Number Theory (IJNT), 11(6): 1961–2017, September 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500852.

Lagarias:2016:PBC

[LM16] Jeffrey C. Lagarias and Harsh Mehta. Products of binomial coefficients and unreduced Farey fractions. International Journal of Number Theory (IJNT), 12(1):57– 91, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042116500044.

Leonetti:2018:SMP

[LM18] Paolo Leonetti and Andrea Marino. Sums of multivariate polynomials in finite subgroups. International Journal of Number Theory (IJNT), 14(2):301–311, March 2018. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500185.

Luca:2019:CTI

[LMO^{+}19] Florian Luca, Pieter Moree, Robert Osburn, Sumaia Saad Eddin, and Alisa Sedunova. Constrained ternary integers. International Journal of Number Theory (IJNT), 15(2):407–431, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119500210.

Lev:2010:ABA

[LMP10] Vsevolod F. Lev, Mikhail E. Muzychuk, and Rom Pinchasi. Additive bases in Abelian groups. International Journal of Number Theory (IJNT), 6(4):799–809, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042110003216.

Le:2010:APT

[LMS10] Daniel Le, Shelly Manber, and Shrenik Shah. On p-adic properties of twisted traces of singular moduli. International Journal of Number Theory (IJNT), 6(3):625–653, May 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003101.

Lorenzini:2022:ID

[LMS^{+}22] Dino Lorenzini, Mentzelos Melistas, Arvind Suresh, Makoto Suwama, and Haiyang Wang. Integer dynamics. International Journal of Number Theory (IJNT), 18(02): 397–415, March 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500245.

Los:2018:RP

[LMT18] Johan Los, Tiemar Mepschen, and Jaap Top. Rational Poncelet. International Journal of Number Theory (IJNT), 14(10):2641– 2655, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501580.

Lin:2016:EPA

[LMW16] Bernard L. S. Lin, Amita Malik, and Andrew Y. Z. Wang. Extensive parity analysis of broken 5-diamond partitions. International Journal of Number Theory (IJNT), 12 (1):219–236, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500135.

Li:2019:MMF

[LN19] Yingkun Li and Michael Neururer. A magnetic modular form. International Journal of Number Theory (IJNT), 15(5):907– 924, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119500489.

Lee:2012:DCN

[LO12] Seok Hyeong Lee and Gyujin Oh. On the distribution of cyclic number fields of prime degree. International Journal of Number Theory (IJNT), 8(6):1463– 1475, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500868.

Louboutin:2015:RZD

[Lou15] Stéphane R. Louboutin. Real zeros of Dedekind zeta functions. International Journal of Number Theory (IJNT), 11(3):843– 848, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042115500463.

Louboutin:2016:DO

[Lou16]  Stéphane R. Louboutin.  Discriminants

of S_{n}-orders. International Journal of Number Theory (IJNT), 12(7):1899–1905, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042116501177.

Lovejoy:2005:TSC

[Lov05] Jeremy Lovejoy. A theorem on seven-colored overpartitions and its applications. International Journal of Number Theory (IJNT), 1(2):215–224, June 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042105000157.

Li:2013:MZS

[LP13] Yuanlin Li and Jiangtao Peng. Minimal zero-sum sequences of length four over finite cyclic groups II. International Journal of Number Theory (IJNT), 9(4):845– 866, June 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042113500012.

Luca:2014:LBO

[LP14] Florian Luca and Carl Pomerance. On the local behavior of the order of appearance in the Fibonacci sequence. International Journal of Number Theory (IJNT), 10(4):915–933, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500079.

Luca:2016:APF

[LP16] Florian Luca and M. Tip Phaovibul. Amicable pairs with few distinct prime factors. International Journal of Number Theory (IJNT), 12(7):1725–1732, November 2016. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116501050.

Lee:2017:ECS

[LP17] Joohee Lee and Yoon Kyung Park. Evaluation of the convolution sums ∑ a_{1}m_{1} + a_{2}m_{2} + a_{3}m_{3} + a_{4}m_{4} = nσ(m_{1})σ(m_{2})σ(m_{3})σ(m_{4}) with lcm(a_{1}, a_{2}, a_{3}, a_{4}) ≤ 4. International Journal of Number Theory (IJNT), 13(8): 2155–2173, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501160.

Li:2020:TPA

[LP20] Zhonghua Li and Ende Pan. Topological properties of q-analogues of multiple zeta values. International Journal of Number Theory (IJNT), 16(5):963–980, June 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500499.

Pankowski:2013:EVF

[łPS13] łukasz Pańkowski and Jörn Steuding. Extreme values of L-functions from the Selberg class. International Journal of Number Theory (IJNT), 9(5):1113–1124, August 2013. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500152.

Li:2021:SRD

[LQ21] Zhonghua Li and Chen Qin. Some relations deduced from regularized double shuffle relations of multiple zeta values. International Journal of Number Theory (IJNT), 17 (01):91–146, February 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500068.

Loh:2006:ACP

[LR06] Po-Ru Loh and Robert C. Rhoades. p-adic and combinatorial properties of modular form coefficients. International Journal of Number Theory (IJNT), 2(2):305–328, June 2006. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042106000590.

Lalin:2017:MMW

[LR17] Matilde N. Lalín and Frank Ramamonjisoa. The Mahler measure of a Weierstrass form. International Journal of Number Theory (IJNT), 13(8):2195–2214, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042117501196.

Leal-Ruperto:2019:MIS

[LR19a] José Luis Leal-Ruperto. On the magnitude of the integer solutions of the semi-diagonal equation ax^{2} + by^{2} + cz^{2} + dxy = 0. International Journal of Number Theory (IJNT), 15 (1):157–165, February 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500040.

Liang:2019:DOO

[LR19b] Ke Liang and Jeremy Rouse. Density of odd order reductions for elliptic curves with a rational point of order 2. International Journal of Number Theory (IJNT), 15(8):1547– 1563, September 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042119500891.

Labihi:2022:SCC

[LR22] O. Labihi and A. Raouj. Sommes courtes de certaines fonctions multiplicatives. International Journal of Number Theory (IJNT), 18(08):1651–1668, September 2022. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042122500841.

Lozano-Robledo:2010:BTE

[LRL10] Álvaro Lozano-Robledo and Benjamin Lundell. Bounds for the torsion of elliptic curves over extensions with bounded ramification. International Journal of Number Theory (IJNT), 6(6):1293–1309, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042110003514.

Larsen:2014:IPR

[LRS14] Acadia Larsen, Alexa Rust, and Holly Swisher. Inequalities for positive rank and crank moments of overpartitions. International Journal of Number Theory (IJNT), 10(8):2115–2133, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042114500717.

Litsyn:2005:IFS

[LS05a] Simon Litsyn and Vladimir Shevelev. Irrational factors satisfying the little Fermat theorem. International Journal of Number Theory (IJNT), 1(4):499–512, December 2005. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042105000339.

Luca:2005:PDL

[LS05b] Florian Luca and Pantelimon Stănică. Prime divisors of Lucas sequences and a conjecture of Skałba. International Journal of Number Theory (IJNT), 1(4):583– 591, December 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042105000285.

Laishram:2006:GCC

[LS06] Shanta Laishram and T. N. Shorey. Grimm’s conjecture on consecutive integers. International Journal of Number Theory (IJNT), 2(2):207–211, June 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042106000498.

Liu:2009:GRT

[LS09a] Yu-Ru Liu and Craig V. Spencer. A generalization of Roth’s theorem in function fields. International Journal of Number Theory (IJNT), 5(7):1149–1154, November 2009. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002602.

Luca:2009:MFP

[LS09b] Florian Luca and Pantelimon Stănică. On Machin’s formula with powers of the Golden section. International Journal of Number Theory (IJNT), 5(6):973–979, September 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042109002493.

Le:2012:CRS

[LS12]  T. A. Le and J. W. Sander.  Con-

volutions of Ramanujan sums and integral circulant graphs. International Journal of Number Theory (IJNT), 8(7):1777– 1788, November 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501023.

Larson:2014:CPT

[LS14] Hannah Larson and Geoffrey Smith. Congruence properties of Taylor coefficients of modular forms. International Journal of Number Theory (IJNT), 10(6):1501– 1518, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500407.

Li:2017:GFP

[LS17a] Xiumei Li and Min Sha. Gauss factorials of polynomials over finite fields. International Journal of Number Theory (IJNT), 13(8): 2039–2054, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501093.

Linowitz:2017:LSO

[LS17b] Benjamin Linowitz and Thomas R. Shemanske. Local selectivity of orders in central simple algebras. International Journal of Number Theory (IJNT), 13(4):853– 884, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042117500452.

Leonetti:2018:NPC

[LS18] Paolo Leonetti and Carlo Sanna. A note on primes in certain residue classes. International Journal of Number Theory (IJNT), 14(8): 2219–2223, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118501336.

Li:2019:PFR

[LS19] Xiumei Li and Min Sha. Polynomial functions in the residue class rings of Dedekind domains. International Journal of Number Theory (IJNT), 15(7):1473–1486, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119500854.

Laslan:2020:THG

[LŞ20] Gonca Kı ZıLaslan and Mehmet Haluk Şengün. Torsion homology growth for noncongruence subgroups of Bianchi groups. International Journal of Number Theory (IJNT), 16(4):787–802, May 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500402.

Luca:2008:DE

[LT08] Florian Luca and Alain Togbé. On the Diophantine equation x^{2} + 2^{a} · 5^{b} = y^{n}. International Journal of Number Theory (IJNT), 4 (6):973–979, December 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042108001791.

Luca:2014:ILS

[LT14] Florian Luca and Yohei Tachiya. Irrationality of Lambert series associated with a periodic sequence. International Journal of Number Theory (IJNT), 10(3):623–636, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042113501121.

Li:2020:DIS

[LTZ20] Junxian Li, Albert Tamazyan, and Alexandru Zaharescu. Ducci iterates and similar ordering of visible points in convex regions. International Journal of Number Theory (IJNT), 16 (1):1–28, February 2020. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500013.

Lu:2020:DPM

[Lu20] Hengfei Lu. The distinction problem for metaplectic case. International Journal of Number Theory (IJNT), 16(6):1161–1183, July 2020. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500591.

Lucht:2010:SRE

[Luc10] Lutz G. Lucht. A survey of Ramanujan expansions. International Journal of Number Theory (IJNT), 6(8):1785–1799, December 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042110003800.

Longo:2012:NCT

[LV12] Matteo Longo and Stefano Vigni. A note on control theorems for quaternionic Hida families of modular forms. International Journal of Number Theory (IJNT), 8(6):1425– 1462, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500856.

Liu:2017:PSC

[LV17] Ji-Cai Liu and Wilberd Van Der Kallen. Proof of Sun’s conjectures on Schröder-like numbers. International Journal of Number Theory (IJNT), 13(6):1627–1638, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117500932.

Dokic:2020:LVD

[– DLV20] Dragan – Dokić, Nikola Lelas, and Ilija Vrećica. Large values of Dirichlet Lfunctions over function fields. International Journal of Number Theory (IJNT), 16(5): 1081–1109, June 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500566.

Lau:2008:LQN

[LW08] Yuk-Kam Lau and Jie Wu. On the least quadratic non-residue. International Journal of Number Theory (IJNT), 4(3):423– 435, June 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042108001432.

Lin:2013:CHF

[LW13] Jan-Li Lin and Chi-Hao Wang. Canonical height functions for monomial maps. International Journal of Number Theory (IJNT), 9(7):1821–1840, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042113500590.

Lalin:2019:RPB

[LW19] Matilde Lalín and Gang Wu. Regulator proofs for Boyd’s identities on genus 2 curves. International Journal of Number Theory (IJNT), 15(5):945–967, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500519.

Li:2020:PAT

[LW20] Runqiao Li and Andrew Y. Z. Wang. Partitions associated with two fifth-order mock theta functions and Beck-type identities. International Journal of Number Theory (IJNT), 16(4):841–855, May 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500438.

Lu:2019:SVS

[LWC19] Xiaodong Lü, Zhiwei Wang, and Bin Chen. On the smooth values of shifted almostprimes. International Journal of Number Theory (IJNT), 15(1):1–9, February 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501683.

Wu:2022:PIR

[lWfS22] Hai liang Wu and Yue feng She. On a polynomial involving roots of unity and its applications. International Journal of Number Theory (IJNT), 18(08):1749– 1763, September 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500890.

Lu:2019:TCF

[LX19] Guangshi Lü and Ping Xi. On triple correlations of Fourier coefficients of cusp forms. II. International Journal of Number Theory (IJNT), 15(4):713–722, May 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042119500374.

Liu:2021:SPF

[LX21] Zhixin Liu and Mengyuan Xue. The sum of a prime and a Fibonacci number. International Journal of Number Theory (IJNT), 17(08): 1815–1823, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500640.

Long:2005:SPM

[LY05] Ling Long and Yifan Yang. A short proof of Milne’s formulas for sums of integer squares. International Journal of Number Theory (IJNT), 1(4):533–551, December 2005. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042105000364.

Liu:2009:SFT

[LY09] Zhi-Guo Liu and Xiao-Mei Yang. On the Schröter formula for theta functions. International Journal of Number Theory (IJNT), 5(8):1477–1488, December 2009. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042109002754.

Luca:2012:LCD

[LY12] Florian Luca and Paul Thomas Young. On the log-concavity of the degenerate Bernoulli numbers. International Journal of Number Theory (IJNT), 8(3):789–800, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112500443.

Li:2019:LMC

[LY19] Jinjun Li and Xiangfeng Yang. On longest matching consecutive subsequence. International Journal of Number Theory (IJNT), 15(8):1745–1758, September 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042119500970.

Lee:2021:ECC

[LY21] Wan Lee and Myungjun Yu. On elliptic curves with complex multiplication and root numbers. International Journal of Number Theory (IJNT), 17(04):827–842, May 2021. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304212150010X.

Languasco:2008:HLP

[LZ08] Alessandro Languasco and Alessandro Zaccagnini. On the Hardy–Littlewood problem in short intervals. International Journal of Number Theory (IJNT), 4(5):715– 723, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304210800164X.

Lebacque:2011:LDZ

[LZ11] Philippe Lebacque and Alexey Zykin. On logarithmic derivatives of zeta functions in families of global fields. International Journal of Number Theory (IJNT), 7(8):2139– 2156, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111005015.

Lee:2012:PPF

[LZ12] Pak-Hin Lee and Alexandr Zamorzaev. Parity of the partition function and traces of singular moduli. International Journal of Number Theory (IJNT), 8(2):395–409, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112500236.

Loeffler:2014:ITA

[LZ14] David Loeffler and Sarah Livia Zerbes. Iwasawa theory and p-adic L-functions over Z_{p}^{2}- extensions. International Journal of Number Theory (IJNT), 10(8):2045–2095, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042114500699.

Liu:2015:TEF

[LZ15] Zhi-Guo Liu and Jiang Zeng. Two expansion formulas involving the Rogers–Szegő polynomials with applications. International Journal of Number Theory (IJNT), 11 (2):507–525, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500268.

Li:2018:HTP

[LZ18a] Jinjiang Li and Min Zhang. Hua’s theorem with the primes in Piatetski–Shapiro prime sets. International Journal of Number Theory (IJNT), 14(1):193–220, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042118500124.

Li:2018:FPM

[LZ18b] Jinjiang Li and Min Zhang. On the fourth power moment of the error term for the divisor problem with congruence conditions. International Journal of Number Theory (IJNT), 14 (6):1525–1546, July 2018. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042118500926.

Li:2018:WGPb

[LZ18c]  Jinjiang Li and Min Zhang.  On the

Waring–Goldbach problem for one square and five cubes. International Journal of Number Theory (IJNT), 14(9):2425–2440, October 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118501476.

Li:2018:WGPa

[LZ18d] Jinjiang Li and Min Zhang. On Waring– Goldbach problem for unlike powers. International Journal of Number Theory (IJNT), 14(6):1669–1687, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501014.

Li:2022:ECI

[LZ22] Yangcheng Li and Yong Zhang. On an elliptic curve involving pairs of triangles and special quadrilaterals. International Journal of Number Theory (IJNT), 18(07): 1517–1533, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500774.

Ma:2017:PPR

[Ma17] Ding Ma. Period polynomial relations of binomial coefficients and binomial realization of formal double zeta space. International Journal of Number Theory (IJNT), 13(3):761–774, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500403.

MachHide:2012:WST

[Mac12] Tomoya MachHide. Weighted sums with two parameters of multiple zeta values and their formulas. International Journal of Number Theory (IJNT), 8(8):1903– 1921, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501084.

Machide:2016:CIR

[Mac16] Tomoya Machide. Congruence identities of regularized multiple zeta values involving a pair of index sets. International Journal of Number Theory (IJNT), 12(2): 409–426, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211650024X.

Macourt:2017:IVE

[Mac17]  Simon Macourt.  An integer-valued ex-

pression of Dedekind sums. International Journal of Number Theory (IJNT), 13(5): 1253–1259, June 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500683.

Madritsch:2012:AFP

[Mad12] Manfred G. Madritsch. q-additive functions on polynomial sequences. International Journal of Number Theory (IJNT), 8 (2):377–393, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500224.

Maga:2013:SAK

[Mag13] Péter Maga. A semi-adelic Kuznetsov formula over number fields. International Journal of Number Theory (IJNT), 9(7):1649– 1681, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500498.

Maharaj:2012:SMF

[Mah12] Hiren Maharaj. Spaces of modular forms and algebraic geometric codes. International Journal of Number Theory (IJNT), 8(6): 1485–1502, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500881.

Mahadevaswamy:2019:AAR

[Mah19] Mahadevaswamy. Analytical approach to Ramanujan type and Ramanujan’s modular equations of degree 7. International Journal of Number Theory (IJNT), 15(7): 1413–1423, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500799.

Mahatab:2020:SNS

[Mah20]  Kamalakshya Mahatab.  Sixth norm of

a Steinhaus chaos. International Journal of Number Theory (IJNT), 16(6):1295– 1306, July 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042120500670.

Maire:2011:PLE

[Mai11] Christian Maire. Plongements locaux et extensions de corps de nombres. (French) [Local embeddings and number field extensions]. International Journal of Number Theory (IJNT), 7(3):721–738, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004332.

Males:2020:FVV

[Mal20] Joshua Males. A family of vector-valued quantum modular forms of depth two. International Journal of Number Theory (IJNT), 16 (1):29–64, February 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500025.

Mallesham:2021:PQF

[Mal21] Kummari Mallesham. On pairs of quadratic forms in five variables. International Journal of Number Theory (IJNT), 17(09): 2055–2070, October 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500767.

Matala-Aho:2006:VCF

[MAM06] Tapani Matala-Aho and Ville Merilä. On the values of continued fractions: q-series II. International Journal of Number Theory (IJNT), 2 (3):417–430, September 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042106000656.

Mammo:2010:DDA

[Mam10]  Behailu Mammo.  On the density of

discriminants of Abelian extensions of a number field. International Journal of Number Theory (IJNT), 6(6):1273–1291, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042110003502.

Mao:2014:ARM

[Mao14] Renrong Mao. Asymptotics for rank moments of overpartitions. International Journal of Number Theory (IJNT), 10(8):2011– 2036, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500675.

Mao:2017:PSC

[Mao17] Guo-Shuai Mao. Proof of some congruences conjectured by Z.-W. Sun. International Journal of Number Theory (IJNT), 13(8): 1983–1993, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117501068.

Marcovecchio:2014:MLP

[Mar14]  Raffaele Marcovecchio.  Multiple Leg-

endre polynomials in Diophantine approximation. International Journal of Number Theory (IJNT), 10(7):1829–1855, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500584.

Mason:2007:VVM

[Mas07] Geoffrey Mason. Vector-valued modular forms and linear differential operators. International Journal of Number Theory (IJNT), 3 (3):377–390, September 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042107000973.

Masri:2008:IFF

[Mas08] Nadia Masri. Infinite families of formulas for sums of integer squares. International Journal of Number Theory (IJNT), 4 (4):613–626, August 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001560.

Matsuno:2008:AII

[Mat08] Kazuo Matsuno. On the 2-adic Iwasawa invariants of ordinary elliptic curves. International Journal of Number Theory (IJNT), 4(3):403–422, June 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001468.

Matomaki:2010:NSN

[Mat10]  Kaisa Matomäki.  A note on smooth

numbers in short intervals. International Journal of Number Theory (IJNT), 6(5): 1113–1116, August 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003381.

Matar:2012:SGG

[Mat12] Ahmed Matar. Selmer groups and generalized class field towers. International Journal of Number Theory (IJNT), 8(4): 881–909, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500522.

Matomaki:2016:ANS

[Mat16] Kaisa Matomäki. Another note on smooth numbers in short intervals. International Journal of Number Theory (IJNT), 12(2): 323–340, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500196.

Matsuno:2017:MWR

[Mat17a] Kazuo Matsuno. Mordell–Weil ranks of elliptic curves in the cyclotomic Z_{2}-extension of the rationals. International Journal of Number Theory (IJNT), 13(2):429–438, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042117500257.

Matthes:2017:SKL

[Mat17b] Roland Matthes. The Saito–Kurokawa lift and Siegel’s theta series. International Journal of Number Theory (IJNT), 13(7): 1679–1693, August 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500968.

Matar:2018:FSG

[Mat18] Ahmed Matar. Fine Selmer groups, Heegner points and anticyclotomic Z_{p}-extensions.

International Journal of Number Theory (IJNT), 14(5):1279–1304, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211850080X.

Matsuda:2019:DES

[Mat19] Kazuhide Matsuda. Differential equations satisfied by modular forms of level 5. International Journal of Number Theory (IJNT), 15 (2):213–250, March 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211950009X.

Maurischat:2012:COS

[Mau12]  Kathrin Maurischat.  Casimir operators

for symplectic groups. International Journal of Number Theory (IJNT), 8(4):923– 932, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112500546.

Maurischat:2017:SOS

[Mau17] Kathrin Maurischat. Sturm’s operator for scalar weight in arbitrary genus. International Journal of Number Theory (IJNT), 13(10): 2677–2686, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501482.

Mayer:2012:SCG

[May12]  Daniel C. Mayer.  The second p-class

group of a number field. International Journal of Number Theory (IJNT), 8(2): 471–505, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211250025X.

Mayer:2014:QRS

[May14] Daniel C. Mayer. Quadratic p-ring spaces for counting dihedral fields. International Journal of Number Theory (IJNT), 10(8): 2205–2242, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500754.

Mayer:2019:DPF

[May19]  Daniel C. Mayer.  Differential princi-

pal factors and Pólya property of pure metacyclic fields. International Journal of Number Theory (IJNT), 15(10):1983– 2025, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119501094.

Mei:2013:APS

[MC13] Shu-Yuan Mei and Yong-Gao Chen. Arithmetic progressions in sumsets and difference sets. International Journal of Number Theory (IJNT), 9(3):601–606, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112501503.

McCarthy:2010:HSP

[McC10] Dermot McCarthy. _{3}F_{2} hypergeometric series and periods of elliptic curves. International Journal of Number Theory (IJNT), 6(3):461– 470, May 2010. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110002946.

McCarthy:2012:EGH

[McC12] Dermot McCarthy. Extending Gaussian hypergeometric series to the p-adic setting. International Journal of Number Theory (IJNT), 8(7):1581–1612, November 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042112500844.

McGown:2012:NEC

[McG12] Kevin J. McGown. Norm-Euclidean cyclic fields of prime degree. International Journal of Number Theory (IJNT), 8(1):227– 254, February 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500133.

McNew:2013:RWL

[McN13] Nathan McNew. Radically weakening the Lehmer and Carmichael conditions. International Journal of Number Theory (IJNT), 9 (5):1215–1224, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500218.

Ma:2019:DPI

[MCW19] Wu-Xia Ma, Yong-Gao Chen, and Bing-Ling Wu. Distribution of the primes involving the ceiling function. International Journal of Number Theory (IJNT), 15(3): 597–611, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500313.

Moree:2017:PTP

[ME17] Pieter Moree and Sumaia Saad Eddin. Products of two proportional primes. International Journal of Number Theory (IJNT), 13(10): 2583–2596, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501445.

Meher:2012:SRR

[Meh12] Jaban Meher. Some remarks on Rankin– Cohen brackets of eigenforms. International Journal of Number Theory (IJNT), 8(8):

2059–2068, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112501175.

Meisner:2018:DPA

[Mei18] Patrick Meisner. Distribution of points on Abelian covers over finite fields. International Journal of Number Theory (IJNT), 14 (5):1375–1401, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500860.

Memic:2017:CER

[Mem17] Nacima Memić. Characterization of ergodic rational functions on the set of 2-adic units. International Journal of Number Theory (IJNT), 13(5):1119–1128, June 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117500609.

Memic:2020:MCL

[Mem20a] Nacima Memić. Mahler coefficients of 1-Lipschitz measure-preserving functions on Z_{p}. International Journal of Number Theory (IJNT), 16(6):1247–1261, July 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042120500645.

Memic:2020:MCU

[Mem20b] Nacima Memić. Mahler coefficients of uniformly differentiable functions modulo p. International Journal of Number Theory (IJNT), 16(09):2113–2127, October 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042120501080.

Mera:2011:ZFR

[Mer11a] Mitsugu Mera. Zero-free regions of a qanalogue of the complete Riemann zeta function. International Journal of Number Theory (IJNT), 7(4):1075–1092, June 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042111004344.

Merila:2011:NLC

[Mer11b] Ville Merilä. A nonvanishing lemma for certain Padé approximations of the second kind. International Journal of Number Theory (IJNT), 7(8):1977–1997, December 2011. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004964.

Merai:2020:VRF

[Mér20] László Mérai. Values of rational functions in small subgroups of finite fields and the identity testing problem from powers. International Journal of Number Theory (IJNT), 16 (2):219–231, March 2020. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500128.

Merriman:2022:GFM

[Mer22] Claire Merriman. Geodesic flows and the mother of all continued fractions. International Journal of Number Theory (IJNT), 18 (04):931–953, May 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304212250049X.

Mestrovic:2012:PCH

[Meš12] Romeo Meštrović. Proof of a congruence for harmonic numbers conjectured by Z.-W. Sun. International Journal of Number Theory (IJNT), 8(4):1081–1085, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112500649.

Mestrovic:2016:PCB

[Meš16] Romeo Meštrović. A primality criterion based on Lucas’ congruence. International Journal of Number Theory (IJNT), 12(5): 1365–1369, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500834.

Mignot:2015:PHB

[Mig15] Teddy Mignot. Points de hauteur bornée sur les hypersurfaces lisses de l’espace triprojectif. (French) [Bounded height points on smooth hypersurfaces of triprojective space]. International Journal of Number Theory (IJNT), 11(3):945–995, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500529.

Miller:2013:WBR

[Mil13] Lance Edward Miller. A Witt–Burnside ring attached to a pro-dihedral group. International Journal of Number Theory (IJNT), 9(3):745–757, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112501588.

Milovic:2018:RNC

[Mil18] Djordjo Z. Milovic. On the 8-rank of narrow class groups of Q(\sqrt{−4pq}), Q(\sqrt{−8pq}), and Q(\sqrt{8pq}). International Journal of Number Theory (IJNT), 14(8):2165–2193, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042118501300.

Minculete:2012:CIA

[Min12] Nicuşor Minculete. On certain inequalities about arithmetic functions which use the exponential divisors. International Journal of Number Theory (IJNT), 8(6):1527– 1535, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500923.

Minelli:2022:SFP

[Min22] Paolo Minelli. On small fractional parts of perturbed polynomials. International Journal of Number Theory (IJNT), 18(08):1669– 1690, September 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500853.

Mishra:2016:GRP

[Mis16] Manish Mishra. Generic representations in L-packets. International Journal of Number Theory (IJNT), 12(6):1613–1624, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042116500986.

Miska:2017:APG

[Mis17] Piotr Miska. Arithmetic properties of generalizations of Schenker sums. International Journal of Number Theory (IJNT), 13 (6):1585–1601, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500890.

Miyazaki:2011:TCE

[Miy11] Takafumi Miyazaki. Terai’s conjecture on exponential Diophantine equations. International Journal of Number Theory (IJNT), 7(4):981–999, June 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004496.

Miyatani:2015:MDC

[Miy15] Kazuaki Miyatani. Monomial deformations of certain hypersurfaces and two hypergeometric functions. International Journal of Number Theory (IJNT), 11(8):2405– 2430, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115501122.

Mizuno:2008:ALS

[Miz08] Yoshinori Mizuno. A p-adic limit of Siegel– Eisenstein series of prime level q. International Journal of Number Theory (IJNT), 4 (5):735–746, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001729.

Muriefah:2009:DE

[MLST09] Fadwa S. Abu Muriefah, Florian Luca, Samir Siksek, and Szabolcs Tengely. On the Diophantine equation x^{2}+C = 2y^{n}. International Journal of Number Theory (IJNT), 5(6): 1117–1128, September 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002572.

Murty:2007:OVF

[MM07] M. Ram Murty and V. Kumar Murty. Odd values of Fourier coefficients of certain modular forms. International Journal of Number Theory (IJNT), 3(3):455– 470, September 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042107001036.

Meemark:2011:DSM

[MM11] Yotsanan Meemark and Nawaphon Maingam. The digraph of the square mapping on quotient rings over the Gaussian integers. International Journal of Number Theory (IJNT), 7(3):835–852, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004459.

McLeman:2013:CNI

[MM13a] Cam McLeman and Dustin Moody. Class numbers via 3-isogenies and elliptic surfaces. International Journal of Number Theory (IJNT), 9(1):125–137, February 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211250128X.

Mourtada:2013:OT

[MM13b] M. Mourtada and V. Kumar Murty. Omega theorems for \frac{L′}{L} (1, χ_{D}). International Journal of Number Theory (IJNT), 9(3):561– 581, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112501485.

Meher:2014:SCF

[MM14] Jaban Meher and M. Ram Murty. Sign changes of Fourier coefficients of half-integral weight cusp forms. International Journal of Number Theory (IJNT), 10(4):905– 914, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500067.

Milas:2008:NTP

[MMO08] Antun Milas, Eric Mortenson, and Ken Ono. Number theoretic properties of Wronskians of Andrews–Gordon series. International Journal of Number Theory (IJNT), 4 (2):323–337, April 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001377.

Medina:2011:IPL

[MMR11] Luis A. Medina, Victor H. Moll, and Eric S. Rowland. Iterated primitives of logarithmic powers. International Journal of Number Theory (IJNT), 7(3):623–634, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211100423X.

Martin:2011:RTF

[MMW11] Kimball Martin, Mark McKee, and Eric Wambach. A relative trace formula for a compact Riemann surface. International Journal of Number Theory (IJNT), 7(2): 389–429, March 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004101.

Martin:2013:NVD

[MN13] Greg Martin and Nathan Ng. Nonzero values of Dirichlet L-functions in vertical arithmetic progressions. International Journal of Number Theory (IJNT), 9(4):813–843, June 2013. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500140.

Ma:2019:APG

[MNZ19] Rong Ma, Yana Niu, and Yulong Zhang. On asymptotic properties of the generalized Dirichlet L-functions. International Journal of Number Theory (IJNT), 15(6): 1305–1321, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500738.

Miezaki:2020:EPZ

[MO20a] Tsuyoshi Miezaki and Manabu Oura. On Eisenstein polynomials and zeta polynomials II. International Journal of Number Theory (IJNT), 16(1):207–218, February 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500116.

Morisawa:2020:FUV

[MO20b] Takayuki Morisawa and Ryotaro Okazaki. Filtrations of units of Viète field. International Journal of Number Theory (IJNT), 16 (5):1067–1079, June 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500554.

Mohammadi:2019:IBG

[Moh19] Ali Mohammadi. Improved bounds on Gauss sums in arbitrary finite fields. International Journal of Number Theory (IJNT), 15(10): 2027–2041, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119501100.

Mokrani:2020:AMM

[Mok20] Youcef Mokrani. Adaptation of Monsky matrices for θ-congruent numbers. International Journal of Number Theory (IJNT), 16 (2):377–396, March 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500207.

Mollin:2005:EAL

[Mol05] R. A. Mollin. On an elementary approach to the Lebesgue–Nagell equation. International Journal of Number Theory (IJNT), 1 (4):553–561, December 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042105000352.

Molteni:2012:RPS

[Mol12] G. Molteni. Representation of a 2-power as sum of k 2-powers: the asymptotic behavior. International Journal of Number Theory (IJNT), 8(8):1923–1963, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042112501096.

Monge:2014:FEP

[Mon14] Maurizio Monge. A family of Eisenstein polynomials generating totally ramified extensions, identification of extensions and construction of class fields. International Journal of Number Theory (IJNT), 10(7):1699– 1727, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500511.

Moore:2019:ZHI

[Moo19] Samantha C. Moore. On the zeroes of halfintegral weight Eisenstein series on Γ_{0}(4). International Journal of Number Theory (IJNT), 15(4):869–882, May 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119500465.

Mortenson:2008:BDP

[Mor08]  Eric Mortenson.  On the broken 1-

diamond partition. International Journal of Number Theory (IJNT), 4(2):199– 218, April 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042108001365.

Mori:2011:PSE

[Mor11] Andrea Mori. Power series expansions of modular forms and their interpolation properties. International Journal of Number Theory (IJNT), 7(2):529–577, March 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042111004095.

Morton:2016:SCF

[Mor16] Patrick Morton. Solutions of the cubic Fermat equation in ring class fields of imaginary quadratic fields (as periodic points of a 3adic algebraic function). International Journal of Number Theory (IJNT), 12(4):853–902, June 2016. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211650055X.

Mertens:2021:MME

[MOR21] Michael H. Mertens, Ken Ono, and Larry Rolen. Mock modular Eisenstein series with nebentypus. International Journal of Number Theory (IJNT), 17(03):683– 697, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304212040028X.

Mori:2022:PSE

[Mor22] Andrea Mori. Power series expansions of modular forms and p-adic interpolation of the square roots of Rankin–Selberg special values. International Journal of Number Theory (IJNT), 18(03):515–555, April 2022. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042122500300.

Marchan:2016:IRW

[MORS16] Luz E. Marchan, Oscar Ordaz, Dennys Ramos, and Wolfgang A. Schmid. Inverse results for weighted Harborth constants. International Journal of Number Theory (IJNT), 12(7): 1845–1861, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116501141.

Moshe:2007:CMR

[Mos07] Yossi Moshe. On a conjecture of McIntosh regarding LP-sequences. International Journal of Number Theory (IJNT), 3(4): 599–610, December 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042107001139.

Marchan:2014:RPM

[MOS14a] Luz E. Marchan, Oscar Ordaz, and Wolfgang A. Schmid. Remarks on the plus–minus weighted Davenport constant. International Journal of Number Theory (IJNT), 10(5): 1219–1239, August 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500225.

Mukhamedov:2014:CEA

[MOS14b] Farrukh Mukhamedov, Bakhrom Omirov, and Mansoor Saburov. On cubic equations over p-adic fields. International Journal of Number Theory (IJNT), 10(5): 1171–1190, August 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500201.

Moscariello:2015:IWR

[Mos15] Alessio Moscariello. On integers which are representable as sums of large squares. International Journal of Number Theory (IJNT), 11(8):2505–2511, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211550116X.

Moustrou:2017:DCL

[Mou17] Philippe Moustrou. On the density of cyclotomic lattices constructed from codes. International Journal of Number Theory (IJNT), 13(5):1261–1274, June 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500695.

Moy:2013:CAP

[Moy13] Richard Moy. Congruences among power series coefficients of modular forms. International Journal of Number Theory (IJNT), 9(6):1447–1474, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042113500371.

Miller:2010:ATN

[MP10] Alison Miller and Aaron Pixton. Arithmetic traces of non-holomorphic modular invariants. International Journal of Number Theory (IJNT), 6(1):69–87, February 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110002818.

Marko:2012:NDD

[MP12] František Marko and Štefan Porubský. A note on density and the Dirichlet condition. International Journal of Number Theory (IJNT), 8(3):823–830, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500479.

Murty:2014:CSV

[MP14] M. Ram Murty and Hector Pasten. Counting squarefree values of polynomials with error term. International Journal of Number Theory (IJNT), 10(7):1743–1760, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500535.

McCarthy:2015:FFH

[MP15] Dermot McCarthy and Matthew A. Papanikolas. A finite field hypergeometric function associated to eigenvalues of a Siegel eigenform. International Journal of Number Theory (IJNT), 11(8):2431– 2450, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115501134.

Martins:2016:EHA

[MP16a] Rodrigo S. V. Martins and Daniel Panario. Erratum: “On the heuristic of approximating polynomials over finite fields by random mappings”. International Journal of Number Theory (IJNT), 12(7):2041– 2042, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116920020. See [MP16b].

Martins:2016:HAP

[MP16b] Rodrigo S. V. Martins and Daniel Panario. On the heuristic of approximating polynomials over finite fields by random mappings. International Journal of Number Theory (IJNT), 12(7):1987–2016, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116501219. See erra-

tum [MP16a].

Murty:2021:NVD

[MP21] M. Ram Murty and Siddhi S. Pathak. A note on values of the Dedekind zeta-function at odd positive integers. International Journal of Number Theory (IJNT), 17(08):1753– 1764, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500603.

Maurischat:2022:TCA

[MP22a] A. Maurischat and R. Perkins. Taylor coefficients of Anderson generating functions and Drinfeld torsion extensions. International Journal of Number Theory (IJNT), 18 (01):113–130, February 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500099.

Muller:2022:HCM

[MP22b] Nicolas Müller and Richard Pink. Hyperelliptic curves with many automorphisms. International Journal of Number Theory (IJNT), 18(04):913–930, May 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500488.

Park:2022:MDA

[mPP22] Cheol min Park and Sun Woo Park. Minimal degrees of algebraic numbers with respect to primitive elements. International Journal of Number Theory (IJNT), 18(03): 485–500, April 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500282.

Monks:2013:CPB

[MPY13] Keenan Monks, Sarah Peluse, and Lynnelle Ye. Congruence properties of Borcherds product exponents. International Journal of Number Theory (IJNT), 9(6):1563– 1578, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500437.

Mu:2018:NDA

[MQ18] Quanwu Mu and Yunyun Qu. A note on Diophantine approximation by unlike powers of primes. International Journal of Number Theory (IJNT), 14(6):1651–1668, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118501002.

Movasati:2006:HSH

[MR06] H. Movasati and S. Reiter. Hypergeometric series and Hodge cycles of four dimensional cubic hypersurfaces. International Journal of Number Theory (IJNT), 2(3): 397–416, September 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042106000632.

Moree:2012:NBT

[MR12] Pieter Moree and Eugenia Roşu. Non–Beiter ternary cyclotomic polynomials with an optimally large set of coefficients. International Journal of Number Theory (IJNT), 8(8): 1883–1902, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112501072.

Maples:2015:BZD

[MR15] Kenneth Maples and Brad Rodgers. Bootstrapped zero density estimates and a central limit theorem for the zeros of the zeta function. International Journal of Number Theory (IJNT), 11(7):2087–2107, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042115500918.

Michaud-rodgers:2022:QPN

[Mr22] Philippe Michaud-rodgers. Quadratic points on non-split Cartan modular curves. International Journal of Number Theory (IJNT), 18(02):245–267, March 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500178.

Maier:2005:CGE

[MS05] H. Maier and A. Sankaranarayanan. On a certain general exponential sum. International Journal of Number Theory (IJNT), 1(2):183–192, June 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042105000224.

Mukhopadhyay:2007:ZDE

[MS07] Anirban Mukhopadhyay and Kotyada Srinivas. A zero density estimate for the Selberg class. International Journal of Number Theory (IJNT), 3(2):263–273, June 2007. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042107000894.

Moree:2009:PPA

[MS09a] P. Moree and B. Sury. Primes in a prescribed arithmetic progression dividing the sequence. International Journal of Number Theory (IJNT), 5(4):641–665, June 2009. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042109002316.

Murty:2009:SVP

[MS09b] M. Ram Murty and N. Saradha. Special values of the polygamma functions. International Journal of Number Theory (IJNT), 5 (2):257–270, March 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002079.

Maier:2010:ESP

[MS10] H. Maier and A. Sankaranarayanan. Exponential sums over primes in residue classes. International Journal of Number Theory (IJNT), 6(4):905–918, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003319.

Mantilla-Soler:2012:NFE

[MS12a] Guillermo Mantilla-Soler. On number fields with equivalent integral trace forms. International Journal of Number Theory (IJNT), 8(7):1569–1580, November 2012. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112500807.

Martinet:2012:BMV

[MS12b] Jacques Martinet and Achill Schürmann. Bases of minimal vectors in lattices, III. International Journal of Number Theory (IJNT), 8(2):551–567, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500303.

Moriya:2013:IDD

[MS13] B. K. Moriya and C. J. Smyth. Indexdependent divisors of coefficients of modular forms. International Journal of Number Theory (IJNT), 9(7):1841–1853, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042113500607.

Masse:2015:FGS

[MS15] Bruno Massé and Dominique Schneider. Fast growing sequences of numbers and the first digit phenomenon. International Journal of Number Theory (IJNT), 11 (3):705–719, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500384.

Mao:2016:TCI

[MS16a] Guo-Shuai Mao and Zhi-Wei Sun. Two congruences involving harmonic numbers with applications. International Journal of Number Theory (IJNT), 12(2):527–539, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042116500330.

Micheli:2016:DCT

[MS16b] Giacomo Micheli and Reto Schnyder. On the density of coprime m-tuples over holomorphy rings. International Journal of Number Theory (IJNT), 12(3):833–839, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500536.

Mu:2018:TMC

[MS18a] Yan-Ping Mu and Zhi-Wei Sun. Telescoping method and congruences for double sums. International Journal of Number Theory (IJNT), 14(1):143–165, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118500100.

Murahara:2018:MZV

[MS18b] Hideki Murahara and Mika Sakata. On multiple zeta values and finite multiple zeta values of maximal height. International Journal of Number Theory (IJNT), 14(4):975–987, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500574.

Morrill:2021:QKT

[MS21] Thomas Morrill and Aleksander Simonič. Quasimodularity of the k th residual cranks. International Journal of Number Theory (IJNT), 17(01):1–13, February 2021. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042121500019.

Mahanta:2022:RSP

[MS22] Pankaj Jyoti Mahanta and Manjil P. Saikia. Refinement of some partition identities of Merca and Yee. International Journal of Number Theory (IJNT), 18(05):1131– 1142, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042122500592.

Meher:2018:CSP

[MSV18] Jaban Meher, Karam Deo Shankhadhar, and G. K. Viswanadham. On the coefficients of symmetric power L-functions. International Journal of Number Theory (IJNT), 14(3):813–824, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500495.

Miyazaki:2012:DE

[MT12] T. Miyazaki and A. Togbé. The Diophantine equation (2am − 1)^{x} + (2m)^{y} = (2am + 1)^{z}. International Journal of Number Theory (IJNT), 8(8):2035–2044, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112501151.

Magyar:2017:APS

[MT17] Ákos Magyar and Tatchai Titichetrakun. Almost prime solutions to diophantine systems of high rank. International Journal of Number Theory (IJNT), 13(6):1491– 1514, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304211750083X.

Munsch:2018:SFP

[MT18] Marc Munsch and Tim Trudgian. Squarefull primitive roots. International Journal of Number Theory (IJNT), 14(4):1013– 1021, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118500604.

Mossinghoff:2021:OWA

[MT21] Michael J. Mossinghoff and Timothy S. Trudgian. Oscillations in weighted arithmetic sums. International Journal of Number Theory (IJNT), 17(07):1697–1716, August 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042121500561.

McCoy:2017:FSC

[MTWZ17] Megan McCoy, Kevin Thielen, Liuquan Wang, and Jianqiang Zhao. A family of super congruences involving multiple harmonic sums. International Journal of Number Theory (IJNT), 13(1):109–128, February 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117500075.

Mu:2017:ODI

[Mu17] Quanwu Mu. One Diophantine inequality with unlike powers of prime variables. International Journal of Number Theory (IJNT), 13(6):1531–1545, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500853.

Matsumoto:2018:VDD

[MU18] Kohji Matsumoto and Yumiko Umegaki. On the value-distribution of the difference between logarithms of two symmetric power L-functions. International Journal of Number Theory (IJNT), 14(7):2045– 2081, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118501245.

Muetzel:2012:NLB

[Mue12] Bjoern Muetzel. A new lower bound for Hermite’s constant for symplectic lattices. International Journal of Number Theory (IJNT), 8(4):1067–1080, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500637.

Muic:2011:NVC

[Mui11] Goran Muić. On the non-vanishing of certain modular forms. International Journal of Number Theory (IJNT), 7(2):351– 370, March 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042111004083.

Muic:2012:ACN

[Mui12] Goran Muić. On the analytic continuation and non-vanishing of L-functions. International Journal of Number Theory (IJNT), 8(8): 1831–1854, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112501059.

Munshi:2010:DPR

[Mun10] Ritabrata Munshi. Density of positive rank fibers in elliptic fibrations, II. International Journal of Number Theory (IJNT), 6 (1):15–23, February 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110002867.

Murru:2015:PWC

[Mur15] Nadir Murru. On the periodic writing of cubic irrationals and a generalization of Rédei functions. International Journal of Number Theory (IJNT), 11(3):779–799, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115500438.

Murahara:2017:DRF

[Mur17] Hideki Murahara. Derivation relations for finite multiple zeta values. International Journal of Number Theory (IJNT), 13(2): 419–427, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500245.

Mehta:2014:QUS

[MV14a] Jay Mehta and G. K. Viswanadham. Quasiuniqueness of the set of “Gaussian prime plus one’s”. International Journal of Number Theory (IJNT), 10(7):1783–1790, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500559.

Moll:2014:PCP

[MV14b] Victor H. Moll and Christophe Vignat. On polynomials connected to powers of Bessel functions. International Journal of Number Theory (IJNT), 10(5):1245–1257, August 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500249.

Mellit:2016:DCC

[MV16] Anton Mellit and Masha Vlasenko. Dwork’s congruences for the constant terms of powers of a Laurent polynomial. International Journal of Number Theory (IJNT), 12(2):313–321, March 2016. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500184.

Murty:2011:TCI

[MW11] M. Ram Murty and Chester J. Weatherby. On the transcendence of certain infinite series. International Journal of Number Theory (IJNT), 7(2):323–339, March 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042111004058.

McNew:2016:ILN

[MW16] Nathan McNew and Thomas Wright. Infinitude of k-Lehmer numbers which are not Carmichael. International Journal of Number Theory (IJNT), 12(7):1863–1869, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042116501153.

Mao:2019:SCI

[MW19] Guo-Shuai Mao and Jie Wang. On some congruences involving Domb numbers and harmonic numbers. International Journal of Number Theory (IJNT), 15(10):2179– 2200, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119501203.

Miyauchi:2013:LNF

[MY13] Michitaka Miyauchi and Takuya Yamauchi. Local newforms and formal exterior square L-functions. International Journal of Number Theory (IJNT), 9(8):1995–2010, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S179304211350070X.

Merca:2021:SPM

[MY21] Mircea Merca and Ae Ja Yee. On the sum of parts with multiplicity at least 2 in all the partitions of n. International Journal of Number Theory (IJNT), 17 (03):665–681, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120400205.

McLaughlin:2022:FRV

[MZ22] James McLaughlin and Peter Zimmer. Further results on vanishing coefficients in infinite products of the form (q^{b}, q^{p} − b; q^{p})_{∞}^{3}(q^{jb}, q^{2p−jb}; q^{2p})_{∞}. International Journal of Number Theory (IJNT), 18(08):1863– 1885, September 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500956.

N:2018:BGP

[N18]  Levent KargıN.  p-Bernoulli and geo-

metric polynomials. International Journal of Number Theory (IJNT), 14(2):595– 613, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118500665.

Najman:2012:EEC

[Naj12] Filip Najman. Exceptional elliptic curves over quartic fields. International Journal of Number Theory (IJNT), 8(5):1231– 1246, August 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112500716.

Nardi:2022:PTC

[Nar22] Jade Nardi. Projective toric codes. International Journal of Number Theory (IJNT), 18 (01):179–204, February 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500142.

Nathanson:2009:HFP

[Nat09] Melvyn B. Nathanson. Heights on the finite projective line. International Journal of Number Theory (IJNT), 5(1):55–65, February 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S179304210900192X.

Nathanson:2011:PAN

[Nat11] Melvyn B. Nathanson. Problems in additive number theory, IV: Nets in groups and shortest length g-adic representations. International Journal of Number Theory (IJNT), 7(8):1999–2017, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042111004940.

Nathanson:2015:FLF

[Nat15] Melvyn B. Nathanson. A forest of linear fractional transformations. International Journal of Number Theory (IJNT), 11(4): 1275–1299, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500694.

Nathanson:2017:FCN

[Nat17] Melvyn B. Nathanson. Forests of complex numbers. International Journal of Number Theory (IJNT), 13(1):15–25, February 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042117500026.

Nebe:2013:AEE

[Neb13] Gabriele Nebe. On automorphisms of extremal even unimodular lattices. International Journal of Number Theory (IJNT), 9(8): 1933–1959, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211350067X.

Nguyen:2011:CDM

[Ngu11] Lan Nguyen. A complete description of maximal solutions of functional equations arising from multiplication of quantum integers. International Journal of Number Theory (IJNT), 7(1):9–56, February 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042111003909.

Nguyen:2019:RUC

[Ngu19] Dong Quan Ngoc Nguyen. Representation of units in cyclotomic function fields. International Journal of Number Theory (IJNT), 15 (7):1385–1401, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500775.

Naika:2017:APR

[NH17] M. S. Mahadeva Naika and B. Hemanthkumar. Arithmetic properties of 5-regular bipartitions. International Journal of Number Theory (IJNT), 13(4):937–956, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211750049X.

Neamah:2020:AOM

[NH20] Ammar Ali Neamah and Titus W. Hilberdink. The average order of the Möbius function for Beurling primes. International Journal of Number Theory (IJNT), 16(5): 1005–1011, June 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500517.

Nicolas:2006:VIF

[Nic06] Jean-Louis Nicolas. Valeurs impaires de la fonction de partition p(n). (French) [Odd values of the partition function p(n)]. International Journal of Number Theory (IJNT), 2(4):469–487, December 2006. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S179304210600067X.

Nitaj:2009:CRCo

[Nit09]  Abderrahmane Nitaj.  Cryptanalysis of

RSA with constrained keys. International Journal of Number Theory (IJNT), 5(2): 311–325, March 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002122.

Nath:2022:LCM

[NJ22] Ayan Nath and Abhishek Jha. On the least common multiple of polynomial sequences at prime arguments. International Journal of Number Theory (IJNT), 18(06): 1227–1237, July 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500622.

Niu:2022:MVG

[NMZJ22] Yana Niu, Rong Ma, Yulong Zhang, and Peilin Jiang. The mean value of the generalized Dirichlet L-functions with the weight of the Gauss sums. International Journal of Number Theory (IJNT), 18(05): 1079–1096, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500567.

Noble:2012:AWD

[Nob12] Rob Noble. Asymptotics of the weighted Delannoy numbers. International Journal of Number Theory (IJNT), 8(1):175– 188, February 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500108.

Nomura:2014:NAB

[Nom14] Jiro Nomura. On non-abelian Brumer and Brumer–Stark conjecture for monomial CMextensions. International Journal of Number Theory (IJNT), 10(4):817–848, June 2014. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500018.

Nowak:2016:LPB

[Now16] Werner Georg Nowak. Lattice points in bodies of rotation with a dent. International Journal of Number Theory (IJNT), 12 (2):567–583, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500366.

Ni:2018:CCG

[NP18] He-Xia Ni and Hao Pan. On a conjectured q-congruence of Guo and Zeng. International Journal of Number Theory (IJNT), 14 (6):1699–1707, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501038.

Nedev:2008:BSV

[NQ08] Zhivko Nedev and Anthony Quas. Balanced sets and the vector game. International Journal of Number Theory (IJNT), 4(3):339–347, June 2008. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304210800133X.

Nguyen:2016:SVE

[NR16] Ngoc Ai Van Nguyen and Damien Roy. A small value estimate in dimension two involving translations by rational points. International Journal of Number Theory (IJNT), 12 (5):1273–1293, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500780.

Nesharim:2020:MDA

[NRS20] Erez Nesharim, René Rühr, and Ronggang Shi. Metric Diophantine approximation with congruence conditions. International Journal of Number Theory (IJNT), 16(09): 1923–1933, October 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500980.

Naika:2018:CAS

[NS18] M. S. Mahadeva Naika and C. Shivashankar. Congruences for Andrews singular overpartition pairs. International Journal of Number Theory (IJNT), 14(4): 989–1008, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500586.

Nowak:2015:HDM

[NSS15] W. Georg Nowak, Klaus Scheicher, and Víctor F. Sirvent. On Hausdorff dimension monotonicity of a family of dynamical subsets of Rauzy fractals. International Journal of Number Theory (IJNT), 11 (4):1089–1098, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211550058X.

Nowak:2014:ANS

[NT14] Werner Georg Nowak and László Tóth. On the average number of subgroups of the group Z_{m} × Z_{n}. International Journal of Number Theory (IJNT), 10(2):363– 374, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S179304211350098X.

Najman:2021:IMP

[Nţ21] Filip Najman and George C. ţurcaş. Irreducibility of modp Galois representations of elliptic curves with multiplicative reduction over number fields. International Journal of Number Theory (IJNT), 17(08):1729– 1738, September 2021. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042121500585.

Nebe:2010:LDS

[NV10] Gabriele Nebe and Boris Venkov. Lowdimensional strongly perfect lattices. III: Dual strongly perfect lattices of dimension

14. International Journal of Number Theory (IJNT), 6(2):387–409, March 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003022.

Niederreiter:2005:ESD

[NW05] Harald Niederreiter and Arne Winterhof. Exponential sums and the distribution of inversive congruential pseudorandom numbers with power of two modulus. International Journal of Number Theory (IJNT), 1(3):431– 438, September 2005. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042105000261.

Nam:2021:LSC

[NY21]  Hayan Nam and Myungjun Yu.  The

largest size of an s s + 1-core partition with parts of the same parity. International Journal of Number Theory (IJNT), 17 (03):699–712, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304212040014X.

Ong:2008:EET

[OEL08] Yao Lin Ong, Minking Eie, and Wen-Chin Liaw. Explicit evaluation of triple Euler sums. International Journal of Number Theory (IJNT), 4(3):437–451, June 2008. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042108001420.

Ong:2013:GWS

[OEL13] Yao Lin Ong, Minking Eie, and Wen-Chin Liaw. On generalizations of weighted sum formulas of multiple zeta values. International Journal of Number Theory (IJNT), 9 (5):1185–1198, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211350019X.

Ogasawara:2014:CEQ

[Oga14] Takeshi Ogasawara. A certain eta-quotient and the class number of an imaginary quadratic field. International Journal of Number Theory (IJNT), 10(6):1485–1499, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042114500390.

Oh:2011:RAP

[Oh11]  Byeong-Kweon Oh.  Representations of

arithmetic progressions by positive definite quadratic forms. International Journal of Number Theory (IJNT), 7(6):1603– 1614, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004915.

Oliveira:2019:WEC

[OLG19] Filipe A. A. Oliveira, Abílio Lemos, and Hemar T. Godinho. Weighted EGZ-constant for p-groups of rank 2. International Journal of Number Theory (IJNT), 15(10):2163– 2177, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119501197.

Olivetto:2014:FCM

[Oli14] René Olivetto. On the Fourier coefficients of meromorphic Jacobi forms. International Journal of Number Theory (IJNT), 10(6): 1519–1540, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500419.

Oliveira:2022:KTR

[Oli22]  André P. Oliveira.  Khintchine’s theo-

rem with rationals coming from neighborhoods in different places. International Journal of Number Theory (IJNT), 18(06): 1367–1396, July 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500695.

Olofsson:2009:LRH

[Olo09] Rikard Olofsson. Local Riemann hypothesis for complex numbers. International Journal of Number Theory (IJNT), 5(7): 1221–1230, November 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002651.

Onodera:2008:BSG

[Ono08] Kazuhiro Onodera. Behavior of some generalized multiple sine functions. International Journal of Number Theory (IJNT), 4 (5):775–796, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001651.

Onozuka:2017:MDP

[Ono17] Tomokazu Onozuka. The multiple Dirichlet product and the multiple Dirichlet series. International Journal of Number Theory (IJNT), 13(8):2181–2193, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042117501184.

Onodera:2021:MHZ

[Ono21] Kazuhiro Onodera. On multiple Hurwitz zeta function of Mordell–tornheim type. International Journal of Number Theory (IJNT), 17(10):2327–2360, November 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042121500913.

Omar:2017:NLC

[OO17] Sami Omar and Raouf Ouni. A note on Li’s coefficients for Hecke L-functions. International Journal of Number Theory (IJNT), 13 (7):1747–1753, August 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117501007.

Oura:2008:TSR

[OPY08] Manabu Oura, Cris Poor, and David S. Yuen. Towards the Siegel ring in genus four. International Journal of Number Theory (IJNT), 4(4):563–586, August 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042108001535.

Orr:2017:CBI

[Orr17] Martin Orr. On compatibility between isogenies and polarizations of abelian varieties. International Journal of Number Theory (IJNT), 13(3):673–704, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117500348.

Odzak:2011:IFG

[OS11] Almasa Odžak and Lejla Smajlović. On interpolation functions for generalized Li coefficients in the Selberg class. International Journal of Number Theory (IJNT), 7(3):771–792, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004356.

OSullivan:2016:APF

[O’S16] Cormac O’Sullivan. Asymptotics for the partial fractions of the restricted partition generating function I. International Journal of Number Theory (IJNT), 12(6):1421– 1474, September 2016. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500895.

Odzak:2017:GES

[OS17] Almasa Odžak and Lejla Smajlović. On the generalized Euler–Stieltjes constants for the Rankin–Selberg L-function. International Journal of Number Theory (IJNT), 13(6): 1363–1379, July 2017. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117500762.

Ohno:2021:RFP

[OS21] Yasuo Ohno and Yoshitaka Sasaki. Recursion formulas for poly-Bernoulli numbers and their applications. International Journal of Number Theory (IJNT), 17(01): 175–189, February 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500081.

Ostafe:2011:MCS

[OSW11] Alina Ostafe, Igor E. Shparlinski, and Arne Winterhof. Multiplicative character sums of a class of nonlinear recurrence vector sequences. International Journal of Number Theory (IJNT), 7(6):1557– 1571, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004484.

Ono:2005:APC

[OT05] Ken Ono and Yuichiro Taguchi. 2-adic properties of certain modular forms and their applications to arithmetic functions. International Journal of Number Theory (IJNT), 1 (1):75–101, March 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042105000066.

Oishi-tomiyasu:2020:PDT

[Ot20]  Ryoko Oishi-tomiyasu.  Positive-definite

ternary quadratic forms with the same representations over Z. International Journal of Number Theory (IJNT), 16(7):1493– 1534, August 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042120500785.

Oura:2009:EPA

[Our09] Manabu Oura. Eisenstein polynomials associated to binary codes. International Journal of Number Theory (IJNT), 5(4): 635–640, June 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002298.

Oyengo:2016:NPC

[Oye16] Michael O. Oyengo. Non-periodic continued fractions for quadratic irrationalities. International Journal of Number Theory (IJNT), 12 (5):1329–1344, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500810.

Ozawa:2017:CTE

[Oza17] Tomomi Ozawa. Constant terms of Eisenstein series over a totally real field. International Journal of Number Theory (IJNT), 13(2):309–324, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500208.

Ozdemir:2021:FPF

[Ozd21] Enver Ozdemir. Factoring polynomials over finite fields. International Journal of Number Theory (IJNT), 17(07):1517–1536, August 2021. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500469.

Pach:2019:IUB

[Pac19]  Péter Pál Pach.  An improved upper

bound for the size of the multiplicative 3-Sidon sets. International Journal of Number Theory (IJNT), 15(8):1721–1729, September 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119500957.

Pal:2007:EID

[Pál07]  Ambrus Pál.  On the Eisenstein ideal

of Drinfeld modular curves. International Journal of Number Theory (IJNT), 3(4): 557–598, December 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042107001115.

Pal:2014:FEC

[Pal14]  Aprameyo Pal.  Functional equation of

characteristic elements of abelian varieties over function fields (ℓ ≠ p). International Journal of Number Theory (IJNT), 10(3):705–735, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113501194.

Pande:2011:DGR

[Pan11] Aftab Pande. Deformations of Galois representations and the theorems of Sato– Tate and Lang–Trotter. International Journal of Number Theory (IJNT), 7(8):2065– 2079, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004939.

Panraksa:2022:RPP

[Pan22] Chatchawan Panraksa. Rational periodic points of x^{d} + c and Fermat–Catalan equations. International Journal of Number Theory (IJNT), 18(05):1111–1129, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042122500580.

Papikian:2011:GAG

[Pap11] Mihran Papikian. On generators of arithmetic groups over function fields. International Journal of Number Theory (IJNT), 7(6): 1573–1587, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004265.

Park:2011:ALG

[Par11] Jeehoon Park. Another look at Gross–Stark units over the number field Q. International Journal of Number Theory (IJNT), 7 (5):1379–1393, August 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004150.

Parks:2015:APA

[Par15a] James Parks. Amicable pairs and aliquot cycles on average. International Journal of Number Theory (IJNT), 11(6):1751– 1790, September 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500761.

Parry:2015:PVM

[Par15b] Daniel Parry. A polynomial variation of Meinardus’ theorem. International Journal of Number Theory (IJNT), 11(1):251– 268, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500153.

Park:2018:MFC

[Par18a] Poo-Sung Park. Multiplicative functions commutable with sums of squares. International Journal of Number Theory (IJNT), 14 (2):469–478, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211850029X.

Park:2018:ECS

[Par18b] Yoon Kyung Park. Evaluation of the convolution sums ∑_{ak+bl+cm=n} σ(k)σ(l)σ(m) with lcm(a, b, c) = 7, 8, or 9. International Journal of Number Theory (IJNT), 14 (6):1637–1650, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500999.

Park:2020:AUP

[Par20] Poo-Sung Park. Additive uniqueness of PRIMES - 1 for multiplicative functions. International Journal of Number Theory (IJNT), 16(6):1369–1376, July 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500724.

Park:2022:RCF

[Par22] Yoon Kyung Park. Ramanujan continued fractions of order sixteen. International Journal of Number Theory (IJNT), 18(05): 1097–1109, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500579.

Pasten:2015:ACA

[Pas15] Hector Pasten. The ABC conjecture, arithmetic progressions of primes and squarefree values of polynomials at prime arguments. International Journal of Number Theory (IJNT), 11(3):721–737, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042115500396.

Patkowski:2010:CGF

[Pat10] Alexander E. Patkowski. On curious generating functions for values of L-functions. International Journal of Number Theory (IJNT), 6(7):1531–1540, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003630.

Patkowski:2015:SPT

[Pat15] Alexander E. Patkowski. A strange partition theorem related to the second Atkin– Garvan moment. International Journal of Number Theory (IJNT), 11(7):2191–2197, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042115500979.

Pathak:2019:DNV

[Pat19a] Siddhi S. Pathak. Distribution and nonvanishing of special values of L-series attached to Erdős functions. International Journal of Number Theory (IJNT), 15(7): 1449–1462, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500829.

Patkowski:2019:SMT

[Pat19b]  Alexander E. Patkowski.  Some Mellin

transforms for the Riemann zeta function in the critical strip. International Journal of Number Theory (IJNT), 15(1): 153–156, February 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500039.

Patane:2020:PHF

[Pat20] Frank Patane. A proof of Hecke’s formula for binary quadratic forms. International Journal of Number Theory (IJNT), 16 (2):233–240, March 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304212050013X.

Pazuki:2013:PMB

[Paz13] F. Pazuki. Polarized morphisms between Abelian varieties. International Journal of Number Theory (IJNT), 9(2):405–411, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112501394.

Pazuki:2019:MII

[Paz19] Fabien Pazuki. Modular invariants and isogenies. International Journal of Number Theory (IJNT), 15(3):569–584, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042119500295.

Cui:2021:ORO

[pCG21] Su ping Cui and Nancy S. S. Gu. On odd ranks of odd Durfee symbols. International Journal of Number Theory (IJNT), 17(07): 1569–1582, August 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500482.

Cui:2020:ROM

[pCGyS20] Su ping Cui, Nancy S. S. Gu, and Chen yang Su. Ranks of overpartitions modulo 4 and 8. International Journal of Number Theory (IJNT), 16(10):2293–2310, November 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042120501171.

Peaucelle:2022:ESI

[Pea22] Baptiste Peaucelle. Explicit small image theorems for residual modular representations. International Journal of Number Theory (IJNT), 18(05):1143–1202, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042122500609.

Pehlivan:2016:AMO

[Peh16] Cihan Pehlivan. The average multiplicative order of a finitely generated subgroup of the rationals modulo primes. International Journal of Number Theory (IJNT), 12(8): 2147–2158, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116501281.

Penniston:2008:ARP

[Pen08] David Penniston. Arithmetic of ℓ-regular partition functions. International Journal of Number Theory (IJNT), 4(2):295– 302, April 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042108001341.

Penniston:2019:RPH

[Pen19]  David Penniston.  11-regular partitions

and a Hecke eigenform. International Journal of Number Theory (IJNT), 15(6): 1251–1259, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500696.

Perucca:2012:MSP

[Per12] Antonella Perucca. The multilinear support problem for products of Abelian varieties and tori. International Journal of Number Theory (IJNT), 8(1):255–264, February 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112500145.

Perucca:2017:ROD

[Per17]  Antonella Perucca.  Reductions of one-

dimensional tori. International Journal of Number Theory (IJNT), 13(6):1473– 1489, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042117500828.

Petsche:2005:QVB

[Pet05] Clayton Petsche. A quantitative version of Bilu’s equidistribution theorem. International Journal of Number Theory (IJNT), 1(2):281–291, June 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042105000145.

Petkov:2016:CTM

[Pet16] Vladislav Petkov. A converse theorem for metaplectic Eisenstein series on the double and triple covers of SL_{2}(Q(\sqrt{−D})). International Journal of Number Theory (IJNT), 12(6):1625–1639, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500998.

Peyrot:2020:LVH

[Pey20] Alexandre Peyrot. Large values of Hecke– Maass L-functions with prescribed argument.

International Journal of Number Theory (IJNT), 16(1):173–206, February 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500104.

Fang:2022:SQS

[pFG22] Jian ping Fang and Victor J. W. Guo. Some q -supercongruences related to Swisher’s (H.3) conjecture. International Journal of Number Theory (IJNT), 18(07):1417–1427, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042122500725.

Pham:2022:SST

[Pha22] Duc Hiep Pham. Schmidt’s subspace theorem for non-subdegenerate families of hyperplanes. International Journal of Number Theory (IJNT), 18(03):557–574, April 2022. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042122500312.

Philippon:2011:AFC

[Phi11] Patrice Philippon. Approximations fonctionnelles des courbes des espaces projectifs. (French) [Functional approximations of the curves of projective spaces]. International Journal of Number Theory (IJNT), 7 (5):1195–1215, August 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004502.

Peng:2019:SSZ

[PHLS19] Jiangtao Peng, Wanzhen Hui, Yuanlin Li, and Fang Sun. On subset sums of zerosum free sets of abelian groups. International Journal of Number Theory (IJNT), 15(3):645–654, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500349.

Pi:2020:CVR

[Pi20] Qinghua Pi. Central values of GL(2) × GL(3) Rankin–Selberg L-functions. International Journal of Number Theory (IJNT), 16(5):941–962, June 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500487.

Pickett:2010:CSD

[Pic10] Erik Jarl Pickett. Construction of selfdual integral normal bases in Abelian extensions of finite and local fields. International Journal of Number Theory (IJNT), 6(7): 1565–1588, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003654.

Pila:2009:EFS

[Pil09] Jonathan Pila. Entire functions sharing arguments of integrality, I. International Journal of Number Theory (IJNT), 5(2): 339–353, March 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002146.

Pitoun:2011:CCM

[Pit11] Frédéric Pitoun. Conoyaux de capitulation et modules d’Iwasawa. (French) [Capitulation cokernels and Iwasawa modules]. International Journal of Number Theory (IJNT), 7(6): 1503–1517, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004861.

Prasad:2013:NPL

[PMM13] K. C. Prasad, Hrishikesh Mahato, and Sudhir Mishra. A new point in Lagrange spectrum. International Journal of Number Theory (IJNT), 9(2):393–403, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112501382.

Poels:2020:CMS

[Poë20] Anthony Poëls. A class of maximally singular sets for rational approximation. International Journal of Number Theory (IJNT), 16(09): 2005–2012, October 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120501031.

Pollack:2011:ESP

[Pol11] Paul Pollack. The exceptional set in the polynomial Goldbach problem. International Journal of Number Theory (IJNT), 7(3):579–591, May 2011. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004423.

Pollack:2014:PSA

[Pol14a] Paul Pollack. Prime splitting in abelian number fields and linear combinations of Dirichlet characters. International Journal of Number Theory (IJNT), 10(4):885– 903, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042114500055.

Pollack:2014:RPD

[Pol14b] Paul Pollack. A remark on prime divisors of partition functions. International Journal of Number Theory (IJNT), 10(1): 125–131, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211350084X.

Pontreau:2006:GLB

[Pon06] Corentin Pontreau. Geometric lower bounds for the normalized height of hypersurfaces. International Journal of Number Theory (IJNT), 2(4):555–568, December 2006. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042106000723.

Pong:2009:LSN

[Pon09] Wai Yan Pong. Length spectra of natural numbers. International Journal of Number Theory (IJNT), 5(6):1089–1102, September 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042109002584.

Pongsriiam:2016:LBN

[Pon16] Prapanpong Pongsriiam. Local behaviors of the number of relatively prime sets. International Journal of Number Theory (IJNT), 12(6):1575–1593, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500962.

Pottmeyer:2018:STA

[Pot18] Lukas Pottmeyer. Small totally p-adic algebraic numbers. International Journal of Number Theory (IJNT), 14(10):2687– 2697, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501622.

Pottmeyer:2021:FGF

[Pot21] Lukas Pottmeyer. Fields generated by finite rank subgroups of Q̄^{∗}. International Journal of Number Theory (IJNT), 17(05): 1079–1089, June 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500251.

Pellarin:2018:VDM

[PP18] Federico Pellarin and Rudolph B. Perkins. Vectorial Drinfeld modular forms over Tate algebras. International Journal of Number Theory (IJNT), 14(6):1729–1783, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118501063.

Pilehrood:2012:CCJ

[PPT12] Kh. Hessami Pilehrood, T. Hessami Pilehrood, and R. Tauraso. Congruences concerning Jacobi polynomials and Apérylike formulae. International Journal of Number Theory (IJNT), 8(7):1789–1811, November 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112501035.

Peng:2014:SZS

[PQSW14] Jiangtao Peng, Guoyou Qian, Fang Sun, and Linlin Wang. On the structure of nzero-sum free sequences over cyclic groups of order n. International Journal of Number Theory (IJNT), 10(8):1991–2009, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500663.

Patsolic:2017:TDQ

[PR17] Jesse Patsolic and Jeremy Rouse. Trinomials defining quintic number fields. International Journal of Number Theory (IJNT), 13 (7):1881–1894, August 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501032.

Pandey:2021:SLC

[PR21] Manish Kumar Pandey and B. Ramakrishnan. Shimura lifts of certain classes of modular forms of half-integral weight. International Journal of Number Theory (IJNT), 17 (06):1391–1404, July 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500421.

Pries:2009:TCL

[Pri09]  Rachel Pries.  The p-torsion of curves

with large p-rank. International Journal of Number Theory (IJNT), 5(6):1103– 1116, September 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002560.

Pineda-Ruelas:2008:EGG

[PRVS08] Mario Pineda-Ruelas and Gabriel D. Villa-Salvador. Explicit Galois group realizations. International Journal of Number Theory (IJNT), 4(4):639–652, August 2008. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042108001584.

Petersen:2011:EAN

[PS11] Kathleen L. Petersen and Christopher D. Sinclair. Equidistribution of algebraic numbers of norm one in quadratic number fields. International Journal of Number Theory (IJNT), 7(7):1841–1861, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042111004666.

Porta:2015:GIM

[PS15] Horacio A. Porta and Kenneth B. Stolarsky. A golden iterated map number system: Results and conjectures. International Journal of Number Theory (IJNT), 11(5):

1617–1631, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211540014X.

Pauli:2017:EEA

[PS17] Sebastian Pauli and Brian Sinclair. Enumerating extensions of π -adic fields with given invariants. International Journal of Number Theory (IJNT), 13(8):2007–2038, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042117501081.

Perucca:2019:KTN

[PS19] Antonella Perucca and Pietro Sgobba. Kummer theory for number fields and the reductions of algebraic numbers. International Journal of Number Theory (IJNT), 15(8): 1617–1633, September 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211950091X.

Pal:2022:BMO

[PS22] Ambrus Pál and Tomer M. Schlank. Brauer– Manin obstruction to the local–global principle for the embedding problem. International Journal of Number Theory (IJNT), 18 (07):1535–1565, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500786.

Paule:2021:HRM

[PsR21] Peter Paule and Cristian silviu Radu. Holonomic relations for modular functions and forms: First guess, then prove. International Journal of Number Theory (IJNT), 17 (03):713–759, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120400278.

Perucca:2020:EKT

[PST20] Antonella Perucca, Pietro Sgobba, and Sebastiano Tronto. Explicit Kummer theory for the rational numbers. International Journal of Number Theory (IJNT), 16(10): 2213–2231, November 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120501146.

Perucca:2021:DKE

[PST21] Antonella Perucca, Pietro Sgobba, and Sebastiano Tronto. The degree of Kummer extensions of number fields. International Journal of Number Theory (IJNT), 17(05): 1091–1110, June 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500263.

Poor:2017:SPF

[PSY17] Cris Poor, Jerry Shurman, and David S. Yuen. Siegel paramodular forms of weight 2 and squarefree level. International Journal of Number Theory (IJNT), 13(10):2627– 2652, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501469.

Poor:2018:PFL

[PSY18] Cris Poor, Ralf Schmidt, and David S. Yuen. Paramodular forms of level 8 and weights 10 and 12. International Journal of Number Theory (IJNT), 14(2):417–467, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118500288.

Polanco:2016:CDA

[PSZ16] Geremías Polanco, Daniel Schultz, and Alexandru Zaharescu. Continuous distributions arising from the three gap theorem. International Journal of Number Theory (IJNT), 12(7):1743–1764, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

                                        worldscientific.com/doi/10.1142/S1793042116501074.

Pacetti:2014:SCL

[PT14] Ariel Pacetti and Gonzalo Tornaría. Shimura correspondence for level p^{2} and the central values of L-series, II. International Journal of Number Theory (IJNT), 10(7):1595– 1635, November 2014. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211450047X.

Pollack:2015:AFC

[PT15] Paul Pollack and Lola Thompson. Arithmetic functions at consecutive shifted primes. International Journal of Number Theory (IJNT), 11(5):1477–1498, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115400023.

Platt:2019:FDC

[PT19] Dave Platt and Tim Trudgian. Fujii’s development on Chebyshev’s conjecture. International Journal of Number Theory (IJNT), 15(3):639–644, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119500337.

Pumplun:2020:DFH

[Pum20] S. Pumplün. Diagonal forms of higher degree over function fields of p-adic curves. International Journal of Number Theory (IJNT), 16 (1):161–172, February 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500098.

Pupyrev:2015:LB

[Pup15] Yury Pupyrev. On a lower bound for ∥(4/3)^{k}∥. International Journal of Number Theory (IJNT), 11(1):299–309, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042115500177.

Purkait:2013:SD

[Pur13] Soma Purkait. On Shimura’s decomposition. International Journal of Number Theory (IJNT), 9(6):1431–1445, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S179304211350036X.

Preparata:2012:RPI

[PV12] Franco P. Preparata and Roberto Vacca. On the representation of the powers of integers in any prime base. International Journal of Number Theory (IJNT), 8(2): 531–541, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500285.

Peruginelli:2016:LPI

[PZ16] Giulio Peruginelli and Paolo Zanardo. The lattice of primary ideals of orders in quadratic number fields. International Journal of Number Theory (IJNT), 12(7):2025–2040, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042117500737.

Pakovich:2018:DZP

[PZ18] Fedor Pakovich and Alexander K. Zvonkin. Davenport–Zannier polynomials over Q. International Journal of Number Theory (IJNT), 14(4):925–974, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042118500562.

Qi:2022:KTP

[QGX22] Jinyun Qi, Victor Zhenyu Guo, and Zhefeng Xu. k =th powers in Piatetski– Shapiro sequences. International Journal of Number Theory (IJNT), 18(08):1791– 1806, September 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500919.

Qu:2016:ITA

[QH16] Yongke Qu and Dongchun Han. An inverse theorem for additive bases. International Journal of Number Theory (IJNT), 12(6): 1509–1518, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500937.

Qu:2017:AB

[QH17] Yongke Qu and Dongchun Han. Additive bases of C_{p} ⊕ C_{p^{n}}. International Journal of Number Theory (IJNT), 13(9): 2453–2459, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117501342.

Qiu:2019:AVS

[QH19] Min Qiu and Shaofang Hong. 2-adic valuations of Stirling numbers of the first kind. International Journal of Number Theory (IJNT), 15 (9):1827–1855, October 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119501021.

Qin:2015:NRP

[QQH15] Xiaoer Qin, Guoyou Qian, and Shaofang Hong. New results on permutation polynomials over finite fields. International Journal of Number Theory (IJNT), 11(2): 437–449, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500220.

Quang:2018:SST

[Qua18] Si Duc Quang. Schmidt’s subspace theorem for moving hypersurfaces in subgeneral position. International Journal of Number Theory (IJNT), 14(1):103–121, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118500082.

Quang:2019:GST

[Qua19] Si Duc Quang. A generalization of the subspace theorem for higher degree polynomials in subgeneral position. International Journal of Number Theory (IJNT), 15(4):775–788, May 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500404.

Qi:2011:MTS

[QY11] Zhi Qi and Chang Yang. Morita’s theory for the symplectic groups. International Journal of Number Theory (IJNT), 7(8): 2115–2137, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004952.

Raji:2009:FCG

[Raj09] Wissam Raji. Fourier coefficients of generalized modular forms of negative weight. International Journal of Number Theory (IJNT), 5 (1):153–160, February 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002006.

Raji:2011:ECT

[Raj11] Wissam Raji. Eichler cohomology theorem for generalized modular forms. International Journal of Number Theory (IJNT), 7 (4):1103–1113, June 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004514.

Ramirez:2017:CCS

[Ram17] Felipe A. Ramírez. Counterexamples, covering systems, and zero-one laws for inhomogeneous approximation. International Journal of Number Theory (IJNT), 13(3): 633–654, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500324.

Raulf:2016:LDC

[Rau16] Nicole Raulf. Limit distribution of class numbers for discriminants in progressions and fundamental discriminants. International Journal of Number Theory (IJNT), 12(5): 1237–1258, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500767.

Ray:2021:DCR

[Ray21] Anwesh Ray. Deformations of certain reducible Galois representations III. International Journal of Number Theory (IJNT), 17 (06):1429–1485, July 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500445.

Ray:2022:DRG

[Ray22] Anwesh Ray. Deformations of reducible Galois representations to Hida-families. International Journal of Number Theory (IJNT), 18 (01):89–112, February 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500087.

Ray:2018:IFC

[RB18] Chiranjit Ray and Rupam Barman. Infinite families of congruences for k-regular overpartitions. International Journal of Number Theory (IJNT), 14(1):19–29, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118500021.

Ruiz-Cabello:2017:ULC

[RC17] Serafín Ruiz-Cabello. On the use of the least common multiple to build a primegenerating recurrence. International Journal of Number Theory (IJNT), 13(4):819–833, May 2017. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500439.

Revesz:2013:GEW

[RD13] Szilárd Gy. Révész and Anne De Roton. Generalization of the effective Wiener–Ikehara theorem. International Journal of Number Theory (IJNT), 9(8):2091–2128, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042113500760.

Reis:2021:MVT

[Rei21] Lucas Reis. Mean value theorems for a class of density-like arithmetic functions. International Journal of Number Theory (IJNT), 17 (04):1013–1027, May 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500214.

Ren:2019:NST

[Ren19] Rufei Ren. Newton slopes for twisted Artin– Schreier–Witt towers. International Journal of Number Theory (IJNT), 15(10):2089– 2105, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119501148.

Ressler:2009:BQF

[Res09] Wendell Ressler. On binary quadratic forms and the Hecke groups. International Journal of Number Theory (IJNT), 5(8):1401– 1418, December 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002730.

Reynya:2013:SHD

[Rey13] M. A. Reynya. Symmetric homogeneous Diophantine equations of odd degree. International Journal of Number Theory (IJNT), 9(4):867–876, June 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500024.

Reynya:2016:RRN

[Rey16] M. A. Reynya. Representations of a rational number as a sum of odd powers. International Journal of Number Theory (IJNT), 12(4):903–911, June 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500561.

Reznick:2021:EST

[Rez21] Bruce Reznick. Equal sums of two cubes of quadratic forms. International Journal of Number Theory (IJNT), 17(03):761– 786, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042120400308.

Roditty-Gershon:2020:VSA

[RGHK20] Edva Roditty-Gershon, Chris Hall, and Jonathan P. Keating. Variance of sums in arithmetic progressions of divisor functions associated with higher degree L-functions in F_{q}[t]. International Journal of Number Theory (IJNT), 16(5):1013–1030, June 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120500529.

Rhoades:2009:SPD

[Rho09] Robert C. Rhoades. Statistics of prime divisors in function fields. International Journal of Number Theory (IJNT), 5(1): 141–152, February 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109001980.

Ribenboim:2011:MPT

[Rib11] Paulo Ribenboim. Multiple patterns of k-tuples of integers. International Journal of Number Theory (IJNT), 7(7):1761– 1779, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004733.

Richard:2013:RGC

[Ric13] Rodolphe Richard. Répartition galoisienne d’une classe d’isogénie de courbes elliptiques. (French) [Galois distribution of an isogeny class of elliptic curves]. International Journal of Number Theory (IJNT), 9 (2):517–543, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501199.

Riffaut:2019:EPS

[Rif19] Antonin Riffaut. Equations with powers of singular moduli. International Journal of Number Theory (IJNT), 15(3):445–468, April 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500234.

Rivoal:2009:AAI

[Riv09] Tanguy Rivoal. Applications arithmétiques de l’interpolation lagrangienne. (French) [Arithmetic applications of Lagrangianp interpolation]. International Journal of Number Theory (IJNT), 5(2):185–208, March 2009. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109001992.

Ruiz:2018:PFS

[RL18] Carlos Alexis Gómez Ruiz and Florian Luca. On prime factors of the sum of two k-Fibonacci numbers. International Journal of Number Theory (IJNT), 14(4): 1171–1195, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500720.

Ruiz:2019:ZMF

[RL19] Carlos Alexis Gómez Ruiz and Florian Luca. On the zero–multiplicity of a fifthorder linear recurrence. International Journal of Number Theory (IJNT), 15(3):585– 595, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119500301.

Rihane:2022:TND

[RLT22] Salah Eddine Rihane, Florian Luca, and Alain Togbé. There are no Diophantine quadruples of Pell numbers. International Journal of Number Theory (IJNT), 18(01): 27–45, February 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304212250004X.

Robertson:2010:MCF

[Rob10] Leanne Robertson. Monogeneity in cyclotomic fields. International Journal of Number Theory (IJNT), 6(7):1589–1607, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042110003666.

Roberts:2011:NPF

[Rob11] David P. Roberts. Nonsolvable polynomials with field discriminant 5^{A}. International Journal of Number Theory (IJNT), 7 (2):289–322, March 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004113.

Roberts:2015:PPB

[Rob15] David P. Roberts. Polynomials with prescribed bad primes. International Journal of Number Theory (IJNT), 11(4):1115– 1148, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304211550061X.

Roberts:2018:NFR

[Rob18] David P. Roberts. PGL_{2}(F_{ℓ}) number fields with rational companion forms. International Journal of Number Theory (IJNT), 14(3):825–845, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500501.

Robert:2020:GTH

[Rob20] Gerardo Gonzalez Robert. Good’s theorem for Hurwitz continued fractions. International Journal of Number Theory (IJNT), 16 (7):1433–1447, August 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500761.

Rodgers:2014:CLT

[Rod14] Brad Rodgers. A central limit theorem for the zeroes of the zeta function. International Journal of Number Theory (IJNT), 10 (2):483–511, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113501054.

Roe:2014:AEB

[Roe14] David Roe. The 3-adic eigencurve at the boundary of weight space. International Journal of Number Theory (IJNT), 10(7): 1791–1806, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500560.

Rolen:2011:GCN

[Rol11] Larry Rolen. A generalization of the congruent number problem. International Journal of Number Theory (IJNT), 7(8):2237– 2247, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111005039.

Rome:2019:PPH

[Rom19] Nick Rome. A positive proportion of Hasse principle failures in a family of Châtelet surfaces. International Journal of Number Theory (IJNT), 15(6):1237–1249, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119500684.

Rosengren:2008:SSE

[Ros08] Hjalmar Rosengren. Sums of squares from elliptic Pfaffians. International Journal of Number Theory (IJNT), 4(6):873–902, December 2008. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001778.

Rosen:2013:MHS

[Ros13] Julian Rosen. Multiple harmonic sums and Wolstenholme’s theorem. International Journal of Number Theory (IJNT), 9(8):2033– 2052, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500735.

Rosenberg:2017:SFV

[Ros17] Shai Rosenberg. Square-free values of multivariate polynomials over function fields in linear sparse sets. International Journal of Number Theory (IJNT), 13(1):77– 108, February 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500063.

Rosen:2018:GFC

[Ros18] Julian Rosen. A general family of congruences for Bernoulli numbers. International Journal of Number Theory (IJNT), 14(7): 1895–1902, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501129.

Rowell:2010:NEL

[Row10] Michael Rowell. A new exploration of the Lebesgue identity. International Journal of Number Theory (IJNT), 6(4):785– 798, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042110003204.

Roy:2005:SAC

[Roy05] Damien Roy. Simultaneous approximation by conjugate algebraic numbers in fields of transcendence degree one. International Journal of Number Theory (IJNT), 1(3): 357–382, September 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042105000212.

Royer:2007:ECS

[Roy07] Emmanuel Royer. Evaluating convolution sums of the divisor function by quasimodular forms. International Journal of Number Theory (IJNT), 3(2):231–261, June 2007. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042107000924.

Roy:2010:SVE

[Roy10] Damien Roy. Small value estimates for the additive group. International Journal of Number Theory (IJNT), 6(4):919– 956, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S179304211000323X.

Rozensztajn:2018:ACR

[Roz18]  Sandra Rozensztajn.  An algorithm for

computing the reduction of 2-dimensional crystalline representations of Gal(Q̄_{p}/Q_{p}). International Journal of Number Theory (IJNT), 14(7):1857–1894, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042118501117.

Robles-Perez:2017:FNS

[RPR17] Aureliano M. Robles-Pérez and José Carlos Rosales. The Frobenius number in the set of numerical semigroups with fixed multiplicity and genus. International Journal of Number Theory (IJNT), 13(4):1003– 1011, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042117500531.

Reznick:2011:STC

[RR11] Bruce Reznick and Jeremy Rouse. On the sums of two cubes. International Journal of Number Theory (IJNT), 7(7):1863– 1882, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004903.

Ramana:2020:VTP

[RR20] D. S. Ramana and O. Ramaré. Variant of the truncated Perron formula and primes in polynomial sets. International Journal of Number Theory (IJNT), 16(2): 309–323, March 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500165.

Rodseth:2006:PPF

[RS06] Øystein J. Rødseth and James A. Sellers. Partitions with parts in a finite set. International Journal of Number Theory (IJNT), 2 (3):455–468, September 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042106000644.

Radu:2011:CPM

[RS11] Silviu Radu and James A. Sellers. Congruence properties modulo 5 and 7 for the pod function. International Journal of Number Theory (IJNT), 7(8):2249– 2259, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111005064.

Radu:2013:CMS

[RS13a] Cristian-Silviu Radu and James A. Sellers. Congruences modulo squares of primes for Fu’s k dots bracelet partitions. International Journal of Number Theory (IJNT), 9(4):939–943, June 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500073.

Ramakrishnan:2013:ECS

[RS13b] B. Ramakrishnan and Brundaban Sahu. Evaluation of the convolution sums ∑_{l+15m=n} σ(l)σ(m) and ∑_{3l+5m=n} σ(l)σ(m) and an application. International Journal of Number Theory (IJNT), 9(3):799–809, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211250162X.

Ramakrishnan:2013:CBJ

[RS13c] B. Ramakrishnan and Karam Deo Shankhadhar. On a correspondence between Jacobi cusp forms and elliptic cusp forms. International Journal of Number Theory (IJNT), 9(4):917–937, June 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500061.

Rogers:2013:SSC

[RS13d] Mathew D. Rogers and Armin Straub. A solution of Sun’s $520 challenge concerning \frac{520}{π}. International Journal of Number Theory (IJNT), 9(5):1273–1288, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042113500267.

Ramakrishnan:2014:NRI

[RS14a] B. Ramakrishnan and Brundaban Sahu. On the number of representations of an integer by certain quadratic forms in sixteen variables. International Journal of Number Theory (IJNT), 10(8):1929–1937, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500638.

Rubinstein-Salzedo:2014:IFC

[RS14b] Simon Rubinstein-Salzedo. Invariants for A_{4} fields and the Cohen–Lenstra heuristics. International Journal of Number Theory (IJNT), 10(5):1259–1276, August 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042114500250.

Ramakrishnan:2018:NRC

[RSS18] B. Ramakrishnan, Brundaban Sahu, and Anup Kumar Singh. On the number of representations by certain octonary quadratic forms with coefficients 1, 2, 3, 4 and 6. International Journal of Number Theory (IJNT), 14(3):751–812, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500483.

Ramare:2020:PPA

[RSS20] Olivier Ramaré, Priyamvad Srivastav, and Oriol Serra. Product of primes in arithmetic progressions. International Journal of Number Theory (IJNT), 16(4):747– 766, May 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042120500384.

Rath:2008:DC

[RST08] P. Rath, K. Srilakshmi, and R. Thangadurai. On Davenport’s constant. International Journal of Number Theory (IJNT), 4 (1):107–115, February 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042108001195.

Royer:2014:SCS

[RSW14] Emmanuel Royer, Jyoti Sengupta, and Jie Wu. Sign changes in short intervals of coefficients of spinor zeta function of a Siegel cusp form of genus 2. International Journal of Number Theory (IJNT), 10(2):327–339, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042113500966.

Reinholz:2018:FEN

[RSY18] Lindsey Reinholz, Blair K. Spearman, and Qiduan Yang. Families of even non-congruent numbers with prime factors in each odd congruence class modulo eight. International Journal of Number Theory (IJNT), 14(3):669–692, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500422.

Ryan:2011:BTC

[RT11] Nathan C. Ryan and Gonzalo Tornaría. A Böcherer-type conjecture for paramodular forms. International Journal of Number Theory (IJNT), 7(5):1395–1411, August 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042111004629.

Rasmussen:2017:ASG

[RT17] Christopher Rasmussen and Akio Tamagawa. Abelian surfaces good away from 2. International Journal of Number Theory (IJNT), 13(4):991–1001, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211750052X.

Ruhl:2010:AIQ

[Rüh10] Klaas-Tido Rühl. Annihilating ideals of quadratic forms over local and global fields. International Journal of Number Theory (IJNT), 6(3):603–624, May 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003095.

Rouse:2016:TCA

[RY16]  Jeremy Rouse and Yilin Yang.  Three

consecutive almost squares. International Journal of Number Theory (IJNT), 12(4): 969–978, June 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500603.

Ru:2022:RVR

[RyW22] Min Ru and Julie Tzu yueh Wang. The Ru–Vojta result for subvarieties. International Journal of Number Theory (IJNT), 18 (01):61–74, February 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500063.

Roehrig:2022:TSQ

[RZ22] Christina Roehrig and Sander Zwegers. Theta series for quadratic forms of signature (n−1, 1) with (spherical) polynomials. International Journal of Number Theory (IJNT), 18(07): 1491–1515, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500762.

Sapar:2009:MEA

[SA09] S. H. Sapar and K. A. Mohd. Atan. A method of estimating the p-adic sizes of common zeros of partial derivative polynomials associated with a quintic form. International Journal of Number Theory (IJNT), 5(3):541–554, May 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002249.

Sadek:2016:CSG

[Sad16] Mohammad Sadek. Character sums, Gaussian hypergeometric series, and a family of hyperelliptic curves. International Journal of Number Theory (IJNT), 12(8):2173– 2187, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116501311.

Saha:2011:PDR

[Sah11] Abhishek Saha. Prime density results for Hecke eigenvalues of a Siegel cusp form. International Journal of Number Theory (IJNT), 7(4):971–979, June 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004642.

Saha:2016:NAF

[Sah16] Biswajyoti Saha. A note on arithmetical functions with absolutely convergent Ramanujan expansions. International Journal of Number Theory (IJNT), 12(6):1595– 1611, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500974.

Sairaiji:2010:FGJ

[Sai10] Fumio Sairaiji. Formal groups of Jacobian varieties of hyperelliptic curves. International Journal of Number Theory (IJNT), 6(7): 1701–1716, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003733.

Sakai:2011:AOP

[Sak11] Yuichi Sakai. The Atkin orthogonal polynomials for the low-level Fricke groups and their application. International Journal of Number Theory (IJNT), 7(6):1637– 1661, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004460.

Sakai:2014:AOP

[Sak14] Yuichi Sakai. The Atkin orthogonal polynomials for the Fricke groups of levels 5 and 7. International Journal of Number Theory (IJNT), 10(8):2243–2255, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500766.

Salle:2009:MPG

[Sal09] Landry Salle. Mild pro-p-groups as Galois groups over global fields. International Journal of Number Theory (IJNT), 5(5): 779–795, August 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002377.

Samuels:2010:FCU

[Sam10] Charles L. Samuels. The finiteness of computing the ultrametric Mahler measure. International Journal of Number Theory (IJNT), 6(8):1731–1753, December 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003745.

Sands:2009:VFM

[San09] Jonathan W. Sands. Values at s = −1 of L-functions for multi-quadratic extensions of number fields, and the fitting ideal of the tame kernel. International Journal of Number Theory (IJNT), 5(3):383–405, May 2009. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109002183.

Sanna:2014:CAP

[San14]  Carlo Sanna.  Covering an arithmetic

progression with geometric progressions and vice versa. International Journal of Number Theory (IJNT), 10(6):1577–1582, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042114500456.

Sanna:2015:NAF

[San15] Carlo Sanna. On the number of arithmetic formulas. International Journal of Number Theory (IJNT), 11(4):1099–1106, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115500591.

Sanna:2017:NDT

[San17] Carlo Sanna. On numbers n dividing the n-th term of a Lucas sequence. International Journal of Number Theory (IJNT), 13(3):725–734, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500373.

Sanna:2018:CBC

[San18a] Carlo Sanna. Central binomial coefficients divisible by or coprime to their indices. International Journal of Number Theory (IJNT), 14(4):1135–1141, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500707.

Sano:2018:GRA

[San18b]  Kaoru Sano.  Growth rate of am-

ple heights and the dynamical Mordell– Lang conjecture. International Journal of Number Theory (IJNT), 14(10):2673– 2685, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501610.

Sanna:2021:CMS

[San21] Carlo Sanna. On the l.c.m. of shifted Fibonacci numbers. International Journal of Number Theory (IJNT), 17(09): 2009–2018, October 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500743.

Sarashina:2022:CRA

[Sar22] Akira Sarashina. Conditional reconstruction of affine open subschemes of elliptic curves in positive characteristic by their geometric fundamental groups. International Journal of Number Theory (IJNT), 18(07): 1465–1490, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500750.

Sasaki:2015:ZMM

[Sas15] Yoshitaka Sasaki. Zeta Mahler measures, multiple zeta values and L-values. International Journal of Number Theory (IJNT), 11(7): 2239–2246, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115501006.

Saunders:2015:SDA

[Sau15] J. C. Saunders. Sums of digits in q-ary expansions. International Journal of Number Theory (IJNT), 11(2):593–611, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115500311.

Sun:2016:ATW

[SC16] Cui-Fang Sun and Zhi Cheng. On the addition of two weighted squares of units mod n. International Journal of Number Theory (IJNT), 12(7):1783–1790, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042116501098.

Scarponi:2017:RDZ

[Sca17] Danny Scarponi. The realization of the degree zero part of the Motivic polylogarithm on abelian schemes in Deligne– Beilinson cohomology. International Journal of Number Theory (IJNT), 13(9):2471– 2485, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117501378.

Schwab:2010:GAF

[Sch10] Emil Daniel Schwab. Generalized arithmetical functions of three variables. International Journal of Number Theory (IJNT), 6(7): 1689–1699, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003721.

Schweiger:2012:DAR

[Sch12] F. Schweiger. A 2-dimensional algorithm related to the Farey–Brocot sequence. International Journal of Number Theory (IJNT), 8 (1):149–160, February 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211250008X.

Schwab:2013:DCB

[Sch13] Emil Daniel Schwab. Dirichlet convolution, bicyclic semigroup and a breaking process. International Journal of Number Theory (IJNT), 9(8):1961–1972, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042113500681.

Schettler:2014:GIR

[Sch14] Jordan Schettler. Generalizations of Iwasawa’s “Riemann–Hurwitz” formula for cyclic p-extensions of number fields. International Journal of Number Theory (IJNT), 10(1): 219–233, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500905.

Schultz:2015:SPE

[Sch15] Dan Schultz. Solving polynomial equations with elliptic modular functions. International Journal of Number Theory (IJNT), 11 (4):1313–1344, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500712.

Schneider:2018:JTP

[Sch18a] Robert Schneider. Jacobi’s triple product, mock theta functions, unimodal sequences and the q-bracket. International Journal of Number Theory (IJNT), 14(7): 1961–1981, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501178.

Schweiger:2018:IMP

[Sch18b]  Fritz Schweiger.  Invariant measures of

piecewise fractional linear maps and piecewise quadratic maps. International Journal of Number Theory (IJNT), 14(6):1559– 1572, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304211850094X.

Schmid:2021:IPM

[Sch21a] Stefan Schmid. Integrality properties in the moduli space of elliptic curves: CM case. International Journal of Number Theory (IJNT), 17(07):1671–1696, August 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304212150055X.

Schonnenbeck:2021:CUG

[Sch21b] Sebastian Schönnenbeck. Computing Sunit groups of orders. International Journal of Number Theory (IJNT), 17(05): 1243–1263, June 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500342.

Sittinger:2020:PIN

[SD20] Brian D. Sittinger and Ryan D. Demoss. The probability that ideals in a number ring are k-wise relatively r-prime. International Journal of Number Theory (IJNT), 16(08): 1753–1765, September 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500918.

Sedunova:2015:AFS

[Sed15] Alisa Sedunova. On the asymptotic formulae for some multiplicative functions in short intervals. International Journal of Number Theory (IJNT), 11(5):1571–1587, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042115400102.

Sengun:2012:THN

[Şen12] Mehmet Haluk Şengün. On the torsion homology of non-arithmetic hyperbolic tetrahedral groups. International Journal of Number Theory (IJNT), 8(2):311–320, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112500182.

Song:2017:LSP

[SFM17] Kunkun Song, Lulu Fang, and Jihua Ma. Level sets of partial maximal digits for Lüroth expansion. International Journal of Number Theory (IJNT), 13(10):2777– 2790, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501536.

Salazar:2017:OES

[SG17] Daniel Barrera Salazar and Shan Gao. Overconvergent Eichler–Shimura isomorphisms for quaternionic modular forms over Q. International Journal of Number Theory (IJNT), 13(10):2687–2715, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117501494.

Schwab:2008:UFC

[SH08] Emil Daniel Schwab and Pentti Haukkanen. A unique factorization in commutative Möbius monoids. International Journal of Number Theory (IJNT), 4(4): 549–561, August 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001523.

Shavgulidze:2009:NRI

[Sha09] Ketevan Shavgulidze. On the number of representations of integers by the sums of quadratic forms. International Journal of Number Theory (IJNT), 5(3):515–525, May 2009. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042109002201.

Sha:2014:BII

[Sha14] Min Sha. Bounding the j-invariant of integral points on certain modular curves. International Journal of Number Theory (IJNT), 10(6):1545–1551, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042114500432.

Sharma:2017:LIG

[Sha17] Devika Sharma. Locally indecomposable Galois representations with full residual image. International Journal of Number Theory (IJNT), 13(5):1191–1211, June 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117500646.

Shevelev:2007:D

[She07] Vladimir Shevelev. On divisibility of \binom{n−i−1}{i−1} by i. International Journal of Number Theory (IJNT), 3(1):119–139, March 2007. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304210700078X.

Shekhar:2014:ECA

[She14] Sudhanshu Shekhar. Euler characteristic of Λ-adic forms over Kummer extensions. International Journal of Number Theory (IJNT), 10(2):401–419, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113501017.

Shen:2016:APR

[She16]  Erin Y. Y. Shen.  Arithmetic proper-

ties of l-regular overpartitions. International Journal of Number Theory (IJNT), 12(3):841–852, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500548.

Shen:2017:CMS

[She17] Erin Y. Y. Shen. Congruences modulo 9 for singular overpartitions. International Journal of Number Theory (IJNT), 13(3):717–724, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500361.

Shen:2022:ZDV

[She22] Qibin Shen. Zero distribution of v-adic multiple zeta values over F_{q}(t). International Journal of Number Theory (IJNT), 18(01): 131–139, February 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500105.

Shiomi:2016:UIF

[Shi16] Daisuke Shiomi. Unit indices of function fields of degree four. International Journal of Number Theory (IJNT), 12(7):1733– 1742, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116501062.

Shkredov:2018:DSS

[Shk18]  I. D. Shkredov.  Differences of sub-

groups in subgroups. International Journal of Number Theory (IJNT), 14(4):1111– 1134, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118500690.

Shlapentokh:2012:ECP

[Shl12] Alexandra Shlapentokh. Elliptic curve points and Diophantine models of Z in large subrings of number fields. International Journal of Number Theory (IJNT), 8(6):1335– 1365, September 2012. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112500789.

Shparlinski:2014:DPG

[Shp14] Igor E. Shparlinski. On the distribution of points on the generalized Markoff–Hurwitz and Dwork hypersurfaces. International Journal of Number Theory (IJNT), 10(1): 151–160, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500863.

Shparlinski:2018:TFD

[Shp18] Igor E. Shparlinski. Trilinear forms with double Kloosterman sums. International Journal of Number Theory (IJNT), 14(8): 2195–2203, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501312.

Shparlinski:2019:STD

[Shp19] Igor E. Shparlinski. The Sato–Tate distribution in thin families of elliptic curves over high degree extensions of finite fields. International Journal of Number Theory (IJNT), 15(3):469–477, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500246.

Siar:2015:SCG

[Şia15] Zafer Şiar. On square classes in generalized Lucas sequences. International Journal of Number Theory (IJNT), 11(2): 661–672, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500414.

Sills:2007:IRR

[Sil07] Andrew V. Sills. Identities of the Rogers– Ramanujan–Slater type. International Journal of Number Theory (IJNT), 3(2):293– 323, June 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042107000912.

Singh:2009:DPS

[Sin09] Jitender Singh. Defining power sums of n and ϕ(n) integers. International Journal of Number Theory (IJNT), 5(1):41–53, February 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304210900189X.

Sinick:2010:RCC

[Sin10]  Jonah Sinick.  Ramanujan congruences

for a class of eta quotients. International Journal of Number Theory (IJNT), 6(4):835–847, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003253.

Siurys:2016:IF

[Šiu16] Jonas Šiurys. On integers not of the form F_{n}±p^{a}. International Journal of Number Theory (IJNT), 12(2):505–512, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500317.

Shahverdian:2013:EFD

[SK13] Ashot Shahverdian and Adem Kilicman. Exact formula for distribution of sequences {ωn}. International Journal of Number Theory (IJNT), 9(1):179–187, February 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112501321.

Skalba:2017:NAF

[Ska17a] M. Skałba. Note on an analogue of Fermat’s two squares theorem for primes p ≡ 3(mod 8). International Journal of Number Theory (IJNT), 13(1):129–132, February 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117500087.

Skalba:2017:PDI

[Ska17b] M. Skałba. Products of distinct integers being high powers. International Journal of Number Theory (IJNT), 13(8):2093– 2096, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117501123.

Skalba:2019:PDB

[Ska19] M. Skałba. Products of disjoint blocks of integers being high powers. International Journal of Number Theory (IJNT), 15(1): 85–88, February 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118501749.

Skalba:2021:RPT

[Ska21] M. Skałba. Rational points on three mutually orthogonal great circles. International Journal of Number Theory (IJNT), 17 (04):925–929, May 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500159.

Smith:2011:VBD

[Smi11] Ethan Smith. A variant of the Barban– Davenport–Halberstam Theorem. International Journal of Number Theory (IJNT), 7(8):2203–2218, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211100499X.

Smilansky:2013:STS

[Smi13] Yotam Smilansky. Sums of two squares — pair correlation and distribution in short intervals. International Journal of Number Theory (IJNT), 9(7):1687–1711, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042113500516.

Smith:2021:LPD

[Smi21] Jerrod Manford Smith. Linear periods and distinguished local parameters. International Journal of Number Theory (IJNT), 17(05): 1067–1077, June 2021. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304212150024X.

Mao:2021:TPC

[sMT21] Guo shuai Mao and Roberto Tauraso. Three pairs of congruences concerning sums of central binomial coefficients. International Journal of Number Theory (IJNT), 17(10): 2301–2314, November 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500895.

Sodaigui:2021:SGR

[Sod21] Bouchaïb Sodaïgui. Sur la structure galoisienne relative de puissances de la différente et idéaux de Stickelberger. International Journal of Number Theory (IJNT), 17(07): 1645–1664, August 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500536.

Sofo:2018:ACE

[Sof18] Anthony Sofo. Alternating cubic Euler sums with binomial squared terms. International Journal of Number Theory (IJNT), 14(5):

1357–1374, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500859.

Somasunderam:2022:LTI

[Som22] Naveen Somasunderam. A LeVeque-type inequality on the ring of p-adic integers. International Journal of Number Theory (IJNT), 18(03):655–671, April 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304212250035X.

Souderes:2010:MDS

[Sou10] Ismael Soudères. Motivic double shuffle. International Journal of Number Theory (IJNT), 6(2):339–370, March 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110002995.

Schulze-pillot:2018:PPB

[SPY18] Rainer Schulze-Pillot and Abdullah Yenirce. Petersson products of bases of spaces of cusp forms and estimates for Fourier coefficients. International Journal of Number Theory (IJNT), 14(8):2277–2290, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042118501385.

Sharma:2019:NAI

[SR19] S. Sharma and M. Rana. A new approach in interpreting some mock theta functions. International Journal of Number Theory (IJNT), 15(7):1369–1383, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119500763.

Schmidt:2007:CLH

[SS07a] Thomas A. Schmidt and Mark Sheingorn. Classifying low height geodesics on Γ^{3} \ H. International Journal of Number Theory (IJNT), 3(3):421–438, September 2007. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042107001012.

Schmidt:2007:LHG

[SS07b] Thomas A. Schmidt and Mark Sheingorn. Low height geodesics on Γ \ H: Height formulas and examples. International Journal of Number Theory (IJNT), 3(3):475– 501, September 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304210700105X.

Sankaranarayanan:2008:ESC

[SS08] A. Sankaranarayanan and N. Saradha. Estimates for the solutions of certain Diophantine equations by Runge’s method. International Journal of Number Theory (IJNT), 4(3):475–493, June 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304210800147X.

Sbeity:2010:CSE

[SS10] Farah Sbeity and Bouchaïb Sodaïgui. Classes de Steinitz d’extensions non abéliennes à groupe de Galois d’ordre 16 ou extraspécial d’ordre 32 et problème de plongement. (French) [Steinitz classes of nonabelian extensions to Galois group of order 16 or extraspecial of order 32 and embedding problem]. International Journal of Number Theory (IJNT), 6(8):1769– 1783, December 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003794.

Saber:2014:EFV

[SS14] Hicham Saber and Abdellah Sebbar. Equivariant functions and vector-valued modular forms. International Journal of Number Theory (IJNT), 10(4):949–954, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042114500092.

Samuels:2015:OFR

[SS15a] Charles L. Samuels and Tanner J. Strunk. Optimal factorizations of rational numbers using factorization trees. International Journal of Number Theory (IJNT), 11(3):739–769, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500402.

Saradha:2015:SBA

[SS15b] N. Saradha and T. N. Shorey. Squares in blocks from an arithmetic progression and Galois group of Laguerre polynomials. International Journal of Number Theory (IJNT), 11 (1):233–250, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500141.

Shparlinski:2017:FEC

[SS17] Igor E. Shparlinski and Andrew V. Sutherland. Finding elliptic curves with a subgroup of prescribed size. International Journal of Number Theory (IJNT), 13(1): 133–152, February 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500099.

Schinzel:2018:CBE

[SS18] A. Schinzel and M. Skałba. On certain biquadratic equations, II. International Journal of Number Theory (IJNT), 14(7):

2095–2105, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501269.

Saha:2019:EMF

[SS19] Biswajyoti Saha and Ayyadurai Sankaranarayanan. On estimates of the Mertens function. International Journal of Number Theory (IJNT), 15(2):327–337, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042119500143.

Schneider:2021:ACP

[SS21] Robert Schneider and Andrew V. Sills. Analysis and combinatorics of partition zeta functions. International Journal of Number Theory (IJNT), 17(03):805–814, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042120400023.

Schwartz:2011:SAP

[SSD11] Ryan Schwartz, József Solymosi, and Frank De Zeeuw. Simultaneous arithmetic progressions on algebraic curves. International Journal of Number Theory (IJNT), 7(4):921–931, June 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004198.

Sourmelidis:2022:VDV

[SSS22] Athanasios Sourmelidis, Teerapat Srichan, and Jörn Steuding. On the vertical distribution of values of L-functions in the Selberg class. International Journal of Number Theory (IJNT), 18(02):277–302, March 2022. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042122500191.

Shokri:2019:GGT

[SST19] Khosro Monsef Shokri, Jafar Shaffaf, and Reza Taleb. Galois groups of Taylor polynomials of some elementary functions. International Journal of Number Theory (IJNT), 15 (6):1127–1141, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119500623.

Scheicher:2014:DSC

[SSTW14] Klaus Scheicher, Paul Surer, Jörg M. Thuswaldner, and Christiaan E. Van De Woestijne. Digit systems over commutative rings. International Journal of Number Theory (IJNT), 10(6):1459–1483, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042114500389.

Schlosser:2021:LCR

[SSU21] Michael J. Schlosser, Koushik Senapati, and Ali K. Uncu. Log-concavity results for a biparametric and an elliptic extension of the q-binomial coefficients. International Journal of Number Theory (IJNT), 17 (03):787–804, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120400187.

Sun:2011:SNC

[ST11] Zhi-Wei Sun and Roberto Tauraso. On some new congruences for binomial coefficients. International Journal of Number Theory (IJNT), 7(3):645–662, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042111004393.

Schwab:2018:GPN

[ST18] Nicholas Schwab and Lola Thompson. A generalization of the practical numbers. International Journal of Number Theory (IJNT), 14(5):1487–1503, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500902.

Sun:2019:AMA

[ST19] Cui-Fang Sun and Tian-Tian Tao. On g-adic minimal asymptotic bases of order h. International Journal of Number Theory (IJNT), 15(2):389–406, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500209.

Steiner:2016:UBS

[Ste16] Raphael S. Steiner. Uniform bounds on supnorms of holomorphic forms of real weight. International Journal of Number Theory (IJNT), 12(5):1163–1185, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500718.

Stockton:2021:CFD

[Sto21] Jacob J. Stockton. On the co-factors of degree 6 Salem number beta expansions. International Journal of Number Theory (IJNT), 17 (05):1175–1200, June 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500317.

Straub:2015:CFN

[Str15] Armin Straub. Congruences for Fishburn numbers modulo prime powers. International Journal of Number Theory (IJNT), 11(5): 1679–1690, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115400175.

Streipel:2022:TMT

[Str22] Jakob Streipel. Twisted moments of GL(3) × GL(2) L-functions. International Journal of Number Theory (IJNT), 18(06): 1301–1334, July 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304212250066X.

Shemanske:2010:CSH

[STW10] T. Shemanske, S. Treneer, and L. Walling. Constructing simultaneous Hecke eigenforms. International Journal of Number Theory (IJNT), 6(5):1117–1137, August 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042110003411.

Su:2016:ESK

[Su16] Ren-He Su. Eisenstein series in the Kohnen plus space for Hilbert modular forms. International Journal of Number Theory (IJNT), 12(3):691–723, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500469.

Suarez:2009:MLC

[Sua09] Ivan Suarez. Modular lattices over CM fields. International Journal of Number Theory (IJNT), 5(5):859–869, August 2009. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042109002420.

Sugiyama:2015:RNS

[Sug15] Rin Sugiyama. Remark on nondegeneracy of simple abelian varieties with many endomorphisms. International Journal of Number Theory (IJNT), 11(4):1217–1226, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115500657.

Sun:2012:ICN

[Sun12]  Zhi-Hong Sun.  Identities and congru-

ences for a new sequence. International Journal of Number Theory (IJNT), 8(1): 207–225, February 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500121.

Sun:2013:CCL

[Sun13]  Zhi-Hong Sun.  Congruences concerning

Legendre polynomials III. International Journal of Number Theory (IJNT), 9(4): 965–999, June 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500097.

Sun:2014:CCL

[Sun14]  Zhi-Hong Sun.  Congruences concern-

ing Lucas sequences. International Journal of Number Theory (IJNT), 10(3):793– 815, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042113501236.

Sun:2015:EDS

[Sun15a] Qingfeng Sun. On effective determination of symmetric-square lifts, level aspect. International Journal of Number Theory (IJNT), 11(1):51–65, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500037.

Sun:2015:CUC

[Sun15b] Zhi-Hong Sun. Congruences for q^{[p/8]}(modp) under the condition 4n^{2}p = x^{2} + qy^{2}. International Journal of Number Theory (IJNT), 11(4):1301–1312, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500700.

Sun:2015:SCC

[Sun15c] Zhi-Hong Sun. Super congruences concerning Bernoulli polynomials. International Journal of Number Theory (IJNT), 11(8): 2393–2404, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115501110.

Sun:2016:CCS

[Sun16a]  Zhi-Hong Sun.  Cubic congruences and

sums involving \binom{3k}{k}. International Journal of Number Theory (IJNT), 12(1):143– 164, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500093.

Sun:2016:SIB

[Sun16b] Zhi-Hong Sun. Supercongruences involving Bernoulli polynomials. International Journal of Number Theory (IJNT), 12(5): 1259–1271, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500779.

Sun:2016:PZF

[Sun16c] Zhi-Wei Sun. On a pair of zeta functions. International Journal of Number Theory (IJNT), 12(8):2323–2342, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042116501396.

Sun:2017:SDP

[Sun17a] Brian Yi Sun. On some divisibility properties of binomial sums. International Journal of Number Theory (IJNT), 13(9): 2265–2276, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211750124X.

Sun:2017:ICC

[Sun17b] Zhi-Hong Sun. Identities and congruences for Catalan–Larcombe–French numbers. International Journal of Number Theory (IJNT), 13(4):835–851, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500440.

Sun:2017:SFP

[Sun17c] Zhi-Hong Sun. Some further properties of even and odd sequences. International Journal of Number Theory (IJNT), 13(6): 1419–1442, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500798.

Sun:2018:CSI

[Sun18]  Zhi-Hong Sun.  Congruences for sums

involving Franel numbers. International Journal of Number Theory (IJNT), 14(1): 123–142, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500094.

Sun:2019:NRL

[Sun19a] Zhi-Hong Sun. The number of representations of n as a linear combination of triangular numbers. International Journal of Number Theory (IJNT), 15(6):1191–1218, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042119500660.

Sun:2019:RTF

[Sun19b] Zhi-Hong Sun. Ramanujan’s theta functions and sums of triangular numbers. International Journal of Number Theory (IJNT), 15(5):969–989, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500520.

Sun:2019:RSF

[Sun19c] Zhi-Wei Sun. Restricted sums of four squares. International Journal of Number Theory (IJNT), 15(9):1863–1893, October 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119501045.

Surer:2020:SNS

[Sur20]  Paul Surer.  Substitutive number sys-

tems. International Journal of Number Theory (IJNT), 16(08):1709–1751, September 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

                                        worldscientific.com/doi/10.1142/S1793042120500906.

Suzuki:2005:RBZ

[Suz05] Masatoshi Suzuki. A relation between the zeros of two different L-functions which have an Euler product and functional equation. International Journal of Number Theory (IJNT), 1 (3):401–429, September 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042105000248.

Suzuki:2017:MVR

[Suz17] Yuta Suzuki. A mean value of the representation function for the sum of two primes in arithmetic progressions. International Journal of Number Theory (IJNT), 13(4):977–990, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500518.

Suzuki:2022:NSI

[Suz22] Koji Suzuki. On the number of semismooth integers. International Journal of Number Theory (IJNT), 18(05):999–1011, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304212250052X.

Schwagenscheidt:2015:LNV

[SV15] Markus Schwagenscheidt and Fabian Völz. Lifting newforms to vector-valued modular forms for the Weil representation. International Journal of Number Theory (IJNT), 11(7):2199–2219, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042115500980.

Stucker:2020:DSA

[SVY20] T. Stucker, A. Vennos, and M. P. Young. Dedekind sums arising from newform Eisenstein series. International Journal of Number Theory (IJNT), 16(10):2129–2139, November 2020. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120501092.

Spearman:2006:DCC

[SW06] Blair K. Spearman and Kenneth S. Williams. On the distribution of cyclic cubic fields with index 2. International Journal of Number Theory (IJNT), 2(2):235–247, June 2006. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042106000541.

Sun:2008:LTC

[SW08] Zhi-Wei Sun and Daqing Wan. Lucastype congruences for cyclotomic ψ-coefficients. International Journal of Number Theory (IJNT), 4(2):155–170, April 2008. CO-DEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042108001286.

Sun:2013:ESC

[SW13] Zhi-Hong Sun and Lin-Lin Wang. An extension of Stern’s congruence. International Journal of Number Theory (IJNT), 9 (2):413–419, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501400.

Sun:2014:DRC

[SW14] Yu Sun and Jun Wu. A dimensional result in continued fractions. International Journal of Number Theory (IJNT), 10(4):849–857, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211450002X.

Shen:2014:ICL

[SX14] Caixia Shen and Li-Meng Xia. On the index-conjecture of length four minimal zerosum sequences II. International Journal of Number Theory (IJNT), 10(3):601– 622, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304211350111X.

Shen:2013:LDP

[SXJ13] Luming Shen, Jian Xu, and Huiping Jing. On the largest degree of the partial quotients in continued fraction expansions over the field of formal Laurent series. International Journal of Number Theory (IJNT), 9 (5):1237–1247, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500231.

Sun:2014:SAC

[SY14] Cui-Fang Sun and Quan-Hui Yang. On the sumset of atoms in cyclic groups. International Journal of Number Theory (IJNT), 10(6): 1355–1363, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042114500328.

Shparlinski:2017:DES

[SY17] Igor E. Shparlinski and Kam Hung Yau. Double exponential sums with exponential functions. International Journal of Number Theory (IJNT), 13(10):2531–2543, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304211750141X.

Seki:2019:NPD

[SY19a] Shin-Ichiro Seki and Shuji Yamamoto. A new proof of the duality of multiple zeta values and its generalizations. International Journal of Number Theory (IJNT), 15 (6):1261–1265, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500702.

Sun:2019:FIB

[SY19b] Haiwei Sun and Yangbo Ye. Further improvement on bounds for L-functions related to GL(3). International Journal of Number Theory (IJNT), 15(7):1487–1517, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119500866.

Shen:2014:NLD

[SYZ14] Luming Shen, Yiying Yu, and Yuxin Zhou. A note on the largest digits in Lüroth expansion. International Journal of Number Theory (IJNT), 10(4):1015–1023, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042114500122.

Sole:2007:MFC

[SZ07]  Patrick Solé and Dmitrii Zinoviev.  A

Macwilliams formula for convolutional codes. International Journal of Number Theory (IJNT), 3(2):191–206, June 2007. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042107000869.

Szalay:2015:DQ

[SZ15]  László Szalay and Volker Ziegler.  S-

Diophantine quadruples with S = {2, q}. International Journal of Number Theory (IJNT), 11(3):849–868, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115500475.

Sellers:2021:PNP

[SZ21] James A. Sellers and Fabrizio Zanello. On the parity of the number of partitions with odd multiplicities. International Journal of Number Theory (IJNT), 17(07): 1717–1728, August 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500573.

Szczuka:2015:CAP

[Szc15] Paulina Szczuka. The closures of arithmetic progressions in Kirch’s topology on the set of positive integers. International Journal of Number Theory (IJNT), 11(3):673–682, May 2015. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500360.

Szczuka:2016:PDT

[Szc16] Paulina Szczuka. Properties of the division topology on the set of positive integers. International Journal of Number Theory (IJNT), 12(3):775–785, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500500.

Sze:2019:ESI

[Sze19] Sandra Sze. Explicit solutions of imaginary quadratic norm equations. International Journal of Number Theory (IJNT), 15(8): 1635–1673, September 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500921.

Steinberger:2013:AFG

[SZZ13] Thomas Steinberger, Alexandru Zaharescu, and Mohammad Zaki. Arithmetic functions on Gaussian integers. International Journal of Number Theory (IJNT), 9(8): 1923–1932, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500668.

Sui:2018:RIC

[SZZ18] Yankun Sui, Wenguang Zhai, and Deyu Zhang. Ω-result for the index of composition of an integer. International Journal of Number Theory (IJNT), 14(2):339– 348, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118500215.

Takahashi:2008:APM

[Tak08] S. Takahashi. p-adic periods of modular elliptic curves and the level-lowering theorem. International Journal of Number Theory (IJNT), 4(1):15–23, February 2008. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042108001183.

Takeshi:2015:ECG

[Tak15] Nao Takeshi. Elliptic curves with good reduction everywhere over cubic fields. International Journal of Number Theory (IJNT), 11(4):1149–1164, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500621.

Takahashi:2021:GGC

[Tak21] Naoya Takahashi. On Greenberg’s generalized conjecture for imaginary quartic fields. International Journal of Number Theory (IJNT), 17(05):1163–1173, June 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500305.

Tamam:2014:FMD

[Tam14] Nattalie Tamam. The fourth moment of Dirichlet L-functions for the rational function field. International Journal of Number Theory (IJNT), 10(1):183–218, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042113500899.

Tanner:2009:SCP

[Tan09]  Noam Tanner.  Strings of consecutive

primes in function fields. International Journal of Number Theory (IJNT), 5(1): 81–88, February 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109001918.

Tanabe:2012:NVD

[Tan12] Naomi Tanabe. Non-vanishing of derivatives of L-functions attached to Hilbert modular forms. International Journal of Number Theory (IJNT), 8(4):1099–1105, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112500662.

Tang:2014:GEF

[Tan14] Min Tang. On a generalization of the Erdős–Fuchs theorem. International Journal of Number Theory (IJNT), 10(4):955– 961, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500109.

Tang:2018:ASD

[Tan18] Yunqing Tang. Algebraic solutions of differential equations over P^{1} − {0, 1, ∞}. International Journal of Number Theory (IJNT), 14(5):1427–1457, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500884.

Tang:2019:VCS

[Tan19] Dazhao Tang. Vanishing coefficients in some q-series expansions. International Journal of Number Theory (IJNT), 15 (4):763–773, May 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500398.

Tauraso:2018:SRI

[Tau18] Roberto Tauraso. Supercongruences related to _{3}F_{2}(1) involving harmonic numbers. International Journal of Number Theory (IJNT), 14(4):1093–1109, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500689.

Taylor:2009:ACN

[Tay09] Karen Taylor. Analytic continuation of nonanalytic vector-valued Eisenstein series. International Journal of Number Theory (IJNT), 5 (5):805–830, August 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002407.

Tayou:2017:IRG

[Tay17] Salim Tayou. Images de représentations galoisiennes associées à certaines formes modulaires de Siegel de genre 2. (French) [Images of Galois representations associated with certain Siegel modular forms of genus 2]. International Journal of Number Theory (IJNT), 13(5):1129–1144, June 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500610.

Takloo-Bighash:2006:RPA

[TB06] Ramin Takloo-Bighash. A remark on a paper of Ahlgren, Berndt, Yee, and Zaharescu. International Journal of Number Theory (IJNT), 2(1):111–114, March 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042106000450.

Templier:2010:AVC

[Tem10] Nicolas Templier. On asymptotic values of canonical quadratic L-functions. International Journal of Number Theory (IJNT), 6(8):1717– 1730, December 2010. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042110003678.

Terracini:2013:WIR

[Ter13] Lea Terracini. A weight independence result for quaternionic Hecke algebras. International Journal of Number Theory (IJNT), 9(8): 1895–1922, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500656.

Thong:2006:CFG

[Tho06] Nguyen Quang Do Thong. Sur la conjecture faible de Greenberg dans le cas abélien p-décomposé. (French) [On the weak conjecture of Greenberg in the abelian p-decomposed case]. International Journal of Number Theory (IJNT), 2(1):49–64, March 2006. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042106000395.

Thompson:2013:D

[Tho13] Lola Thompson. On the divisors of x^{n} − 1 in F_{p}[x]. International Journal of Number Theory (IJNT), 9(2):421–430, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112501412.

Thunder:2008:PBH

[Thu08] Jeffrey Lin Thunder. Points of bounded height on Schubert varieties. International Journal of Number Theory (IJNT), 4(5): 757–765, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001742.

Tanti:2010:ECS

[TK10] Jagmohan Tanti and S. A. Katre. Euler’s criterion for septic nonresidues. International Journal of Number Theory (IJNT), 6(6): 1329–1347, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003563.

Tan:2022:RCS

[TLZ22] Xiaoyan Tan, Jia Liu, and Zhenliang Zhang. Relative convergence speed of Lüroth expansions and Hausdorff dimension. International Journal of Number Theory (IJNT), 18(05):977–997, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500518.

Toh:2008:GTO

[Toh08] Pee Choon Toh. Generalized m-th order Jacobi theta functions and the Macdonaldcg identities. International Journal of Number Theory (IJNT), 4(3):461– 474, June 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042108001456.

Toh:2013:RFN

[Toh13] Pee Choon Toh. On representations by figurate numbers: a uniform approach to the conjectures of Melham. International Journal of Number Theory (IJNT), 9(4): 1055–1071, June 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500127.

Tolev:2006:DTS

[Tol06] D. I. Tolev. On the distribution of r-tuples of squarefree numbers in short intervals. International Journal of Number Theory (IJNT), 2(2):225–234, June 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304210600053X.

Toth:2013:TGB

[Tót13] László Tóth. Two generalizations of the Busche–Ramanujan identities. International Journal of Number Theory (IJNT), 9(5): 1301–1311, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500280.

Toth:2018:MTI

[Tót18] László Tóth. Menon-type identities concerning Dirichlet characters. International Journal of Number Theory (IJNT), 14(4): 1047–1054, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211850063X.

Toulmonde:2009:CAV

[Tou09] Vincent Toulmonde. Comportement au voisinage de 1 de la fonction de répartition de ϕ(n)/n. (French) [Behavior in the neighborhood of 1 of the partition function ϕ(n)/n]. International Journal of Number Theory (IJNT), 5(8):1347–1384, December 2009. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109001414.

Tou:2019:ACT

[Tou19] Erik R. Tou. Asymptotic counting theorems for primitive juggling patterns. International Journal of Number Theory (IJNT), 15(5):1037–1050, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500568.

Towsley:2013:HPP

[Tow13] Adam Towsley. A Hasse principle for periodic points. International Journal of Number Theory (IJNT), 9(8):2053–2068, December 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042113500747.

Tran:2017:CDS

[Tra17] Ha Thanh Nguyen Tran. Computing dimensions of spaces of Arakelov divisors of number fields. International Journal of Number Theory (IJNT), 13(2):487–512, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117500270.

Tretkoff:2015:SAP

[Tre15a] Paula Tretkoff. K3 surfaces with algebraic period ratios have complex multiplication. International Journal of Number Theory (IJNT), 11(5):1709–1724, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042115400217.

Trevino:2015:BIL

[Tre15b] Enrique Treviño. The Burgess inequality and the least k-th power non-residue. International Journal of Number Theory (IJNT), 11 (5):1653–1678, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115400163.

Troupe:2017:ORE

[Tro17] Lee Troupe. Orders of reductions of elliptic curves with many and few prime factors. International Journal of Number Theory (IJNT), 13(8):2115–2134, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117501147.

Tsang:2017:RCS

[Tsa17] Sin Yi Cindy Tsang. On the realizable classes of the square root of the inverse different in the unitary class group. International Journal of Number Theory (IJNT), 13(4):913–932, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500476.

Tripathi:2020:AHS

[TSB20] Mohit Tripathi, Neelam Saikia, and Rupam Barman. Appell’s hypergeometric series over finite fields. International Journal of Number Theory (IJNT), 16(4):673–692, May 2020. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500347.

Tsukerman:2015:EDS

[Tsu15] Emmanuel Tsukerman. Equality of Dedekind sums mod Z, 2Z and 4Z. International Journal of Number Theory (IJNT), 11(6): 1735–1738, September 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500748.

Tsumura:2018:SIA

[Tsu18] Hirofumi Tsumura. On series identities arising from Jacobi’s identity of the theta function. International Journal of Number Theory (IJNT), 14(5):1317–1327, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118500823.

Tretkoff:2009:TSV

[TT09] Marvin D. Tretkoff and Paula Tretkoff. Transcendence of special values of Pochhammer functions. International Journal of Number Theory (IJNT), 5(4):667–677, June 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304210900233X.

Taniguchi:2014:EEC

[TT14] Takashi Taniguchi and Frank Thorne. An error estimate for counting S_{3}-sextic number fields. International Journal of Number Theory (IJNT), 10(4):935–948, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042114500080.

Tanaka:2018:AIV

[TT18a] Taka-Aki Tanaka and Yusuke Tanuma. Algebraic independence of the values of the Hecke–Mahler series and its derivatives at algebraic numbers. International Journal of Number Theory (IJNT), 14(9):2369– 2384, October 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118501440.

Tran:2018:SFC

[TT18b] Ha Thanh Nguyen Tran and Peng Tian. The size function for cyclic cubic fields. International Journal of Number Theory (IJNT), 14(2):399–415, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500276.

Tu:2011:ONA

[Tu11] Fang-Ting Tu. On orders of M(2, K) over a non–Archimedean local field. International Journal of Number Theory (IJNT), 7 (5):1137–1149, August 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004654.

Tengely:2020:DEF

[TU20] Szabolcs Tengely and MacIej Ulas. The Diophantine equation F_{n} = P (x). International Journal of Number Theory (IJNT), 16(09): 2095–2111, October 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120501079.

Turkelli:2011:CMC

[Tür11] Seyfi Türkelli. Counting multisections in conic bundles over a curve defined over F_{q}. International Journal of Number Theory (IJNT), 7(6):1663–1680, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211100485X.

Turkelli:2018:NTC

[Tür18]  Seyfi Türkelli.  A note on a theorem

of Cadoret and Tamagawa. International Journal of Number Theory (IJNT), 14(2): 349–353, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500227.

Tang:2019:RS

[TW19]  Min Tang and Meng-Ting Wei.  Re-

stricted sumsets in Z_{n}. International Journal of Number Theory (IJNT), 15(10):2057– 2065, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119501124.

Tang:2016:IMA

[TX16] Hengcai Tang and Xuanxuan Xiao. Integral moments of automorphic L-functions. International Journal of Number Theory (IJNT), 12(7):1827–1843, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211650113X.

Tasaka:2013:SMZ

[TY13] Koji Tasaka and Shuji Yamamoto. On some multiple zeta-star values of one-two-three indices. International Journal of Number Theory (IJNT), 9(5):1171–1184, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042113500188.

Tong:2016:MLC

[TYZ16] Xin Tong, Yueli Yu, and Yanfen Zhao. On the maximal length of consecutive zero digits of β-expansions. International Journal of Number Theory (IJNT), 12(3):625– 633, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042116500408.

Tanigawa:2009:FPM

[TZ09] Yoshio Tanigawa and Wenguang Zhai. On the fourth power moment of Δx and E(x) in short intervals. International Journal of Number Theory (IJNT), 5(2):355– 382, March 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042109002055.

Tsai:2012:SCIa

[TZ12a] Mu-Tsun Tsai and Alexandru Zaharescu. On the sum of consecutive integers in sequences. International Journal of Number Theory (IJNT), 8(3):643–652, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112500364.

Tsai:2012:SCIb

[TZ12b] Mu-Tsun Tsai and Alexandru Zaharescu. On the sum of consecutive integers in sequences II. International Journal of Number Theory (IJNT), 8(5):1281–1299, August 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112500753.

Tang:2017:SSP

[TZ17] Liang Tang and Ting Zhong. Some stochastic properties on the GCF_{ϵ} expansion when ϵ(k_{n}) = ck_{n}^{τ} + ck_{n}^{τ−1}. International Journal of Number Theory (IJNT), 13(4): 1071–1081, May 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500579.

Tan:2021:DTP

[TZ21]  Bo Tan and Qinglong Zhou.  Dimen-

sion theory of the product of partial quotients in Lüroth expansions. International Journal of Number Theory (IJNT), 17(05): 1139–1154, June 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500287.

Ulas:2010:VHT

[Ula10] Maciej Ulas. Variations on higher twists of pairs of elliptic curves. International Journal of Number Theory (IJNT), 6(5): 1183–1189, August 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003472.

Ulas:2012:APS

[Ula12] Maciej Ulas. Arithmetic properties of the sequence of degrees of Stern polynomials and related results. International Journal of Number Theory (IJNT), 8(3):669–687, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112500388.

Ulas:2019:AVC

[Ula19] Maciej Ulas. 2-adic valuations of coefficients of certain integer powers of formal power series. International Journal of Number Theory (IJNT), 15(4):723–762, May 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042119500386.

Ulas:2013:NES

[US13] Maciej Ulas and Andrzej Schinzel. A note on Erdős–Straus and Erdős–Graham divisibility problems (with an appendix by Andrzej Schinzel). International Journal of Number Theory (IJNT), 9(3):583–599, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042112501497.

Uygul:2014:ARS

[UW14] Faruk Uygul and Kenneth S. Williams. Analogues of Ramanujan’s 24 squares formula. International Journal of Number Theory (IJNT), 10(5):1099–1119, August 2014. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500171.

Vaananen:2015:RAC

[Vää15] Keijo Väänänen. On rational approximations of certain Mahler functions with a connection to the Thue–Morse sequence. International Journal of Number Theory (IJNT), 11 (2):487–493, March 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500244.

Vaananen:2016:AIC

[Vää16] Keijo Väänänen. Algebraic independence of certain Mahler numbers. International Journal of Number Theory (IJNT), 12(8): 2159–2166, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116501293.

Vaananen:2018:PAG

[Vää18]  Keijo Väänänen.  On Padé approxima-

tions and global relations of some Eulertype series. International Journal of Number Theory (IJNT), 14(8):2303–2315, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042118501403.

Vallieres:2014:RRI

[Val14] Daniel Vallières. A remark on a result of Iwasawa on Hecke characters of type A_{0}. International Journal of Number Theory (IJNT), 10(8):1939–1953, December 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304211450064X.

VanWamelen:2007:NEM

[Van07] Paul Van Wamelen. New explicit multiplicative relations between Gauss sums. International Journal of Number Theory (IJNT), 3(2):275–292, June 2007. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042107000900.

VanGorder:2012:GTP

[Van12a]  Robert A. Van Gorder.  Glaisher-type

products over the primes. International Journal of Number Theory (IJNT), 8(2): 543–550, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500297.

VanOrder:2012:QAF

[Van12b] Jeanine Van Order. On the quaternionic p-adic L-functions associated to Hilbert modular eigenforms. International Journal of Number Theory (IJNT), 8(4):1005– 1039, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/

S1793042112500601. See erratum [Van16a].

VanOrder:2016:EQA

[Van16a] Jeanine Van Order. Erratum: On the quaternionic p-adic L-functions associated to Hilbert modular eigenforms. International Journal of Number Theory (IJNT), 12(1): 305–311, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116920019. See [Van12b].

Vandehey:2016:DPC

[Van16b] Joseph Vandehey. Diophantine properties of continued fractions on the Heisenberg group. International Journal of Number Theory (IJNT), 12(2):541–560, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042116500342.

VanBommel:2018:IGP

[Van18] Raymond Van Bommel. Inverse Galois problem for ordinary curves. International Journal of Number Theory (IJNT), 14(5): 1305–1315, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500811.

VanHille:2021:FMF

[Van21a] Siegfried Van Hille. On a family of mild functions. International Journal of Number Theory (IJNT), 17(06):1379–1390, July 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S179304212150041X.

VanLangen:2021:SFP

[Van21b] Joey M. Van Langen. On the sum of fourth powers in arithmetic progression. International Journal of Number Theory (IJNT), 17 (01):191–221, February 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500093.

Vargas:2022:ACP

[Var22] João Campos Vargas. On Artin’s conjecture for pairs of diagonal forms. International Journal of Number Theory (IJNT), 18 (03):613–628, April 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500336.

Vaughan:2015:SAQ

[Vau15] Robert C. Vaughan. Squares: Additive questions and partitions. International Journal of Number Theory (IJNT), 11(5): 1367–1409, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115400096.

Vasuki:2019:CQR

[VB19] K. R. Vasuki and E. N. Bhuvan. “Cases? for ϕ also” a question raised by Ramanujan. International Journal of Number Theory (IJNT), 15(1):183–188, February 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042119500076.

Vega:2011:HFF

[Veg11] M. Valentina Vega. Hypergeometric functions over finite fields and their relations to algebraic curves. International Journal of Number Theory (IJNT), 7(8):2171– 2195, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004976.

Vepkhvadze:2017:NRP

[Vep17] Teimuraz Vepkhvadze. On the number of representations of positive integers by the diagonal quadratic forms in nine variables with coefficients that are ones and twos. International Journal of Number Theory (IJNT), 13(8): 2135–2154, September 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501159.

Verrill:2010:CRM

[Ver10]  H. A. Verrill.  Congruences related to

modular forms. International Journal of Number Theory (IJNT), 6(6):1367–1390, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042110003587.

Viada:2010:LBN

[Via10] Evelina Viada. Lower bounds for the normalized height and non-dense subsets of subvarieties of Abelian varieties. International Journal of Number Theory (IJNT), 6(3):471–499, May 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003010.

Vienney:2010:NCA

[Vie10]  Mathieu Vienney.  A new construction

of p-adic Rankin convolutions in the case of positive slope. International Journal of Number Theory (IJNT), 6(8):1875– 1900, December 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003782.

Viguie:2012:GUM

[Vig12] Stéphane Viguié. Global units modulo elliptic units and ideal class groups. International Journal of Number Theory (IJNT), 8(3):569– 588, May 2012. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500315.

Villacorta:2018:DFI

[Vil18a]  Jorge Luis Cimadevilla Villacorta.  Di-

visor function, inequalities and arithmetic progressions. International Journal of Number Theory (IJNT), 14(8):2225–2238, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042118501348.

Villacorta:2018:JPR

[Vil18b]  Jorge Luis Cimadevilla Villacorta.  On

Jackson’s proof of Ramanujan’s _{1}ψ_{1} summation formula. International Journal of Number Theory (IJNT), 14(2):313–328, March 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118500197.

Vinh:2014:TVE

[Vin14]  Le Anh Vinh.  On three-variable ex-

panders over finite fields. International Journal of Number Theory (IJNT), 10 (3):689–703, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113501182.

Vinas:2019:SCC

[Viñ19] Víctor Julio Ramírez Viñas. A simple criterion for the class number of a quadratic number field to be one. International Journal of Number Theory (IJNT), 15(9): 1857–1862, October 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119501033.

Volkov:2010:ASS

[Vol10]  Maja Volkov.  Abelian surfaces with

supersingular good reduction and nonsemisimple Tate module. International Journal of Number Theory (IJNT), 6(4): 811–818, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003228.

Veras:2020:EV

[VR20] Daiane S. Veras and Paulo H. A. Rodrigues. Exact values of Γ^{∗}(10, p). International Journal of Number Theory (IJNT), 16(3):639–649, April 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500323.

Villa-Salvador:2010:EPC

[VS10] Gabriel Villa-Salvador. An elementary proof of the conductor–discriminant formula. International Journal of Number Theory (IJNT), 6 (5):1191–1197, August 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003459.

Viu-sos:2021:SCR

[Vs21] Juan Viu-sos. A semi-canonical reduction for periods of Kontsevich–zagier. International Journal of Number Theory (IJNT), 17 (01):147–174, February 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304212150007X.

Vsemirnov:2014:CGK

[Vse14] Maxim Vsemirnov. On a conjecture of Guo and Krattenthaler. International Journal of Number Theory (IJNT), 10(6):1541– 1543, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500420.

Vaserstein:2010:PPP

[VSF10] Leonid Vaserstein, Takis Sakkalis, and Sophie Frisch. Polynomial parametrization of Pythagorean tuples. International Journal of Number Theory (IJNT), 6(6):1261– 1272, September 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042110003496.

Vulakh:2006:DA

[Vul06] L. Ya. Vulakh. Diophantine approximation in Q(\sqrt{−5}) and Q(\sqrt{−5}). International Journal of Number Theory (IJNT), 2(1):25–48, March 2006. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042106000462.

Vulakh:2009:MSC

[Vul09] L. Ya. Vulakh. The Markov spectra for cocompact Fuchsian groups. International Journal of Number Theory (IJNT), 5(4): 679–718, June 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002341.

Vulakh:2010:DAI

[Vul10a] L. Ya. Vulakh. Diophantine approximation in imaginary quadratic fields. International Journal of Number Theory (IJNT), 6(4):731–766, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003137.

Vulakh:2010:HPM

[Vul10b] L. Ya Vulakh. Hermitian points in Markov spectra. International Journal of Number Theory (IJNT), 6(4):713–730, June 2010. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003186.

Vulakh:2010:MBI

[Vul10c] L. Ya. Vulakh. Minima of binary indefinite Hermitian forms. International Journal of Number Theory (IJNT), 6(2): 411–435, March 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003034.

Vatwani:2017:PPC

[VW17] Akshaa Vatwani and Peng-Jie Wong. Patterns of primes in Chebotarev sets. International Journal of Number Theory (IJNT), 13 (7):1651–1677, August 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500956.

Vilardi:2018:DEL

[VX18] Trevor Vilardi and Hui Xue. Distinguishing eigenforms of level one. International Journal of Number Theory (IJNT), 14(1): 31–36, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118500033.

Voutier:2013:LCS

[VY13] Paul Voutier and Minoru Yabuta. Lang’s conjecture and sharp height estimates for the elliptic curves y^{2} = x^{3} + ax. International Journal of Number Theory (IJNT), 9 (5):1141–1170, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500176.

Voloch:2014:RPS

[VZ14] José Felipe Voloch and Michael E. Zieve. Rational points on some Fermat curves and surfaces over finite fields. International Journal of Number Theory (IJNT), 10(2): 319–325, March 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500954.

Wakabayashi:2012:NSQ

[Wak12a] Isao Wakabayashi. Number of solutions for quartic simple Thue equations. International Journal of Number Theory (IJNT), 8(6): 1367–1386, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500790.

Wakabayashi:2012:HRM

[Wak12b] Noriko Wakabayashi. On Hoffman’s relation for multiple zeta-star values. International Journal of Number Theory (IJNT), 8(8): 1971–1976, December 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501114.

Wakabayashi:2017:DSH

[Wak17] Noriko Wakabayashi. Double shuffle and Hoffman’s relations for interpolated multiple zeta values. International Journal of Number Theory (IJNT), 13(9):2245–2251, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042117501226.

Walling:2006:AHO

[Wal06] Lynne H. Walling. Action of Hecke operators on Siegel theta series I. International Journal of Number Theory (IJNT), 2(2):169–186, June 2006. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042106000516.

Walling:2008:AHO

[Wal08] Lynne H. Walling. Action of Hecke operators on Siegel theta series, II. International Journal of Number Theory (IJNT), 4 (6):981–1008, December 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001845.

Waldherr:2013:AMH

[Wal13] Matthias Waldherr. Asymptotics for moments of higher ranks. International Journal of Number Theory (IJNT), 9(3):675– 712, May 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042112501552.

Walling:2017:HER

[Wal17a] Lynne H. Walling. Hecke eigenvalues and relations for Siegel Eisenstein series of arbitrary degree, level, and character. International Journal of Number Theory (IJNT), 13 (2):325–370, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211750021X.

Walling:2017:HOH

[Wal17b] Lynne H. Walling. Hecke operators on half-integral weight Siegel Eisenstein series. International Journal of Number Theory (IJNT), 13(9):2335–2372, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117501287.

Walji:2021:CRS

[Wal21] Nahid Walji. A conjectural refinement of strong multiplicity one for GL(n). International Journal of Number Theory (IJNT), 17 (07):1583–1598, August 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500494.

Wang:2012:HPF

[Wan12] Julie Tzu-Yueh Wang. Hensley’s problem for function fields. International Journal of Number Theory (IJNT), 8(2):507– 524, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042112500261.

Wang:2015:APP

[Wan15a] Liuquan Wang. Arithmetic properties of partition triples with odd parts distinct. International Journal of Number Theory (IJNT), 11(6):1791–1805, September 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042115500773.

Wang:2015:MHD

[Wan15b] Shaozu Wang. Maximal height of divisors of x^{pq^{b}}−1. International Journal of Number Theory (IJNT), 11(1):67–79, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042115500049.

Wang:2016:AIC

[Wan16] Liuquan Wang. Arithmetic identities and congruences for partition triples with 3cores. International Journal of Number Theory (IJNT), 12(4):995–1010, June 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500627.

Wang:2017:IQL

[Wan17] Ruiqing Wang. On indecomposable quadratic lattices over global function fields. International Journal of Number Theory (IJNT), 13 (6):1611–1616, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500919.

Wang:2018:KSP

[Wan18]  Liyuan Wang.  The Korselt set of a

power of a prime. International Journal of Number Theory (IJNT), 14(1):233– 240, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500148.

Wang:2020:AFC

[Wan20] Biao Wang. An analogue of a formula for Chebotarev densities. International Journal of Number Theory (IJNT), 16(7): 1557–1565, August 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500815.

Wang:2021:SLI

[Wan21] Guoqing Wang. Structure of long idempotentsum-free sequences over finite cyclic semigroups. International Journal of Number Theory (IJNT), 17(04):871–886, May 2021. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042121500123.

Watt:2008:BMV

[Wat08] Nigel Watt. Bounds for a mean value of character sums. International Journal of Number Theory (IJNT), 4(2):249– 293, April 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042108001328.

Wang:2011:API

[WC11] Xinna Wang and Yingchun Cai. An additive problem involving Piatetski–Shapiro primes. International Journal of Number Theory (IJNT), 7(5):1359–1378, August 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042111004630.

Wang:2019:STF

[WC19] Chun Wang and Shane Chern. Some qtransformation formulas and Hecke type identities. International Journal of Number Theory (IJNT), 15(7):1349–1367, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042119500751.

Wear:2022:PER

[Wea22]  Peter Wear.  Properties of extended

Robba rings. International Journal of Number Theory (IJNT), 18(08):1765–1789, September 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042122500907.

Wei:2015:EST

[Wei15a] Bin Wei. Exponential sums twisted by Fourier coefficients of automorphic cusp forms for SL(2, Z). International Journal of Number Theory (IJNT), 11(1):39– 49, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042115500025.

Weitzer:2015:CAS

[Wei15b] Mario Weitzer. Characterization algorithms for shift radix systems with finiteness property. International Journal of Number Theory (IJNT), 11(1):211–232, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211550013X.

Weidner:2019:CRP

[Wei19] Matthew Weidner. On conjectural rank parities of quartic and sextic twists of elliptic curves. International Journal of Number Theory (IJNT), 15(9):1895–1918, October 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119501057.

Weingartner:2020:CFA

[Wei20] Andreas Weingartner. The constant factor in the asymptotic for practical numbers. International Journal of Number Theory (IJNT), 16(3):629–638, April 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500311.

Williams:2011:SFO

[WG11] H. C. Williams and R. K. Guy. Some fourthorder linear divisibility sequences. International Journal of Number Theory (IJNT), 7 (5):1255–1277, August 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042111004587.

Wang:2019:EBC

[WHZ19] Haoli Wang, Jun Hao, and Lizhen Zhang. The Erdős–Burgess constant of the multiplicative semigroup of a factor ring of F_{q}[x]. International Journal of Number Theory (IJNT), 15 (1):131–136, February 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119500015.

Widmer:2011:NFN

[Wid11] Martin Widmer. On number fields with nontrivial subfields. International Journal of Number Theory (IJNT), 7(3):695– 720, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042111004204.

Wiese:2009:MSC

[Wie09] Gabor Wiese. On modular symbols and the cohomology of Hecke triangle surfaces. International Journal of Number Theory (IJNT), 5(1):89–108, February 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042109001967.

Williams:2005:CS

[Wil05] Kenneth S. Williams. The convolution sum ∑_{m<n/9} σ(m)σ(n − 9m). International Journal of Number Theory (IJNT), 1(2):193–205, June 2005. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042105000091.

Williams:2010:UIC

[Wil10] Gerald Williams. Unimodular integer circulants associated with trinomials. International Journal of Number Theory (IJNT), 6(4):869–876, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003289.

Williams:2012:FSC

[Wil12] Kenneth S. Williams. Fourier series of a class of eta quotients. International Journal of Number Theory (IJNT), 8(4): 993–1004, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500595.

Williams:2019:VVE

[Wil19] Brandon Williams. Vector-valued Eisenstein series of small weight. International Journal of Number Theory (IJNT), 15(2): 265–287, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500118.

Wittmann:2007:PDC

[Wit07] Christian Wittmann. l-parts of divisor class groups of cyclic function fields of degree l. International Journal of Number Theory (IJNT), 3(2):171–190, June 2007. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042107000857.

Witno:2010:EPP

[Wit10] Amin Witno. On elite primes of period four. International Journal of Number Theory (IJNT), 6(3):667–671, May 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042110003149.

Witte:2020:NCI

[Wit20] Malte Witte. Non-communtative Iwasawa main conjecture. International Journal of Number Theory (IJNT), 16(09):2041– 2094, October 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120501067.

Wang:2020:RSR

[WJ20] Yujie Wang and Chungang Ji. Ramanujan’s sum in the ring of integers of an algebraic number field. International Journal of Number Theory (IJNT), 16(1):65– 76, February 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042120500037.

Jiang:2021:NLB

[wJShY21] Xing wang Jiang, Csaba Sándor, and Quan hui Yang. A note on the lower bound of representation functions. International Journal of Number Theory (IJNT), 17(10): 2243–2250, November 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304212150086X.

Wang:2022:DMT

[WLWY22] Weiliang Wang, Lu Li, Xiujian Wang, and Qin Yue. Doubly metric theory and simultaneous shrinking target problem in Cantor series expansion. International Journal of Number Theory (IJNT), 18(08):1807– 1821, September 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500920.

Wong:2017:VHC

[Won17] Peng-Jie Wong. A variant of Heilbronn characters. International Journal of Number Theory (IJNT), 13(6):1547–1570, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117500865.

Wong:2018:AGT

[Won18] Peng-Jie Wong. Applications of group theory to conjectures of Artin and Langlands. International Journal of Number Theory (IJNT), 14(3):881–898, April 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500537.

Wong:2020:UFE

[Won20] Tian An Wong. Universal Fourier expansions of Bianchi modular forms. International Journal of Number Theory (IJNT), 16(4):731–746, May 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500372.

Wu:2010:RGD

[WS10] Qingquan Wu and Renate Scheidler. The ramification groups and different of a compositum of Artin–Schreier extensions. International Journal of Number Theory (IJNT), 6(7): 1541–1564, November 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110003617.

Wang:2016:NRL

[WS16] Min Wang and Zhi-Hong Sun. On the number of representations of n as a linear combination of four triangular numbers. International Journal of Number Theory (IJNT), 12(6): 1641–1662, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116501001.

Wang:2017:NRL

[WS17] Min Wang and Zhi-Hong Sun. On the number of representations of n as a linear combination of four triangular numbers II. International Journal of Number Theory (IJNT), 13(3):593–617, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117500300.

Wang:2019:DRF

[WS19] Chen Wang and Zhi-Wei Sun. Divisibility results on Franel numbers and related polynomials. International Journal of Number Theory (IJNT), 15(2):433–444, March 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119500222.

Sun:2020:QRQ

[wS20a] Zhi wei Sun. Quadratic residues and quartic residues modulo primes. International Journal of Number Theory (IJNT), 16(08): 1833–1858, September 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500955.

Wu:2020:ABS

[WS20b] Liuying Wu and Sanying Shi. On the asymptotic behavior of sums ∑_{n≤x} f(n){x/n}^{k}. International Journal of Number Theory (IJNT), 16(6):1263–1274, July 2020. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042120500657.

Wu:2017:NSA

[Wu17]  Han Wu.  A note on spectral anal-

ysis in automorphic representation theory for GL_{2}: I. International Journal of Number Theory (IJNT), 13(10):2717– 2750, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501500.

Wu:2018:AUM

[Wu18] Hai-Liang Wu. Almost universal mixed sums of squares and polygonal numbers. International Journal of Number Theory (IJNT), 14(8):2239–2256, September 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211850135X.

Wakhare:2021:SPM

[WV21] Tanay Wakhare and Christophe Vignat. Structural properties of multiple zeta values. International Journal of Number Theory (IJNT), 17(08):1873–1897, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042121500676.

Wei:2018:WTS

[WW18] Chuanan Wei and Xiaoxia Wang. Whippletype _{3}F_{2}-series and summation formulae involving generalized harmonic numbers. International Journal of Number Theory (IJNT), 14(9):2385–2407, October 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118501452.

Wang:2020:ASV

[WY20] Xiaoxia Wang and Mingbing Yue. A qanalogue of the (A.2) supercongruence of Van Hamme for any prime p ≡ 3 mod 4. International Journal of Number Theory (IJNT), 16(6):1325–1335, July 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500694.

Washington:2021:ARL

[WY21] Lawrence C. Washington and Ambrose Yang. Analogues of the Robin–Lagarias criteria for the Riemann hypothesis. International Journal of Number Theory (IJNT), 17 (04):843–870, May 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500111.

Weston:2009:PRF

[WZ09] Tom Weston and Elena Zaurova. Power residues of Fourier coefficients of elliptic curves with complex multiplication. International Journal of Number Theory (IJNT), 5(1): 109–124, February 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109001955.

Wang:2012:CNF

[WZ12] Q. H. Wang and J. J. Zhuang. On the critical number of finite groups of order pq. International Journal of Number Theory (IJNT), 8 (5):1271–1279, August 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112500741.

Wang:2016:NDC

[WZWQ16] Haoli Wang, Lizhen Zhang, Qinghong Wang, and Yongke Qu. Note on the Davenport constant of the multiplicative semigroup of the quotient ring \frac{F_{p}[x]}{⟨f(x)⟩}. International Journal of Number Theory (IJNT), 12(3):663–669, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500433.

Xia:2010:BNC

[XC10] Binzhou Xia and Tianxin Cai. Bernoulli numbers and congruences for harmonic sums. International Journal of Number Theory (IJNT), 6(4):849–855, June 2010. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042110003265.

Xia:2013:ICL

[Xia13]  Li-Meng Xia.  On the index-conjecture

on length four minimal zero-sum sequences. International Journal of Number Theory (IJNT), 9(6):1505–1528, September 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042113500401.

Xiong:2011:TII

[Xio11] Xinhua Xiong. Two identities involving the cubic partition function. International Journal of Number Theory (IJNT), 7(6):1615– 1626, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004757.

Xiong:2016:OFM

[Xio16] Xinhua Xiong. Overpartition function modulo 16 and some binary quadratic forms. International Journal of Number Theory (IJNT), 12 (5):1195–1208, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500731.

Xiong:2017:SVC

[Xio17]  Xinhua Xiong.  Small values of coeffi-

cients of a half Lerch sum. International Journal of Number Theory (IJNT), 13(9): 2461–2470, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042117501366.

Xiao:2022:SJI

[XLD22] Guanju Xiao, Lixia Luo, and Yingpu Deng. Supersingular j-invariants and the class number of Q(\sqrt{−p}). International Journal of Number Theory (IJNT), 18(05): 1065–1078, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500555.

Ma:2022:PRC

[xMgC22] Wu xia Ma and Yong gao Chen. On a problem on restricted k-colored partitions. International Journal of Number Theory (IJNT), 18 (03):467–472, April 2022. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042122500269.

Xia:2016:NRI

[XMT16] Ernest X. W. Xia, Y. H. Ma, and L. X. Tian. On the number of representations of integers by sums of mixed numbers. International Journal of Number Theory (IJNT), 12(4):945–954, June 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500585.

Xia:2014:ECS

[XTY14] Ernest X. W. Xia, X. L. Tian, and Olivia X. M. Yao. Evaluation of the convolution sum ∑_{i+25j=n} σ(i)σ(j). International Journal of Number Theory (IJNT), 10(6): 1421–1430, September 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042114500365.

Xu:2019:AER

[Xu19a]  Aimin Xu.  Asymptotic expansion re-

lated to the generalized Somos recurrence constant. International Journal of Number Theory (IJNT), 15(10):2043–2055, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042119501112.

Xu:2019:SEI

[Xu19b] Ce Xu. Some evaluations of infinite series involving parametric harmonic numbers. International Journal of Number Theory (IJNT), 15(7):1531–1546, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211950088X.

Xia:2011:DRG

[XY11] Ernest X. W. Xia and X. M. Yao. The 8-dissection of the Ramanujan–Göllnitz– Gordon continued fraction by an iterative method. International Journal of Number Theory (IJNT), 7(6):1589–1593, September 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042111004824.

Xia:2013:RIC

[XY13] Ernest X. W. Xia and Olivia X. M. Yao. On the representations of integers by certain quadratic forms. International Journal of Number Theory (IJNT), 9(1):189– 204, February 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501333.

Xia:2014:PKC

[XY14] Ernest X. W. Xia and Olivia X. M. Yao. A proof of Keith’s conjecture for 9-regular partitions modulo 3. International Journal of Number Theory (IJNT), 10(3):669–674, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113501169.

Xia:2019:PSC

[XY19] Ernest X. W. Xia and Zhang Yan. Proofs of some conjectures of Sun on the relations between sums of squares and sums of triangular numbers. International Journal of Number Theory (IJNT), 15(1):189– 212, February 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500088.

Xu:2017:EEQ

[XYZ17] Ce Xu, Yingyue Yang, and Jianwen Zhang. Explicit evaluation of quadratic Euler sums. International Journal of Number Theory (IJNT), 13(3):655–672, April 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117500336.

Xue:2021:LZC

[XZ21] Hui Xue and Daozhou Zhu. Location of the zeros of certain cuspforms. International Journal of Number Theory (IJNT), 17(08): 1899–1903, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500688.

Yamamoto:2015:SFM

[Yam15]  Shuji Yamamoto.  A sum formula of

multiple L-values. International Journal of Number Theory (IJNT), 11(1):127–137, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042115500074.

Yamagishi:2016:DAP

[Yam16] Shuntaro Yamagishi. Diophantine approximation of polynomials over F_{q}[t] satisfying a divisibility condition. International Journal of Number Theory (IJNT), 12(5): 1371–1390, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500846.

Yang:2013:GCT

[Yan13]  Quan-Hui Yang.  A generalization of

Chen’s theorem on the Erdős–Turán conjecture. International Journal of Number Theory (IJNT), 9(7):1683–1686, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042113500504.

Yan:2017:NSS

[Yan17] Qiming Yan. A note on Schmidt’s subspace theorem. International Journal of Number Theory (IJNT), 13(7):1639–1649, August 2017. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500944.

Yan:2019:PSN

[Yan19] Xiao-Hui Yan. Partitions of the set of nonnegative integers with identical representation functions. International Journal of Number Theory (IJNT), 15(10):1969–1975, November 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119501070.

Yao:2015:APF

[Yao15] Olivia X. M. Yao. Arithmetic properties for Fu’s 9 dots bracelet partitions. International Journal of Number Theory (IJNT), 11(4): 1063–1072, June 2015. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115500566.

Yao:2018:BTT

[Yao18]  Yanjun Yao.  Bombieri-type theorems

of nonlinear Weyl sums over primes in short intervals. International Journal of Number Theory (IJNT), 14(10):2571–2581, November 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042118501531.

Yasuda:2013:KGT

[Yas13] Masaya Yasuda. Kummer generators and torsion points of elliptic curves with bad reduction at some primes. International Journal of Number Theory (IJNT), 9(7): 1743–1752, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500541.

Yasuda:2015:RKE

[Yas15] Masaya Yasuda. Ramification of the Kummer extension generated from torsion points of elliptic curves. International Journal of Number Theory (IJNT), 11(6):1725–1734, September 2015. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115500736.

Yasuda:2016:FRM

[Yas16] Seidai Yasuda. Finite real multiple zeta values generate the whole space Z. International Journal of Number Theory (IJNT), 12(3):787–812, May 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042116500512.

Yang:2017:SCH

[YC17] Peng Yang and Tianxin Cai. Super congruence on harmonic sums. International Journal of Number Theory (IJNT), 13(10): 2569–2582, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501433.

Ye:2015:ECS

[Ye15a] Dongxi Ye. Evaluation of the convolution sums ∑_{l+36m=n} σ(l)σ(m) and ∑_{4l+9m=n} σ(l)σ(m). International Journal of Number Theory (IJNT), 11(1):171–183, February 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042115500104.

Ye:2015:RCB

[Ye15b] Dongxi Ye. Representations of certain binary quadratic forms as a sum of Lambert series and eta-quotients. International Journal of Number Theory (IJNT), 11(4): 1073–1088, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500578.

Ye:2016:QF

[Ye16] Dongxi Ye. On the quaternary form x^{2} + xy + 7y^{2} + z^{2} + zt + 7t^{2}. International Journal of Number Theory (IJNT), 12(7): 1791–1800, November 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116501104.

Yee:2009:BPT

[Yee09] Ae Ja Yee. Bijective proofs of a theorem of Fine and related partition identities. International Journal of Number Theory (IJNT), 5(2):219–228, March 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042109002043.

Yang:2015:UBL

[YF15] Hai Yang and Ruiqin Fu. An upper bound for least solutions of the exponential Diophantine equation D_{1}x^{2} − D_{2}y^{2} = λk^{z}. International Journal of Number Theory (IJNT), 11 (4):1107–1114, June 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042115500608.

Yang:2018:EDE

[YF18] Hai Yang and Ruiqin Fu. The exponential diophantine equation x^{y} + y^{x} = z^{2} via a generalization of the Ankeny–Artin–Chowla conjecture. International Journal of Number Theory (IJNT), 14(5):1223–1228, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042118500756.

Yang:2019:CSS

[YF19] Hai Yang and Ruiqin Fu. Complete solutions of the simultaneous Pell’s equations x^{2} − (a^{2} − 1)y^{2} = 1 and y^{2} − pz^{2} = 1. International Journal of Number Theory (IJNT), 15(5):1069–1074, June 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042119500593.

Kang:2021:DPF

[yK21] Soon yi Kang. Divisibility properties of the Fourier coefficients of (mock) modular functions and Ramanujan. International Journal of Number Theory (IJNT), 17 (03):577–589, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120400333.

Yuan:2016:LUZ

[YL16] Pingzhi Yuan and Yuanlin Li. Long unsplittable zero-sum sequences over a finite cyclic group. International Journal of Number Theory (IJNT), 12(4):979–993, June 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500615.

Young:2012:TTF

[You12]  Justin Young.  The twisted tensor L-

function of GSp_{4}. International Journal of Number Theory (IJNT), 8(2):411– 470, March 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042112500248.

You:2016:NRF

[You16a] Xu You. A note on Ramanujan’s factorial approximation. International Journal of Number Theory (IJNT), 12(5):1187– 1194, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S179304211650072X.

Young:2016:BPB

[You16b] Paul Thomas Young. Bernoulli and poly-Bernoulli polynomial convolutions and identities of p-adic Arakawa–Kaneko zeta functions. International Journal of Number Theory (IJNT), 12(5):1295–1309, August 2016. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042116500792.

Young:2017:PZF

[You17] Paul Thomas Young. Polylogarithmic zeta functions and their p-adic analogues. International Journal of Number Theory (IJNT), 13(10):2751–2768, November 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117501512.

Yu:2011:EMO

[Yu11] Chia-Fu Yu. On the existence of maximal orders. International Journal of Number Theory (IJNT), 7(8):2091–2114, December 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042111005003.

Yao:2012:NRI

[YX12] Olivia X. M. Yao and Ernest X. W. Xia. The number of representations of an integer as a sum involving generalized pentagonal numbers. International Journal of Number Theory (IJNT), 8(4):1041–1056, June 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112500613.

Yao:2014:ECS

[YX14] Olivia X. M. Yao and Ernest X. W. Xia. Evaluation of the convolution sums ∑_{i+3j=n} σ(i)σ_{3}(j) and ∑_{3i+j=n} σ(i)σ_{3}(j). International Journal of Number Theory (IJNT), 10(1):115–123, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042113500838.

Yao:2013:EFF

[YXJ13] Olivia X. M. Yao, Ernest X. W. Xia, and J. Jin. Explicit formulas for the Fourier coefficients of a class of eta quotients. International Journal of Number Theory (IJNT), 9(2):487–503, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S179304211250145X.

You:2022:RPI

[YZ22] Yuanyi You and Yichao Zhang. Rationality of the Petersson Inner Product of Cohen’s kernels. International Journal of Number Theory (IJNT), 18(07):1445–1463, August 2022. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042122500749.

Zaman:2016:LPI

[Zam16] Asif Zaman. On the least prime ideal and Siegel zeros. International Journal of Number Theory (IJNT), 12(8):2201– 2229, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116501335.

Zantout:2016:CHG

[Zan16]  Dania Zantout.  Cuspidality in higher

genus. International Journal of Number Theory (IJNT), 12(8):2043–2060, December 2016. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116300015.

Zhao:2017:AGM

[ZC17a] Xiao-Peng Zhao and Zhen-Fu Cao. Another generalization of Menon’s identity. International Journal of Number Theory (IJNT), 13 (9):2373–2379, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501299.

Zhong:2017:LPL

[ZC17b] Hao Zhong and Tianxin Cai. On the Lucas property of linear recurrent sequences. International Journal of Number Theory (IJNT), 13(6):1617–1625, July 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117500920.

Zheng:2022:GSD

[ZCX22] Zhiyong Zheng, Man Chen, and Jie Xu. On Gauss sums over Dedekind domains. International Journal of Number Theory (IJNT), 18 (05):1047–1063, June 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500543.

Zelinsky:2019:UBS

[Zel19] Joshua Zelinsky. Upper bounds on the second largest prime factor of an odd perfect number. International Journal of Number Theory (IJNT), 15(6):1183–1189, July 2019. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042119500659.

Zemel:2017:NCG

[Zem17] Shaul Zemel. Normalizers of congruence groups in SL_{2}(R) and automorphisms of lattices. International Journal of Number Theory (IJNT), 13(5):1275–1300, June 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042117500701.

Zhao:2020:ECP

[ZG20] Xiaodong Zhao and Wenxu Ge. Eight cubes of primes and 204 powers of 2. International Journal of Number Theory (IJNT), 16 (7):1547–1555, August 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500803.

Zhang:2016:NII

[ZH16a] Wenpeng Zhang and Di Han. A new identity involving the classical Kloosterman sums and 2-dimensional Kloosterman sums. International Journal of Number Theory (IJNT), 12 (1):111–119, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S179304211650007X.

Zhang:2016:WBL

[ZH16b] Zhizheng Zhang and Junli Huang. A WP– Bailey lattice and its applications. International Journal of Number Theory (IJNT), 12 (1):189–203, February 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500111.

Zhao:2016:CGG

[ZH16c] Zhengjun Zhao and Wanbao Hu. On l-class groups of global function fields. International Journal of Number Theory (IJNT), 12 (2):341–356, March 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042116500202.

Zhang:2005:EET

[Zha05] Liang-Cheng Zhang. Explicit evaluations of two Ramanujan–Selberg continued fractions. International Journal of Number Theory (IJNT), 1(4):593–601, December 2005. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042105000297.

Zhao:2008:WTT

[Zha08] Jianqiang Zhao. Wolstenholme type theorem for multiple harmonic sums. International Journal of Number Theory (IJNT), 4 (1):73–106, February 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001146.

Zhabitskaya:2012:CFO

[Zha12] E. N. Zhabitskaya. Continued fractions with odd partial quotients. International Journal of Number Theory (IJNT), 8(6):1541– 1556, September 2012. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042112500947.

Zhao:2014:LBC

[Zha14] Feng-Zhen Zhao. The log-behavior of the Catalan–Larcombe–French sequence. International Journal of Number Theory (IJNT), 10 (1):177–182, February 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042113500887.

Zhang:2015:EBP

[Zha15a]  Wen-Bin Zhang.  Extensions of Beurl-

ing’s prime number theorem. International Journal of Number Theory (IJNT), 11(5): 1589–1616, August 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042115400126.

Zhang:2015:ESS

[Zha15b] Zhen-Liang Zhang. On the exceptional sets in Sylvester continued fraction expansion. International Journal of Number Theory (IJNT), 11(8):2369–2380, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115501092.

Zhang:2016:DPW

[Zha16] Yichao Zhang. Divisibility properties for weakly holomorphic modular forms with sign vectors. International Journal of Number Theory (IJNT), 12(8):2107–2123, December 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042116501268.

Zhang:2017:RFT

[Zha17a] Bin Zhang. Remarks on the flatness of ternary cyclotomic polynomials. International Journal of Number Theory (IJNT), 13 (2):529–547, March 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501354.

Zhang:2017:LCT

[Zha17b] Qing Zhang. A local converse theorem for U(1, 1). International Journal of Number Theory (IJNT), 13(8):1931–1981, September 2017. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501056.

Zhang:2017:SRS

[Zha17c] Qing Zhang. Stability of Rankin–Selberg gamma factors for Sp(2n), S̃p(2n) and U(n, n). International Journal of Number Theory (IJNT), 13(9):2393–2432, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042117501317.

Zhang:2017:DE

[Zha17d] Zhongfeng Zhang. On the Diophantine equation (x − d)^{4} + x^{4} + (x + d)^{4} = y^{n}. International Journal of Number Theory (IJNT), 13 (9):2229–2243, October 2017. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042117501214.

Zhang:2021:IES

[Zha21] Mengjie Zhang. On the intersections of exceptional sets in Borel’s normal number theorem and Erdős–Rényi limit theorem. International Journal of Number Theory (IJNT), 17(04):959–972, May 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042121500172.

Zhang:2022:FGS

[Zha22a] Helen W. J. Zhang. Further generalizations of some Hecke-type identities. International Journal of Number Theory (IJNT), 18 (02):361–387, March 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500221.

Zhao:2022:GDR

[Zha22b]  Alice X. H. Zhao.  Generalizations of

Dyson’s rank on overpartitions. International Journal of Number Theory (IJNT), 18 (04):743–755, May 2022. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042122500397.

Zhelezov:2015:IBA

[Zhe15] Dmitrii Zhelezov. Improved bounds for arithmetic progressions in product sets. International Journal of Number Theory (IJNT), 11(8):2295–2303, December 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042115501043.

Zhou:2018:PHO

[Zho18] Nian Hong Zhou. Primes in higher-order progressions on average. International Journal of Number Theory (IJNT), 14(7): 1943–1959, August 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118501166.

Zhou:2021:NPP

[Zho21] Nian Hong Zhou. Note on partitions into polynomials with number of parts in an arithmetic progression. International Journal of Number Theory (IJNT), 17(09):1951–1963, October 2021. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042121500718.

Zhu:2018:BAP

[Zhu18] Wenbin Zhu. A binary additive problem with fractional powers. International Journal of Number Theory (IJNT), 14(1): 221–232, February 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042118500136.

Zhu:2022:DAP

[Zhu22] Guangyan Zhu. On the divisibility among power GCD and power LCM matrices on gcdclosed sets. International Journal of Number Theory (IJNT), 18(07):1397–1408, August 2022. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042122500701.

Zieve:2008:SFP

[Zie08] Michael E. Zieve. Some families of permutation polynomials over finite fields. International Journal of Number Theory (IJNT), 4 (5):851–857, October 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042108001717.

Ziegler:2011:AUS

[Zie11] Volker Ziegler. The additive S-unit structure of quadratic fields. International Journal of Number Theory (IJNT), 7(3):635– 644, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042111004216.

Zhao:2019:ZTA

[ZJ19] Zhengjun Zhao and Qingzhong Ji. Zsigmondy theorem for arithmetic dynamics induced by a Drinfeld module. International Journal of Number Theory (IJNT), 15(6):1111–1125, July 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119500611.

Zhang:2018:WGP

[ZL18a]  Min Zhang and Jinjiang Li.  Waring–

Goldbach problem for unlike powers with almost equal variables. International Journal of Number Theory (IJNT), 14(9):2441– 2472, October 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042118501488.

Zhang:2018:MTF

[ZL18b] Zhizheng Zhang and Xiaoqian Li. Mock theta functions in terms of q-hypergeometric double sums. International Journal of Number Theory (IJNT), 14(6):1715–1728, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042118501051.

Zhang:2019:DIF

[ZL19]  Min Zhang and Jinjiang Li.  A Dio-

phantine inequality with four prime variables. International Journal of Number Theory (IJNT), 15(9):1759–1770, October 2019. CODEN ???? ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119500982.

Zou:2016:CES

[ZLL16] Yuru Zou, Wenxia Li, and Jian Lu. On the cardinality of β-expansions of some numbers. International Journal of Number Theory (IJNT), 12(6):1497–1507, September 2016. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042116500925.

Zhao:2010:EDP

[ZLZ10] Yusheng Zhao, Wei Li, and Xianke Zhang. Effective determination of prime decompositions of cubic function fields. International Journal of Number Theory (IJNT), 6 (2):437–448, March 2010. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042110002983.

Zhang:2013:MVI

[ZLZ13] Deyu Zhang, Meimei Lü, and Wenguang Zhai. On the mean value of the index of composition of an integer II. International Journal of Number Theory (IJNT), 9 (2):431–445, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042112501424.

Zhong:2015:MPE

[ZMS15] Ting Zhong, Quanwu Mu, and Luming Shen. Metrical property for GCF_{ϵ} expansion with the parameter function ϵ(k) = c(k + 1). International Journal of Number Theory (IJNT), 11(7):2065–2072, November 2015. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S179304211550089X.

Zoeteman:2019:UCP

[Zoe19] M. Zoeteman. Uniformly counting primes with a given primitive root and in an arithmetic progression. International Journal of Number Theory (IJNT), 15(10):2115– 2134, November 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042119501161.

Zorn:2011:EDI

[Zor11] Christian Zorn. Explicit doubling integrals for Sp_{2}(F) and S̃p_{2}(F) using “good test vectors”. International Journal of Number Theory (IJNT), 7(2):449–527, March 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042111004046.

Zorin:2013:ZOE

[Zor13] Evgeniy Zorin. Zero order estimates for analytic functions. International Journal of Number Theory (IJNT), 9(2):333– 392, March 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042112501370.

Zhu:2018:HTD

[ZS18] Bao-Xuan Zhu and Zhi-Wei Sun. Hankeltype determinants for some combinatorial sequences. International Journal of Number Theory (IJNT), 14(5):1265–1277, June 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

                                        /www.worldscientific.com/doi/10.1142/
                                        S1793042118500793.

Zhang:2020:SFH

[ZS20] Zhizheng Zhang and Hanfei Song. Some further Hecke-type identities. International Journal of Number Theory (IJNT), 16(09): 1945–1967, October 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120501006.

Zhong:2013:SDC

[ZT13] Ting Zhong and Liang Tang. The sets of different continued fractions with the same partial quotients. International Journal of Number Theory (IJNT), 9(7):1855– 1863, November 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042113500619.

Zubrilina:2020:LRP

[Zub20] Nina Zubrilina. Left–right pairs and complex forests of infinite rooted binary trees. International Journal of Number Theory (IJNT), 16 (7):1637–1642, August 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120500864.

Zuevsky:2018:CPE

[Zue18]  Alexander Zuevsky.  Complex powers

of eta-function. International Journal of Number Theory (IJNT), 14(6):1619– 1625, July 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042118500975.

Zumalacarregui:2011:CPM

[Zum11] Ana Zumalacárregui. Concentration of points on modular quadratic forms. International Journal of Number Theory (IJNT), 7(7): 1835–1839, November 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111004897.

Zhang:2014:EGZ

[ZW14a] Haiyan Zhang and Guoqing Wang. On the Erdős–Ginzburg–Ziv invariant and zero-sum Ramsey number for intersecting families. International Journal of Number Theory (IJNT), 10(7):1637–1647, November 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/
S1793042114500481.

Zhang:2014:RTC

[ZW14b] Wenlong Zhang and Chun Wang. Ramanujantype congruences for several partition functions. International Journal of Number Theory (IJNT), 10(3):641–652, May 2014. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042113501145.

Zhang:2008:ACS

[ZY08] Lingrui Zhang and Qin Yue. Another case of a Scholz’s theorem on class groups. International Journal of Number Theory (IJNT), 4(3):495–501, June 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042108001493.

Zhang:2012:DE

[ZY12] Zhongfeng Zhang and Pingzhi Yuan. On the Diophantine equation ax^{y} + by^{z} + cz^{x} = 0. International Journal of Number Theory (IJNT), 8(3):813–821, May 2012. ISSN 1793-0421 (print), 1793-7310 (electronic).

                                        URL https://www.worldscientific.com/
                                        doi/10.1142/S1793042112500467.

Zhang:2020:PTD

[ZY20]  Yong Zhang and Peisen Yuan.  Proof

of three divisibilities of Franel numbers and binomial coefficients. International Journal of Number Theory (IJNT), 16(7): 1535–1546, August 2020. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

                                        //www.worldscientific.com/doi/10.1142/
                                        S1793042120500797.

Zywina:2011:RKC

[Zyw11] David Zywina. A refinement of Koblitz’s conjecture. International Journal of Number Theory (IJNT), 7(3):739–769, May 2011. ISSN 1793-0421 (print), 1793-7310 (electronic).

URL https://www.worldscientific.com/
doi/10.1142/S1793042111004411.

Zhang:2011:FPM

[ZZ11] Deyu Zhang and Wenguang Zhai. On the fifth-power moment of Δ(x). International Journal of Number Theory (IJNT), 7 (1):71–86, February 2011. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042111003922.

Zeilberger:2021:ADI

[ZZ21] Doron Zeilberger and Wadim Zudilin. Automatic discovery of irrationality proofs and irrationality measures. International Journal of Number Theory (IJNT), 17(03): 815–825, April 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:

//www.worldscientific.com/doi/10.1142/
S1793042120400230.